A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1970–1974

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.10

Title word cross-reference

#5824 [1784].

(3 + 1) [1104]. (: ϕ^{4}: )_{1+1} [1573, 1574]. (∇^{2})ϕ = −ϕ − ϕ^{3} [432]. (∂/∂t+a(t, x, y)∂/∂x+b(t, x, y)∂/∂y+c(t, x, y)+d(t, x, y)∂^{2}/∂x∂y)u = f(t, xy) [251]. (∂^{2}/(∂x∂y) + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0 [84, 155]. (∂^{2}/∂x∂y + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0 [108]. (φ^{4})_{3} [1243]. + [290]. −f^{2}/r^{4} [1126]. 0 [290]. 1 [290]. 1/2 [440, 1068, 1374]. 12 [1664]. 15 [1664]. 1 ≤ m ≤ n [1754]. 2 × N [828, 1502, 1457, 381]. 3 [1672]. 3 + 2 [566, 544]. 6 [925]. 9 [1148, 805, 1297]. ^{−} [1129]. ^{4} [234]. _{2} [868, 861]. _{3} [1395]. _{3}F_{2} [763]. sY_{lm}(θ, ϕ) [682]. A_{n} [68]. α [714]. B_{n} [68]. C∗ [962, 1644, 1645, 1646]. C^{∞} [1232]. C_{n} [68]. : ϕ^{4}: [272]. D [400, 1198]. D^{j} [909]. D_{4} [1124]. D_{n} [68]. δ [772]. E(2) [436]. E2 [200]. f [364, 757, 1773, 1441]. F_{4} [486]. F_{D} [1010]. f_{tt} − f_{ss} = γ^{2}f [1780, 1598]. g [1773, 1441]. G(0, 1) [404]. G_{2} [372]. Γ_{jk;l}^{i} = 0, g_{ij;k} = 0 [926]. GL(n) [729]. GL(N) ⊃ Σ_{m} [1465]. H [380, 1054]. H = −(1/2)d^{2}/dx^{2} + m^{2}x^{2}/2 + g/x^{2} [1624]. I [420].

I^{m}[S_{p}(n) ⊗ S_{p}(m)] ⇒ S_{p}(n, m) [1754]. I^{m}[SO(n) ⊗ SO(m)] ⇒ SO(n, m) [1754]. I^{m}[U(n) ⊗ U(m)] ⇒ U(n, m) [1754]. ISO_{n} [462]. IU_{n} [1043]. iU_{t} + U_{xx} − c/x^{2}U = 0 [1699]. iU_{t} + U_{xx} = 0 [1699]. IUSp(2ν_{1}, 2ν_{2}) [226]. J [615, 1672, 1148, 1664, 805, 1297, 925]. k [1713]. k_{ν} [333]. k ≥ 4 [1713]. L [695, 1626]. L^{2} [1089, 523, 803]. L^{∞} [1745]. λ: ϕ_{2}^{4}(x): [1080, 1079]. λφ_{2}^{4} [1036]. λx^{2m} [1290]. μ [1751]. N [185, 490, 1123, 1784, 565, 822, 917, 1041, 767, 1228, 874, 1155, 1030, 401, 1329, 1157, 1212, 1719, 243, 407]. N/D [16, 17, 327]. N/V [1436, 1649]. O [1124]. O(1, 1) ⊗ O(2) [995]. O(2, 1) [555, 475, 1410, 267]. O(2, 1) × O(3) [964]. O(3) [1708, 734]. O(3, 1) [89]. O(3, 1) ⊃ O(2, 1) ⊃ O(1, 1) [497]. O(4) [290, 89, 627, 1414]. O(5) [137, 152, 635, 1559]. O(n) [729, 895]. P (φ)_{2} [1230]. P (x, y, t = 0) = Φ(x, y) [84]. ∂t^{2}f − Δf = 0 [130]. ϕ^{4} [1446]. ϕ^{4}: i + 1 [721]. πN → γN [615]. πN → πB_{s} [578]. πN → πN [615]. ∏_{i=1}^{N}(∇^{2} + k_{i}^{2})Ψ = 0 [1242]. Q(4, 2) [361]. R [1683, 1684]. R_{jkl}^{i} ≠ 0 [926]. R_{4} [317]. R_{5} [511, 879]. S [429, 1045, 637, 520, 73, 121, 1505, 523, 1482, 1626, 1629, 682, 1700]. S_{N} [1157]. SL(2, C) [200, 132, 812, 336, 1363, 1344, 538, 539]. SL(3, C) [1219, 224]. SL(n, C) [138]. SL(n, C) · R(n^{2}) [354]. SO(2, 1) [1203, 1633]. SO(2, C) [1533]. SO(3) [1633]. SO(3, 2) [361]. SO(4, 1) [361]. SO(4, 2) [537]. SO(5) [1166]. SO(N) [1429, 728, 649]. SO(n, 1) [1421]. SO(p, 1) [1528, 659]. SO(p, 1) ⊃ SO(p − 1, 1) [724, 1183]. SO(p, q) [671, 1725]. SO^{+}(n, 1) [171]. SO_{0}(n, 1) [451]. SO_{0}(p, 1) [1195]. SO_{0}(p, q) [209]. SO_{0}(p − 1, q − 1)[multiplicationinrighthalfcircle]T_{p+q−2} [209]. SO_{n} [462]. SO_{n,1} [462]. SO_{n,1} ⊃ SO_{1,1} ⊗ SO_{n−1} [1514]. Sp(4) [1167, 1166]. SP(n) [729]. Sp(n + 2) [186]. SU(1, 1) [1408, 267, 1488]. SU(1, 1) ⊗ SU(2) [821]. SU(2) [296, 1408]. SU(2, 1) [1648]. SU(2, 2) [821, 481]. SU(3) [985, 1708, 122, 1302, 1259, 224, 240, 518, 1427]. SU(3) ⊗ SU(2) ⊂ SU(6) [1623]. SU(4) [1167]. SU(4) ⊃ SU(2) ⊗ SU(2) [402]. SU(6) [1623]. SU(6) ⊃ SU(3) ⊗ SU(2) [240]. SU(n) [1138, 889, 1028, 1198, 817, 491]. SU(n − 2) × SU(2) × U(1) [817]. SU(p, q) [671, 1648]. T [597, 768, 755, 178, 1628, 1778, 627, 1414, 1154, 1340, 1626, 1341, 1127]. T(2n)\frac{←}{+}Sp(n) [186]. U [1111, 772]. U(3) [1094, 1095, 1009]. U(4) ⊃ U(2) ⊗ U(2) [184]. U(k) [931]. U(m) [752]. u(n) [458, 1759, 881, 1604, 301, 1157]. U(n) ⊃ SO(n) [183]. U(N) ⊃ U(M) [1216]. U(n + m) ↓ U(n) × U(m) [346, 766]. U(p, 1) ⊃ U(p) [1141]. U_{2n} [28]. U_{2n} ⊃ U_{n}+U_{n} [28]. U_{n,1} [1043]. USP (2ν_{1} + 2, 2ν_{2}) [226]. USp(2ν_{1}, 2ν_{2} + 2) [226]. V [1111]. V θ [958]. X [1214]. X − Y [1233]. XY [440]. Y [1214]. Z [1219]. |ϕ|^{2N} [600].

-algebraic [962, 1644, 1645, 1646]. -body [1212, 490, 767, 1123, 1784, 1155]. -calculation [138]. -dependent [1089]. -dimensional [364, 757, 185, 874]. -fold [565]. -formulation [537]. -function [772]. -functions [380]. -invariant [171]. -matrix [520, 1505, 1684, 1482, 755, 178]. -operator [1219]. -Parametrization [200]. -particle [243, 917]. -perturbations [1232]. -point [407, 1030]. -representability [822, 401]. -representable [1041]. -soliton [1228]. -space [1751]. -spin [420]. -Symbol [925]. -Theorem [1054]. -type [1198, 17]. -wave [1700, 695, 73].

1 [515]. 1-D [515]. 11 [857]. 12 [1784]. 13 [1785]. 14 [1466]. 1D [1524].

2- [1041, 419]. 2-Dimensional [305]. 2-Form [662].

3 [775]. 3-J [775]. 3-parameter [419]. 3-Space [920].

43 [1784].

6 [775]. 6-J [775].

Abnormal [635, 672, 152]. Absence [1675, 25, 1056, 1226, 1614, 1552, 954]. absolute [559]. absorption [337]. absorptive [727]. Abstract [569, 1135]. AC [562]. Accelerated [585]. Accelerating [707]. Acceleration [944]. accelerations [434]. accelerator [499]. accidental [1208, 1209]. According [228]. Accuracy [50, 270]. accurate [700]. Action [809, 1512, 1692]. action-at-a-distance [1692]. Actions [845]. Adapted [628, 1157]. Addendum [833, 636, 1614, 1537, 1311]. Addition [5]. additively [158]. Adiabatic [50, 748, 412, 496, 1190, 166]. adjacent [1098]. Adjoint [602, 301, 649, 825]. adjointness [1574]. admit [1621, 1566, 199]. Admitting [765, 14, 1295, 1131]. advance [1474]. advanced [1753, 1779]. Advection [722]. Agranovitch [1358]. Albedo [220]. Algebra [873, 510, 1484, 937, 1235, 872, 1429, 88, 1572, 1436, 1708, 1625, 1449, 1649, 1613, 1754, 1121]. Algebraic [1039, 143, 769, 1055, 735, 958, 1335, 367, 1246, 829, 1488, 962, 1078, 1644, 1645, 1646]. Algebraically [1142, 1484, 1609, 1511]. Algebras [873, 764, 819, 676, 916, 186, 376, 628, 887, 1354, 1386, 1656, 575, 1523, 1152, 495, 231, 921, 922, 1617, 158, 848, 1044, 226, 1632, 744, 677]. Alisauskas [1297]. Almost [505, 831]. along [386]. Alternative [76]. Amati [1464]. ambiguity [1356]. among [1184, 1207, 862]. Amplitude [969, 349, 1371, 223, 1657, 1600, 1601]. Amplitudes [615, 181, 142, 164, 513, 39, 64, 1356, 135, 520, 1713, 1515, 344, 475, 971, 727, 1718, 654, 716, 717]. Analogs [38]. Analysis [873, 230, 156, 620, 146, 325, 945, 872, 924, 232, 1079, 1516, 1493, 1173, 127, 1144, 1246]. Analytic [909, 308, 476, 956, 557, 934, 961, 1118, 667, 919, 661, 594, 1463, 220, 504, 1255, 1035, 1356, 945, 1731, 153, 1371, 1300, 512, 993, 1349, 1636, 1264, 1602, 1420, 1280, 1626, 1627, 1629, 1775, 997, 1161, 515]. Analytical [106, 860]. Analyticity [223, 1461, 547, 755, 1084, 1626]. and/or [490, 1654, 1784]. angles [874]. Angular [531, 437, 876, 796, 594, 858, 859, 1610, 418]. Anharmonic [999, 100, 1290, 1130, 1549]. Anholonomic [275]. anisotropic [1248, 1721, 1117, 1403, 198, 1107, 1208, 280, 58]. Anisotropy [41]. annular [522]. Anomalous [1133, 1479]. antifermion [500]. Antiferromagnetic [57, 58]. antiunitary [847, 1695]. Any [1023, 496]. Appell [203]. applicability [1333]. applicable [380]. Application [714, 800, 795, 125, 1662, 480, 222, 1323, 1359, 1360, 300, 441, 1586, 1774, 208, 254, 1185, 1681, 1148, 1171, 775, 258, 315, 1694, 442, 1525, 1179, 873]. Applications [986, 995, 1081, 1501, 324, 1585, 814, 319, 1378, 1101, 1458, 21, 629, 1222, 1553, 205, 278, 896, 712]. applied [1298, 996, 893, 1524, 51, 1623]. Approach [849, 584, 457, 289, 571, 816, 439, 362, 778, 1366, 1548, 1313, 1146, 877, 1210, 1508, 1423, 1697, 516, 1498, 1664, 374, 1124, 1545, 860, 1477, 1064, 823, 1488, 580, 389, 1519, 910, 805, 1085, 1120, 711, 712]. approaches [1500, 1676]. approximant [1145]. approximants [1146, 1134, 1288]. Approximate [317, 714, 265, 478, 1048, 1413, 270, 551, 902, 1630, 359]. approximately [775]. Approximation [241, 707, 124, 1014, 762, 1234, 254, 993, 1349, 1500, 1323, 1359, 1360, 1732, 893, 941, 1180, 579, 243]. Approximations [55, 751, 69, 558, 314, 448, 844]. arbitrarily [1346, 1347, 1262]. Arbitrary [944, 504, 536, 26, 1263, 182, 434, 664, 1666, 1688, 1551, 1539, 1015, 942]. Arbitrary-Spin [536]. Argument [173]. Arising [815, 159]. array [1088, 1502]. arrays [1457, 381]. Ashkin [1389]. aspect [1335]. Aspects [484, 321, 1709, 1489]. assemblies [1518]. Associated [873, 1060, 1074, 872, 1392, 1149]. assume [1689]. asymmetric [619]. Asymptotic [1387, 888, 527, 1143, 802, 742, 552, 30, 1433, 44, 1729, 1747, 698, 1042, 172, 443, 831, 798, 1368, 1237, 26, 489, 929, 1361, 222, 314, 1258, 816, 818, 442, 1058]. Asymptotically [777, 1652, 950, 832, 1074, 156, 607, 1492, 1082]. asymptotics [914]. atmosphere [337]. atmospheres [1240, 614]. atom [639, 796, 31, 179, 548, 1162, 1163, 1743, 742, 1024, 768, 315, 654, 1504]. atom-rigid [1024]. Atomic [343, 1387, 161, 704]. atomicity [1047]. Atoms [335, 700]. Attractive [294, 392, 1700]. Automorphisms [510, 1656]. autonomous [1021]. Average [770, 159]. Averages [1011, 648, 774]. axial [994]. axially [1650, 1777, 1032]. Axiom [1003]. axiomatic [1047, 1337, 365, 1602]. axiomatics [1769]. Axioms [764, 918].

b.c.c [680]. background [1100]. Bäcklund [1760, 1383]. backscattering [1529]. Baker [1752]. Balescu [1115]. Bargmann [1664, 805]. Barnes [1186, 546]. barrier [963]. Baryon [615]. based [1238]. Bases [27, 28, 676, 817, 137, 677, 275]. Basic [1046]. Basis [769, 457, 995, 400, 1009, 2, 525, 628, 538, 261, 812, 566, 1234, 689, 1363, 1203, 1257, 633, 1106, 1200, 1438, 766, 224, 1410, 1642, 1643, 1685, 1686, 1278, 1503, 931, 539]. Basis-Dependent [457]. Basis-Independent [457]. bcc [561, 1015]. Be [453, 255, 1539]. beam [1301, 1282, 1110]. Behavior [944, 888, 802, 1619, 900, 563, 570, 1387, 987, 1390, 1723, 1258, 1729, 818, 1368]. Bel [1120]. Bertocchi [1464]. Bertotti [1768]. Bessel [460, 550]. best [1264]. Bethe [635, 857, 152, 639, 1187, 90, 54, 500, 1055, 651, 656]. between [1060, 89, 1702, 340, 1027, 731, 321, 484, 583, 1199, 16, 67, 1405, 1438, 523, 932, 189, 504, 1069, 15]. biased [562]. Biaxial [859]. Biedenharn [1333]. Bifurcate [454]. bifurcation [1199, 315]. biharmonic [1103]. bilinear [968]. Binary [590, 1592]. binding [371]. binomial [4, 282]. binomial-coefficient [282]. biorthogonal [1515]. Birkhoff [113]. Bivector [765]. black [691, 1673, 1577, 820, 870, 1334]. Bloch [991, 1458, 1745]. Block [291]. bodies [20]. Body [308, 490, 142, 767, 531, 86, 687, 276, 653, 332, 470, 471, 85, 778, 1123, 1654, 1784, 1170, 1212, 1678, 629, 467, 1666, 1155, 746, 1325, 1467, 1653]. body-centered [629, 467, 1666]. Boltzmann [1151, 29, 493, 509]. Bonded [394]. Bondi [1072]. Born [1033, 69, 1618]. Bose [1424, 669, 670, 1364, 1451, 622, 737, 665, 1766]. Boson [122, 1153, 1548, 1714, 696, 1405, 1318, 1196]. boson-fermion [1548]. bosons [559, 1588, 1257]. both [1408, 1474]. Bound [1014, 192, 798, 1387, 1123, 216, 756, 500, 1557, 996, 318, 1385, 1711, 727, 1078, 448, 844]. Bound-State [1014, 216]. boundaries [1640, 701, 846]. Boundary [586, 257, 374, 830, 459, 1248, 195, 190, 35, 713, 1637, 955, 1636, 686, 1478, 106, 923]. boundary-initial [713]. Boundary-Value [586, 257, 374, 35, 955, 106]. Bounded [761, 784, 1682, 1388]. bounding [1145]. Bounds [1443, 1738, 3, 577, 943, 770, 1076, 814, 1, 642, 314, 1670, 824, 825, 1403]. Box [809, 834, 162]. bra [1326]. Bracket [219]. Brackets [640]. Branching [183, 1216, 197, 908, 1465, 187, 409, 395]. Brans [551, 902, 1630]. breakdown [1671]. Breaking [230, 208, 960, 537, 1289]. Breit [135, 1515]. Breit-system [135]. Brillouin [1509]. Broglie [398, 650]. Bronzan [1111]. Brownian [1518]. BS [290]. Bubble [912]. Bubble-Time [912]. Bulk [3, 577].

calculated [489]. calculating [1109, 384]. Calculation [1251, 561, 399, 107, 56, 395, 925, 1428, 829, 774, 1519, 138]. Calculations [815, 1631]. calculus [1094, 1238, 1476, 696, 1405, 533, 910, 441]. Campbell [1752]. Can [453, 1673, 1539]. Cancellation [1557]. Cannot [255]. Canonical [1298, 681, 457, 72, 912, 907, 301, 1208, 1209, 935, 685, 249, 1206, 342, 526, 330, 217, 348, 1639, 1755, 937, 1094, 1095, 135, 10, 799, 848, 12, 1318, 684, 1272, 1375, 952, 278]. Canonically [286]. Canonization [747]. capacity [1428]. Cartesian [126, 202]. case [551, 836]. cases [176]. Casimir [236, 1604]. Cauchy [1731, 884, 826]. Causal [1640, 414, 1174]. Causality [1567, 345]. Cavity [1071, 204, 328, 820]. Cell [410, 1669, 529, 1177, 1756]. Center [815, 135]. center-of-mass- [135]. centered [1544, 629, 467, 1666, 1762]. Centers [1007]. Central [1222, 798, 820]. centrally [1388]. Certain [925, 1076, 391, 380, 1147, 1256, 648, 1760, 1404, 801, 1595, 818, 1293, 147, 448, 844]. Chain [28, 201, 1514, 27, 641, 1519]. Chains [673, 1520, 756, 996, 1057, 1137, 1586]. Change [782]. Channel [177, 887, 1366]. chaotic [1119]. character [1069]. Characteristic [729, 1225]. characteristics [1728]. characterization [960]. characterized [461, 1095]. Characters [326, 1287]. Charge [449, 12, 252, 1023, 691, 499, 420, 1526, 1765, 1783]. Charged [123, 1007, 348, 632, 1156, 1693, 1647, 1558, 408, 902, 631, 1530, 596, 654, 1375, 604]. Charges [1296, 537]. chemistry [1506]. Chiang [679]. chiral [93]. Choquet [1740]. Circular [204, 1229, 1680]. Class [806, 927, 661, 431, 517, 353, 570, 866, 965, 1251, 361, 380, 575, 1594, 835, 262, 892, 826, 1124, 1258, 733, 523, 1293, 1576, 1499, 1629]. Classes [130, 744]. Classical [773, 219, 379, 534, 228, 482, 807, 864, 92, 292, 213, 363, 918, 748, 1427, 438, 1386, 94, 1170, 1229, 1147, 552, 236, 1556, 227, 1659, 774, 109, 1748, 1373, 1277, 928, 1489, 187, 165]. Classification [1169, 284, 610, 1044, 249, 1663, 1523]. Clebsch [132, 133, 1343, 233, 202, 400, 1302, 894, 920, 1642, 1643, 1685, 1686, 1744, 1694, 1695, 267]. Clifford [785]. Closed [1546, 304, 941, 1129, 155, 1284]. Closed-Form [155, 1546, 1129]. closure [669]. Cluster [1424, 1204, 48, 1312, 772, 1690]. coarse [1644, 1645, 1646]. coated [18]. coaxial [1680]. codons [1716]. Coefficient [142, 4, 1500, 282, 225]. Coefficients [881, 400, 894, 920, 377, 396, 591, 132, 133, 1343, 1251, 233, 1672, 1273, 520, 1424, 1395, 929, 436, 1759, 1166, 1559, 1664, 458, 1302, 763, 1149, 1642, 1643, 1685, 1686, 1198, 1744, 1694, 1695, 772, 1196, 267, 805, 1297]. Coherence [37, 1060]. Coherent [1717, 1617, 1634, 1190]. cohomology [1615]. coinciding [365]. Colineation [1485]. Collective [668]. Colliding [837]. Collineations [307, 103, 199]. collision [941, 1024, 493]. collisions [1299, 418, 654, 1129]. Colorings [98, 385]. comb [701, 846, 499]. Combinatorial [491, 1138]. Combined [781]. Comment [1786, 679, 1443]. Comments [944, 1466]. Commutation [312, 313, 583, 249, 10, 799, 1632]. commutator [1468]. Commutators [1052, 77, 514, 223, 407]. Commuting [848, 289, 1708]. Commuting-Operator [289]. compact [393, 1603, 1625, 1172, 1355, 1305]. compactness [1121]. Comparison [333, 1500, 1376, 1764, 1373]. compensation [1714]. Complementary [466]. Complete [393, 531, 1708, 899, 1514, 793, 1655, 560]. Completely [456, 1763]. Completeness [1155, 832, 1410]. Complex [284, 662, 1674, 1650, 1777, 605, 963, 1221, 305, 1060, 1430, 1416, 594, 251, 1113, 1415, 1639, 1755]. Complexification [1609]. complexified [1105]. component [1249, 171, 492]. Components [583, 1011, 636, 256, 430]. Composite [86, 687, 1342]. Composition [1634]. compound [1671]. compressible [1328]. computable [1178]. Computation [772, 462]. computing [680]. Concavity [99]. Concealed [255]. concept [1501]. Concepts [1046]. Concerning [1732, 1541, 1610]. condensation [87, 1364, 1451]. Condition [84, 1333, 720, 863, 384]. Conditions [312, 482, 1026, 1102, 195, 1460, 190, 822, 1360, 1078, 1478, 459]. Conducting [532, 1000, 522, 621, 851, 601, 18]. conduction [116, 1543]. conductivity [672, 1532]. Cone [871, 1143, 956, 1042]. cones [1212, 992]. configuration [7]. configurations [316, 161]. confluent [1488]. conform [1741]. Conformal [574, 449, 633, 1106, 1200, 1494, 1480, 1596, 368, 1770, 794, 1485]. Conformally [1168]. Congruence [958, 1495]. congruences [1074]. conical [860]. Conjecture [239, 1435, 1345]. conjectures [792]. conjugate [286].

Conjugation [449]. Connected [284, 1354, 675, 344]. connectedness [543]. connecting [396]. Connection [16, 1702, 1027, 619, 1069, 275]. connections [1612]. consequence [743, 573]. Consequences [519, 1513, 298]. Conservation [206, 214, 120, 760, 977, 373, 1207, 1530, 1541, 1499]. conservative [1295]. Conserved [479, 1285]. Considerations [43, 782, 1546]. consistent [1234]. consisting [61]. constant [1621]. Constants [845, 423, 1274, 640, 1428, 1295, 1659, 1176, 51, 1223]. Constitutive [1002]. constrained [1392]. constraint [1355]. constraints [357]. Constructing [928, 391, 426, 372, 486]. Construction [598, 1300, 68, 628, 1356]. Containing [1443, 909, 1, 941, 160, 1474, 280, 824]. content [717]. Continua [1065]. continuation [594, 1264]. continued [1509]. Continuity [510]. Continuous [787, 788, 542, 571, 1029, 1307, 659, 1562, 688, 633, 1642, 1643, 1685, 1686, 816]. Continuum [1356, 20, 1557]. contour [1220, 1204, 963]. contours [1709]. Contractions [819]. Contributions [615, 983]. Convection [830]. Convergence [1151, 1033, 1288, 1154, 1340, 645]. convergences [1132]. convergent [721, 1698]. Converging [1076]. convex [1786, 1041, 1568, 1587]. Convolution [243, 1237]. Cook [692, 1116]. Coordinate [1074, 1107, 1570, 1256, 465, 1678, 1221]. Coordinates [767, 668, 559, 12, 1525]. Core [778, 25, 1341, 1127]. corepresentations [1742]. correction [254]. corrections [1091, 36, 225, 509]. Correlation [309, 507, 723, 904, 753, 1233, 637, 355, 1377, 292, 1067, 299, 1015, 399, 839, 1519, 160]. Correlations [78, 128, 57, 842, 1284, 58, 702, 1627]. corresponding [17, 1522, 627, 1414]. corrugated [701, 846, 1568, 1587]. Coset [1140]. Cosmological [853, 280, 1232]. Cosmologies [808, 304, 38, 941]. cosmology [1565]. Coulomb [1118, 1414, 23, 1473, 984, 150, 961, 364, 757, 1055, 1516, 950, 935, 177, 832, 768, 1606, 147, 1077, 627]. counting [565]. Coupled [1366, 981, 507, 965, 884, 1549, 1518, 1057, 1359, 1758, 1021, 1365]. couples [1773, 1441]. Coupling [1616, 501, 272, 666, 1023, 132, 136, 1428, 174, 1461, 1309, 51, 1283, 1112, 1504]. couplings [1182, 1408]. Covariance [1285, 1638]. Covariant [736, 1252, 170, 1035, 975, 256, 1197]. covering [544]. Criteria [270]. criterion [1732]. Critical [570, 136, 1723, 127]. Critical-Point [570]. Critical-region [127]. criticality [1687]. cross [438, 891, 1258, 1129]. Crossing [135, 646, 39, 475, 64, 17, 357]. crossing-symmetric [357]. crystal [776, 1771, 852, 237, 672, 1532]. crystallization [208]. Crystals [984, 66, 394, 858, 859, 1543]. Cubic [413, 973, 546, 776, 191, 1186, 1544, 629, 467, 1666, 1762, 560]. cumulant [1662]. curiosity [1610]. Current [583, 1052, 514, 1152, 231, 1285, 575, 559, 9, 8, 7, 495, 1436, 1649, 1613, 1422, 1121]. Current-Stress [514]. Currents [807, 76, 252, 1524, 1197]. Curvature [307, 103, 398, 461, 199]. Curved [726, 386, 786, 780, 1091, 465, 752, 535, 351, 323]. Curves [660]. cutoff [1420]. Cutoffs [1036, 1454]. cuts [1371]. Cutting [75]. Cylinder [1017, 18].

Cylindrical [422, 834, 461, 1734, 1735, 162, 1756]. cylindrically [472].

D [515]. Damped [424]. damping [489]. dangerous [1714]. data [1356, 216, 1264]. Death [1565]. Debye [1418]. Decay [789, 1307]. Decision [936]. Decomposition [724, 821, 886, 1310, 1183, 1724, 746, 1008, 1222, 1623, 1051, 1690, 187, 1168]. decreasing [223]. Deep [1362]. defined [943, 221]. Definite [13, 435, 343, 521]. Definiteness [1184]. definition [256, 1538]. Deformable [813, 1002, 20, 851]. Deformation [1533]. Deformations [1386, 1754]. Degeneracy [23, 115, 240, 1208, 1209, 1622, 1715, 1270]. Degenerate [724, 209, 1528, 659, 1183, 1206, 966, 1726, 1200]. degeneration [1274]. delta [1185, 232, 1588]. delta-function [1185, 1588]. Democracy [767]. densities [420]. Density [559, 540, 1031, 521, 1751, 1184, 575, 1786, 822, 1284, 1398, 1180, 1207, 1750, 1018, 1269, 911, 702, 711, 712]. density-matrix [1284]. Dependence [60, 360]. Dependent [168, 554, 908, 457, 950, 29, 538, 887, 1772, 718, 1317, 1473, 355, 1089, 1319, 1677, 1431, 539, 790]. Depolarization [1346, 1347, 1455]. Derivation [206, 646, 769, 302, 1266, 1267, 1268, 192, 909, 739, 48, 311, 1458]. Derivative [1006, 1005]. Derivatives [736, 371, 216]. Derive [409]. derived [1592]. Deriving [76]. described [941]. describing [391, 1291]. Description [13, 435, 1137, 61, 1599, 943, 221]. destroy [1673]. Detection [1119]. determinant [375, 932]. determinants [70]. determination [942, 1196]. Determining [636, 1649, 430]. deterministic [1489]. development [1225, 703]. developments [1318]. deviations [719]. diagram [1657, 1291, 1357]. Diagrammatic [426, 383, 1312, 1085]. diagrams [1562, 1714, 291]. Diatomic [988, 704]. diced [1534]. Dicke [551, 902, 1630]. dielectric [193, 994, 148, 212]. Dielectrics [47]. difference [843, 1474, 862]. differences [473]. different [1428, 622, 1655]. differentiable [1174]. Differential [1158, 620, 84, 257, 1550, 1103, 258, 965, 155, 1242, 391, 612, 1510, 1658, 1661, 67, 516, 231, 1637, 982, 108, 431, 1150, 1701, 587, 1501, 299, 733, 1149, 300, 399, 839, 160, 1474, 1499, 1093, 1311, 1160, 825, 1607, 1058]. differential-difference [1474]. Diffin [1378]. Diffraction [461, 726, 499, 1728, 1088, 193, 204, 288, 18, 1497, 314, 522, 350, 896]. Diffusion [1020, 1379, 954, 830, 1758]. dilatation [1420]. dilatations [1075]. dilute [547]. Dimension [276, 470, 471, 1562, 1444, 1352, 669, 670, 245, 1376, 1764, 1688, 1588, 1040, 1415]. Dimensional [655, 736, 185, 284, 490, 1434, 244, 466, 264, 141, 819, 1737, 528, 730, 305, 394, 1330, 1248, 1251, 1151, 1396, 1597, 1654, 1784, 1214, 1201, 31, 179, 548, 1162, 1163, 33, 1424, 720, 25, 355, 512, 419, 1185, 852, 1520, 1377, 993, 1349, 799, 874, 364, 757, 32, 1398, 1477, 1476, 1442, 901, 998, 1304, 1622, 801, 328, 1479, 1136, 1642, 1643, 1685, 1686, 932, 1133, 1519, 237, 672, 139, 562, 1267, 1268, 1710, 1504, 923]. Dimensionality [930, 445, 1397, 1291]. Dimensions [483, 1751, 776, 114, 1117, 1133, 1454, 1452]. Dimers [110]. dipole [136, 389]. Dirac [1039, 23, 884, 43, 216, 219, 903, 1241, 1009, 433, 785, 678, 63]. Direct [240, 342, 526, 1323, 1359, 1360, 1008, 154]. Directed [660]. directions [1238]. Directors [1065]. Discontinuity [181, 339]. discovered [233]. Discrete [544, 1245, 1193, 1367, 1444, 1716, 1106, 663]. disk [522, 604]. disks [1680]. Disordered [673, 201, 1543, 505, 237, 672, 159, 1532, 1586]. Dispersion [609, 969, 1590, 1280, 1626, 1775]. dispersive [1717, 1220, 527, 1093, 1311, 1160, 911]. displacement [613]. dissipative [969, 1147, 987, 1417, 1328]. Distance [504, 1692, 1438, 15, 1690]. distinct [163]. distorted [1745]. distributed [7, 515]. Distribution [616, 348, 1576, 56, 1531, 565, 9, 8, 437, 392, 1619, 1700, 1522, 810, 1175, 36, 690, 509, 443, 243]. Distributions [452, 609, 1346, 1347, 1636, 414, 967, 1255, 1368]. disturbances [1252]. Divergence [312]. divergent [1097]. divisible [1603]. Domain [60, 29, 7, 251]. Domains [919, 1527]. dominance [1143, 1121]. Dominated [808, 853]. Doppler [188]. Double [713, 1713]. drainhole [1114, 1507]. Dressing [948]. driven [962]. drop [596]. Dual [492]. Duality [1556, 1291, 750, 1068]. Due [834, 162, 902, 711, 712]. Duffin [1429]. dumbbells [1457, 381, 828, 1622, 1502]. Dust [808, 949, 353]. Dust-Filled [353]. dyadic [375, 1107]. dyadics [791]. Dynamic [47]. Dynamical [1659, 875, 1596, 1401, 336, 1053, 182, 1523, 88, 1723, 1295, 1392, 818, 359, 1075]. Dynamics [1229, 413, 1743, 996, 1005, 918, 191, 1707, 245, 907, 962, 1504, 1506]. Dynamo [378, 651, 809, 693, 1323, 1359, 1360]. Dynamos [699]. Dyson [239, 651].

each [420]. earth [351]. Edge [726]. Effect [335, 493, 596]. Effective [3, 577, 148, 1738, 1177, 1403, 1763]. Effects [1758]. eigenenergies [1309]. Eigenfunctions [604, 1546, 295]. eigenrays [1049, 1785]. Eigenvalue [588, 694, 905, 1540, 503, 197, 1683, 976, 1576]. Eigenvalues [22, 1290, 1443, 496, 295, 889, 862, 443, 1388, 824, 825, 1058, 1368]. Eight [1542, 593]. Eight-vertex [1542]. Eikonal [241]. Einstein [975, 153, 662, 1674, 403, 738, 468, 1022, 1379, 769, 877, 549, 854, 419, 1364, 1451, 347, 1202, 1521, 498, 930, 472, 1733]. Elastic [1589, 47, 320, 1090, 1535, 791, 1357, 1215, 1568, 1587, 1176, 1064, 1067, 727, 442]. Electric [9, 1346, 1347, 23, 1316, 477, 193]. electrical [1738, 1177, 1403, 1763]. Electrodynamic [332]. Electrodynamics [77, 324, 585, 119, 1692, 1748, 1489]. Electromagnetic [1262, 1263, 322, 927, 522, 406, 621, 804, 1680, 212, 780, 882, 138, 603, 461, 1346, 1347, 14, 1039, 1339, 1616, 250, 791, 1000, 1471, 1594, 1462, 1621, 1215, 20, 1558, 869, 1275, 907, 858, 859, 193, 1171, 328, 188, 1025, 1635, 596, 1768, 1394, 1704, 851, 1733, 760, 759, 180, 351, 601, 1120]. Electron [1339, 317, 213, 815, 798, 1039, 991, 1298, 700, 499, 1506, 1477, 359, 704, 1137]. Electron-Ion [815]. electron-oscillators [1506]. electronic [1064]. Electrons [980, 150, 1606]. electrostatic [994]. Elementary [170, 898, 293, 81, 1087]. Elementary-Particle [170]. Elements [151, 685, 628, 462, 1043, 1184, 659, 796, 150, 426, 1546, 1203, 1207, 178]. Elimination [1678]. elliptic [793, 560, 1529]. embedding [1560]. emission [1119, 31]. Employing [769]. Empty [103]. encountered [74, 303]. ended [350, 896]. Energies [661, 1033, 371]. Energy [313, 583, 1443, 417, 228, 1506, 29, 563, 1016, 538, 521, 1388, 192, 1076, 1084, 1356, 814, 175, 176, 1, 598, 101, 1428, 637, 1000, 254, 607, 934, 1561, 749, 1089, 1600, 373, 1258, 1393, 1601, 1750, 899, 539, 760, 1244, 1537, 824, 1341, 1127]. Energy-Dependent [29, 538, 539]. Energy-Momentum [313, 583, 417, 1393]. ensemble [1522]. Ensembles [806, 503, 552, 383]. entity [1087]. Entropy [886, 1321, 782, 48, 642, 1402, 941]. envelope [1227]. envelope-soliton [1227]. enveloping [1708]. Equal [312, 313, 583, 414]. Equal-Time [312, 313, 583, 414]. Equation [131, 246, 939, 635, 356, 339, 265, 652, 84, 586, 857, 880, 130, 54, 120, 855, 651, 29, 885, 40, 732, 678, 155, 798, 1307, 290, 1242, 1702, 285, 1084, 152, 1342, 23, 1151, 190, 612, 1658, 1139, 1591, 1201, 1199, 43, 17, 102, 73, 1147, 1448, 1313, 174, 929, 854, 667, 496, 1361, 993, 1349, 619, 232, 1227, 1228, 295, 1126, 1187, 90, 708, 1598, 1780, 500, 594, 1716, 1265, 108, 1475, 229, 1323, 299, 1149, 97, 1488, 300, 399, 251, 656, 1677, 464, 1345, 1260, 1756, 647, 315, 493, 432, 237, 741, 1061, 1614, 1552]. equation [443, 1093, 1311, 1160, 106, 604, 825, 923]. Equations [634, 715, 833, 849, 662, 287, 16, 181, 620, 241, 769, 281, 981, 272, 257, 995, 398, 1082, 266, 524, 506, 883, 536, 965, 134, 705, 706, 1351, 391, 1158, 639, 1285, 1027, 731, 1419, 1384, 119, 975, 153, 1674, 1510, 210, 380, 1367, 613, 884, 1661, 800, 403, 17, 33, 125, 474, 863, 877, 1723, 1458, 1391, 1322, 549, 67, 1486, 892, 903, 1703, 1637, 982, 198, 826, 1699, 1275, 431, 1055, 1550, 1760, 1655, 693, 1359, 1360, 1521, 1442, 1150, 1701, 1344, 1103, 901, 1501, 171, 813, 733, 1149, 839, 926, 1747, 803, 1175, 1113, 36]. equations [169, 785, 1325, 1467, 843, 311, 160, 96, 139, 1474, 663, 942, 279, 1283, 1112, 242, 862, 1726, 1499, 1427, 1058, 1237]. equilibration [756]. Equilibrium [1641, 362, 1217, 970, 31, 179, 548, 1162, 1163, 719, 552, 1423, 292, 1504, 1758]. Equivalence [144, 948, 457, 622, 639, 552, 1555, 401, 86]. Equivalent [1247, 1564, 1271, 1253]. Erdahl [1786]. Ergodic [1299, 1470]. Errata [1045, 1507, 1281, 1535, 1536, 857, 834, 1412, 1414]. Erratum [1781, 634, 715, 879, 939, 635, 1183, 576, 1784, 1779, 1073, 413, 1777, 1226, 1508, 1118, 880, 1780, 1782, 1785, 577, 524, 1340, 1282, 1467, 1776, 525, 1711, 1413, 526, 1783, 1283, 1310, 1773, 1537, 1311, 1774, 1775, 1778, 1341]. Erroneous [1078]. Error [314]. estimates [968]. estimation [1143]. Ether [1114, 1507]. Euclidean [1548, 873, 208, 1744]. Euler [874, 671]. Eulerian [571]. Evaluation [881, 572, 62, 207, 648, 172, 147, 63, 1216, 1161, 1210, 1508, 1453, 1698, 1204, 1465, 1551, 816]. evolution [877, 1377, 1760, 1175]. Exact [210, 1123, 1654, 718, 1170, 1428, 867, 33, 1457, 1316, 1227, 1228, 756, 629, 198, 809, 381, 730, 998, 1304, 1622, 1715, 169, 71, 600, 883, 110, 778, 1448, 961, 1118, 1400, 693, 1519, 1502, 1773, 1441, 351, 1162, 1163, 348]. Exactly [529, 980, 332, 122, 803, 616]. Example [797]. examples [928]. exceptional [1252, 836]. Exchange [1073, 906, 1028]. Excitation [1071, 193]. Excitations [898]. Excited [757, 1346, 1347]. Exciton [959]. excitons [502]. execute [1539]. Exhibiting [900, 1479, 1710]. Existence [17, 1358, 408, 1668, 832, 1355, 464, 1061, 1753, 1779, 863, 1505, 1538, 926, 1026, 647, 741, 862]. Exotic [1207]. expanding [212]. Expansion [425, 576, 1594, 440, 226, 186, 504, 407, 923, 7, 721, 1178, 375, 929, 550, 961, 1118, 1185, 1097, 1509, 1722, 1315, 1438, 383, 690, 48, 1312, 1353, 627, 1414, 1720]. Expansion-free [1594]. Expansions [277, 714, 933, 1515, 595, 452, 89, 296, 1749, 489, 1462, 222, 21, 1177, 1314, 122, 475, 1438, 1309, 1180, 1154, 1340, 357, 1042, 1048, 1413, 15, 1266, 1267, 1268]. Expectation [417, 624, 384]. Expected [163]. Explicit [1559, 746, 1772, 895, 766, 1196, 216, 1631]. explicit-time-dependent [1772]. Explicitly [538, 539]. explored [1084]. Explosive [221]. Exponential [79, 1057, 155, 1440, 956, 597, 1700]. Exponentially [1600, 1601, 223, 262, 515]. expressions [216, 1546, 1432, 1058]. extended [1672, 1693, 1580, 1581, 1326, 1213, 1445]. Extension [382, 1172, 1094, 1125, 1417, 291, 1695]. Extensions [873, 1059, 872, 1615, 512, 1276]. Exterior [769, 1577]. Exterior-Algebraic [769]. external [1596, 1156, 1097, 1288, 99, 823, 1422, 279, 1667]. Extra [1439, 109]. extraneous [1538]. extrapolation [1371]. Extremal [370]. Extremum [131, 358, 432].

face [1544, 1762]. face-centered [1762]. Facial [1335]. Factorizability [1718]. Factorizable [1603]. factorization [639, 671]. factorized [796]. Factors [879, 1023, 511, 1102, 138]. Faddeev [33, 1486, 1442, 901]. Faddeev-like [1486]. family [210, 198]. Far [1587]. Fast [29]. fcc [561]. feedback [26]. Fermi [1189, 87, 1180, 390, 63, 1645, 1646]. Fermion [294, 630, 1751, 1548, 992, 500, 1479, 952, 87]. fermion-antifermion [500]. Fermions [1281, 1588, 1257, 122, 1224, 1096]. ferroelectric [469]. Ferromagnet [310, 99]. ferromagnetic [58]. Ferromagnetism [1073, 906]. ferromagnets [309, 1385, 1711, 1270]. Feynman [1158, 1053, 1254, 411, 1185, 53, 1563, 344, 638]. Field [168, 554, 972, 113, 13, 435, 769, 417, 1036, 410, 657, 457, 807, 366, 398, 650, 651, 605, 606, 581, 643, 519, 1006, 883, 866, 754, 56, 325, 412, 798, 167, 1346, 1262, 1734, 260, 261, 1039, 1339, 1384, 991, 1460, 1597, 1235, 1434, 336, 1156, 9, 8, 31, 179, 548, 1162, 1163, 1462, 868, 223, 319, 1650, 1777, 1587, 150, 946, 1097, 1288, 99, 924, 218, 1527, 30, 1757, 1294, 416, 1398, 408, 994, 1337, 256, 1067, 823, 1487, 328, 1463, 1175, 258, 1133, 579, 1345, 1541, 360, 1025, 359, 71, 600, 1116, 1602, 1230, 1765]. field [1783, 147, 1267, 396, 851, 1365, 1733, 1690, 280, 1446, 977, 1427, 1306, 1504, 1120, 1720, 350, 813]. field-free [396]. Field-Source [519]. Field-Theoretic [13, 435]. Fields [608, 335, 736, 123, 593, 618, 919, 927, 834, 625, 1005, 882, 603, 274, 781, 521, 831, 1060, 167, 1382, 14, 1100, 681, 1214, 476, 1567, 940, 1273, 1621, 1741, 20, 1673, 903, 1316, 1411, 10, 30, 1558, 199, 292, 1202, 1294, 1608, 1782, 907, 858, 859, 477, 1651, 162, 1609, 1583, 1584, 1483, 1494, 1131, 1032, 1171, 1425, 573, 1530, 360, 1635, 472, 279, 1394, 1704, 535, 1480]. fifth [1345]. Filled [353]. Films [1056, 1226]. Finding [22, 415, 96]. Finite [284, 1220, 272, 979, 518, 628, 330, 1272, 1531, 985, 827, 1083, 1162, 720, 1152, 847, 171, 1670, 1410, 744, 1742, 110, 1694, 1695, 1645, 1327, 179, 1714]. finite-component [171]. Finite-contour [1220]. Finite-Dimensional [284, 720]. Finite-Sized [979]. first [101, 1091, 1732, 1103]. first-order [1091]. Five [1477, 1476]. Five-dimensional [1477, 1476]. fixed [1356, 598, 101, 749, 1089, 899, 1626]. fixed-energy [1356]. FKG [1595]. flanged [869]. Flat [777, 156, 1492, 1082, 248]. Flow [673, 1114, 1507, 1218, 1328]. flows [1525]. Fluctuation [651, 252]. fluctuations [1758, 711, 712]. Fluid [957, 304, 517, 1547, 1781, 1676, 823, 1409, 533, 280]. Fluids [392, 885, 1018, 1631]. flux [739]. Fock [79, 1257, 76, 947, 692, 1116]. Fokker [1723]. fold [565]. folding [967]. foliations [1213]. Föppl [131]. force [1175]. forced [1539]. Forces [541, 294, 1123, 1486, 1294, 996]. Form [662, 1023, 217, 348, 155, 1102, 742, 218, 1546, 648, 895, 698, 1129, 138, 1297, 278]. Formal [639, 319, 1409]. Formalism [16, 450, 912, 1219, 814, 1679, 161, 961, 1118, 1185, 1205, 1580, 1581, 1326, 836, 692, 1116, 1206, 952, 1644, 1645, 1646, 711, 297]. formed [700, 1728]. forms [1136, 258, 1306]. Formula [409, 1379, 1191, 1030]. Formulas [143, 745, 1752, 1457, 10, 1291]. Formulation [86, 687, 37, 285, 537, 1053, 126, 367, 1650, 1777, 1692, 569, 1284, 1107, 1665, 631, 365, 1250, 1776, 1730, 638]. forward [520, 1280, 1775]. Foundations [797, 271]. Four [531, 930, 776]. Four-Body [531]. Four-Dimensionality [930]. Fourier [460, 710, 1143, 1057, 645, 1042, 15]. fourth [1658]. fourth-power [1658]. fractions [1509]. Frames [324]. Fredholm [1342, 1419, 613, 653, 932, 755, 1726]. Free [879, 881, 1692, 192, 511, 1074, 1076, 1298, 1594, 1152, 934, 1433, 1475, 1483, 698, 396, 1646]. Free-particle-like [1692]. Frequency [532, 32, 55, 1211, 188, 493, 1529, 1586]. frequency-modulated [1211]. friction [996]. Friedmann [607]. Froissart [609]. Fubini [1464]. fugacity [169]. Full [1263, 340, 1455, 1735, 1720, 1347]. full-wave [340]. fully [1249]. Function [308, 246, 939, 917, 542, 545, 546, 894, 920, 173, 723, 988, 904, 770, 555, 1430, 565, 1249, 986, 1471, 1185, 331, 232, 1126, 793, 1534, 364, 757, 1186, 1544, 392, 629, 1315, 374, 467, 1666, 1557, 1588, 1398, 1700, 1404, 913, 1762, 1712, 561, 680, 1015, 1756, 772, 469, 509, 1179, 790, 1420, 243, 404, 1504, 407]. function-square [680]. Functional [761, 787, 788, 83, 885, 533, 37, 1661, 1313, 1321, 416, 1550, 1165, 1613, 237, 910, 1197, 1048, 1413]. functionals [1649, 1271, 968]. Functions [556, 333, 145, 531, 207, 92, 857, 272, 592, 400, 1010, 203, 107, 504, 640, 463, 1035, 56, 491, 1387, 393, 945, 558, 380, 1753, 1779, 1201, 637, 355, 91, 550, 1377, 314, 21, 1013, 1187, 90, 1138, 473, 664, 292, 1619, 1515, 1055, 1554, 1449, 475, 1030, 1438, 299, 1261, 81, 560, 810, 399, 839, 1350, 932, 36, 189, 1519, 1744, 505, 547, 772, 910, 15, 616, 1058]. Further [4, 229, 1318]. future [1128]. Fuzzy [1450].

G [861]. Gårding [919]. galactic [1099]. Galilean [873, 80, 1705]. Galilei [1523, 1222]. Gamma [573]. Gas [264, 753, 665, 918, 1249, 1424, 669, 670, 331, 1688, 737, 1766, 1727]. Gases [590, 169, 547, 1592, 772]. Gauge [453, 1096, 1281, 1705, 1375, 274, 1051, 1310, 521, 1616, 1025, 1746, 275, 1499, 1427]. Gauss [51]. Gaussian [1282, 145, 565, 1110]. Gel’fand [812, 1009, 1623]. Gel’fand-pattern [1623]. General [949, 375, 111, 955, 104, 1432, 1171, 810, 305, 315, 485, 274, 781, 479, 409, 1169, 1547, 1781, 167, 14, 391, 1401, 1303, 1640, 175, 176, 1053, 1448, 1114, 1507, 49, 1090, 1535, 1256, 218, 619, 1660, 42, 1521, 746, 1107, 247, 1612, 1494, 758, 1425, 1355, 589, 1018, 836, 1694, 1695, 1687, 469, 1270, 1168, 175, 872, 126, 1580, 631, 1266, 1644]. Generalization [874, 692, 491, 1308, 1536, 1138, 229, 1175, 1438]. Generalizations [652, 120, 1009, 498, 144]. Generalized [1734, 100, 316, 230, 5, 940, 567, 143, 474, 1189, 769, 1638, 207, 164, 42, 1483, 1192, 1331, 1332, 173, 832, 568, 1059, 428, 1704, 915, 1348, 1247, 460, 1158, 195, 1564, 1358, 964, 1352, 1276, 447, 622, 229, 971, 97, 1028, 1026, 843, 311, 645, 968, 1233, 242, 976, 750, 716, 717, 1120, 1346, 1262]. Generalizing [1381]. generated [1396, 9, 8]. Generating [565, 1554, 1649, 1197, 549, 854, 924, 921, 922, 1613]. Generation [1202, 1557]. generator [1574]. generators [729, 1556, 1296]. geodesic [434, 1049, 1785]. Geodesics [864, 1448, 1445]. Geometric [516, 535, 877, 1308, 1536, 1748, 1541]. Geometrical [206, 340, 1091, 1025]. geometrical-optical [340]. geometries [1687, 441]. Geometrization [605, 606, 1583, 1584]. Geometrodynamics [912, 1545]. Geometry [149, 586, 427, 19, 118, 231, 1381, 1105, 95, 1607]. geons [876]. Gibbs [952]. given [1460]. Glasser [1466]. Glauber [654, 1129]. Global [334, 93, 1173, 190, 1560]. Goldstein [54]. Gordan [132, 133, 1343, 233, 202, 400, 1302, 894, 920, 1642, 1643, 1685, 1686, 1744, 1694, 1695, 267]. Gordon [1139, 17, 73, 477, 40]. governing [1175]. Grad [1611]. Gradient [266, 524]. graining [1644, 1645, 1646]. grand [331, 952]. graphical [1709, 418]. Graphs [873, 872, 1620]. Gravitation [1050, 573, 1393, 508, 1767, 1189, 1400, 1571]. Gravitational [307, 113, 489, 268, 593, 207, 75, 366, 581, 563, 1016, 837, 781, 1099, 336, 876, 1273, 1448, 1650, 1741, 1777, 434, 946, 1757, 1608, 1782, 416, 907, 1651, 1609, 1131, 1301, 579, 1345, 1733, 811, 278]. Gravity [468, 1562]. Green [939, 1779, 857, 246, 1753, 791, 917, 1471, 91, 207, 92, 793, 1013, 1534, 1187, 364, 757, 90, 1186, 1544, 272, 629, 374, 467, 545, 546, 664, 1666, 1557, 1055, 1762, 1261, 561, 560, 680, 988, 107, 1756, 547, 1179]. Gribov [609]. Griffiths [1389, 165]. Ground [416, 470, 1, 992, 234, 1284]. ground-state [1]. Grounded [834, 162]. Group [879, 845, 574, 1020, 317, 244, 466, 1253, 230, 646, 709, 388, 995, 289, 1009, 1072, 894, 427, 1059, 487, 628, 861, 909, 293, 1217, 511, 1570, 133, 1447, 1691, 1596, 991, 566, 1514, 1597, 1287, 689, 688, 114, 437, 27, 835, 1164, 326, 436, 752, 1703, 1363, 1533, 544, 1625, 1124, 42, 1770, 1081, 1222, 1483, 1719, 1038, 1412, 633, 475, 1106, 1200, 1579, 1068, 794, 154, 224, 1136, 1642, 1643, 1685, 1686, 1463, 189, 889, 1278, 1153, 1503, 768, 1744, 95, 931, 1157, 1348, 1754, 717, 404]. Group-Theoretic [646]. Group-Theoretical [230]. grouping [1267, 1268]. Groups [574, 284, 480, 235, 379, 534, 864, 199, 347, 400, 897, 920, 451, 11, 909, 1001, 1343, 1094, 1095, 1386, 194, 1083, 1615, 776, 182, 1523, 1603, 1675, 236, 419, 495, 227, 516, 847, 157, 1633, 658, 696, 1406, 1318, 671, 205, 1648, 1725, 1541, 744, 1742, 1694, 1695, 1069, 1305, 1296, 1485, 187, 1607, 1327]. guided [1164].

H [1366, 1129]. hadronic [960]. Hafnians [70]. Half [425, 576, 320, 1098, 435]. Half-Integer [435]. Half-Range [425, 576]. Half-Space [320]. half-spaces [1098]. Hamel [1328]. Hamilton [1147, 1417, 785]. Hamiltonian [411, 652, 630, 1573, 687, 1036, 1279, 648, 53, 1599, 746, 1165, 1613, 1469, 1355, 1274, 109, 1070, 836, 1565, 1230]. Hamiltonians [1663, 868, 570]. Hankel [314]. Hard [385, 778, 846, 25]. Hard-Core [778, 25]. Hardy [880, 708, 982]. Harmonic [945, 673, 894, 555, 639, 296, 718, 1234, 7, 852, 1518, 1057, 1543, 258, 237, 672, 1532, 1388]. harmonically [756, 996, 1758]. Harmonics [635, 5, 6, 152, 682, 277, 627, 1414, 278]. Harrison [610]. Hartree [315]. Hausdorff [1752]. Having [266, 524, 101]. Heat [673, 1543, 1702, 116, 1428, 642, 852, 739, 851]. Heisenberg [1045, 429, 309, 564, 723, 810, 904, 1015, 1385, 1711, 469, 165]. helical [1359]. Helicity [937, 39]. Helium [333, 700, 315]. Heliumlike [145]. Helmholtz [106]. Hencky [131]. Hermite [452]. Hertz [7]. Heterogeneous [3, 577, 1738, 1177, 1403, 1763, 1176]. Heun [1278]. Hexagonal [98]. Hidden [567, 59, 415]. Hidden-Variable [59]. High [167, 24, 661, 440, 1180, 1529, 56, 1084, 1738, 1600, 1258, 1601, 1635]. High-Field [56]. High-frequency [1529]. High-Temperature [24, 440]. Higher [1562, 13, 435, 1006, 1005, 520, 1771, 247, 1712, 448, 844]. Higher-Derivative [1006, 1005]. higher-order [1712, 448, 844]. Hilbert [1776, 1679, 1300, 446, 79, 1580, 1581, 1326, 1250, 938]. Hill [1201, 496]. hole [691, 1577, 401, 1334]. holes [1673, 870]. Holography [288]. Homogeneous [995, 352, 582, 167, 393, 133, 1597, 1448, 1558, 1703, 804, 596, 280]. honeycomb [1013, 1542]. Hooft [1435]. Hooke [1536, 1308]. Hopf [97]. Horizons [512]. horospheric [1363]. Hulthén [725]. Hydrodynamic [853, 311]. Hydrodynamics [957, 80]. Hydrogen [335, 394, 639, 796, 742, 768, 396, 654]. Hydromagnetic [699]. hyperbolic [1093, 1311, 1160]. hyperboloids [1720]. Hypercontractivity [1224]. Hypergeometric [473, 971, 173, 491, 21, 1138, 1554, 1449, 1488]. Hyperplane [349, 196]. hyperspheres [945].

Ice [394, 934]. Ideal [86, 169]. Identical [1443, 1, 61, 824]. Identities [929, 1066, 894, 920, 4, 729, 1468, 1415, 282]. Identity [1009, 458, 519, 1198, 728, 1524, 568]. II [1508, 1311, 705, 554, 1188, 260, 484, 632, 1095, 210, 176, 1372, 1245, 873, 689, 1320, 126, 16, 181, 435, 161, 1273, 179, 846, 990, 62, 182, 1212, 543, 1587, 1458, 787, 997, 670, 323, 49, 854, 86, 1210, 298, 534, 1349, 10, 922, 1403, 1187, 1138, 1186, 392, 1451, 1295, 1080, 1764, 859, 896, 1359, 1600, 1701, 1209, 1584, 1331, 1106, 1002, 1031, 1581, 920, 318, 904, 1643, 1684, 758, 353, 1601, 1207, 1632, 1766, 1338, 1631, 237, 159, 1695, 741, 644, 1266, 539, 1394, 1272, 127, 348, 1160, 778, 1755, 1645, 717, 1586, 1774, 1626]. II [844, 712]. III [833, 706, 261, 1084, 153, 662, 700, 202, 548, 788, 687, 1453, 30, 1763, 1714, 1360, 1318, 1332, 1200, 1015, 1410, 1685, 353, 418, 506, 110, 57, 1267, 1324, 1646, 481, 1627]. image [909]. imaginary [1315]. imbedded [136]. imbedding [1342, 1419, 613, 1726]. impedance [461, 180]. impenetrable [1040]. implementation [1096]. implementations [1281]. Implemented [453]. implications [1318]. important [1623]. improper [1404]. improved [1393]. Impurities [655, 1364, 1451]. impurity [32]. incident [698]. include [1417]. including [182]. Incoherent [959]. Incompressible [1359, 533, 1323]. indecomposable [1579]. Indefinite [574, 602, 1491]. Independence [349]. Independent [457, 950, 146, 1317, 1647, 1107, 256, 963]. indicatrices [624]. induced [1447]. inelastic [1362]. Inequalities [507, 1031, 203, 1184, 1220, 309, 1550, 1389, 1192, 1331, 1332, 1595, 589, 1590, 1233]. Infilling [117]. Infinite [736, 244, 466, 1397, 564, 723, 904, 306, 827, 1354, 1163, 1391, 163, 1361, 799, 1088, 996, 1015, 1756, 611, 1646, 1585, 1586, 1774]. Infinite-Dimensional [736, 244, 466, 799]. Infinite-Spin [564]. infinitely [1603]. Infinitesimal [330, 93, 1272]. infinity [1128, 946]. influence [1301, 1504]. information [1738]. Inhomogeneous [850, 123, 186, 293, 136, 1346, 1347, 1262, 1263, 1251, 1386, 1214, 527, 474, 1240, 226, 1503, 1365]. Initial [84, 251, 306, 830, 713, 1472, 1345, 811, 1168]. Initial-Boundary-Value [830]. Initial-Value [306, 251, 1168]. inner [1465, 187, 395]. instabilities [221]. Instantaneous [245, 1692]. insulated [1390]. insulators [851]. Integer [435, 13]. Integral [145, 751, 368, 6, 1391, 272, 546, 504, 463, 285, 1084, 1027, 1342, 1419, 41, 1151, 1396, 210, 380, 613, 62, 125, 961, 1118, 1185, 218, 10, 232, 793, 1186, 198, 826, 473, 1295, 1275, 1323, 1359, 1392, 560, 464, 189, 303, 1048, 1413, 1726, 443, 1459, 604, 448, 844, 1058, 1237, 800]. Integral-Transform [145, 463]. Integrals [411, 455, 1165, 786, 815, 172, 63, 1158, 572, 986, 1204, 53, 1551, 963, 914, 816, 258, 967, 1042, 704, 147]. integrand [1069]. integration [1053, 1254, 1564, 156]. Integro [257, 1637]. Integro-Differential [257, 1637]. Intensity [651]. Interacting [980, 559, 1599, 122, 737, 390, 279, 952, 943, 1137]. Interaction [1037, 1073, 906, 268, 630, 257, 66, 1733, 999, 956, 1573, 20, 302, 245, 1126, 392, 1688, 1323, 1359, 1360, 573, 704, 127, 851, 1653, 1341, 1127]. Interactions [16, 413, 984, 1749, 191, 1424, 25, 924, 1397, 641, 1758, 1325, 1467, 1431, 1244, 1537]. interest [102]. interface [112, 1437]. interior [1371, 1630]. Intermediate [329]. Internal [779, 255, 305, 841, 861, 960, 1596, 860, 45, 666, 840]. interpolating [1294].

Interpretation [233, 1006, 438, 891, 520, 1748]. interpretative [1513]. intertwining [1081]. interval [1361]. intrinsic [587]. Introduction [442, 545, 1144]. Invariance [366, 536, 991, 1772, 368, 1289, 516, 45, 171, 1541, 139, 663, 942, 1499]. Invariant [773, 1342, 1419, 153, 613, 584, 992, 423, 207, 513, 1408, 1310, 1726, 1315, 1397, 1344, 1222, 895, 171, 889, 109, 166, 1375, 1051, 1168, 1377]. Invariants [1351, 795, 578, 1618, 236, 1190]. Inverse [953, 264, 216, 695, 422, 789, 1496, 1658, 1654, 1245, 1193, 1367, 1444, 277, 101, 102, 73, 405, 749, 1089, 1472, 1582, 899, 1093, 1311, 1160, 1456, 1628, 1778]. Inversely [490, 1784]. Investigation [632]. involutional [485]. involving [572, 929, 690, 1408]. Ion [815]. Ionization [830]. ionized [1249]. Irreducible [284, 1083, 873, 776, 126, 202, 28, 1625, 636, 451, 354, 1484, 194, 872, 821, 27, 68, 658, 746, 1008, 794, 430, 1742, 599, 644, 481, 10]. irregular [1658, 1437]. Irreversible [1390, 358, 1717]. Irrotational [808, 853]. Ising [1248, 655, 1430, 642, 1520, 99, 913, 492, 810, 1519, 1592, 562, 310, 1092, 57, 58, 469, 1270, 1266, 1267, 1268, 127, 750, 1374]. Ising-Model [57, 58]. iso [481]. iso-Poincaré [481]. isometries [1731, 419]. isoperimetric [1192, 1331, 1332]. Isoscalar [879, 511]. Isospin [865]. Isospinors [1239]. isotopic [32]. isotopically [852, 1543, 237, 672, 1532]. Isotropic [1029, 1323, 1209, 57]. Isotropy [283]. Iterated [463]. iteration [62, 125]. Iterative [1496]. IV [588, 1162, 652, 1036, 1686, 58, 1268]. Iwamoto [1312].

J [1784, 1466, 857, 1785, 775]. Jacobi [1147, 785]. Jahn [697]. Jeffery [1328]. Jost [1358, 1700, 1350, 932, 790, 1629]. Jucys [1297]. JWKB [707].

Källén [958]. Kato [1231]. Kemmer [1378, 1429]. Kepler [537, 964, 1117, 1277]. Kernel [989, 990, 542, 10, 1342, 1396, 232, 1590, 966, 443, 1237]. kernels [613, 675, 1300, 1726]. Kerr [1589, 1594, 1173, 1032, 1221, 1495, 1105, 1445, 870, 1334]. ket [1326]. Killing [14, 765, 1566, 1223, 1480]. Kinematic [378, 651, 809, 699, 693, 1323, 1359, 1360]. kinematics [1354]. Kinetic [1330, 1384, 1365, 1017, 1727, 1303, 1423, 941, 1086, 686]. kinetics [1729]. Kinks [678, 977]. Kirkwood [1199, 1265, 443]. Klauder [1605]. Klein [17, 73, 477]. KMS [370, 1682]. Kondo [1551]. Kontorowich [460]. Korteweg [1307, 652, 120, 732]. Kraichnan [1323, 1359, 1360]. Kramers [1019]. Kronecker [194]. Kronig [1019]. Kruskal [341]. KS [1773, 1441]. KS-related [1773, 1441].

label [1514]. labeled [158]. Labeling [346, 372, 486, 861]. ladder [1546]. ladder-operator [1546]. Lagrangian [588, 694, 905, 1540, 924, 1545, 571, 816, 1018, 836, 1525]. Lagrangians [1247, 119, 1564, 865, 1206, 1108]. Lame [1278, 475]. Landau [1298, 404]. Lapiedra [1120]. Laplace [41, 713, 1438, 1756, 1255]. Large [651, 1309, 1568, 1587, 1719, 51, 160]. Large-Scale [651]. Lattice [98, 385, 264, 1012, 413, 191, 954, 997, 973, 793, 1013, 1534, 1186, 1544, 467, 545, 546, 664, 1666, 1688, 590, 269, 560, 988, 107, 57, 55, 1320, 1428, 1572, 1161, 1210, 1508, 1453, 331, 629, 32, 998, 828, 1304, 1622, 1068, 561, 680, 1015, 672, 110, 1592, 159, 58, 1542, 1452]. Lattice-Spin [269]. Lattices [413, 683, 959, 730, 988, 65, 191, 1228, 1013, 664, 1762, 1261, 561, 560, 680, 58, 1267, 1268]. lattices-field [1267]. lattices-temperature [1268]. Lauricella [1010]. Law [264, 1308, 1536, 218, 395]. Laws [206, 120, 1207, 1530, 1541, 1499, 187]. layer [218]. Lebedev [460]. Lebesgue [880, 708, 982, 382]. Lee [1372, 1111, 88, 247, 803, 1420]. Legendre [277, 1134, 21, 189]. Lemma [382, 1333, 1181]. Lenard [1115]. Length [753, 328]. levels [404]. Levinson [1027]. Lie [1780, 845, 1484, 1354, 284, 1386, 1656, 872, 873, 819, 676, 1523, 1429, 1572, 236, 864, 1617, 157, 848, 1598, 1633, 1699, 1449, 1044, 1010, 1488, 226, 186, 677, 404, 1607]. Lie-algebra [1572]. Lie-algebraic [1488]. Lienard [940, 1425]. Light [425, 576, 871, 345, 565, 1143, 956, 740, 1301, 1042]. Light-Cone [871]. lightlike [1156, 688]. like [1486, 1692]. Limit [1528, 564, 1229, 355, 1436, 496, 1619, 1397, 1719, 1649, 1269]. Limitable [182]. Limited [252]. Limits [1757, 1710, 414, 1470]. line [1145, 499, 1593]. Linear [705, 706, 833, 1396, 380, 141, 124, 586, 1190, 1031, 684, 641, 1236, 1386, 1510, 710, 67, 222, 374, 955, 42, 1736, 1150, 1747, 968, 166, 1196, 1137, 1639, 1653, 1585, 1586, 1774, 825]. linearized [975, 416, 1115, 493]. linearly [1067]. linked [1204]. Liouville [1496, 888, 714, 403]. Lippmann [1467, 1325, 647, 741]. Liquid [707, 700, 234]. little [1447]. Litvin [1181]. Lobachevski [1315]. Local [972, 975, 453, 779, 1241, 926, 1682, 1292, 1445, 274, 831, 403, 868, 719, 45, 420, 1649, 1613, 1746, 71, 238, 1765, 1783, 750, 1341, 1127, 1627]. Locality [1102, 80]. localizable [476]. Localized [196]. locally [1603]. located [1388]. Logarithmically [263]. Logic [484, 321]. logics [1135, 1338]. Long [413, 984, 1056, 191, 25, 1226, 1228, 392, 1294, 1493, 1758, 1325, 1467, 127]. Long-Range [413, 1056, 191, 25, 1226, 392, 1294, 1325, 1467, 127]. Long-Wavelength [984]. longitudinal [1568, 1587]. loop [1158, 1713]. Lorentz [1412, 133, 1447, 773, 566, 244, 466, 689, 43, 871, 1574, 207, 388, 1533, 544, 995, 366, 1081, 1038, 920, 1172, 1410, 1136, 1642, 1643, 1685, 1686, 170, 189, 1503, 768, 536, 663, 1035, 1727]. Lorentz-Covariant [170]. Lorentz-Invariant [207]. Lorentzian [1416]. losses [499]. Low [532, 819, 637, 737, 1266, 1267, 1268, 1668]. Low-Dimensional [819]. Low-Frequency [532]. low-temperature [737, 1266, 1267, 1268]. Lower [1, 1443, 192, 824, 825]. Lowering [879, 511]. Lowest [1469, 1576]. LSZ [1205]. Lucas [447, 843]. Lyra [1050, 951].

Mach [217]. Mackey [1615]. Magnet [723, 904, 310, 1015]. Magnetic [8, 123, 115, 509, 1439, 1346, 1347, 537, 23, 991, 1214, 1738, 1458, 903, 1411, 1177, 1403, 1763, 477, 1530, 1742, 1365, 280]. magnetically [813, 1002]. magnetization [99]. magnon [1385, 1711]. Majorana [140]. Mandelstam [259, 62]. Manifold [1050, 508, 1693, 1172, 1445]. Manifolds [951, 1370, 1652, 1355, 1168]. Manuals [1338]. Many [483, 308, 294, 317, 181, 86, 687, 83, 470, 471, 85, 778, 87, 492, 359, 702]. many- [87]. Many-Body [308, 86, 687, 470, 471, 85, 778]. many-component [492]. Many-Electron [317]. Many-Fermion [294]. Many-Particle [181, 83, 702]. Mapping [811]. mappings [1392]. Marchenko [1367, 16, 1358]. Markoffian [1729]. Markov [711, 712]. Markovian [1179]. Martin [720]. Mass [143, 348, 175, 176, 1434, 135, 876, 297, 298, 1490, 1491, 902, 1579, 938, 599, 644]. mass-energy [175, 176]. masses [1641]. Massive [606, 256, 1004, 892]. Massless [13, 435, 605]. Master [849, 731, 474, 229, 923]. matched [489, 222]. matching [1171]. Materials [3, 577, 1669, 1738, 1177, 1403, 1763, 1064]. Math [1784, 1466, 857, 1785]. Mathematical [634, 715, 833, 484, 61, 433, 1683, 960, 134, 705, 706, 321, 1125, 1336, 869, 211, 1580, 1581, 1684, 1489, 711]. Mathematics [327, 674]. Mathieu [287, 1126]. Matrices [503, 1031, 1410, 170, 558, 1786, 822, 720, 1041, 874, 1344, 383, 1522, 1670, 1576, 1368]. Matrix [725, 1045, 659, 151, 557, 806, 886, 150, 1697, 685, 628, 462, 1043, 909, 118, 1122, 1188, 1184, 429, 210, 796, 637, 520, 597, 426, 121, 1505, 1546, 447, 1203, 1284, 1314, 1259, 839, 1683, 1684, 523, 1482, 1154, 1340, 1207, 372, 486, 768, 755, 178, 1626, 1629, 627, 1414, 1341, 1127]. Matter [1303, 941]. Maximal [1213, 1095, 1648, 1725, 624]. Maxwell [524, 972, 884, 738, 1022, 375, 1202, 1521, 266, 1113, 258, 472, 1733]. Maxwellian [690]. may [318]. Mayer [1493, 910]. Mean [866, 1151, 1433, 1475]. means [1137, 395]. Measure [1132]. Measurement [111, 743, 674]. Measurements [764, 867]. measures [827, 799]. mechanical [1076, 1772, 572, 964, 548, 1147, 569, 292, 61, 1676, 1402, 823, 739, 311]. mechanically [962]. Mechanics [608, 936, 219, 271, 657, 228, 864, 128, 69, 915, 293, 1702, 1450, 1247, 1351, 827, 1003, 1513, 1399, 970, 800, 1564, 182, 367, 1379, 552, 978, 1556, 1641, 1130, 1697, 921, 922, 1706, 1204, 1047, 1191, 1337, 1581, 589, 109, 962, 1470, 1373, 1286, 506, 1087, 1746, 238, 611, 1293, 928, 1048, 1413, 1431, 1375, 674, 783, 494]. mechanism [1723]. Media [850, 324, 571, 1029, 136, 1251, 250, 527, 986, 1357, 1176, 1107, 813, 1002, 801, 816, 820]. Medium [406, 345, 118, 340, 1721, 791, 621, 804, 1398, 1655, 911]. Mellin [1186, 546, 1553]. melting [529]. Meson [519, 1121]. mesonic [700]. metal [1529]. Method [308, 22, 391, 1109, 50, 989, 990, 618, 327, 76, 762, 341, 462, 41, 1562, 743, 1125, 1254, 214, 549, 854, 1185, 1722, 1698, 993, 1349, 222, 1681, 1472, 1295, 1563, 1405, 1578, 1171, 300, 690, 110, 1694, 1695, 1021, 1324, 1459, 459, 395]. Methods [415, 82, 653, 105, 215, 814, 710, 1357, 1703, 1264, 418, 1312, 1048, 1413].

Metric [574, 602, 949, 1589, 1142, 334, 902, 1173, 1032]. metrics [610, 1049, 1785, 1221]. Metzner [1072]. Microelectromagnetism [253, 1665]. Micromagnetism [623]. Microscopic [632, 631]. Mills [1310, 1004, 1235, 408, 883, 1051]. Milne [113]. Minimal [1276, 784]. minimizing [878]. minimum [956]. Minkowski [488, 792, 1113]. Minton [592]. misrepresentation [1078]. missing [1514]. Mixed [1203, 400, 316, 713, 161]. Mixed-Basis [400]. mixing [174]. mixture [1569, 1588]. Mode [981]. Model [655, 385, 1073, 906, 444, 410, 83, 440, 980, 564, 362, 898, 390, 570, 57, 146, 137, 1531, 1382, 1330, 1248, 1464, 144, 1380, 1399, 1430, 1679, 1548, 1372, 1099, 1214, 1111, 530, 1182, 182, 1114, 1507, 987, 1243, 88, 529, 1299, 1377, 499, 231, 1321, 1080, 1079, 1389, 122, 1442, 247, 1030, 1362, 318, 1487, 81, 810, 1479, 803, 938, 1696, 562, 1116, 1092, 58, 1233, 1266, 1267, 1268, 127, 1446, 1420, 1137, 1232, 1374, 1542, 165]. Model-Independent [146]. Modeled [957]. Models [954, 997, 853, 352, 582, 866, 1135, 1352, 934, 1505, 1681, 30, 492, 1592, 469, 280, 750]. Modes [984, 1071, 740]. Modification [658]. Modified [893, 1325, 1467, 508]. modulated [1211, 1453]. modulation [1324]. Modulus [3, 577]. Molecular [66, 418, 493]. molecules [739]. Moment [1413, 1655, 774, 1149, 96, 441, 1048]. Moments [248, 323, 503, 299, 733, 839, 101, 214, 1426, 1398, 1522, 967]. Momentum [1531, 313, 583, 417, 1354, 368, 437, 876, 757, 594, 1610, 1393, 418, 760, 790, 1645, 1646]. momentum-dependent [790]. monochromatic [1339]. Monopole [115]. Montserrat [1120]. Moore [1615]. Morse [72]. Moshinsky [1141]. Most [724, 209, 1528, 659, 1183]. Motion [845, 1156, 1411, 1558, 477, 758, 643, 1353, 1384, 119, 1316, 1295, 1659, 1518, 1082, 389, 1274, 506, 139, 1223]. Moving [406, 489, 250, 499, 621, 804, 359, 759, 911, 18]. MR [1784]. multi [1069]. multi-Veneziano [1069]. multibin [1538]. multibody [1486]. Multichannel [1250, 1776, 1317, 1473, 1647]. multicomponent [632, 631, 1696]. Multidimensional [172, 91, 1739]. Multigroup [974, 1667, 966]. multilayered [1346, 1347, 1262, 1263, 1455, 1734, 1735]. multilevel [1743]. Multiparticle [1014, 146, 675, 520, 1178, 1718]. multiperipheral [1464]. Multiple [740, 1240, 1446, 702, 373, 365, 442]. multiplets [158]. Multiplicative [1054, 1723, 983, 1101, 1662, 1458]. Multiplicities [379, 534, 227, 187]. Multiplicity [879, 881, 511, 346, 395]. Multiplicity-Free [879, 881, 511]. multiply [210]. multipoint [1145, 1146]. Multipole [1462, 248, 323, 892, 1426]. multipoles [579]. Multiregion [1687]. multisection [351]. Must [1689]. Myriotic [608].

n [485]. Nagel [1141]. Naimark [812]. Natural [164, 1438]. near [434]. Nearest [730, 1688, 515, 998, 1622, 1715, 641]. Nearest-Neighbor [730]. Nearly [722, 699, 1264]. nearly-best-possible [1264]. Necessary [822]. necessity [1264]. neglect [1404]. Neighbor [730, 1688, 515, 998, 1622, 1715, 641]. network [318, 81, 1137]. networks [1739]. Neumann [1003, 916, 1438, 1024, 887]. neutral [632, 700, 742, 631]. Neutrino [781, 1638, 1477, 1583, 1584, 535]. Neutrino-Gravitational [781]. Neutron [195, 974, 1020, 306, 1283, 1112, 112, 41, 1151, 1396, 380, 91, 74, 1433, 1475, 1578, 95]. Neutron-Transport [306, 74]. neutron-wave [112]. Newman [156, 423, 870]. Newton [720]. Newtonian [1692]. next [515, 1715, 641]. next-nearest [515, 641]. Nice [747]. node [318]. Noether [1285, 439]. Noise [273]. Non [803, 1397, 1581, 1649, 1613]. Non- [803]. non-relativistic [1581, 1613]. non-translationally [1397]. noncentral [1316]. noncharacteristic [1731]. Noncluster [778]. Noncommutation [360]. Noncommutative [916]. Noncompact [158, 400, 1488, 157, 1410]. nonderivative [1108]. Nonequilibrium [753, 962, 506]. Nonexistence [190, 120]. Nonhomogeneous [586, 190]. Nonideal [665, 1766]. Noninertial [324]. Noninteracting [917]. Noninvariance [205]. noniterative [1275]. nonlightlike [1731]. Nonlinear [168, 554, 856, 356, 119, 1139, 1372, 77, 620, 241, 1134, 619, 900, 424, 1062, 160, 965, 830, 325, 345, 34, 391, 1084, 612, 969, 681, 197, 835, 987, 1697, 1227, 1228, 93, 826, 1760, 1359, 96, 1021, 166, 1324, 1485, 943, 1244, 1537, 1499, 106, 977]. Nonlocal [246, 939, 257, 1215, 1557, 1280, 1626, 1627, 1628, 1629, 1775, 1778, 1341, 1127]. nonlocalizable [1567]. Nonmetrical [1034]. nonnearest [515]. nonnegative [1469]. Nonnegativity [1279, 1070]. Nonorthogonal [714, 2, 525]. nonparallel [340]. Nonphysical [121]. nonplanar [1170]. Nonpolynomial [865, 1108]. Nonquantal [1641]. Nonradiative [1511]. Nonrelativistic [495, 1436, 146, 887, 1229, 1294]. Nonspherical [553, 967]. nonstandard [1080, 1079, 1144]. nonstrictly [476]. Nonuniform [335, 501, 1455, 1734, 1735, 195, 978, 1269]. Nonuniqueness [254]. nonunitary [1343]. nonvanishing [101]. nonzero [1750]. Nordstrom [1693, 441]. Normal [984, 155, 1107, 1470]. Normal-Ordering [155]. Normalization [448, 844, 1459]. Note [777, 876, 102, 70, 1621, 895, 1261, 585, 1062, 1499, 591, 4, 1231, 282, 1452, 1121]. Noumenon [1087]. novel [1519]. nth [418]. Nuclear [714, 668, 137, 1681]. nuclei [1284, 359]. Null [14, 1052, 765, 1635, 1074, 1095, 1238]. Number [1508, 908, 636, 1481, 163, 1257, 430, 1210]. Number-theoretic [1210]. numbers [1337, 785]. Numerical [415, 232]. numerically [1084]. NUT [334, 117, 1173].

object [1568, 1587]. Oblique [180]. Observables [657, 784, 1450, 1556, 1132, 1241, 1555, 238]. observations [1445, 165, 1624]. observing [1091]. Obstacles [532, 759]. obtaining [1295, 1694, 1695]. Occupation [908, 828, 1502, 87, 1257]. Occupation-Branching [908]. occupation-number [1257]. occurring [314, 1344]. occurs [420]. octonions [1369]. odd [139]. odd-dimensional [139]. Off [178, 627, 1414, 1341, 1127]. Off-energy-shell [1341, 1127]. Off-shell [178, 627, 1414]. once [420]. One [574, 490, 264, 707, 779, 528, 1442, 276, 815, 470, 471, 1330, 118, 1251, 1158, 1654, 1784, 35, 1214, 1444, 1201, 31, 179, 548, 1162, 1163, 33, 1424, 1352, 25, 355, 669, 670, 1185, 852, 1520, 302, 993, 1349, 245, 32, 1376, 1764, 1688, 1588, 1398, 1475, 901, 1040, 998, 1304, 1622, 801, 328, 1479, 1133, 1519, 237, 672, 562, 599, 644, 1710, 1504, 923]. one- [118]. One-Center [815]. one-dimension [1764]. One-Dimensional [490, 264, 1442, 1330, 1251, 1654, 1784, 1214, 1201, 31, 179, 548, 1162, 1163, 33, 1424, 25, 355, 1185, 852, 1520, 993, 1349, 32, 1398, 901, 998, 1304, 1622, 801, 328, 1479, 1133, 1519, 237, 672, 562, 1710, 1504, 923]. one-loop [1158]. One-Parameter [574]. One-Particle [779, 599, 644]. One-Speed [528, 35, 1475]. only [420]. onto [811]. open [350, 896, 1261, 962, 506]. open-ended [350, 896]. Operational [1046, 1338]. operations [1338]. Operator [634, 715, 833, 812, 827, 764, 289, 134, 705, 706, 1366, 1219, 814, 1679, 1246, 1546, 622, 1124, 1736, 1150, 1701, 1524, 1115, 303, 1305, 638, 1628, 1778, 1624]. Operators [879, 856, 151, 411, 602, 908, 457, 779, 301, 595, 832, 428, 841, 155, 511, 1001, 937, 1094, 1095, 1663, 728, 1083, 1358, 1517, 1246, 1481, 286, 1140, 302, 1132, 1527, 1604, 387, 878, 848, 1708, 53, 1155, 211, 1081, 1432, 1318, 895, 1104, 1258, 889, 1028, 1141, 1026, 1292, 599, 644, 1745, 1407, 1236, 649, 1327]. Optical [1555, 1060, 340]. optics [1091, 987]. Optimal [1371]. Orbit [595, 1229, 749, 1283, 1112]. Orbit-Orbit [595]. Orbitals [164, 704]. Orbits [1691, 482, 900, 419]. Order [620, 1056, 707, 1663, 1562, 1771, 1091, 1738, 1226, 67, 496, 1732, 1103, 1712, 1747, 418, 750, 1585, 1586, 1774, 825, 448, 844]. ordered [956, 1197]. Ordering [155, 638]. orderings [1165]. Ordinary [620, 431, 300, 399]. Orientation [338]. origin [216]. Orlicz [800]. Orthogonal [1761, 235, 628, 1591, 572, 1675, 874, 658, 696, 1406, 1168]. Orthogonality [74, 670]. Orthogonalization [215, 1302]. Orthogonalized [456]. orthorhombic [793]. Oscillations [424, 391, 580]. Oscillator [1217, 100, 900, 894, 34, 555, 999, 639, 718, 1253, 1322, 1130, 1697, 756, 569, 878, 1057, 211, 1208, 1209, 935, 166, 1388]. Oscillators [507, 1290, 681, 1549, 1518, 1506, 1021]. Oscillatory [412]. other [1551, 278]. outer [1216]. outgoing [929]. Overcomplete [2, 525]. Overlap [164]. overlapping [842].

p [1129]. packets [1353]. Padé [1145, 1146, 1134, 1288]. Pair [490, 87, 723, 904, 1784, 543, 1549, 1359, 1015]. pairs [286, 1238]. pairwise [459]. paper [1443, 679, 626, 1466]. parafermi [1286]. parallel [1728, 199, 1088, 828]. Parameter [574, 1084, 419, 473]. parameters [1419, 842, 67, 160, 750]. Parametrization [200, 64]. parametrized [344]. Parastatistics [1380, 1304]. Parastochastics [52]. Parity [13, 435]. Parker [1360]. part [727]. Partial [249, 965, 1242, 89, 516, 1501, 1149, 357, 1499, 1093, 1311, 1160]. Partial-Tensor-Product [249]. partial-wave [89, 357]. partially [210]. Particle [917, 181, 890, 779, 83, 834, 170, 643, 820, 479, 668, 1384, 1234, 1156, 742, 1114, 1507, 1693, 892, 1707, 1706, 1692, 392, 1500, 1659, 1599, 162, 1205, 369, 389, 1175, 1482, 36, 698, 1530, 647, 401, 741, 928, 599, 644, 654, 1218, 243, 702]. particle-distribution [36]. particle-hole [401]. particle-hydrogen [742]. Particlelike [168, 554]. Particles [773, 1443, 123, 13, 435, 86, 687, 281, 581, 293, 1119, 632, 1, 1434, 216, 1229, 1647, 297, 298, 1490, 1491, 1641, 245, 1558, 392, 61, 907, 477, 1040, 631, 698, 942, 279, 1375, 824, 494]. Particular [574, 176]. partition [1430, 331, 913, 1712, 469]. past [499]. Path [411, 455, 786, 914, 1053, 1254, 1564, 961, 1118, 1185, 53, 963, 1459]. Paths [386, 1433, 1475]. pattern [1094, 1623]. Pauli [958]. peaking [1099]. penetration [963]. Penrose [423]. pentagon [1657]. Peratization [262, 263]. Percolation [543]. Perfect [304, 517, 1018]. Perfect-Fluid [304]. Perfectly [532]. perform [1539]. Period [732]. Periodic [612, 482, 900, 701, 846, 1677, 505, 1353]. permanents [70]. permittivity [1177, 1403, 1763, 148]. Perpendicular [1520]. persistent [1524]. Perturbation [856, 294, 500, 1265, 829, 424, 745, 51, 665, 1766, 105, 1230, 341, 748, 1324, 1108, 783, 34, 814, 718, 681, 961, 1118, 1097, 1681, 1563, 431, 383, 1578, 580, 418, 1231, 968, 1021, 1585, 1586, 1774, 1568, 1177, 1763]. Perturbations [207, 952, 1142, 1509, 914, 1682, 1334, 1232, 1036]. Perturbative [849, 1749, 1362]. perturbed [1384]. Petiau [1378, 1429]. Petrov [153, 662, 1329]. Petrov-type [1329]. Pfaffians [70]. Phase [264, 557, 620, 761, 787, 788, 590, 1750, 1696, 1407, 1217, 1669, 1199, 1517, 1164, 25, 286, 87, 752, 569, 878, 211, 1479, 159, 390, 1709, 1233, 371, 1710, 943, 750, 221, 448, 844, 1453]. phase-integral [448, 844]. Phase-Plane [620]. Phases [370]. phenomenon [1717, 1605]. Phonon [66]. Phonons [444]. photoelectric [565]. Photoelectron [273]. Photon [450, 1750, 1407]. photons [632, 1422]. photosynthetic [502]. Phys [1784, 1466, 857, 1785]. Physical [181, 339, 1101, 1458, 86, 771, 1769, 1006, 1740, 298, 1539, 475, 1553]. Physical-Region [181, 339]. physically [1623]. Physics [634, 715, 833, 484, 129, 668, 134, 705, 706, 321, 1638, 982, 1079]. picture [205]. piezoelectric [1771]. planar [193, 1620]. Planck [1723]. Plane [1052, 620, 456, 433, 837, 1039, 1430, 1462, 1448, 1568, 1587, 1722, 594, 1521, 180]. plane-wave [1039, 1722]. planes [1728]. Plasma [953, 501, 78, 422, 1017, 391, 1211, 322, 1575, 188, 1179, 1365, 601, 18]. Plasmas [979, 1328]. plate [1529]. plethysms [1216, 1287, 1465]. Poincaré [1412, 942, 1602, 1447, 937, 1597, 873, 689, 688, 835, 326, 1038, 633, 1106, 1200, 1579, 794, 154, 1463, 599, 644, 481]. Poincaré-irreducible [599, 644]. Point [691, 707, 570, 85, 778, 94, 776, 956, 1000, 1424, 1707, 387, 1080, 1124, 902, 1030, 318, 1694, 459, 407]. Points [707, 504, 1658, 1199, 1371, 664, 1666, 1438, 1747, 15]. pointwise [1677]. Poisson [1103, 640]. Polariton [66]. Polariton-Phonon [66]. Polarization [1606, 1097, 1288]. Polarized [425, 576, 1064, 210]. polaron [1454]. pole [12, 389]. pole-dipole [389]. Poles [615, 871, 365, 1718]. Polish [1615]. polycrystalline [148]. Polynomial [676, 120, 690, 376, 1632, 677, 999, 728, 21, 1149]. Polynomials [170, 1611, 1761, 1591, 572, 447, 74, 848, 1318, 1278, 843, 138]. Polytype [838]. Position [1246, 1104, 841, 1432, 369]. Positive [521, 99, 599, 644]. Positive-Definite [521]. Positivity [1352, 357]. possibility [1091]. possible [1723, 1264, 717]. Potential [725, 751, 259, 553, 950, 82, 798, 1384, 1658, 1313, 1650, 1777, 405, 597, 961, 1118, 1185, 1588, 1700, 1600, 1319, 893, 1309, 51, 523, 1601, 1271, 225, 178, 790, 604, 448, 844]. Potentials [490, 264, 263, 683, 661, 950, 832, 1310, 23, 1654, 1784, 216, 598, 101, 262, 1391, 749, 1089, 12, 1557, 1493, 1418, 1270, 275, 1051, 1745, 1353, 1280, 1626, 1627, 1628, 1629, 1775, 1778, 627, 1414, 1341, 1127]. Power [264, 1376, 1764, 1658, 1553, 1086]. Powers [447]. pp [307]. Prequantization [1526]. Presence [115, 903, 1364, 1451]. pressure [1191]. principal [132, 659, 821, 633, 1106]. principal-series [132]. Principle [129, 1714, 1512, 1335, 1417, 48, 217]. Principles [131, 1066, 358, 1638, 432, 371]. Probabilistic [438, 891]. Probability [916, 1398]. Problem [308, 953, 588, 905, 378, 695, 422, 767, 531, 86, 687, 586, 789, 54, 201, 276, 571, 306, 653, 332, 470, 471, 830, 778, 340, 1497, 1496, 1540, 1731, 1514, 1654, 1193, 1367, 1444, 884, 1125, 1336, 1170, 216, 101, 102, 964, 73, 1650, 1777, 1117, 222, 1678, 749, 1089, 1637, 569, 1472, 1155, 774, 1302, 1582, 346, 823, 1171, 1642, 1643, 1685, 1686, 251, 1683, 1668, 1593, 686, 1277, 401, 899, 674, 976, 1093, 1311, 1160, 106, 1653, 1628, 1778, 811, 1168]. Problems [694, 490, 886, 257, 528, 220, 118, 1076, 1251, 537, 1139, 1784, 35, 1245, 986, 197, 1212, 254, 713, 314, 374, 955, 1433, 935, 690, 1730, 1459, 1667]. Procedure [636, 156, 680, 430]. Procedures [936, 1003, 1161]. process [1442]. Processes [1054, 124, 358, 1014, 830, 1119, 827, 543, 983, 1101, 1662, 1723, 1458, 446]. Product [480, 455, 1257, 249, 240, 1059, 956, 878, 154, 159, 1348]. Product-vector [1257]. production [1321, 941]. products [194, 1287, 1675, 1276, 658, 1008, 1292, 1197]. profile [911]. Projected [687]. Projection [402, 1305, 1003, 1150, 1701, 763]. Projective [1219, 1370, 840, 1128, 847]. Prolate [595, 1058]. Prolate-Spheroidal [595]. Proof [863, 239, 1402, 1765, 1783, 1620, 587, 1595]. Propagation [1721, 1211, 185, 850, 1012, 571, 1029, 603, 351, 345, 1717, 340, 1262, 1263, 1734, 1735, 112, 1251, 116, 322, 527, 986, 1320, 1472, 1376, 1764, 1655, 801, 816, 911]. Propagators [625, 1178]. Proper [338, 1707, 1706, 397, 663]. Properties [1669, 675, 530, 557, 24, 1573, 234, 771, 1259, 2, 525, 1035, 1145, 632, 1679, 1772, 476, 1517, 1738, 934, 1299, 1490, 1177, 1403, 1763, 693, 211, 344, 1344, 1115, 1684, 1409, 1750, 1631, 237, 472, 147, 1690, 535]. Property [997, 433, 669, 1492, 913, 1667]. pseudo [1675, 226]. pseudo-orthogonal [1675]. pseudo-unitary [226]. PUA [1603]. Pulse [911, 1717]. pulses [1211]. Pure [370].

Quadratic [490, 630, 1784, 1295, 1701]. quantal [718]. quantities [1067, 1553]. Quantization [1401, 1053, 1254, 482, 618, 72, 875, 1720, 975, 256, 205, 1194, 384, 261]. Quantized [343, 1384]. Quantum [608, 936, 1597, 797, 1000, 271, 417, 1036, 683, 852, 657, 1130, 919, 128, 564, 318, 81, 328, 1530, 1005, 311, 470, 471, 638, 225, 242, 1306, 293, 1382, 1450, 891, 1119, 827, 1003, 1513, 1689, 1772, 1434, 1562, 681, 1317, 572, 970, 1335, 1569, 964, 31, 179, 548, 1162, 1163, 1135, 182, 319, 1246, 987, 214, 80, 1556, 1549, 10, 1498, 30, 569, 1047, 1500, 848, 292, 1716, 61, 1402, 1337, 1580, 1581, 631, 1463, 36, 169, 589, 1133, 1470, 1746, 547, 238, 1602, 1230, 616, 1765, 1783, 674, 783, 1446, 977, 1504, 494]. quantum-mechanical [548, 61, 1402]. Quantum-Statistical [128, 31, 548, 1162, 1163, 1504]. Quantumlike [1730]. Quark [1369, 1380]. Quarter [528]. quasi [1525]. quasi-Lagrangian [1525]. quasiaverages [1109]. Quasigroups [465]. Quasilocal [510]. Quasiparticle [161]. Quasiperiodic [1677]. Quenching [959]. Question [217]. questions [1061].

Racah [316, 1395, 1759, 592, 458, 377, 395, 591]. Radial [796, 1693, 1350, 967, 798, 555, 254, 1411, 1126, 935, 1260, 147, 614, 1755]. radial-angular [796]. Radiating [979]. Radiation [136, 1100, 320, 268, 738, 607, 75, 213, 386, 44, 780, 210, 489, 1099, 94, 31, 179, 548, 1162, 1163, 892, 499, 869, 804, 1506, 596, 1236, 1504, 278]. radiation-field [1163]. Radiative [337, 614, 1098, 380, 74, 740, 1240, 1376, 1764, 1582]. radiative-transfer [74]. Radio [1734, 1735]. radius [1460]. Radon [716]. Rainich [587]. raising [1663]. Rajnak [568]. Random [134, 634, 715, 833, 655, 1012, 24, 83, 515, 503, 1011, 643, 65, 838, 603, 705, 706, 731, 1721, 986, 1320, 791, 1357, 1471, 1313, 1416, 163, 1177, 1403, 1763, 1321, 1398, 1655, 1404, 1067, 383, 1522, 801, 1175, 1519, 160, 1525, 1576, 1085, 1437, 1368, 923]. Random-Walk [83, 1321]. Randomized [882, 1394]. randomly [1384, 1211, 1478]. Range [425, 576, 413, 1056, 191, 1391, 25, 1226, 392, 1294, 1493, 1551, 1758, 1325, 1467, 127]. Ranged [468]. Rank [933, 302, 1566, 1157]. rate [645]. ratios [968]. Ray [386, 350, 896, 776, 1107]. Rayleigh [337]. Rayleigh-scattering [337]. Reactions [146]. reactors [26]. Real [819, 1060, 1145, 1656, 1337]. Realization [735, 404, 829, 1153]. realizations [260, 835, 848]. rearrangement [1442]. reciprocity [1618]. Recombination [830]. Rectangular [730, 664, 561]. Recurrence [227, 1030, 775, 395]. Recursion [133, 381, 1502]. Recursive [462, 1681]. Reduced [151, 1786, 822, 1041, 757]. Reduction [1447, 361, 209, 709, 826, 1038, 1412, 154, 202, 497, 346, 766, 976, 481]. reductions [1203]. reference [667]. reflection [1251, 1125, 1336, 171, 180, 1098]. Reflectors [306]. refraction [860]. regained [1545]. regardless [190]. Regge [1464, 233, 405, 1572, 1148, 661, 1309, 441]. Region [181, 339, 223, 121, 127]. Regions [900, 136]. Regular [673, 16, 1509, 594, 580]. regularity [624]. regularization [1434]. regularized [1749]. Reissner [1693, 441]. Related [797, 1066, 1295, 1392, 1744, 1773, 1441]. relating [70]. Relation [1060, 293, 89, 313, 1405, 523, 932, 731, 1759, 587, 381, 1068, 1410, 840, 395, 1280, 1626, 1775]. Relations [1184, 312, 583, 39, 1207, 249, 133, 609, 1004, 135, 286, 67, 227, 10, 799, 1019, 775, 189, 1632, 1502, 862, 1407]. Relationship [340, 484, 398, 460, 321, 1199]. relationships [1246]. Relative [369]. Relativistic [1611, 253, 20, 618, 995, 104, 855, 162, 834, 1071, 1065, 38, 1075, 1746, 663, 759, 537, 1303, 1640, 1253, 1170, 1229, 598, 1246, 150, 1090, 1535, 1641, 245, 1432, 1064, 1665, 369, 1581, 1649, 1613, 803, 690, 785, 1748, 71, 928, 1765, 1783, 760, 280]. Relativistically [47, 109, 1315, 1599]. Relativity [949, 897, 305, 274, 781, 479, 1169, 1547, 1781, 167, 14, 175, 176, 1114, 1507, 1370, 80, 49, 1256, 218, 619, 1660, 1476, 1612, 1494, 758, 1425, 1355, 1018, 1704, 1168]. Relaxation [444, 179, 548, 719, 31, 1162, 1163, 1458, 1504]. Relaxing [1063]. remain [794]. Remark [592, 1702, 1762]. Remarkable [1069]. Remarks [1037, 575, 688, 1605, 626, 72, 647, 139, 741, 94, 502]. Renormalization [938, 1415, 956, 1737, 1719, 1312]. Renormalized [34, 1481, 1322, 1150, 1701, 1362]. replication [1667]. representability [822, 401]. representable [1041]. Representation [1045, 662, 259, 908, 289, 504, 462, 1043, 1607, 429, 153, 1416, 1657, 1167, 858, 859, 369, 1291, 485, 1092, 1077]. Representations [1570, 991, 284, 724, 184, 244, 466, 209, 28, 676, 709, 235, 542, 1739, 157, 995, 420, 189, 140, 1062, 249, 240, 427, 451, 611, 538, 734, 1629, 1060, 1217, 132, 354, 1528, 1484, 1219, 361, 659, 1183, 194, 1674, 575, 688, 776, 821, 1462, 27, 1603, 1378, 1572, 1713, 1675, 68, 495, 446, 10, 1363, 544, 847, 622, 658, 1081, 1008, 1406, 1483, 763, 1579, 1259, 1649, 794, 1488, 154, 684, 224, 1136, 372, 486, 1278, 1503, 744, 518, 1157, 677, 539, 840, 1421, 481, 1626]. Reproducing [989, 990, 1300, 1590]. repulsion [392]. repulsive [1588, 1155, 1493, 893]. requirements [360]. residues [530]. resolvent [1342, 1419, 721]. resonance [842, 1718]. resonances [842]. Resonant [580, 345]. Respect [209, 361, 1038, 1412, 481]. Response [92, 412]. restricted [821, 1081, 1383, 794, 414, 1255]. Restrictions [1052]. resulting [1541]. results [1399, 1786, 629, 1270]. retardation [1474]. retarded [1753, 1779, 7]. reversal [1289]. Revisited [79]. Ricci [1273]. Riemann [382, 398, 1468, 275]. Riemannian [743, 1381, 199, 508, 1418, 1168]. Rigged [1580, 1581]. rigid [116, 1568, 1587, 1707, 1024, 1351]. Rigorous [978, 1270, 1399]. ring [1429, 499, 994]. Riseman [443]. Robinson [1768]. Role [1488]. Rosenbaum [4, 282]. Rotating [582, 941, 596]. Rotation [304, 920, 909, 1691, 1273, 1278]. Rotational [343]. rotator [1351]. rotor [1024]. Rough [373, 1357, 1471, 1478, 1085]. Rough-surface [373]. Rowlinson [1696]. RPA [976]. Rule [556, 394, 1765, 1783, 1620]. Rules [183, 646, 409, 1439, 1513, 1094, 1216, 658, 1465, 1019].

Sachs [1072]. Salam [1773, 1441]. Salpeter [635, 857, 152, 639, 1187, 90, 54, 500, 1055, 651, 656]. Salsburg [1199, 1265].

Sandwich [104, 1345]. satisfied [458, 1198]. Sato [1757]. saturated [813, 1002]. Scalar [726, 468, 1050, 605, 606, 38, 1577, 508, 217, 348, 94, 1767, 1400, 1315, 656, 1085, 1437]. Scalar-Tensor [1050, 38, 508, 217, 348, 1767, 1400]. scalars [1708]. Scale [651]. scales [1446]. Scaling [563, 1171, 1553]. scatter [373]. scattered [210]. Scattering [483, 1382, 532, 751, 1591, 250, 714, 259, 1249, 695, 181, 738, 701, 846, 1357, 1212, 1568, 1587, 349, 553, 1007, 1319, 247, 1014, 950, 276, 82, 177, 947, 815, 1422, 617, 1745, 1017, 887, 1085, 1437, 438, 891, 1366, 1084, 1356, 1159, 41, 1589, 1654, 1245, 1193, 1367, 1444, 1761, 1317, 1473, 1371, 216, 101, 102, 520, 1471, 223, 125, 742, 1391, 214, 669, 670, 929, 1647, 302, 749, 1089, 198, 740, 1240, 1275, 1600, 1516, 1205, 373, 475, 1680, 1258, 1362, 1040, 1264, 1350, 698, 1601, 303, 212, 1325, 1250, 1467, 1776, 1593, 1606, 1271, 899, 727]. scattering [337, 442, 1614, 1552, 760, 759, 178, 1456, 711, 712]. SCF [333]. Schild [1594]. Schrödinger [1658, 1313, 880, 1716, 285, 1201, 174, 929, 993, 1349, 1126, 387, 708, 594, 1677, 279, 1745, 798]. Schrödinger-type [387]. Schwarzschild [1100, 691, 1142, 434, 1221, 1260, 1577, 1213, 1105, 441]. Schwinger [1467, 1325, 647, 741]. Scintillation [882]. Scintillations [1394]. screened [1606]. Search [114]. Second [343, 620, 707, 737, 1663, 1096, 1281, 67, 205, 1747, 579, 225, 825]. Second-Order [620, 707, 1663, 67, 1747, 825]. second-quantization [205]. section [438, 891, 1258]. sections [1129]. Sector [958, 1209, 656]. sectors [1111, 88, 247]. Secular [287, 596]. segment [1593]. selection [1439]. Self [268, 602, 999, 1234, 1574, 924, 573, 1590, 952, 825]. Self-Adjoint [602, 825]. self-adjointness [1574]. self-consistent [1234]. self-generating [924]. self-interacting [952]. Self-Interaction [268, 999, 573]. self-reproducing [1590]. semi [1548, 996, 1607]. semi-Euclidean [1548]. semi-infinite [996]. Semiclassical [418, 1751]. Semiconductors [56]. Semidirect [480, 1059, 1276, 1348]. semigroups [783]. Semisimple [284, 676, 11, 1447, 157, 1044, 677]. seniority [137]. Separable [558, 279, 1314, 1154, 1340, 627, 1414]. Separation [50, 1703, 1598, 1633, 1699, 1780]. Series [294, 333, 466, 831, 132, 1145, 659, 1195, 1033, 821, 319, 1134, 1698, 1563, 633, 1106, 1200, 1024, 645, 442, 267]. sesquilinear [1306]. Set [918, 1041, 74, 1655]. sets [393, 1708, 1625]. seven [1597]. seven-dimensional [1597]. several [318]. shadowing [373]. shallow [1228]. shallow-water [1228]. Shape [944]. shear [1074, 1525]. shear-free [1074]. shearing [1049, 1785]. sheath [18]. shell [161, 1000, 1681, 1284, 614, 178, 627, 1414, 1341, 1127]. Shift [761, 787, 788, 188, 938]. shifts [371]. Shock [1016]. Short [468, 239, 392, 1360]. short-range [392]. Short-Ranged [468]. short-sudden [1360]. Shot [273]. sided [41]. Similarity [1328]. Simple [185, 406, 546, 269, 636, 430, 362, 909, 55, 236, 1299, 1186, 1652, 1044, 1595, 365, 165, 1607]. Simplex [1374]. simplexes [970, 1740]. Simplified [587, 363, 234]. Simultaneous [867]. Sine [40, 1139]. Sine-Gordon [40, 1139]. Single [668, 1234, 1713, 177, 887]. Single-Particle [668, 1234].

Singleton [544]. Singular [142, 263, 707, 1578, 303, 1027, 1749, 1658, 718, 262, 1256, 740, 1323, 1359, 893, 914, 1120]. Singularities [520, 808, 871, 853, 334, 121, 523, 758, 703, 1042]. Singularity [414, 956, 1678]. sites [163]. Sitter [293, 260, 261, 566, 544, 1104, 666]. Six [6]. Sized [979]. Slab [586, 306, 118, 195, 91, 1398, 212, 601]. Slater [704]. Slater-type [704]. Sliced [872, 873]. slit [1728]. slowly [489, 322, 1698]. small [1433, 1475, 1600, 1601, 160, 1690]. smooth [351]. Smoothed [1478]. Smoothing [1527, 1732, 300]. soft [701]. solid [1067]. solids [116, 1215]. soliton [1227, 1228]. Soluble [980, 122, 803, 616]. Solution [490, 1784, 287, 714, 1147, 88, 84, 708, 880, 901, 276, 958, 306, 843, 155, 798, 1242, 118, 1496, 1039, 1514, 210, 1654, 718, 102, 964, 548, 1162, 1163, 33, 1448, 1587, 174, 929, 1361, 516, 1400, 198, 594, 1265, 108, 1275, 1433, 1521, 893, 551, 902, 1630, 1149, 97, 1469, 1024, 464, 899, 1092, 943, 106, 277, 1360, 237]. Solutions [168, 554, 1084, 635, 662, 1020, 378, 265, 981, 130, 1475, 220, 883, 965, 348, 1307, 340, 1347, 1263, 1455, 1735, 391, 152, 731, 1151, 190, 612, 153, 1170, 1229, 1358, 197, 1594, 863, 1568, 549, 854, 67, 1505, 667, 892, 903, 1227, 1228, 1126, 1103, 299, 733, 46, 399, 839, 926, 656, 1747, 1668, 523, 803, 1355, 1677, 647, 493, 741, 1773, 1441, 1629, 811, 1058, 1237]. Solvable [444, 332, 529, 318, 81, 1648, 1725]. solved [1125]. Solving [1510, 1014, 775]. Some [321, 484, 1513, 944, 1399, 1037, 94, 1490, 1466, 1047, 995, 130, 693, 211, 269, 344, 771, 1344, 1670, 782, 589, 472, 502, 1474, 1709, 63, 165, 1624, 1058, 118, 144, 1123, 776, 1786, 675, 197, 1352, 67, 30, 839, 1750, 1489, 298]. Sound [571, 1029, 737, 816, 1543, 686]. Source [266, 524, 519, 7]. source-current [7]. Sources [406, 1034, 1346, 1347, 94, 621, 1082, 360]. Space [320, 103, 802, 584, 765, 1022, 1128, 338, 248, 323, 283, 454, 880, 117, 79, 397, 650, 128, 528, 930, 445, 920, 488, 517, 305, 1034, 421, 947, 252, 1169, 1751, 1731, 1095, 1597, 194, 1679, 1640, 743, 368, 1616, 986, 1135, 156, 1091, 1621, 1352, 434, 49, 1238, 607, 1724, 512, 1256, 446, 1492, 465, 1426, 1566, 752, 364, 757, 1703, 708, 982, 1315, 199, 1716, 1257, 1652, 1104, 828, 1304, 1622, 1580, 1581, 792, 1113, 1260, 1326, 1250, 1776, 1213, 1744, 938, 1694, 692, 1454, 1415, 666, 275, 1511, 535, 840, 1646, 1174, 1720]. Space-Charge-Limited [252]. Space-Temperature [128]. Spacelike [1503, 688, 223]. Spaces [777, 602, 86, 347, 1131, 786, 780, 393, 800, 1300, 68, 1082, 998, 1635, 139, 1329, 1418, 1485, 1098, 420]. spacing [1368]. spatial [527, 946]. Spatially [352, 582, 280]. Special [1660, 894, 920, 1142, 1609, 1744, 1069, 1511]. Specific [642]. Specification [1034]. Spectral [1517, 1657, 1115, 1713, 1077]. Spectrality [1690]. Spectrum [444, 868, 890, 236, 417, 228, 921, 922, 1307, 55, 667, 387, 32, 858, 859, 1736, 1553, 1586]. specular [1098]. Speed [528, 35, 1475]. Sphere [707, 388, 1547, 1781]. Spherical [175, 176, 264, 5, 6, 296, 144, 682, 277, 7, 1000, 550, 551, 739, 95, 820, 1614, 1552, 614, 1236]. Spherically [35, 1566, 832, 348, 955, 902, 1667]. Spheroidal [1516, 595, 295, 1319, 1058]. Spin [615, 13, 435, 1182, 450, 890, 24, 235, 164, 908, 440, 855, 269, 735, 581, 595, 1068, 564, 723, 904, 170, 519, 536, 57, 1023, 290, 1691, 937, 1634, 682, 688, 182, 642, 749, 1397, 1319, 913, 829, 1432, 763, 1064, 420, 1712, 1015, 641, 942, 58, 1270, 1283, 1112, 275, 1078, 977, 1374, 494]. Spin- [440, 1068, 615, 290, 682, 1374]. Spin-2 [519]. Spin-Dependent [908, 1319]. Spin-Orbit [595, 749, 1283, 1112]. Spin-Orbitals [164]. Spin-Pair [723, 904, 1015]. Spin-Spin [595]. Spin-Wave [735]. Spin-Zero [855]. Spinning [213, 907]. Spinor [49, 1476, 1406, 1612, 421, 682, 619, 1494, 656, 1120]. spinor-spinor [656]. Spinorial [898, 260]. Spinors [149]. splitting [1094]. Spontaneous [230, 1671, 31]. Square [385, 1654, 405, 664, 680, 110, 225, 1341, 1127]. square-well [225]. Squares [385, 194]. Srivastava [626]. Stability [168, 554, 1661, 707, 900, 979, 26, 1139, 969, 403, 596]. stabilizers [393]. staggered [1728]. standard [1080]. Stark [335, 396]. State [1020, 86, 1014, 958, 1029, 470, 1387, 1119, 1, 216, 992, 234, 757, 1716, 1284, 416, 36, 169, 1078]. States [356, 540, 370, 784, 878, 196, 491, 1217, 1691, 1634, 1123, 970, 1740, 437, 1569, 174, 670, 1377, 1617, 1138, 1557, 408, 1222, 1190, 318, 372, 486, 1408, 1682, 1750, 1470, 71, 600, 1385, 1711, 611, 599, 644, 952, 1407, 448, 844]. Static [593, 1273, 579, 348, 512, 1492, 1566, 551, 902, 1704]. Stationary [356, 927, 1032, 421, 1741, 1724, 1426, 1566, 1202, 1608, 1782, 1049, 1785, 1250, 1776, 359, 71, 600]. Statistical [608, 437, 1177, 1403, 1763, 128, 1676, 69, 1286, 159, 1431, 1060, 1076, 1399, 970, 800, 31, 179, 548, 1162, 1163, 367, 1738, 552, 978, 1641, 1204, 648, 1191, 693, 823, 739, 631, 589, 962, 1470, 311, 506, 238, 611, 1293, 1048, 1413, 616, 783, 1504]. Statistically [809, 1029, 1568, 1587, 1403]. Statistics [1046, 329, 730, 629, 1376, 1764, 1688, 998, 828, 1338, 1502, 110]. Steady [1020, 1029, 1475]. Steady-State [1020, 1029]. steps [515]. Stieltjes [1145, 1288]. Stochastic [849, 1054, 1289, 446, 900, 571, 1218, 494, 1702, 1159, 1510, 1661, 983, 1101, 1662, 1723, 1458, 1550, 1150, 1701, 299, 733, 300, 399, 839, 816, 464, 1086, 1282, 1110, 160, 96, 1730, 562]. Stokes [1103]. straight [1593]. Stranghellini [1464]. stratified [340, 1211]. strength [1461, 1501]. strengthened [1513]. Stress [77, 514, 1561, 739]. stress-energy [1561]. Stress-Tensor [77]. striated [180]. strips [1088]. Strong [541, 174, 1402]. Structure [700, 970, 946, 1138, 1664, 387, 897, 1072, 491, 716, 717, 805, 1297, 1464, 1094, 1095, 316, 743, 1253, 1786, 701, 846, 1041, 49, 1369, 1318, 971, 1068, 810, 758, 414, 1255]. Structures [838, 479, 461, 1346, 1347, 1262, 1263, 1455, 1734, 1735, 1674, 193, 1174]. structures-full [1455, 1735]. structures-Generalized [1734]. Studies [425, 576, 962, 617]. Study [1063, 982, 1679, 31, 1498, 1681, 1563, 1524, 159, 275]. Sturm [1496, 888, 714]. subadditivity [1402]. subalgebras [1523, 1378, 1044]. Subgroup [628, 1447, 821, 1533, 1203, 481]. Subgroups [574, 11, 794, 1648, 1725, 1157]. Subject [84]. subrepresentations [1675]. substituted [852]. sudden [1360]. Sugawara [231]. Suggestive [359]. Sum [556, 646, 973, 1124, 1019, 627, 1414, 407]. Summability [294, 105, 1288].

Summation [1749, 1759, 961, 1118]. sums [1161, 1210, 1508, 1453, 159, 1452]. Superconductivity [623, 1512]. Superfluidity [487]. Superposition [129, 1689, 1335]. superpropagator [1440]. Superpropagators [1636, 1122, 1188]. superselection [1765, 1783]. supplementary [1195]. support [1255]. Surface [218, 461, 1357, 1471, 1568, 1587, 373, 180, 1085]. surfaces [1225, 1771, 1107, 703, 291, 1478, 1480]. susceptibility [1520, 562]. Symbol [925, 1148]. symbols [775]. Symmetric [219, 1009, 832, 240, 64, 722, 699, 348, 1001, 1351, 260, 261, 1342, 1287, 35, 17, 1650, 1777, 1566, 955, 475, 902, 1032, 357, 472, 1157, 1270, 1168, 1667]. Symmetries [1672, 779, 771, 377, 841, 1303, 1671, 1739, 921, 922, 1295, 1659, 45, 1075, 840]. symmetrization [1140]. Symmetrized [19, 194]. Symmetry [317, 230, 343, 584, 1022, 415, 592, 255, 347, 1392, 1277, 628, 591, 960, 290, 167, 133, 537, 1596, 233, 175, 176, 336, 114, 208, 1289, 752, 1148, 1521, 122, 551, 768, 359, 1635, 1705, 1746, 1329, 1296, 1157, 147, 666, 1614, 1552, 614, 1236, 441]. Symplectic [186, 27, 658, 746, 696, 226, 1153]. synchronously [1021]. System [714, 531, 652, 83, 290, 1401, 632, 135, 559, 1053, 33, 863, 1390, 1299, 1377, 1604, 826, 61, 1506, 1539, 901, 1501, 369, 1068, 739, 631, 96, 1269, 616, 1137, 702]. Systems [588, 694, 905, 145, 1443, 888, 24, 269, 900, 585, 1074, 1439, 1540, 1151, 1772, 489, 1, 1123, 1740, 675, 179, 1162, 1163, 1147, 474, 1243, 642, 992, 25, 355, 978, 1152, 87, 1090, 1535, 516, 1498, 1295, 1659, 500, 1417, 1714, 1599, 1392, 913, 1432, 1580, 1581, 1712, 1274, 1086, 1530, 1758, 962, 818, 1286, 506, 505, 836, 390, 611, 1179, 1710, 952, 943, 1645, 1646, 1585, 1586, 1774, 824].

t [1314, 1435, 725]. tableaux [42]. Tachyon [949]. Tachyons [1571]. Tail [75]. Takahashi [1004]. Taub [334, 1173]. TCP [139, 942]. Technique [1014, 504, 96, 125, 426, 1079, 1150, 1701, 775, 1623]. Techniques [402, 572, 1662, 1305]. Teller [697, 1389]. Temperature [24, 128, 440, 723, 904, 1430, 1152, 1714, 737, 1015, 1266, 1267, 1268]. temperatures [1531]. temporal [527]. Tensor [313, 583, 151, 682, 77, 514, 1050, 398, 301, 498, 930, 38, 636, 249, 508, 19, 217, 348, 278, 1327, 1001, 1094, 1095, 1083, 1767, 1741, 1675, 1724, 419, 749, 1703, 1400, 1318, 739, 1393, 1468, 430, 360, 599, 644, 1157, 1773, 1441, 649]. tensorial [1625, 260]. tensors [316, 126, 202, 1561, 1566, 746, 1223, 1168]. Terms [546, 216, 550, 793, 1186, 1404, 560, 1306]. tests [1539, 1538]. tetragonal [680]. th [1157]. Their [423, 398, 452, 731, 216, 1096, 1281, 446, 1318, 205, 684, 816, 1338]. Theorem [403, 5, 567, 1054, 513, 104, 645, 1027, 697, 637, 214, 1295, 1555, 1392, 315]. Theorems [425, 576, 6, 593, 439, 296, 1047, 165]. Theoretic [646, 13, 435, 1508, 1132, 717, 1210]. Theoretical [230, 1703, 1079]. Theories [168, 554, 410, 59, 807, 170, 762, 1006, 1548, 223, 1771, 1189, 74, 431, 1477, 1541, 1373, 1705, 71, 600, 977]. Theory [261, 1020, 317, 141, 738, 468, 142, 797, 660, 795, 319, 253, 1215, 417, 1036, 297, 298, 1050, 92, 111, 72, 289, 1176, 1284, 528, 651, 44, 735, 1538, 424, 1010, 324, 1590, 885, 947, 665, 519, 105, 252, 508, 487, 754, 748, 217, 348, 1017, 325, 288, 887, 37, 1060, 34, 1382, 1330, 260, 112, 1004, 1303, 1689, 632, 41, 1597, 814, 380, 175, 1434, 681, 35, 710, 1761, 1591, 872, 1317, 1473, 527, 1335, 800, 437, 161, 637, 868, 125, 1767, 1246, 208, 1146, 1370, 983, 1647, 1490, 852, 1423, 20, 924, 1641, 222, 1241, 30, 1167]. theory [1177, 1403, 1763, 869, 1400, 1598, 1633, 1699, 1780, 374, 740, 955, 1240, 1294, 1376, 1764, 515, 860, 350, 896, 1323, 1359, 1360, 1404, 1493, 551, 902, 1630, 1575, 1258, 941, 1002, 1665, 1580, 1581, 1578, 318, 1487, 631, 81, 328, 51, 1463, 365, 1683, 1684, 523, 1350, 36, 464, 1086, 1133, 303, 1409, 962, 1325, 1250, 1467, 1776, 418, 1766, 1025, 686, 505, 533, 502, 95, 1116, 1231, 755, 1618, 966, 1602, 968, 1765, 1783, 1266, 1108, 1365, 242, 1745, 1078, 1690, 1773, 1441, 1446, 351, 1727, 1644, 1585, 1586, 1774, 1306, 1667, 1120, 494]. Thermal [1532, 1217, 1738, 719, 1177, 1403, 1763, 292, 774, 672]. thermally [1390]. Thermodynamic [355, 1619, 1269]. Thermodynamical [660, 370]. Thermodynamics [1248, 957, 1588, 363, 471]. thickness [1346, 1347, 1262, 1263]. thickness- [1262]. Thin [1056, 104, 1345, 701, 1226, 522, 846]. Thin-Sandwich [104]. Third [142, 1500]. Thomas [1180]. Three [385, 1037, 141, 142, 276, 653, 1751, 1151, 1396, 1123, 1654, 1170, 1428, 956, 1678, 1040, 1136, 1642, 1643, 1685, 1686, 1747, 932, 1482, 698, 1744, 1653]. Three-Body [142, 276, 653, 1123, 1654, 1170, 1678, 1653]. Three-Colorings [385]. Three-Dimensional [141, 1151, 1396, 1136, 1642, 1643, 1685, 1686, 932]. three-particle [1482, 698]. three-point [956]. Three-Waves [1037]. Tidal [434]. tightly [500]. Time [168, 554, 312, 313, 583, 1473, 367, 1647, 1377, 912, 950, 723, 904, 887, 1772, 718, 743, 1616, 1317, 7, 956, 584, 1022, 1091, 338, 1289, 355, 1238, 607, 1256, 465, 650, 1652, 445, 941, 963, 1015, 1677, 414, 1034, 1260, 160, 1197, 1415, 666, 1446, 535, 840, 441, 1174, 1720]. Time-Dependent [168, 554, 950, 887, 1473, 718, 1317, 355, 1677]. Time-Independent [950, 1317]. time-ordered [956, 1197]. Timelike [831, 14, 434, 291, 1480]. times [1169, 1731, 1640, 103, 802, 765, 156, 1128, 1621, 49, 1724, 512, 1256, 454, 1492, 1426, 1566, 199, 517, 421, 1511]. Tomimatsu [1757]. Tomonaga [1531]. Tonin [1464]. Tool [617]. top [1351]. Topological [924, 325, 1615, 344, 1709]. Topologies [478, 792]. Topology [488]. trace [1258]. trace-class [1258]. Trajectories [661, 405, 1693, 1309, 1218]. Trajectory [123, 1464]. Transfer [118, 886, 380, 74, 740, 1240, 1506, 1582, 1086, 337, 966, 614, 1236, 1244, 1537, 1098]. Transfer-Matrix [886]. transfiguration [1565]. Transform [145, 463, 460, 710, 1716]. Transformation [958, 1570, 1234, 1358, 713, 1539, 1383, 1553, 1373, 1025, 1196, 1733, 1328, 459, 1581]. Transformations [453, 948, 423, 685, 342, 526, 330, 985, 1285, 1396, 1298, 368, 1616, 1096, 1281, 465, 93, 1760, 1208, 1209, 492, 684, 935, 1410, 1221, 1277, 518, 485, 663, 1293, 1272, 1485, 1374]. Transforms [85, 778, 609, 460, 1346, 1262, 1734, 277, 1143, 1057, 1255, 1639, 1755, 716]. transient [9, 8, 7]. Transition [1317, 44, 732, 529, 1479, 1747, 1233, 1710]. Transitions [264, 590, 1199, 25, 87, 515, 1696, 390, 750]. translation [1358, 45]. translationally [1397]. translations [367]. Transmission [237, 601, 1251, 1543]. Transport [974, 141, 124, 586, 386, 528, 306, 118, 195, 41, 1396, 380, 35, 710, 1591, 1189, 91, 852, 222, 74, 374, 955, 1376, 1764, 1433, 1736, 1475, 1578, 188, 1409, 1631, 95, 820, 966, 1283, 1112, 1667, 711, 712]. Transverse [193]. Trapped [703, 1225]. trapping [502]. Traps [65]. treated [1116]. Treatment [421, 812, 613, 700, 961, 1118, 1763, 500, 704]. Tree [954, 997]. Trial [463]. Triangular [57, 1428, 1013, 58]. tridiagonal [387]. trilinear [1408]. trivial [1675]. true [337]. tuned [1021]. Tung [261]. turbulence [1676, 1323, 1359, 1575, 464, 533, 1179]. turbulent [693, 1218]. Turning [707]. Twistors [1770]. Two [655, 1020, 707, 1500, 528, 1007, 730, 723, 904, 1015, 220, 140, 504, 815, 332, 1385, 1711, 95, 394, 118, 1248, 1339, 41, 558, 1663, 1669, 1658, 1123, 1654, 1249, 4, 114, 512, 1520, 1117, 1377, 1566, 245, 392, 1728, 1599, 746, 1680, 1438, 369, 1325, 1250, 1467, 1776, 1373, 647, 279, 741, 1454, 15, 282, 1021, 1773, 1441, 291, 1452, 1480, 1267, 1268]. two- [1123]. Two-Body [332, 1654, 746, 1325, 1467]. two-component [1249]. Two-Dimensional [655, 528, 730, 394, 1248, 512, 1377, 1267, 1268]. Two-Electron [815]. Two-Group [1020, 95]. two-Hilbert [1776, 1250]. Two-magnon [1385, 1711]. two-medium [118]. Two-particle [1500, 392, 1599, 369, 647, 741]. two-phase [1669]. two-sided [41]. two-surfaces [291, 1480]. two-tensor [1773, 1441]. Two-Time [723, 904, 1015]. Type [662, 287, 546, 517, 925, 153, 17, 1096, 1281, 387, 1186, 1479, 1198, 1329, 704].

U.I.R. [1447]. ultraspherical [138]. Unbounded [981, 238, 1555]. Uncertainty [784, 286, 878]. underdense [188]. Uniaxial [858]. Unidirectional [1244, 1537]. Unification [1596]. unifications [1348]. Unified [704, 1487]. Uniform [798, 994]. Uniformly [1361, 250, 804]. unimodular [909, 889]. Unique [53]. Uniqueness [1256, 403, 17, 102, 992, 1668, 1355, 1768, 1061, 870, 120]. Unit [301, 1011, 173, 649, 502]. Unitarity [265, 863, 1670, 1061, 1614, 1552]. Unitary [354, 574, 343, 1363, 995, 224, 451, 1421, 481, 1001, 1356, 1094, 1095, 821, 27, 62, 720, 1675, 495, 752, 847, 658, 1318, 1539, 1488, 684, 226, 889, 1153, 1069, 931, 1296]. universal [114, 544]. Universe [113, 898, 1768]. Universes [353]. Unrenormalizable [625]. unrestricted [1689]. Unstable [890, 281, 1205]. Unti [156]. Upper [825]. Ursell [910]. Use [384, 1539, 371]. Useful [680, 1490]. Using [714, 489, 1738, 755, 1415, 147]. utility [1314].

V [694, 277, 1163, 709, 120]. Vacuum [802, 765, 417, 927, 624, 1731, 1724, 1097, 1288, 1492, 1049, 1785, 551, 1773, 1441]. valid [1361]. validity [1732]. Value [586, 257, 306, 830, 770, 35, 713, 1637, 1472, 374, 955, 251, 1345, 686, 106, 811, 1168]. valued [827]. Values [417, 216, 1636, 624, 384]. vanish [1460]. Variable [59, 461, 1419, 297, 298, 1490, 1641, 328]. variable-length [328]. Variables [567, 747, 558, 1663, 336, 1416, 1598, 1633, 1699, 1780, 109]. variance [163]. variation [814]. variation-perturbation [814]. Variational [814, 1066, 1665, 82, 754, 968, 613, 1146, 1271, 371]. various [518]. varying [1346, 1347, 1262, 322]. Vector [728, 6, 7, 1135, 656, 1460, 776, 1257, 256, 1485, 1121, 1480]. vector-meson [1121]. Vector-scalar [656]. Vectors [919, 1011, 958, 2, 525, 14, 199, 1463, 1602]. Velocity [808, 853, 1164, 1359, 1431]. Velocity-Dominated [808, 853]. Veltman [1435]. Veneziano [1679, 530, 1182, 971, 1030, 1069, 716, 717]. version [1193, 1367]. versions [792]. vertex [1399, 1542]. vertices [1713]. VI [905, 1504]. via [411, 53]. Vibration [397]. vicinity [691]. view [1080]. viewpoint [1761]. VII [1540]. Virial [142, 520, 1500, 225]. visited [163]. Volkov [46]. Volterra [1361]. volume [1645]. Vries [1307, 652, 120, 732].

Waals [1409, 1631]. Walk [83, 163, 1321]. Walks [65, 838, 731, 515, 923]. Ward [1004, 1415]. water [1228]. Watson [460, 973]. Wave [246, 939, 356, 850, 1012, 1320, 695, 281, 257, 130, 398, 650, 1655, 855, 735, 801, 832, 762, 536, 1387, 89, 1351, 340, 1347, 1262, 1263, 1455, 1734, 1735, 112, 1251, 1039, 116, 190, 814, 322, 527, 986, 73, 1462, 1448, 1568, 1587, 1722, 302, 892, 1227, 1228, 295, 1472, 1376, 1764, 1155, 1700, 1107, 1344, 171, 1524, 803, 1026, 1260, 686, 357, 663, 942, 1324, 242, 1078, 180, 943, 1244, 1537, 1353, 1085]. wave-operator [1524]. wave-wave [943, 1244, 1537]. Wavefunction [550, 700, 742, 698]. Wavefunctions [343, 270, 669, 81, 396, 1077, 448, 844]. waveguides [869, 350, 896]. Wavelength [984]. wavenumbers [1655]. Wavepacket [46]. Wavepackets [1491]. Waves [483, 307, 185, 944, 1037, 501, 456, 1063, 433, 1016, 837, 461, 1346, 1347, 1339, 1252, 969, 250, 791, 1164, 1228, 1275, 193, 829, 188, 1577, 1733, 1745, 759, 351, 601, 18, 1437]. Weak [1651, 1194, 392, 1493]. Weaker [792]. Weakly [123, 34, 969, 1753, 1779, 737, 1365]. Weight [511, 879, 68]. weights [395]. Weinberg [261]. Weisskopf [179, 548, 1162, 1163, 1743, 1504]. well [225, 943, 221, 448, 844, 1341, 1127]. well-defined [943, 221]. Weyl [149, 1741, 1243, 1724, 1483, 794, 931, 951]. where [420]. which [1566, 199, 1539, 794, 318, 1058]. Whose [640, 81]. Widom [1696]. widths [1522]. Wiechert [940, 1425]. Wiener [97]. Wiesskopf [31]. Wightman [1690]. Wigner [881, 31, 179, 548, 1162, 1163, 1743, 1395, 436, 1509, 1166, 1559, 1515, 805, 1504]. Wind [954, 997]. Wind-Tree [954, 997]. within [1625]. without [537, 1003, 1036, 12, 1454, 394, 750]. WKB [1125, 254, 893, 384]. World [352, 582]. Wybourne [568].

Yamada [1312]. Yang [1310, 1004, 1235, 408, 883, 1051]. Young [42, 1291]. Yukawa [868, 1352, 1279, 1309, 1469, 1070].

Zeeman [792]. Zero [855, 1434, 559, 1000, 1182, 1579]. zero-mass [1434]. zero-point [1000]. Zeroes [1430, 1700]. Zeros [331, 1712, 469, 314, 913]. Zhelobenko [1405]. Zwanzig [1620].

References

Calogero:1969:LBG

[1] F. Calogero and C. Marchioro. Lower bounds to the ground-state energy of systems containing identical particles. Journal of Mathematical Physics, 10(4):562–569, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1443].

Rowe:1969:PON

[2] D. J. Rowe. Properties of overcomplete and nonorthogonal basis vectors. Journal of Mathematical Physics, 10(9):1774–1777, September 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1774_

s1. See erratum [525].

Miller:1969:BEB

[3] Melvin N. Miller. Bounds for effective bulk modulus of heterogeneous materials. Journal of Mathematical Physics, 10(11):2005–2013, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v10/i11/p2005_s1. See erratum [577].

Chiang:1969:FNT

[4] David Chiang. Further note on two binomial coefficient identities of Rosenbaum. Journal of Mathematical Physics, 10(11):2098, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [679].

Clapp:1970:GAT

[5] Roger E. Clapp. Generalized addition theorem for spherical harmonics. Journal of Mathematical Physics, 11(1):1–4, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/p1_s1.

Clapp:1970:SIT

[6] Roger E. Clapp and H. Tung Li. Six integral theorems for vector spherical harmonics. Journal of Mathematical Physics, 11(1):4–9, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i1/p4_s1.

Clapp:1970:VSH

[7] Roger E. Clapp, Luc Huang, and H. Tung Li. Vector spherical harmonic expansion in the time domain of the retarded Hertz vector for a distributed, transient source-current configuration. Journal of Mathematical Physics, 11(1):9–13, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i1/p9_s1.

Clapp:1970:MFG

[8] Roger E. Clapp and Luc Huang. Magnetic field generated by a transient current distribution. Journal of Mathematical Physics, 11(1):14– 16, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i1/p14_s1.

Clapp:1970:EFG

[9] Roger E. Clapp, Luc Huang, and H. Tung Li. Electric field generated by a transient current distribution. Journal of Mathematical Physics, 11(1): 16–20, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i1/p16_s1.

Hegerfeldt:1970:KIF

[10] Gerhard C. Hegerfeldt. Kernel integral formulas for the canonical commutation relations of quantum fields. II. Irreducible representations. Journal of Mathematical Physics, 11(1):21–29, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/p21_s1.

Stone:1970:SSS

[11] Anthony P. Stone. Semisimple subgroups of semisimple groups. Journal of Mathematical Physics, 11(1):29–38, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i1/p29_s1.

Kerner:1970:CPC

[12] Edward H. Kerner. Charge and pole: canonical coordinates without potentials. Journal of Mathematical Physics, 11(1):39–46, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/
p39_s1.

Cukierda:1970:FTD

[13] T. Cukierda and J. Lukierski. Field-theoretic description of massless particles with higher spin and definite parity. I. Integer spin. Journal of Mathematical Physics, 11(1):46–50, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i1/p46_s1.

Banerjee:1970:NEF

[14] A. Banerjee. Null electromagnetic fields in general relativity admitting timelike or null Killing vectors. Journal of Mathematical Physics, 11(1): 51–55, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i1/p51_s1.

Sourour:1970:FEF

[15] A. R. Sourour and A. A. Ashour. Fourier expansions of functions of the distance between two points. Journal of Mathematical Physics, 11(1): 56–60, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i1/p56_s1.

Cornille:1970:CBM

[16] Henri Cornille. Connection between Marchenko formalism and N/D equations: Regular interactions. II. Journal of Mathematical Physics, 11(1):61–78, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i1/p61_s1.

Cornille:1970:EUC

[17] H. Cornille. Existence and uniqueness of crossing symmetric N/D-type equations corresponding to the Klein–Gordon equation. Journal of Mathematical Physics, 11(1):79–98, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i1/p79_s1.

Yeh:1970:DWC

[18] C. Yeh. Diffraction of waves by a conducting cylinder coated with a moving plasma sheath. Journal of Mathematical Physics, 11(1):99–104, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i1/p99_s1.

Tompkins:1970:STG

[19] Donald R. Tompkins. Symmetrized tensor geometry. Journal of Mathematical Physics, 11(1):105–108, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i1/p105_s1.

Grot:1970:RCT

[20] R. A. Grot. Relativistic continuum theory for the interaction of electromagnetic fields with deformable bodies. Journal of Mathematical Physics, 11(1):109–113, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i1/p109_s1.

Holdeman:1970:LPE

[21] Jonas T. Holdeman, Jr. Legendre polynomial expansions of hypergeometric functions with applications. Journal of Mathematical Physics, 11(1):114–117, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i1/p114_s1.

Anderson:1970:NMF

[22] J. T. Anderson. New method for finding eigenvalues. Journal of Mathematical Physics, 11(1):117–124, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i1/p117_s1.

Berrondo:1970:DDE

[23] Manuel Berrondo and Harold V. McIntosh. Degeneracy of the Dirac equation with electric and magnetic Coulomb potentials. Journal of Mathematical Physics, 11(1):125–141, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i1/p125_s1.

Gallavotti:1970:HTP

[24] G. Gallavotti. High-temperature properties of random spin systems. Journal of Mathematical Physics, 11(1):141–146, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/p141_s1.

Gallavotti:1970:APT

[25] G. Gallavotti and S. Miracle-Sole. Absence of phase transitions in hardcore one-dimensional systems with long-range interactions. Journal of Mathematical Physics, 11(1):147–154, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i1/p147_s1.

Akcasu:1970:ASR

[26] A. Z. Akcasu and P. Akhtar. On the asymptotic stability of reactors with arbitrary feedback. Journal of Mathematical Physics, 11(1):155– 162, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i1/p155_s1.

Devi:1970:BIRa

[27] V. Syamala Devi. Bases for irreducible representations of the unitary group in the symplectic group chain. Journal of Mathematical Physics, 11(1):162–168, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i1/p162_s1.

Devi:1970:BIRb

[28] V. Syamala Devi and T. Venkatarayudu. Bases for irreducible representations of U_{2n} in the chain U_{2n} ⊃ U_{n} ˙+U_{n}. Journal of Mathematical Physics, 11(1):169–173, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i1/p169_s1.

Nicolaenko:1970:EDB

[29] Basil Nicolaenko. Energy-dependent Boltzmann equation in the fast domain. Journal of Mathematical Physics, 11(1):174–184, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/
p174_s1.

Hoegh-Krohn:1970:AFS

[30] Raphael Høegh-Krohn. Asymptotic fields in some models of quantum field theory III. Journal of Mathematical Physics, 11(1):185–188, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i1/p185_s1.

Davidson:1970:RQSa

[31] Russell Davidson and John J. Kozak. On the relaxation to quantumstatistical equilibrium of the Wigner–Wiesskopf atom in a onedimensional radiation field. I. A study of spontaneous emission. Journal of Mathematical Physics, 11(1):189–202, January 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i1/p189_s1.

Kinzer:1970:FSO

[32] Earl T. Kinzer. Frequency spectrum of a one-dimensional lattice with an isotopic impurity. Journal of Mathematical Physics, 11(1):203–206, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i1/p203_s1.

Dodd:1970:ESF

[33] L. R. Dodd. Exact solution of the Faddeev equations for a onedimensional system. Journal of Mathematical Physics, 11(1):207–213, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i1/p207_s1.

Aks:1970:RPT

[34] S. Ø. Aks and R. A. Carhart. Renormalized perturbation theory for the weakly nonlinear oscillator. Journal of Mathematical Physics, 11(1):214– 222, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i1/p214_s1.

Case:1970:SSB

[35] K. M. Case, R. Zelazny, and M. Kanal. Spherically symmetric boundaryvalue problems in one-speed transport theory. Journal of Mathematical Physics, 11(1):223–239, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i1/p223_s1.

Nienhuis:1970:TQC

[36] G. Nienhuis. On the theory of quantum corrections to the equations of state and to the particle-distribution functions. Journal of Mathematical Physics, 11(1):239–243, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i1/p239_s1.

Zardecki:1970:FFC

[37] Andrzej Zardecki. Functional formulation of the coherence theory. Journal of Mathematical Physics, 11(1):244–250, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/p244_s1.

McIntosh:1970:RAS

[38] C. B. G. McIntosh. Relativistic analogs of scalar-tensor cosmologies. Journal of Mathematical Physics, 11(1):250–252, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/p250_s1.

Hara:1970:CRH

[39] Yasuo Hara. On crossing relations for helicity amplitudes. Journal of Mathematical Physics, 11(1):253–257, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i1/p253_s1.

Rubinstein:1970:SGE

[40] Julio Rubinstein. Sine-Gordon equation. Journal of Mathematical Physics, 11(1):258–266, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i1/p258_s1.

Boffi:1970:ASN

[41] V. C. Boffi. Anisotropy of scattering in neutron transport theory by the two-sided Laplace integral method. Journal of Mathematical Physics, 11(1):267–279, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i1/p267_s1.

King:1970:GYT

[42] R. C. King. Generalized Young tableaux and the general linear group. Journal of Mathematical Physics, 11(1):280–293, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/p280_s1.

Cohn:1970:CLD

[43] Jack Cohn. Considerations on the Lorentz–Dirac equation. Journal of Mathematical Physics, 11(1):294–295, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i1/p294_s1.

Lewis:1970:ATT

[44] Robert M. Lewis and Jack K. Cohen. Asymptotic theory of transition radiation. Journal of Mathematical Physics, 11(1):296–305, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i1/p296_s1.

Landau:1970:TIL

[45] Lawrence J. Landau and Eyvind H. Wichmann. On the translation invariance of local internal symmetries. Journal of Mathematical Physics, 11(1):306–311, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i1/p306_s1.

Meyer:1970:WVS

[46] James W. Meyer. Wavepacket Volkov solutions. Journal of Mathematical Physics, 11(1):312–317, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i1/p312_s1.

Bragg:1970:RDE

[47] Lincoln E. Bragg. Relativistically dynamic elastic dielectrics. Journal of Mathematical Physics, 11(1):318–338, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i1/p318_s1.

Ramanathan:1970:EPD

[48] G. V. Ramanathan, J. M. Dawson, and M. D. Kruskal. Entropy principle for the derivation of a new cluster expansion. Journal of Mathematical Physics, 11(1):339–342, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i1/p339_s1.

Geroch:1970:SSS

[49] Robert Geroch. Spinor structure of space–times in general relativity. II. Journal of Mathematical Physics, 11(1):343–348, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i1/p343_s1.

Caswell:1970:AAS

[50] Randall S. Caswell and Michael Danos. On the accuracy of the adiabatic separation method. Journal of Mathematical Physics, 11(2):349– 354, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p349_s1.

Muller:1970:PTL

[51] Harald J. W. Müller. Perturbation theory for large coupling constants applied to the Gauss potential. Journal of Mathematical Physics, 11(2):355– 364, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p355_s1.

Frisch:1970:P

[52] U. Frisch and R. Bourret. Parastochastics. Journal of Mathematical Physics, 11(2):364–390, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p364_s1.

Kerner:1970:UHO

[53] Edward H. Kerner and William G. Sutcliffe. Unique Hamiltonian operators via Feynman path integrals. Journal of Mathematical Physics, 11(2):391–393, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i2/p391_s1.

Keam:1970:BSE

[54] R. F. Keam. Bethe–Salpeter equation and the Goldstein problem. Journal of Mathematical Physics, 11(2):394–403, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p394_s1.

Baroody:1970:AFS

[55] E. M. Baroody. Approximations for the frequency spectrum of a simple lattice. Journal of Mathematical Physics, 11(2):404–407, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i2/p404_s1.

Vassell:1970:CHF

[56] M. O. Vassell. Calculation of high-field distribution functions in semiconductors. Journal of Mathematical Physics, 11(2):408–412, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i2/p408_s1.

Stephenson:1970:IMSa

[57] John Stephenson. Ising-model spin correlations on the triangular lattice. III. Isotropic antiferromagnetic lattice. Journal of Mathematical Physics, 11(2):413–419, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p413_s1.

Stephenson:1970:IMSb

[58] John Stephenson. Ising-model spin correlations on the triangular lattice. IV. Anisotropic ferromagnetic and antiferromagnetic lattices. Journal of Mathematical Physics, 11(2):420–431, February 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i2/p420_s1.

Gudder:1970:HVT

[59] Stanley P. Gudder. On hidden-variable theories. Journal of Mathematical Physics, 11(2):431–436, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p431_s1.

Geroch:1970:DD

[60] Robert Geroch. Domain of dependence. Journal of Mathematical Physics, 11(2):437–449, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p437_s1.

Kummer:1970:MDS

[61] H. Kummer. Mathematical description of a system consisting of identical quantum-mechanical particles. Journal of Mathematical Physics, 11(2):449–474, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i2/p449_s1.

Dilley:1970:EUI

[62] James Dilley. Evaluation of a unitary integral. II. Mandelstam iteration. Journal of Mathematical Physics, 11(2):475–476, February 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p475_
s1.

Wasserman:1970:ESF

[63] A. Wasserman, T. J. Buckholtz, and H. E. DeWitt. Evaluation of some Fermi–Dirac integrals. Journal of Mathematical Physics, 11(2):477–482, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p477_s1.

Roskies:1970:PCS

[64] Ralph Z. Roskies. Parametrization of crossing symmetric amplitudes. Journal of Mathematical Physics, 11(2):482–486, February 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p482_
s1.

Rosenstock:1970:RWL

[65] Herbert B. Rosenstock. Random walks on lattices with traps. Journal of Mathematical Physics, 11(2):487–490, February 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i2/p487_s1.

Mavroyannis:1970:PPI

[66] C. Mavroyannis. Polariton-phonon interaction in molecular crystals. Journal of Mathematical Physics, 11(2):491–507, February 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p491_
s1.

Goel:1970:SRB

[67] Narendra S. Goel and Frank P. Buff. On some relations between the solutions and the parameters of second-order linear differential equations. Journal of Mathematical Physics, 11(2):508–512, February 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p508_
s1.

Gilmore:1970:CWS

[68] Robert Gilmore. Construction of weight spaces for irreducible representations of A_{n}, B_{n}, C_{n}, D_{n}. Journal of Mathematical Physics, 11(2):513– 523, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p513_s1.

Marchand:1970:BAS

[69] Jean-Paul Marchand. Born approximations in statistical mechanics. Journal of Mathematical Physics, 11(2):523–529, February 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p523_
s1.

DeLuca:1970:NRP

[70] A. De Luca, L. M. Ricciardi, and R. Vasudevan. Note on relating Pfaffians and Hafnians with determinants and permanents. Journal of Mathematical Physics, 11(2):530–535, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i2/p530_s1.

Rosen:1970:ESS

[71] Gerald Rosen. Exact stationary states in local relativistic field theories. Journal of Mathematical Physics, 11(2):536–538, February 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p536_
s1.

Hurt:1970:RMT

[72] Norman E. Hurt. Remarks on Morse theory in canonical quantization. Journal of Mathematical Physics, 11(2):539–551, February 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p539_
s1.

Degasperis:1970:IPK

[73] Antonio Degasperis. On the inverse problem for the Klein–Gordon Swave equation. Journal of Mathematical Physics, 11(2):551–567, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i2/p551_s1.

Inonu:1970:OSP

[74] Erdal Inönü. Orthogonality of a set of polynomials encountered in neutron-transport and radiative-transfer theories. Journal of Mathematical Physics, 11(2):568–577, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i2/p568_s1.

Hallidy:1970:GRC

[75] William H. Hallidy and Allen I. Janis. Gravitational radiation: Cutting the tail. Journal of Mathematical Physics, 11(2):578–589, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i2/p578_s1.

Ott:1970:AMD

[76] R. H. Ott. Alternative method for deriving the Fock currents. Journal of Mathematical Physics, 11(2):590–595, February 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i2/p590_s1.

Deser:1970:STC

[77] S. Deser and L. K. Morrison. Stress-tensor commutators in nonlinear electrodynamics. Journal of Mathematical Physics, 11(2):596–602, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i2/p596_s1.

Dasgupta:1970:CP

[78] Brahmananda Dasgupta and Padmanabha Dasgupta. Correlations in plasma. I. Journal of Mathematical Physics, 11(2):603–609, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i2/p603_s1.

Klauder:1970:EHS

[79] John R. Klauder. Exponential Hilbert space: Fock space revisited. Journal of Mathematical Physics, 11(2):609–630, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/p609_s1.

Garrison:1970:GRL

[80] John C. Garrison, Harry L. Morrison, and Jack Wong. Galilean relativity, locality, and quantum hydrodynamics. Journal of Mathematical Physics, 11(2):630–634, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p630_s1.

Montroll:1970:QTN

[81] Elliott W. Montroll. Quantum theory on a network. I. A solvable model whose wavefunctions are elementary functions. Journal of Mathematical Physics, 11(2):635–648, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p635_s1.

Morawitz:1970:VMP

[82] Hans Morawitz. Variational methods in potential scattering. Journal of Mathematical Physics, 11(2):649–656, February 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i2/p649_s1.

Hosokawa:1970:FRW

[83] Iwao Hosokawa. Functional random-walk model of the many-particle system. Journal of Mathematical Physics, 11(2):657–666, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i2/
p657_s1.

Goldstein:1970:SDE

[84] M. E. Goldstein. Solution of the differential equation (∂^{2}/(∂x∂y) + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0 subject to the initial condition P (x, y, t = 0) = Φ(x, y). Journal of Mathematical Physics, 11(2):667– 668, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p667_s1.

Witriol:1970:MBP

[85] Norman M. Witriol. Many-body point transforms. Journal of Mathematical Physics, 11(2):669–680, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i2/p669_s1.

Girardeau:1970:FMB

[86] M. D. Girardeau. Formulation of the many-body problem for composite particles. II. Equivalence of physical and ideal state spaces. Journal of Mathematical Physics, 11(2):681–683, February 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i2/p681_s1.

Girardeau:1970:POF

[87] M. D. Girardeau. Pair occupation, Fermi condensation and phase transitions in many- Fermion systems. Journal of Mathematical Physics, 11(2):684–698, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i2/p684_s1.

Fivel:1970:SLM

[88] Daniel I. Fivel. Solution of the Lee model in all sectors by dynamical algebra. Journal of Mathematical Physics, 11(2):699–705, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i2/p699_s1.

Akyeampong:1970:RBP

[89] D. A. Akyeampong, J. F. Boyce, and M. A. Rashid. Relation between the O(4) and O(3, 1) partial-wave expansions. Journal of Mathematical Physics, 11(2):706–714, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p706_s1.

Huang:1970:GFB

[90] Justin C. Huang and Brian DeFacio. On the Green’s functions for the Bethe–Salpeter equation. Journal of Mathematical Physics, 11(2):715– 724, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p715_s1. See errata [857].

Garrettson:1970:GFM

[91] G. Garrettson and A. Leonard. Green’s functions for multidimensional neutron transport in a slab. Journal of Mathematical Physics, 11(2):725– 740, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p725_s1.

Herzel:1970:CGF

[92] John C. Herzel. Classical Green’s functions in response theory. Journal of Mathematical Physics, 11(3):741–747, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i3/p741_s1.

Joseph:1970:GIN

[93] A. Joseph and A. I. Solomon. Global and infinitesimal nonlinear chiral transformations. Journal of Mathematical Physics, 11(3):748–761, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i3/p748_s1.

Cawley:1970:SRC

[94] Robert G. Cawley. Some remarks on classical scalar radiation from point sources. Journal of Mathematical Physics, 11(3):761–765, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i3/p761_s1.

Schnatz:1970:TGN

[95] T. W. Schnatz and C. E. Siewert. Two-group neutron transport theory in spherical geometry. Journal of Mathematical Physics, 11(3):766– 771, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p766_s1.

Sancho:1970:TFM

[96] N. G. F. Sancho. Technique for finding the moment equations of a nonlinear stochastic system. Journal of Mathematical Physics, 11(3):771– 774, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p771_s1.

Mittra:1970:SGW

[97] R. Mittra and S. W. Lee. On the solution of a generalized Wiener– Hopf equation. Journal of Mathematical Physics, 11(3):775–783, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i3/p775_s1.

Baxter:1970:CHL

[98] R. J. Baxter. Colorings of a hexagonal lattice. Journal of Mathematical Physics, 11(3):784–789, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i3/p784_s1.

Griffiths:1970:CMI

[99] Robert B. Griffiths, C. A. Hurst, and S. Sherman. Concavity of magnetization of an Ising ferromagnet in a positive external field. Journal of Mathematical Physics, 11(3):790–795, March 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i3/p790_s1.

Bender:1970:GAO

[100] Carl M. Bender. Generalized anharmonic oscillator. Journal of Mathematical Physics, 11(3):796–804, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i3/p796_s1.

Cox:1970:ISP

[101] Joseph R. Cox and Kerris W. Thompson. On the inverse scattering problem at fixed energy for potentials having nonvanishing first moments. Journal of Mathematical Physics, 11(3):805–814, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i3/p805_s1.

Cox:1970:NUS

[102] Joseph R. Cox and Kerris W. Thompson. Note on the uniqueness of the solution of an equation of interest in the inverse scattering problem. Journal of Mathematical Physics, 11(3):815–817, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i3/p815_s1.

Collinson:1970:CCE

[103] C. D. Collinson. Curvature collineations in empty space–times. Journal of Mathematical Physics, 11(3):818–819, March 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i3/p818_s1.

Komar:1970:GRT

[104] Arthur Komar. General relativistic thin-sandwich theorem. Journal of Mathematical Physics, 11(3):820–821, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i3/p820_s1.

Reeken:1970:SMP

[105] M. Reeken. Summability methods in perturbation theory. Journal of Mathematical Physics, 11(3):822–824, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i3/p822_s1.

Wilhelm:1970:ASB

[106] H. E. Wilhelm. Analytical solution of the boundary-value problem for the nonlinear Helmholtz equation. Journal of Mathematical Physics, 11(3): 824–826, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p824_s1.

Oitmaa:1970:CLG

[107] J. Oitmaa. Calculation of lattice Green’s functions. Journal of Mathematical Physics, 11(3):827–829, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i3/p827_s1.

Kolsrud:1970:SDE

[108] M. Kolsrud. On the solution of the differential equation (∂^{2}/∂x∂y + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0. Journal of Mathematical Physics, 11(3):829–831, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i3/p829_s1.

Pearle:1970:RIC

[109] Philip Mark Pearle. Relativistically invariant classical Hamiltonian mechanics with extra variables. Journal of Mathematical Physics, 11(3): 832–841, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p832_s1.

Runnels:1970:EFM

[110] L. K. Runnels. Exact finite method of lattice statistics. III. Dimers on the square lattice. Journal of Mathematical Physics, 11(3):842–850, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i3/p842_s1.

Horst:1970:TGT

[111] David R. Horst. Toward a general theory of measurement. I. Journal of Mathematical Physics, 11(3):851–855, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i3/p851_s1.

Baldonado:1970:TNW

[112] O. C. Baldonado and R. C. Erdmann. The theory of neutron-wave propagation through an interface. Journal of Mathematical Physics, 11(3): 855–866, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p855_s1.

Boyling:1970:BGF

[113] J. B. Boyling. Birkhoff gravitational field of Milne’s universe. Journal of Mathematical Physics, 11(3):867–870, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i3/p867_s1.

Cisneros:1970:SUS

[114] Arturo Cisneros and Harold V. McIntosh. Search for a universal symmetry group in two dimensions. Journal of Mathematical Physics, 11(3): 870–895, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p870_s1.

McIntosh:1970:DPM

[115] Harold V. McIntosh and Arturo Cisneros. Degeneracy in the presence of a magnetic monopole. Journal of Mathematical Physics, 11(3):896– 916, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p896_s1.

Bogy:1970:HCW

[116] D. B. Bogy and P. M. Naghdi. On heat conduction and wave propagation in rigid solids. Journal of Mathematical Physics, 11(3):917– 923, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p917_s1.

Jackson:1970:INS

[117] J. C. Jackson. Infilling for NUT space. Journal of Mathematical Physics, 11(3):924–930, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i3/p924_s1.

Aronson:1970:TMS

[118] Raphael Aronson. Transfer matrix solution of some one- and twomedium transport problems in slab geometry. Journal of Mathematical Physics, 11(3):931–940, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i3/p931_s1.

Boillat:1970:NEL

[119] Guy Boillat. Nonlinear electrodynamics: Lagrangians and equations of motion. Journal of Mathematical Physics, 11(3):941–951, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i3/
p941_s1.

Kruskal:1970:KVE

[120] Martin D. Kruskal, Robert M. Miura, Clifford S. Gardner, and Norman J. Zabusky. Korteweg–de Vries equation and generalizations. V. Uniqueness and nonexistence of polynomial conservation laws. Journal of Mathematical Physics, 11(3):952–960, March 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i3/p952_s1.

Goddard:1970:NRS

[121] Peter Goddard. Nonphysical region singularities of the S matrix. Journal of Mathematical Physics, 11(3):960–974, March 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i3/p960_s1.

Li:1970:BEE

[122] Shou Yung Li, Abraham Klein, and R. M. Dreizler. Boson expansions for an exactly soluble model of interacting fermions with SU(3) symmetry. Journal of Mathematical Physics, 11(3):975–985, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i3/p975_s1.

Cook:1970:TCP

[123] David M. Cook. Trajectory of charged particles in weakly inhomogeneous magnetic fields. Journal of Mathematical Physics, 11(3):986–994, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i3/p986_s1.

Hejtmanek:1970:ALT

[124] H. J. Hejtmanek. Approximation of linear transport processes. Journal of Mathematical Physics, 11(3):995–1000, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i3/p995_s1.

Doleschall:1970:AIT

[125] P. Doleschall and J. Révai. Application of iteration technique to the integral equations of scattering theory. Journal of Mathematical Physics, 11(3):1001–1002, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i3/p1001_s1.

Coope:1970:ICTa

[126] J. A. R. Coope and R. F. Snider. Irreducible Cartesian tensors. II. General formulation. Journal of Mathematical Physics, 11(3):1003– 1017, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p1003_s1.

Thompson:1970:IML

[127] Colin J. Thompson, Arnold J. F. Siegert, and David J. Vezzetti. On the Ising model with long-range interaction. II. Critical-region analysis. Journal of Mathematical Physics, 11(3):1018–1028, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i3/p1018_s1.

Kunkin:1970:STC

[128] William Kunkin and Jerome K. Percus. Space-temperature correlations in quantum-statistical mechanics. Journal of Mathematical Physics, 11(3):1029–1036, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i3/p1029_s1.

Gudder:1970:SPP

[129] Stanley P. Gudder. A superposition principle in physics. Journal of Mathematical Physics, 11(3):1037–1040, March 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i3/p1037_s1.

Kastrup:1970:SCS

[130] H. A. Kastrup and D. H. Mayer. On some classes of solutions of the wave equation ∂t^{2}f − Δf = 0. Journal of Mathematical Physics, 11(3):1041– 1047, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p1041_s1.

Anderson:1970:EPF

[131] N. Anderson and A. M. Arthurs. Extremum principles for the Föppl– Hencky equation. Journal of Mathematical Physics, 11(3):1048–1049, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p1048_s1.

Anderson:1970:CGC

[132] R. L. Anderson, R. Raczka, M. A. Rashid, and P. Winternitz. Clebsch– Gordan coefficients for the coupling of SL(2, C) principal-series representations. Journal of Mathematical Physics, 11(3):1050–1058, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i3/p1050_s1.

Anderson:1970:RSR

[133] R. L. Anderson, R. Raczka, M. A. Rashid, and P. Winternitz. Recursion and symmetry relations for the Clebsch–Gordan coefficients of the homogeneous Lorentz group. Journal of Mathematical Physics, 11(3):1059– 1068, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p1059_s1.

Adomian:1970:ROE

[134] G. Adomian. Random operator equations in mathematical physics. I. Journal of Mathematical Physics, 11(3):1069–1084, March 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i3/p1069_

s1. See erratum [634, 715].

Chakrabarti:1970:CRC

[135] A. Chakrabarti. Crossing relations for center-of-mass- and Breit-system canonical amplitudes. Journal of Mathematical Physics, 11(3):1085– 1104, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i3/p1085_s1.

Bahar:1970:RDI

[136] E. Bahar. Radiation from a dipole imbedded in an inhomogeneous media with critical coupling regions. Journal of Mathematical Physics, 11(3):1104–1111, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i3/p1104_s1.

Ahmed:1970:BNS

[137] K. Ahmed and R. T. Sharp. O(5) bases for nuclear seniority model. Journal of Mathematical Physics, 11(3):1112–1117, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i3/p1112_s1.

Thornber:1970:EFF

[138] N. S. Thornber. Electromagnetic form factors in SL(n, C)-calculation of ultraspherical polynomials. Journal of Mathematical Physics, 11 (3):1118–1122, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i3/p1118_s1.

Santhanam:1970:REM

[139] T. S. Santhanam. Remarks on the equations of motion in odddimensional spaces and TCP invariance. Journal of Mathematical Physics, 11(3):1122, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i3/p1122_s1.

Pavkovic:1970:TMR

[140] Marko I. Pavković. On the two Majorana representations. Journal of Mathematical Physics, 11(4):1123–1125, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i4/p1123_s1.

Case:1970:TDL

[141] K. M. Case and R. D. Hazeltine. Three-dimensional linear transport theory. Journal of Mathematical Physics, 11(4):1126–1135, April 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1126_
s1.

Dashen:1970:STB

[142] Roger Dashen and Shang keng Ma. Singular three-body amplitudes in the theory of the third virial coefficient. Journal of Mathematical Physics, 11(4):1136–1143, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1136_s1.

Dickens:1970:GAM

[143] E. D. Dickens, Jr. and A. D. Levine. Generalized algebraic mass formulas. I. Journal of Mathematical Physics, 11(4):1144–1146, April 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1144_
s1.

Bettoney:1970:ESG

[144] G. V. Bettoney and R. M. Mazo. Equivalence of some generalizations on the spherical model. Journal of Mathematical Physics, 11(4):1147– 1149, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i4/p1147_s1.

Bishop:1970:ITG

[145] David M. Bishop and R. L. Somorjai. Integral-transform Gaussian functions for heliumlike systems. Journal of Mathematical Physics, 11(4):1150–1154, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1150_s1.

Warren:1970:MIA

[146] Russell E. Warren and William H. Klink. Model-independent analysis of nonrelativistic multiparticle reactions. Journal of Mathematical Physics, 11(4):1155–1164, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1155_s1.

Swamy:1970:ECR

[147] N. V. V. J. Swamy, R. G. Kulkarni, and L. C. Biedenharn. Evaluation of certain radial Coulomb integrals using symmetry properties of the Coulomb field. Journal of Mathematical Physics, 11(4):1165–1171, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i4/p1165_s1.

Molyneux:1970:EPP

[148] John E. Molyneux. Effective permittivity of a polycrystalline dielectric. Journal of Mathematical Physics, 11(4):1172–1184, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1172_s1.

Adler:1970:SWG

[149] Ronald J. Adler. Spinors in a Weyl geometry. Journal of Mathematical Physics, 11(4):1185–1191, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1185_s1.

Gargaro:1970:MER

[150] W. W. Gargaro and D. S. Onley. Matrix elements of relativistic electrons in a Coulomb field. Journal of Mathematical Physics, 11(4):1191– 1197, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i4/p1191_s1.

Briggs:1970:RME

[151] John S. Briggs. Reduced matrix elements of tensor operators. Journal of Mathematical Physics, 11(4):1198–1204, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i4/p1198_s1.

Basu:1970:HAS

[152] Debabrata Basu and S. N. Biswas. O(5) harmonics and abnormal solutions in the Bethe–Salpeter equation. Journal of Mathematical Physics, 11(4):1204–1209, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1204_s1. See erratum [635].

Brans:1970:IRA

[153] Carl H. Brans. Invariant representation of all analytic Petrov type III solutions to the Einstein equations. Journal of Mathematical Physics, 11(4):1210–1218, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1210_s1.

Moses:1970:RDP

[154] H. E. Moses. Reduction of the direct product of representations of the Poincaré group. Journal of Mathematical Physics, 11(4):1219–1234, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i4/p1219_s1.

Wilcox:1970:CFS

[155] Ralph Wilcox. Closed-form solution of the differential equation (∂^{2}/(∂x∂y) + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0 by normal-ordering exponential operators. Journal of Mathematical Physics, 11(4):1235– 1237, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i4/p1235_s1.

Dixon:1970:ANU

[156] W. G. Dixon. Analysis of the Newman–Unti integration procedure for asymptotically flat space–times. Journal of Mathematical Physics, 11(4):1238–1248, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1238_s1.

Joseph:1970:RNS

[157] David W. Joseph. Representations of noncompact semisimple Lie groups. Journal of Mathematical Physics, 11(4):1249–1271, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1249_s1.

Joseph:1970:NAA

[158] D. W. Joseph and C. J. Hieggelke. Noncompact algebras with additively labeled multiplets. Journal of Mathematical Physics, 11(4):1272– 1278, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i4/p1272_s1.

Sah:1970:SAP

[159] Priyamvada Sah. Statistical average of a product of phase sums arising in the study of disordered lattice. II. Journal of Mathematical Physics, 11(4):1279–1282, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1279_s1.

Sancho:1970:NSD

[160] N. G. F. Sancho. Nonlinear stochastic differential equations containing random parameters with small and large correlation time. Journal of Mathematical Physics, 11(4):1283–1287, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i4/p1283_s1.

Cunningham:1970:QFA

[161] M. J. Cunningham and B. G. Wybourne. Quasiparticle formalism and atomic shell theory. II. Mixed configurations. Journal of Mathematical Physics, 11(4):1288–1295, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1288_s1.

Liboff:1970:RFD

[162] Richard L. Liboff. Relativistic fields due to a particle in a grounded cylindrical box. Journal of Mathematical Physics, 11(4):1295–1306, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v11/i4/p1295_s1. See errata [834].

Gillis:1970:END

[163] Joseph E. Gillis and George H. Weiss. Expected number of distinct sites visited by a random walk with an infinite variance. Journal of Mathematical Physics, 11(4):1307–1312, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i4/p1307_s1.

Goscinski:1970:NSO

[164] O. Goscinski and P. Lindner. Natural spin-orbitals and generalized overlap amplitudes. Journal of Mathematical Physics, 11(4):1313–1317, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i4/p1313_s1.

Yeh:1970:SSO

[165] Robert H. T. Yeh. Some simple observations on Griffiths’ theorems for the classical Heisenberg model. Journal of Mathematical Physics, 11(4):1317–1319, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1317_s1.

Symon:1970:AIL

[166] K. R. Symon. The adiabatic invariant of the linear or nonlinear oscillator. Journal of Mathematical Physics, 11(4):1320–1330, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1320_s1.

Aichelburg:1970:HSF

[167] P. C. Aichelburg. High symmetry fields and the homogeneous field in general relativity. Journal of Mathematical Physics, 11(4):1330–1335, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i4/p1330_s1.

Anderson:1970:STD

[168] D. L. T. Anderson and G. H. Derrick. Stability of time-dependent particlelike solutions in nonlinear field theories. I. Journal of Mathematical Physics, 11(4):1336–1346, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1336_s1.

Nieto:1970:ESF

[169] Michael Martin Nieto. Exact state and fugacity equations for the ideal quantum gases. Journal of Mathematical Physics, 11(4):1346–1354, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i4/p1346_s1.

Nelson:1970:LCM

[170] T. J. Nelson and R. H. Good, Jr. Lorentz-covariant matrices for elementary-particle theories as polynomials in the spin matrices. Journal of Mathematical Physics, 11(4):1355–1359, April 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1355_s1.

Mathews:1970:IFC

[171] P. M. Mathews. SO^{+}(n, 1)-invariant finite-component wave equations and reflection invariance. Journal of Mathematical Physics, 11(4):1360– 1364, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i4/p1360_s1.

Sirovich:1970:AEM

[172] Lawrence Sirovich. Asymptotic evaluation of multidimensional integrals. Journal of Mathematical Physics, 11(4):1365–1374, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1365_s1.

Minton:1970:GHF

[173] Barry M. Minton. Generalized hypergeometric function of unit argument. Journal of Mathematical Physics, 11(4):1375–1376, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1375_s1.

Garrido:1970:SCS

[174] L. M. Garrido and L. Navarro. Strong coupling solution to Schrödinger equation: The mixing of states. Journal of Mathematical Physics, 11(4):1377–1381, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1377_s1.

Cahill:1970:SSMa

[175] Michael E. Cahill and G. C. McVittie. Spherical symmetry and massenergy in general relativity. I. General theory. Journal of Mathematical Physics, 11(4):1382–1391, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1382_s1.

Cahill:1970:SSMb

[176] Michael E. Cahill and G. C. McVittie. Spherical symmetry and massenergy in general relativity II. Particular cases. Journal of Mathematical Physics, 11(4):1392–1401, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i4/p1392_s1.

Mulherin:1970:CSS

[177] Denis Mulherin and I. I. Zinnes. Coulomb scattering. I. Single channel. Journal of Mathematical Physics, 11(4):1402–1408, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1402_s1.

Underhill:1970:SME

[178] J. Underhill. Off-shell T -matrix elements in potential scattering. Journal of Mathematical Physics, 11(4):1409–1419, April 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1409_s1.

Davidson:1970:RQSb

[179] Russell Davidson and John J. Kozak. Relaxation to quantum statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation field. II. Finite systems. Journal of Mathematical Physics, 11(4):1420–1436, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1420_s1.

Wait:1970:ORE

[180] James R. Wait. Oblique reflection of an electromagnetic plane wave from a striated impedance surface. Journal of Mathematical Physics, 11(4):1437–1440, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1437_s1.

Coster:1970:PRDa

[181] Joseph Coster and Henry P. Stapp. Physical-region discontinuity equations for many-particle scattering amplitudes. II. Journal of Mathematical Physics, 11(4):1441–1463, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i4/p1441_s1.

Doebner:1970:LDG

[182] H. D. Doebner and O. Melsheimer. Limitable dynamical groups in quantum mechanics. II. A model including arbitrary spin. Journal of Mathematical Physics, 11(4):1463–1471, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i4/p1463_s1.

Brunet:1970:BR

[183] M. Brunet and M. Resnikoff. Branching rules for U(n) ⊃ SO(n). Journal of Mathematical Physics, 11(4):1471–1473, April 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1471_s1.

Brunet:1970:R

[184] M. Brunet and M. Resnikoff. The representations of U(4) ⊃ U(2)⊗U(2). Journal of Mathematical Physics, 11(4):1474–1481, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1474_s1.

Boillat:1970:SWD

[185] Guy Boillat. Simple waves in N-dimensional propagation. Journal of Mathematical Physics, 11(4):1482–1483, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i4/p1482_s1.

Nagel:1970:EIS

[186] John G. Nagel and K. Tahir Shah. Expansion of the inhomogeneous symplectic Lie algebras T(2n)\frac{←}{+}Sp(n) to Sp(n + 2). Journal of Mathematical Physics, 11(4):1483–1488, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i4/p1483_s1.

Wong:1970:BLI

[187] M. K. F. Wong. Branching laws, inner multiplicities, and decomposition of classical groups. Journal of Mathematical Physics, 11(4):1489– 1495, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i4/p1489_s1.

Peacher:1970:DSF

[188] Jerry L. Peacher and Kenneth M. Watson. Doppler shift in frequency in the transport of electromagnetic waves through an underdense plasma. Journal of Mathematical Physics, 11(4):1496–1504, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i4/p1496_s1.

Pajas:1970:RLG

[189] P. Pajas and P. Winternitz. Representations of the Lorentz group: new integral relations between Legendre functions. Journal of Mathematical Physics, 11(5):1505–1510, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1505_s1.

Bownds:1970:NGS

[190] John M. Bownds. Nonexistence of global solutions to the nonhomogeneous wave equation, regardless of boundary conditions. Journal of Mathematical Physics, 11(5):1511–1512, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i5/p1511_s1.

Davies:1970:LDC

[191] John A. Davies. Lattice dynamics of cubic lattices with long-range interactions. Journal of Mathematical Physics, 11(5):1513–1520, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v11/i5/p1513_s1. See erratum [413].

Yeh:1970:DLB

[192] Robert H. T. Yeh. Derivation of a lower bound on the free energy. Journal of Mathematical Physics, 11(5):1521, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i5/p1521_s1.

Lehman:1970:DEW

[193] G. W. Lehman. Diffraction of electromagnetic waves by planar dielectric structures. I. Transverse electric excitation. Journal of Mathematical Physics, 11(5):1522–1535, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1522_s1.

Bradley:1970:KPS

[194] C. J. Bradley and B. L. Davies. Kronecker products and symmetrized squares of irreducible representations of space groups. Journal of Mathematical Physics, 11(5):1536–1552, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i5/p1536_s1.

Belleni-Morante:1970:NTN

[195] A. Belleni-Morante. Neutron transport in a nonuniform slab with generalized boundary conditions. Journal of Mathematical Physics, 11(5): 1553–1558, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i5/p1553_s1.

McDonald:1970:LSH

[196] S. C. McDonald. Localized states on a hyperplane. Journal of Mathematical Physics, 11(5):1558–1566, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i5/p1558_s1.

Dean:1970:BSS

[197] Edward T. Dean and Paul L. Chambré. Branching of solutions to some nonlinear eigenvalue problems. Journal of Mathematical Physics, 11(5): 1567–1574, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i5/p1567_s1.

Kagiwada:1970:ESF

[198] H. Kagiwada and R. Kalaba. Exact solution of a family of integral equations of anisotropic scattering. Journal of Mathematical Physics, 11(5): 1575–1578, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i5/p1575_s1.

Katzin:1970:GCC

[199] Gerald H. Katzin, Jack Levine, and William R. Davis. Groups of curvature collineations in Riemannian space–times which admit fields of parallel vectors. Journal of Mathematical Physics, 11(5):1578–1580, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i5/p1578_s1.

Iverson:1970:P

[200] G. J. Iverson and G. Mack. Ē2-parametrization of SL(2, C). Journal of Mathematical Physics, 11(5):1581–1584, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i5/p1581_s1.

Lin:1970:PDC

[201] Ta-Feng F. Lin. Problem of the disordered chain. Journal of Mathematical Physics, 11(5):1584–1590, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i5/p1584_s1.

Coope:1970:ICTb

[202] J. A. R. Coope. Irreducible Cartesian tensors. III. Clebsch–Gordan reduction. Journal of Mathematical Physics, 11(5):1591–1612, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i5/
p1591_s1.

Minton:1970:IAF

[203] B. M. Minton. Inequalities for Appell functions. Journal of Mathematical Physics, 11(5):1613–1616, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1613_s1.

Ludwig:1970:DCC

[204] Donald Ludwig. Diffraction by a circular cavity. Journal of Mathematical Physics, 11(5):1617–1630, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1617_s1.

Moshinsky:1970:NGS

[205] M. Moshinsky and C. Quesne. Noninvariance groups in the secondquantization picture and their applications. Journal of Mathematical Physics, 11(5):1631–1639, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1631_s1.

Alderman:1970:GDC

[206] Jerry L. Alderman and Otto Bergmann. Geometrical derivation of the conservation laws. Journal of Mathematical Physics, 11(5):1639–1644, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i5/p1639_s1.

Goenner:1970:LIG

[207] Hubert Goenner. Lorentz-invariant gravitational perturbations and the evaluation of generalized Green’s functions. Journal of Mathematical Physics, 11(5):1645–1655, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1645_s1.

Emch:1970:BES

[208] Gérard G. Emch, Hubert J. F. Knops, and Edward J. Verboven. Breaking of Euclidean symmetry with an application to the theory of crystallization. Journal of Mathematical Physics, 11(5):1655–1668, May 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i5/p1655_
s1.

Castell:1970:RMD

[209] L. Castell. Reduction of the most degenerate representations of SO_{0}(p, q) with respect to SO_{0}(p − 1, q − 1)[multiplicationinrighthalfcircle]T_{p+q−2}. Journal of Mathematical Physics, 11(5):1669–1672, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i5/p1669_s1.

Buell:1970:ESF

[210] June Buell, John Casti, Robert Kalaba, and Sueo Ueno. Exact solution of a family of matrix integral equations for multiply scattered partially polarized radiation. II. Journal of Mathematical Physics, 11(5):1673– 1678, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i5/p1673_s1.

Lerner:1970:SMP

[211] E. C. Lerner, H. W. Huang, and G. E. Walters. Some mathematical properties of oscillator phase operators. Journal of Mathematical Physics, 11(5):1679–1684, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1679_s1.

Pogorzelski:1970:ESE

[212] Ronald J. Pogorzelski. Electromagnetic scattering from an expanding dielectric slab. Journal of Mathematical Physics, 11(5):1685–1689, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i5/p1685_s1.

Kolsrud:1970:RCS

[213] Marius Kolsrud. Radiation from a “classical spinning electron”. Journal of Mathematical Physics, 11(5):1690, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i5/p1690_s1.

Galindo:1970:CTM

[214] V. Galindo, N. Amitay, and Y. C. Lee. Conservation theorem and the method of moments in quantum scattering. Journal of Mathematical Physics, 11(5):1691–1692, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1691_s1.

Schweinler:1970:OM

[215] H. C. Schweinler and E. P. Wigner. Orthogonalization methods. Journal of Mathematical Physics, 11(5):1693–1694, May 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i5/p1693_s1.

Corbella:1970:ISP

[216] O. D. Corbella. Inverse scattering problem for Dirac particles. Explicit expressions for the values of the potentials and their derivatives at the origin in terms of the scattering and bound-state data. Journal of Mathematical Physics, 11(5):1695–1713, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i5/p1695_s1.

Toton:1970:STTa

[217] Edward T. Toton. Scalar-tensor theory in canonical form. I. The question of Mach’s Principle. Journal of Mathematical Physics, 11(5):1713– 1724, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i5/p1713_s1.

Gupta:1970:SLG

[218] Amritava Gupta. Surface layer in general relativity and integral form of field law. Journal of Mathematical Physics, 11(5):1725–1729, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i5/p1725_s1.

Franke:1970:SDB

[219] Wilhelm H. Franke and Andrés J. Kálnay. Symmetric Dirac bracket in classical mechanics. Journal of Mathematical Physics, 11(5):1729– 1734, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i5/p1729_s1.

Noerdlinger:1970:AST

[220] Peter D. Noerdlinger. Analytic solutions to two albedo problems. Journal of Mathematical Physics, 11(5):1735–1737, May 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i5/p1735_s1.

Wilhelmsson:1970:EIW

[221] H. Wilhelmsson, L. Stenflo, and F. Engelmann. Explosive instabilities in the well-defined phase description. Journal of Mathematical Physics, 11(5):1738–1742, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1738_s1.

Hendry:1970:AMM

[222] W. L. Hendry. Application of the method of matched asymptotic expansions to a problem in linear transport theory. Journal of Mathematical Physics, 11(5):1743–1749, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i5/p1743_s1.

Dobson:1970:ASA

[223] Peter N. Dobson, Jr. Analyticity of the scattering amplitude in field theories with exponentially decreasing commutators in the spacelike region. Journal of Mathematical Physics, 11(5):1750–1758, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i5/p1750_s1.

Mukunda:1970:URG

[224] N. Mukunda. Unitary representations of the group SL(3, C), in an SU(3) basis. Journal of Mathematical Physics, 11(5):1759–1771, May 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i5/p1759_
s1.

Theumann:1970:QCS

[225] Alba Theumann. Quantum corrections to the second virial coefficient for a square-well potential. Journal of Mathematical Physics, 11(5):1772– 1778, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i5/p1772_s1.

Nagel:1970:EIP

[226] John G. Nagel. Expansion of the inhomogeneous pseudo-unitary symplectic Lie algebras IUSp(2ν_{1}, 2ν_{2}) to USP (2ν_{1} + 2, 2ν_{2}) or to USp(2ν_{1}, 2ν_{2} + 2). Journal of Mathematical Physics, 11(5):1779–1782, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i5/p1779_s1.

Gruber:1970:RRM

[227] B. Gruber. Recurrence relations for the multiplicities in the classical groups. Journal of Mathematical Physics, 11(5):1783–1790, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i5/
p1783_s1.

Gutzwiller:1970:ESA

[228] Martin C. Gutzwiller. Energy spectrum according to classical mechanics. Journal of Mathematical Physics, 11(6):1791–1806, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1791_s1.

Leaf:1970:FGG

[229] Boris Leaf. Further generalization of the generalized master equation. Journal of Mathematical Physics, 11(6):1806–1814, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1806_s1.

Chan:1970:GGT

[230] Choi-Lai Chan. Generalized group-theoretical analysis of spontaneous symmetry breaking. Journal of Mathematical Physics, 11(6):1815–1824, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i6/p1815_s1.

Hermann:1970:CAS

[231] Robert Hermann. Current algebras, the Sugawara model, and differential geometry. Journal of Mathematical Physics, 11(6):1825–1829, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i6/p1825_s1.

Hetherington:1970:NAI

[232] J. H. Hetherington. Numerical analysis of an integral equation with a delta function in the kernel. Journal of Mathematical Physics, 11(6): 1829–1834, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i6/p1829_s1.

Bincer:1970:ISC

[233] Adam M. Bincer. Interpretation of the symmetry of the Clebsch–Gordan coefficients discovered by Regge. Journal of Mathematical Physics, 11(6): 1835–1844, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i6/p1835_s1.

Girardeau:1970:PSL

[234] M. D. Girardeau. Properties of a simplified liquid ^{4}He ground state. Journal of Mathematical Physics, 11(6):1845–1852, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1845_s1.

Gilmore:1970:SRO

[235] Robert Gilmore. Spin representations of the orthogonal groups. Journal of Mathematical Physics, 11(6):1853–1854, June 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1853_s1.

Gilmore:1970:SCI

[236] Robert Gilmore. Spectrum of Casimir invariants for the simple classical Lie groups. Journal of Mathematical Physics, 11(6):1855–1856, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i6/p1855_s1.

Rubin:1970:TPI

[237] Robert J. Rubin. Transmission properties of an isotopically disordered one-dimensional harmonic crystal. II. Solution of a functional equation. Journal of Mathematical Physics, 11(6):1857–1867, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1857_s1.

Sewell:1970:ULO

[238] G. L. Sewell. Unbounded local observables in quantum statistical mechanics. Journal of Mathematical Physics, 11(6):1868–1884, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/
p1868_s1.

Good:1970:SPC

[239] I. J. Good. Short proof of a conjecture by Dyson. Journal of Mathematical Physics, 11(6):1884, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i6/p1884_s1.

Resnikoff:1970:DDP

[240] M. Resnikoff. Degeneracy of the SU(3) direct product and the symmetric representations of SU(6) ⊃ SU(3) ⊗ SU(2). Journal of Mathematical Physics, 11(6):1885–1893, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i6/p1885_s1.

Emery:1970:EAN

[241] V. J. Emery. Eikonal approximation for nonlinear equations. Journal of Mathematical Physics, 11(6):1893–1900, June 1970. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i6/p1893_s1.

Tripathy:1970:QTG

[242] Kishor C. Tripathy. Quantum theory of the generalized wave equations. I. Journal of Mathematical Physics, 11(6):1901–1912, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1901_s1.

Wu:1970:CAP

[243] F. Y. Wu and M. K. Chien. Convolution approximation for the n-particle distribution function. Journal of Mathematical Physics, 11(6):1912–1916, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i6/p1912_s1.

Carmeli:1970:IDR

[244] Moshe Carmeli. Infinite-dimensional representations of the Lorentz group. Journal of Mathematical Physics, 11(6):1917–1918, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/
p1917_s1.

Hill:1970:IIR

[245] Robert Nyden Hill. Instantaneous interaction relativistic dynamics for two particles in one dimension. Journal of Mathematical Physics, 11(6): 1918–1937, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i6/p1918_s1.

Baraff:1970:GFN

[246] G. A. Baraff. Green’s function for the nonlocal wave equation. Journal of Mathematical Physics, 11(6):1938–1941, June 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1938_s1. See er-

ratum [939].

Liu:1970:SGH

[247] Tsu-Huei Liu and Robert L. Zimmerman. Scattering in general higher sectors of the Lee model. Journal of Mathematical Physics, 11(6):1941– 1954, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i6/p1941_s1.

Geroch:1970:MMF

[248] Robert Geroch. Multipole moments. I. Flat space. Journal of Mathematical Physics, 11(6):1955–1961, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i6/p1955_s1.

Reents:1970:PTP

[249] G. Reents. Partial-tensor-product representations of the canonical commutation relations: Classification. Journal of Mathematical Physics, 11(6):1961–1967, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i6/p1961_s1.

Censor:1970:SEW

[250] Dan Censor. Scattering of electromagnetic waves in uniformly moving media. Journal of Mathematical Physics, 11(6):1968–1976, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i6/p1968_s1.

Multhei:1970:IVP

[251] Heinrich N. Mülthei. Initial-value problem for the equation (∂/∂t + a(t, x, y)∂/∂x + b(t, x, y)∂/∂y + c(t, x, y) + d(t, x, y)∂^{2}/∂x∂y)u = f(t, xy) in the complex domain. Journal of Mathematical Physics, 11(6):1977– 1980, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i6/p1977_s1.

Rosen:1970:FTS

[252] Gerald Rosen and Gene Eckstut. Fluctuation theory for space-charge-limited currents. Journal of Mathematical Physics, 11(6):1981–1984, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i6/p1981_s1.

Eringen:1970:RTM

[253] A. Cemal Eringen and Charles B. Kafadar. Relativistic theory of microelectromagnetism. Journal of Mathematical Physics, 11(6):1984–1991, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i6/p1984_s1.

Engelke:1970:NEC

[254] Ray Engelke and Charles L. Beckel. Nonuniqueness of the energy correction in application of the WKB approximation to radial problems. Journal of Mathematical Physics, 11(6):1991–1994, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1991_s1.

Kantor:1970:ISC

[255] Paul B. Kantor. Internal symmetry cannot be concealed. Journal of Mathematical Physics, 11(6):1995–1996, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i6/p1995_s1.

Mathews:1970:MVF

[256] P. M. Mathews. Massive vector field: covariant definition of independent components and quantization. Journal of Mathematical Physics, 11(6): 1997–1999, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i6/p1997_s1.

Kalaba:1970:BVP

[257] Robert Kalaba. Boundary-value problems for the integro-differential equations of nonlocal wave interaction. I. Journal of Mathematical Physics, 11(6):1999–2004, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i6/p1999_s1.

Ohkuro:1970:DFM

[258] Shigeru Ohkuro. Differential forms and Maxwell’s field: An application of harmonic integrals. Journal of Mathematical Physics, 11(6):2005– 2012, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i6/p2005_s1.

Chela-Flores:1970:MRP

[259] Julian Chela-Flores. Mandelstam representation in potential scattering. Journal of Mathematical Physics, 11(7):2013–2015, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2013_s1.

Bakri:1970:SSFa

[260] M. M. Bakri, A. F. Ghaleb, and M. I. Hessein. De Sitter symmetric field theory. II. Tensorial and spinorial realizations. Journal of Mathematical Physics, 11(7):2015–2026, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i7/p2015_s1.

Bakri:1970:SSFb

[261] M. M. Bakri. De Sitter symmetric field theory. III. Quantization in the Tung–Weinberg basis. Journal of Mathematical Physics, 11(7):2027– 2041, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i7/p2027_s1.

Frank:1970:PCE

[262] William M. Frank and David J. Land. Peratization of a class of exponentially singular potentials. Journal of Mathematical Physics, 11(7): 2041–2057, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i7/p2041_s1.

Frank:1970:PLS

[263] William M. Frank and David J. Land. Peratization of logarithmically singular potentials. Journal of Mathematical Physics, 11(7):2058–2062, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i7/p2058_s1.

Carpenter:1970:ODP

[264] Kenneth H. Carpenter and Joseph L. Strecker. One-dimensional phase transitions in the spherical lattice gas from inverse power law potentials. Journal of Mathematical Physics, 11(7):2063–2070, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2063_s1.

Eftimiu:1970:ASU

[265] C. Eftimiu. Approximate solutions of the unitarity equation. Journal of Mathematical Physics, 11(7):2070–2074, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i7/p2070_s1.

Muraskin:1970:MEH

[266] M. Muraskin. Maxwell’s equations having a gradient as source. Journal of Mathematical Physics, 11(7):2075–2077, July 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2075_s1. See er-

ratum [524].

Wang:1970:CGS

[267] Kuo-Hsiang Wang. Clebsch–Gordan series and the Clebsch–Gordan coefficients of O(2, 1) and SU(1, 1). Journal of Mathematical Physics, 11(7): 2077–2095, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i7/p2077_s1.

Couch:1970:SIG

[268] W. E. Couch and R. J. Torrence. Self-interaction of gravitational radiation. Journal of Mathematical Physics, 11(7):2096–2114, July 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2096_
s1.

Lieh:1970:SSL

[269] Jong-Won Lieh. Some simple lattice-spin systems. Journal of Mathematical Physics, 11(7):2114–2127, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i7/p2114_s1.

Weinhold:1970:CAA

[270] Frank Weinhold. Criteria of accuracy of approximate wavefunctions. Journal of Mathematical Physics, 11(7):2127–2138, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2127_s1.

Giles:1970:FQM

[271] Robin Giles. Foundations for quantum mechanics. Journal of Mathematical Physics, 11(7):2139–2161, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i7/p2139_s1.

Johnson:1970:FIE

[272] Rodney W. Johnson. Finite integral equations for Green’s functions for : ϕ^{4}: coupling. I. Journal of Mathematical Physics, 11(7):2161–2165, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i7/p2161_s1.

Picinbono:1970:PSN

[273] B. Picinbono, C. Bendjaballah, and J. Pouget. Photoelectron shot noise. Journal of Mathematical Physics, 11(7):2166–2176, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2166_s1.

Treat:1970:LGF

[274] Richard P. Treat. Local gauge fields in general relativity. Journal of Mathematical Physics, 11(7):2176–2187, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i7/p2176_s1.

Treat:1970:ABS

[275] Richard P. Treat. Anholonomic bases in the study of the spin connection and gauge potentials in a Riemann space. Journal of Mathematical Physics, 11(7):2187–2193, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i7/p2187_s1.

Marchioro:1970:STB

[276] C. Marchioro. Solution of a three-body scattering problem in one dimension. Journal of Mathematical Physics, 11(7):2193–2196, July 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2193_
s1.

Chambers:1970:ESH

[277] Ll. G. Chambers and R. A. Sack. Expansions in spherical harmonics. V. Solution of inverse Legendre transforms. Journal of Mathematical Physics, 11(7):2197–2202, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i7/p2197_s1.

Zerilli:1970:THC

[278] Frank J. Zerilli. Tensor harmonics in canonical form for gravitational radiation and other applications. Journal of Mathematical Physics, 11(7): 2203–2208, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i7/p2203_s1.

Seiler:1970:SSE

[279] R. Seiler and R. H. Pratt. Separable Schrödinger equations for two interacting particles in external fields. Journal of Mathematical Physics, 11(7):2208–2211, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i7/p2208_s1.

Vajk:1970:SHA

[280] J. Peter Vajk and Peter G. Eltgroth. Spatially homogeneous anisotropic cosmological models containing relativistic fluid and magnetic field. Journal of Mathematical Physics, 11(7):2212–2222, July 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i7/p2212_s1.

Humi:1970:WEU

[281] Mayer Humi. Wave equations for unstable particles. Journal of Mathematical Physics, 11(7):2222–2224, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i7/p2222_s1.

Srivastava:1970:ANT

[282] H. M. Srivastava. Another note on two binomial-coefficient identities of Rosenbaum. Journal of Mathematical Physics, 11(7):2225, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i7/
p2225_s1.

Gorini:1970:IS

[283] Vittorio Gorini and Antonio Zecca. Isotropy of space. Journal of Mathematical Physics, 11(7):2226–2230, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i7/p2226_s1.

Bose:1970:CFD

[284] A. K. Bose and J. Patera. Classification of finite-dimensional irreducible representations of connected complex semisimple Lie groups. Journal of Mathematical Physics, 11(7):2231–2234, July 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i7/p2231_s1.

Baird:1970:NIF

[285] L. C. Baird. New integral formulation of the Schrödinger equation. Journal of Mathematical Physics, 11(8):2235–2242, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2235_s1.

Garrison:1970:CCP

[286] John C. Garrison and Jack Wong. Canonically conjugate pairs, uncertainty relations, and phase operators. Journal of Mathematical Physics, 11(8):2242–2249, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2242_s1.

Chan:1970:SMT

[287] Yau W. Chan. Solution of Mathieu’s type of secular equations. Journal of Mathematical Physics, 11(8):2250–2253, August 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2250_s1.

Wolf:1970:DTH

[288] E. Wolf and J. R. Shewell. Diffraction theory of holography. Journal of Mathematical Physics, 11(8):2254–2267, August 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2254_s1.

Killingbeck:1970:COA

[289] J. Killingbeck. Commuting-operator approach to group representation theory. Journal of Mathematical Physics, 11(8):2268–2272, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/
p2268_s1.

Abo-Shaeer:1970:SBE

[290] Muhsin A. R. Abo-Shaeer. O(4) symmetry and the BS equation for a spin-1 + spin-0 system. Journal of Mathematical Physics, 11(8):2272– 2279, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i8/p2272_s1.

Walker:1970:BDE

[291] Martin Walker. Block diagrams and the extension of timelike twosurfaces. Journal of Mathematical Physics, 11(8):2280–2286, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i8/p2280_s1.

Keller:1970:CQM

[292] Joseph B. Keller. Classical and quantum mechanical correlation functions of fields in thermal equilibrium. Journal of Mathematical Physics, 11(8):2286–2296, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2286_s1.

Aghassi:1970:RIS

[293] J. J. Aghassi, P. Roman, and R. M. Santilli. Relation of the inhomogeneous de Sitter group to the quantum mechanics of elementary particles. Journal of Mathematical Physics, 11(8):2297–2301, August 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2297_
s1.

Baker:1970:SMF

[294] George A. Baker, Jr. Summability of the many-fermion perturbation series with attractive forces. Journal of Mathematical Physics, 11(8):2302– 2308, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i8/p2302_s1.

Hodge:1970:EES

[295] D. B. Hodge. Eigenvalues and eigenfunctions of the spheroidal wave equation. Journal of Mathematical Physics, 11(8):2308–2312, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i8/p2308_s1.

Beers:1970:NTA

[296] Brian Lee Beers and Alex J. Dragt. New theorems about spherical harmonic expansions and SU(2). Journal of Mathematical Physics, 11(8):2313–2328, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2313_s1.

Greenberger:1970:TPVa

[297] Daniel M. Greenberger. Theory of particles with variable mass. I. Formalism. Journal of Mathematical Physics, 11(8):2329–2341, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/
p2329_s1.

Greenberger:1970:TPVb

[298] Daniel M. Greenberger. Theory of particles with variable mass. II. Some physical consequences. Journal of Mathematical Physics, 11(8):2341– 2347, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i8/p2341_s1.

McKenna:1970:MCF

[299] J. McKenna and J. A. Morrison. Moments and correlation functions of solutions of a stochastic differential equation. Journal of Mathematical Physics, 11(8):2348–2360, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i8/p2348_s1.

Morrison:1970:ASM

[300] J. A. Morrison and J. McKenna. Application of the smoothing method to a stochastic ordinary differential equation. Journal of Mathematical Physics, 11(8):2361–2367, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i8/p2361_s1.

Louck:1970:CUA

[301] James D. Louck and L. C. Biedenharn. Canonical unit adjoint tensor operators in U(n). Journal of Mathematical Physics, 11(8):2368– 2415, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i8/p2368_s1.

Grubb:1970:DWS

[302] M. Anthea Grubb and D. B. Pearson. Derivation of the wave and scattering operators for an interaction of rank one. Journal of Mathematical Physics, 11(8):2415–2424, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i8/p2415_s1.

Pearson:1970:SIO

[303] D. B. Pearson. Singular integral operator encountered in scattering theory. Journal of Mathematical Physics, 11(8):2425–2431, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/
p2425_s1.

Matzner:1970:RCP

[304] Richard A. Matzner. Rotation in closed perfect-fluid cosmologies. Journal of Mathematical Physics, 11(8):2432–2440, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2432_s1.

Oliveira:1970:CDI

[305] Colber G. Oliveira. Complex 2-dimensional internal space in general relativity. Journal of Mathematical Physics, 11(8):2440–2444, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i8/p2440_s1.

Newman:1970:SIV

[306] Perry A. Newman and Robert L. Bowden. Solution of the initial-value neutron-transport problem for a slab with infinite reflectors. Journal of Mathematical Physics, 11(8):2445–2458, August 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2445_s1.

Aichelburg:1970:CCG

[307] P. C. Aichelburg. Curvature collineations for gravitational pp waves. Journal of Mathematical Physics, 11(8):2458–2462, August 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2458_
s1.

Anderson:1970:AFM

[308] J. T. Anderson. Analytic function method for the many-body problem. Journal of Mathematical Physics, 11(8):2463–2472, August 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2463_
s1.

Hurst:1970:CIH

[309] C. A. Hurst and S. Sherman. Correlation inequalities for Heisenberg ferromagnets. Journal of Mathematical Physics, 11(8):2473–2480, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i8/p2473_s1.

Sherman:1970:WIM

[310] S. Sherman. When is an Ising magnet a ferromagnet? Journal of Mathematical Physics, 11(8):2480–2481, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i8/p2480_s1.

Robertson:1970:QSM

[311] Baldwin Robertson. Quantum statistical mechanical derivation of generalized hydrodynamic equations. Journal of Mathematical Physics, 11(8):2482–2489, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2482_s1.

Bergstrom:1970:ETCa

[312] K. Bergström. Equal-time commutation relations and divergence conditions. Journal of Mathematical Physics, 11(8):2489–2498, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/
p2489_s1.

Bergstrom:1970:ETCb

[313] K. Bergström. Equal-time commutation relation for the energymomentum tensor. Journal of Mathematical Physics, 11(8):2498–2501, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i8/p2498_s1.

Hethcote:1970:EBA

[314] Herbert W. Hethcote. Error bounds for asymptotic approximations of zeros of Hankel functions occurring in diffraction problems. Journal of Mathematical Physics, 11(8):2501–2504, August 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2501_s1.

Reeken:1970:GTB

[315] M. Reeken. General theorem on bifurcation and its application to the Hartree equation of the helium atom. Journal of Mathematical Physics, 11(8):2505–2512, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2505_s1.

Butler:1970:GRT

[316] P. H. Butler and B. G. Wybourne. Generalized Racah tensors and the structure of mixed configurations. Journal of Mathematical Physics, 11(8):2512–2518, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2512_s1.

Butler:1970:GAS

[317] P. H. Butler and B. G. Wybourne. Is the group R_{4} an approximate symmetry for many-electron theory? Journal of Mathematical Physics, 11(8):2519–2524, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2519_s1.

Mills:1970:QTN

[318] R. G. J. Mills and E. W. Montroll. Quantum theory on a network. II. A solvable model which may have several bound states per node point. Journal of Mathematical Physics, 11(8):2525–2538, August 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2525_
s1.

Dubois-Violette:1970:TFS

[319] M. Dubois-Violette. Theory of formal series with applications to quantum field theory. Journal of Mathematical Physics, 11(8):2539–2546, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i8/p2539_s1.

Case:1970:ERH

[320] K. M. Case and R. D. Hazeltine. Elastic radiation in a half-space. Journal of Mathematical Physics, 11(8):2546–2552, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2546_s1.

Benioff:1970:SAR

[321] Paul A. Benioff. Some aspects of the relationship between mathematical logic and physics. I. Journal of Mathematical Physics, 11(8):2553– 2569, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i8/p2553_s1.

Chakraborty:1970:EWP

[322] B. Chakraborty. Electromagnetic wave propagation through slowly varying plasma. Journal of Mathematical Physics, 11(8):2570–2579, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i8/p2570_s1.

Geroch:1970:MMI

[323] Robert Geroch. Multipole moments. II. Curved space. Journal of Mathematical Physics, 11(8):2580–2588, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i8/p2580_s1.

Mo:1970:TEM

[324] Tse Chin Mo. Theory of electrodynamics in media in noninertial frames and applications. Journal of Mathematical Physics, 11(8):2589–2610, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i8/p2589_s1.

Williams:1970:TAN

[325] J. G. Williams. Topological analysis of a nonlinear field theory. Journal of Mathematical Physics, 11(8):2611–2616, August 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i8/p2611_s1.

Fuchs:1970:CPG

[326] Gérard Fuchs and Pierre Renouard. Characters of the Poincaré group. Journal of Mathematical Physics, 11(9):2617–2645, September 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2617_
s1.

Lyth:1970:MM

[327] D. H. Lyth. Mathematics of the N/D method. Journal of Mathematical Physics, 11(9):2646–2678, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i9/p2646_s1.

Moore:1970:QTE

[328] Gerald T. Moore. Quantum theory of the electromagnetic field in a variable-length one-dimensional cavity. Journal of Mathematical Physics, 11(9):2679–2691, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i9/p2679_s1.

Katsura:1970:IS

[329] Shigetoshi Katsura, Katsuji Kaminishi, and Sakari Inawashiro. Intermediate statistics. Journal of Mathematical Physics, 11(9):2691–2697, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i9/p2691_s1.

Testa:1970:FIC

[330] Francis J. Testa. Finite and infinitesimal canonical transformations. Journal of Mathematical Physics, 11(9):2698–2701, September 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2698_
s1.

Heilmann:1970:ZGP

[331] Ole J. Heilmann. Zeros of the grand partition function for a lattice gas. Journal of Mathematical Physics, 11(9):2701–2704, September 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2701_
s1.

Rudd:1970:ESE

[332] R. A. Rudd and Robert Nyden Hill. Exactly solvable electrodynamic two-body problem. Journal of Mathematical Physics, 11(9):2704–2710, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i9/p2704_s1.

Bishop:1970:CSF

[333] David M. Bishop and Barbara E. Schneider. Comparison of SCF and k_{ν} functions for the helium series. Journal of Mathematical Physics, 11 (9):2711–2713, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i9/p2711_s1.

Feinblum:1970:GST

[334] David A. Feinblum. Global singularities and the Taub–NUT metric. Journal of Mathematical Physics, 11(9):2713–2721, September 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2713_
s1.

Bekenstein:1970:SEH

[335] J. D. Bekenstein and J. B. Krieger. Stark effect in hydrogen atoms for nonuniform fields. Journal of Mathematical Physics, 11(9):2721–2727, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i9/p2721_s1.

Carmeli:1970:SGF

[336] Moshe Carmeli. SL(2, C) symmetry of the gravitational field dynamical variables. Journal of Mathematical Physics, 11(9):2728–2732, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i9/p2728_s1.

Schnatz:1970:RTR

[337] T. W. Schnatz and C. E. Siewert. Radiative transfer in a Rayleighscattering atmosphere with true absorption. Journal of Mathematical Physics, 11(9):2733–2739, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i9/p2733_s1.

Enosh:1970:POS

[338] M. Enosh. Proper orientation of space–time. Journal of Mathematical Physics, 11(9):2739–2742, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i9/p2739_s1.

Coster:1970:PRDb

[339] Joseph Coster and Henry P. Stapp. Physical-region discontinuity equation. Journal of Mathematical Physics, 11(9):2743–2763, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i9/p2743_s1.

Bahar:1970:RBG

[340] E. Bahar. Relationship between geometrical-optical and full-wave solutions to the problem of propagation over a nonparallel stratified medium. Journal of Mathematical Physics, 11(9):2764–2775, September 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2764_
s1.

Stern:1970:KPM

[341] David P. Stern. Kruskal’s perturbation method. Journal of Mathematical Physics, 11(9):2771–2775, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i9/p2771_s1.

Stern:1970:DCT

[342] David P. Stern. Direct canonical transformations. Journal of Mathematical Physics, 11(9):2776–2788, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i9/p2776_s1. See erratum [526].

Cunningham:1970:SQA

[343] M. J. Cunningham. Second quantized atomic wavefunctions with definite unitary and rotational symmetry. Journal of Mathematical Physics, 11 (9):2781–2788, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i9/p2781_s1.

Liu:1970:STP

[344] C. J. Liu and Yutze Chow. Some topological properties connected to the parametrized Feynman amplitudes. Journal of Mathematical Physics, 11 (9):2789–2799, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i9/p2789_s1.

daCosta:1970:NLP

[345] R. C. T. Da Costa. Nonlinear light propagation in a resonant medium and causality. Journal of Mathematical Physics, 11(9):2799–2803, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i9/p2799_s1.

Mickelsson:1970:LMP

[346] Jouko Mickelsson. Labeling and multiplicity problem in the reduction U(n + m) ↓ U(n) × U(m). Journal of Mathematical Physics, 11(9):2803– 2806, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i9/p2803_s1.

Kerr:1970:ESS

[347] R. P. Kerr and G. C. Debney, Jr. Einstein spaces with symmetry groups. Journal of Mathematical Physics, 11(9):2807–2817, September 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2807_
s1.

Toton:1970:STTb

[348] Edward T. Toton. Scalar-tensor theory in canonical form. II. Exact solutions for a static spherically symmetric charged mass distribution. Journal of Mathematical Physics, 11(9):2818–2825, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2818_s1.

Fleming:1970:HIS

[349] Gordon N. Fleming and James T. Miller, Jr. Hyperplane independence of the scattering amplitude. Journal of Mathematical Physics, 11(9):2826– 2829, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i9/p2826_s1.

Lee:1970:RTD

[350] S. W. Lee. Ray theory of diffraction by open-ended waveguides I. Field in waveguides. Journal of Mathematical Physics, 11(9):2830–2850, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i9/p2830_s1.

Wait:1970:PEW

[351] James R. Wait. Propagation of electromagnetic waves over a smooth multisection curved earth — an exact theory. Journal of Mathematical Physics, 11(9):2851–2860, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i9/p2851_s1.

Ozsvath:1970:SHW

[352] István Ozsváth. Spatially homogeneous world models. Journal of Mathematical Physics, 11(9):2860–2870, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i9/p2860_s1.

Ozsvath:1970:DFU

[353] István Ozsváth. Dust-filled universes of class II and class III. Journal of Mathematical Physics, 11(9):2871–2883, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2871_s1.

Angelopoulos:1970:UIR

[354] E. Angelopoulos. Unitary irreducible representations of SL(n, C)·R(n^{2}). I. Journal of Mathematical Physics, 11(9):2883–2897, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/
p2883_s1.

Gallavotti:1970:TLT

[355] G. Gallavotti, O. E. Lanford, and Joel L. Lebowitz. Thermodynamic limit of time-dependent correlation functions for one-dimensional systems. Journal of Mathematical Physics, 11(9):2898–2905, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i9/p2898_s1.

Berger:1970:SSN

[356] Melvyn S. Berger. Stationary states for a nonlinear wave equation. Journal of Mathematical Physics, 11(9):2906–2912, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2906_s1.

Roskies:1970:PCC

[357] Ralph Roskies. Positivity constraints on crossing-symmetric partial-wave expansions. Journal of Mathematical Physics, 11(9):2913–2919, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i9/p2913_s1.

Keller:1970:EPI

[358] Joseph B. Keller. Extremum principles for irreversible processes. Journal of Mathematical Physics, 11(9):2919–2931, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i9/p2919_s1.

Ray:1970:SAD

[359] Supriya Ray. Suggestive approximate dynamical symmetry of an electron moving in the field of many stationary nuclei. Journal of Mathematical Physics, 11(9):2931–2934, September 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i9/p2931_s1.

Raman:1970:NRF

[360] K. Raman. Noncommutation requirements and field dependence of sources of tensor fields. Journal of Mathematical Physics, 11(10):2935– 2938, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i10/p2935_s1.

Barut:1970:RCR

[361] A. O. Barut and A. Böhm. Reduction of a class of Q(4, 2) representations with respect to SO(4, 1) and SO(3, 2). Journal of Mathematical Physics, 11(10):2938–2945, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i10/p2938_s1.

Radin:1970:AES

[362] Charles Radin. Approach to equilibrium in a simple model. Journal of Mathematical Physics, 11(10):2945–2955, October 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i10/p2945_s1.

Rastall:1970:CTS

[363] Peter Rastall. Classical thermodynamics simplified. Journal of Mathematical Physics, 11(10):2955–2966, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i10/p2955_s1.

Hostler:1970:CGF

[364] Levere C. Hostler. Coulomb Green’s function in f-dimensional space. Journal of Mathematical Physics, 11(10):2966–2970, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i10/
p2966_s1.

Nakanishi:1970:AFT

[365] Noboru Nakanishi. An axiomatic formulation of the theory of coinciding simple poles and multiple poles. Journal of Mathematical Physics, 11(10):2970–2982, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i10/p2970_s1.

Kirkwood:1970:LIG

[366] Robert L. Kirkwood. Lorentz invariance in a gravitational field. Journal of Mathematical Physics, 11(10):2983–2989, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i10/p2983_s1.

Dubin:1970:TTA

[367] D. A. Dubin and G. L. Sewell. Time translations in the algebraic formulation of statistical mechanics. Journal of Mathematical Physics, 11 (10):2990–2998, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i10/p2990_s1.

Castell:1970:ITM

[368] L. Castell. Integral transformations in momentum space and conformal invariance. Journal of Mathematical Physics, 11(10):2999–3001, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i10/p2999_s1.

McDonald:1970:RPR

[369] S. C. McDonald. Relative position representation for a relativistic twoparticle system. Journal of Mathematical Physics, 11(10):3001–3008, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i10/p3001_s1.

Emch:1970:PTP

[370] Gérard G. Emch and Hubert J. F. Knops. Pure thermodynamical phases as extremal KMS states. Journal of Mathematical Physics, 11(10):3008– 3018, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i10/p3008_s1.

Tikochinsky:1970:DPS

[371] Y. Tikochinsky. Derivatives of phase shifts and binding energies by use of variational principles. Journal of Mathematical Physics, 11(10):3019– 3026, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i10/p3019_s1.

Patera:1970:LSC

[372] J. Patera. Labeling states and constructing matrix representations of G_{2}. Journal of Mathematical Physics, 11(10):3027–3032, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i10/
p3027_s1.

Lynch:1970:RSS

[373] P. J. Lynch and R. J. Wagner. Rough-surface scattering: shadowing, multiple scatter, and energy conservation. Journal of Mathematical Physics, 11(10):3032–3042, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i10/p3032_s1.

Kanal:1970:BVP

[374] Madhoo Kanal. Boundary-value problems of linear transport theory — Green’s function approach. Journal of Mathematical Physics, 11 (10):3042–3052, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i10/p3042_s1.

Gelman:1970:GED

[375] Harry Gelman. General expansion of the determinant of the Maxwell dyadic. Journal of Mathematical Physics, 11(10):3053–3054, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i10/p3053_s1.

Raghavacharyulu:1970:PA

[376] I. V. V. Raghavacharyulu and Nalini B. Menon. Polynomial algebras. Journal of Mathematical Physics, 11(10):3055–3060, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i10/
p3055_s1.

Minton:1970:SRC

[377] B. M. Minton. Symmetries of the Racah coefficients. Journal of Mathematical Physics, 11(10):3061–3062, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i10/p3061_s1.

Childress:1970:NSK

[378] Stephen Childress. New solutions of the kinematic dynamo problem. Journal of Mathematical Physics, 11(10):3063–3076, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i10/
p3063_s1.

Gruber:1970:MCG

[379] B. Gruber. Multiplicities in the classical groups. I. Journal of Mathematical Physics, 11(10):3077–3090, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i10/p3077_s1.

Burniston:1970:LIE

[380] E. E. Burniston and C. E. Siewert. Linear integral equations for a certain class of H-functions applicable to the theory of neutron transport and radiative transfer. Journal of Mathematical Physics, 11(10):3091– 3095, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i10/p3091_s1.

McQuistan:1970:ERR

[381] R. B. McQuistan and S. J. Lichtman. Exact recursion relation for 2 × N arrays of dumbbells. Journal of Mathematical Physics, 11(10):3095– 3099, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i10/p3095_s1.

Kantor:1970:ERL

[382] Paul B. Kantor. Extension of the Riemann–Lebesgue lemma. Journal of Mathematical Physics, 11(10):3099–3103, October 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i10/p3099_s1.

McDonald:1970:DPE

[383] J. F. McDonald and L. D. Favro. Diagrammatic perturbation expansion for ensembles of random matrices. Journal of Mathematical Physics, 11(10):3103–3110, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i10/p3103_s1.

Siebert:1970:UWQ

[384] E. Siebert and J. B. Krieger. Use of the WKB quantization condition for calculating expectation values. Journal of Mathematical Physics, 11(10):3111–3115, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i10/p3111_s1.

Baxter:1970:TCS

[385] R. J. Baxter. Three-colorings of the square lattice: A hard squares model. Journal of Mathematical Physics, 11(10):3116–3124, October 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i10/p3116_s1.

Lau:1970:RTA

[386] Chau-Wa Lau and Kenneth M. Watson. Radiation transport along curved ray paths. Journal of Mathematical Physics, 11(11):3125–3137, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3125_s1.

Ifantis:1970:SPS

[387] Evangelos K. Ifantis. Structure of the point spectrum of Schrödinger-type tridiagonal operators. Journal of Mathematical Physics, 11(11):3138– 3144, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3138_s1.

Held:1970:LGS

[388] A. Held, E. T. Newman, and R. Posadas. The Lorentz group and the sphere. Journal of Mathematical Physics, 11(11):3145–3154, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i11/p3145_s1.

Newman:1970:NAM

[389] E. T. Newman and Roger Young. A new approach to the motion of a pole-dipole particle. Journal of Mathematical Physics, 11(11):3154– 3158, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3154_s1.

Scott:1970:MPT

[390] H. L. Scott. Model of phase transitions for interacting Fermi systems. Journal of Mathematical Physics, 11(11):3159–3165, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i11/
p3159_s1.

Barrett:1970:MCG

[391] E. B. Barrett. Method of constructing general solutions of certain nonlinear differential equations describing plasma oscillations. Journal of Mathematical Physics, 11(11):3165–3167, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i11/p3165_s1.

Jalickee:1970:FPS

[392] John B. Jalickee, Arnold J. F. Siegert, and David J. Vezzetti. Fluids of particles with short-range repulsion and weak long-range attractive interaction. II. The two-particle distribution function. Journal of Mathematical Physics, 11(11):3168–3176, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i11/p3168_s1.

Anderson:1970:CSF

[393] Robert Leonard Anderson and Kurt Bernardo Wolf. Complete sets of functions on homogeneous spaces with compact stabilizers. Journal of Mathematical Physics, 11(11):3176–3183, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i11/p3176_s1.

Sutherland:1970:TDH

[394] Bill Sutherland. Two-dimensional hydrogen bonded crystals without the ice rule. Journal of Mathematical Physics, 11(11):3183–3186, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i11/p3183_s1.

Wong:1970:CIM

[395] M. K. F. Wong and B. Gruber. Calculation of the inner multiplicity of weights by means of the branching law method and by Racah’s recurrence relation. Journal of Mathematical Physics, 11(11):3187–3192, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3187_s1.

Tarter:1970:CCS

[396] C. Bruce Tarter. Coefficients connecting the Stark and field-free wavefunctions for hydrogen. Journal of Mathematical Physics, 11(11):3192– 3195, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3192_s1.

Kulhanek:1970:PVS

[397] J. Kulhánek. The proper vibration of the space. Journal of Mathematical Physics, 11(11):3195–3200, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i11/p3195_s1.

Kulhanek:1970:EBW

[398] J. Kulhánek. Equations of the de Broglie wave field and their relationship to Riemann’s curvature tensor. Journal of Mathematical Physics, 11 (11):3199–3200, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i11/p3199_s1.

Morrison:1970:CCF

[399] J. A. Morrison. Calculation of correlation functions of solutions of a stochastic ordinary differential equation. Journal of Mathematical Physics, 11(11):3200–3209, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i11/p3200_s1.

Klink:1970:MBF

[400] William H. Klink. Mixed-basis D functions and Clebsch–Gordan coefficients of noncompact groups. Journal of Mathematical Physics, 11 (11):3210–3218, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i11/p3210_s1.

Ruskai:1970:RPP

[401] Mary Beth Ruskai. N-representability problem: particle-hole equivalence. Journal of Mathematical Physics, 11(11):3218–3224, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i11/p3218_s1.

Draayer:1970:PT

[402] J. P. Draayer. SU(4) ⊃ SU(2) ⊗ SU(2) projection techniques. Journal of Mathematical Physics, 11(11):3225–3238, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i11/p3225_s1.

Choquet-Bruhat:1970:TUL

[403] Yvonne Choquet-Bruhat. Theorem of uniqueness and local stability for Liouville–Einstein equations. Journal of Mathematical Physics, 11 (11):3238–3243, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i11/p3238_s1.

Yanagawa:1970:RLG

[404] S. Yanagawa and T. Moriya. Realization of the Lie group G(0, 1) by the function of Landau levels. Journal of Mathematical Physics, 11(11): 3244, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3244_s1.

Ferreira:1970:RTI

[405] Erasmo Ferreira and Javier Sesma. Regge trajectories for the inverse square potential. Journal of Mathematical Physics, 11(11):3245–3250, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3245_s1.

Johannsen:1970:ESM

[406] Günther Johannsen. Electromagnetic sources in moving simple medium. Journal of Mathematical Physics, 11(11):3251–3253, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i11/
p3251_s1.

Zagier:1970:EPF

[407] Don B. Zagier. Expansion of an n-point function as a sum of commutators. Journal of Mathematical Physics, 11(11):3253–3257, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i11/p3253_s1.

Loos:1970:ECS

[408] Hendricus G. Loos. Existence of charged states of the Yang–Mills field. Journal of Mathematical Physics, 11(11):3258–3274, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v11/i11/
p3258_s1.

vanDaele:1970:GFD

[409] A. van Daele. General formula to derive branching rules. Journal of Mathematical Physics, 11(11):3275–3282, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i11/p3275_s1.

Grimmer:1970:CMF

[410] Hans Grimmer. Cell model field theories. Journal of Mathematical Physics, 11(11):3283–3295, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i11/p3283_s1.

Cohen:1970:HOF

[411] Leon Cohen. Hamiltonian operators via Feynman path integrals. Journal of Mathematical Physics, 11(11):3296–3297, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i11/p3296_s1.

Young:1970:ARO

[412] Ralph H. Young and Walter J. Deal, Jr. Adiabatic response to an oscillatory field. Journal of Mathematical Physics, 11(11):3298–3306, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i11/p3298_s1.

Davies:1970:ELD

[413] John A. Davies. Erratum: “Lattice Dynamics of Cubic Lattices with Long-Range Interactions”. Journal of Mathematical Physics, 11(11): 3306, November 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i11/p3306_s1. See [191].

Otterson:1970:SSC

[414] P. J. Otterson and J. Sucher. Singularity structure of causal distributions and restricted equal-time limits. Journal of Mathematical Physics, 11 (12):3307–3316, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i12/p3307_s1.

Heilmann:1970:FHS

[415] Ole J. Heilmann. Finding hidden symmetry by numerical methods. Journal of Mathematical Physics, 11(12):3317–3321, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i12/p3317_s1.

Kuchar:1970:GSF

[416] Karel Kuchar. Ground state functional of the linearized gravitational field. Journal of Mathematical Physics, 11(12):3322–3334, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i12/p3322_s1.

Glimm:1970:EMS

[417] James Glimm and Arthur Jaffe. Energy-momentum spectrum and vacuum expectation values in quantum field theory. Journal of Mathematical Physics, 11(12):3335–3338, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i12/p3335_s1.

Rabitz:1970:SPT

[418] H. A. Rabitz and R. Gordon. Semiclassical perturbation theory of molecular collisions III. Graphical angular momentum methods and the nth order. Journal of Mathematical Physics, 11(12):3339–3357, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i12/p3339_s1.

Goenner:1970:ETP

[419] Hubert Goenner and John Stachel. Einstein tensor and 3-parameter groups of isometries with 2- dimensional orbits. Journal of Mathematical Physics, 11(12):3358–3370, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i12/p3358_s1.

Mendes:1970:RLS

[420] Rui V. Mendes and Yuval Ne’eman. Representations of local I-spin charge densities. I. Spaces where each I occurs only once. Journal of Mathematical Physics, 11(12):3371–3382, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i12/p3371_s1.

Perjes:1970:STS

[421] Z. Perjés. Spinor treatment of stationary space–times. Journal of Mathematical Physics, 11(12):3383–3391, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i12/p3383_s1.

Degasperis:1970:IPC

[422] Antonio Degasperis. Inverse problem for a cylindrical plasma. Journal of Mathematical Physics, 11(12):3392–3400, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i12/p3392_s1.

Glass:1970:NPC

[423] Edward N. Glass and Joshua N. Goldberg. Newman–Penrose constants and their invariant transformations. Journal of Mathematical Physics, 11(12):3400–3413, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i12/p3400_s1.

Mendelson:1970:PTD

[424] Kenneth S. Mendelson. Perturbation theory for damped nonlinear oscillations. Journal of Mathematical Physics, 11(12):3413–3415, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i12/p3413_s1.

Burniston:1970:HRE

[425] E. E. Burniston and C. E. Siewert. Half-range expansion theorems in studies of polarized light. Journal of Mathematical Physics, 11(12):3416– 3420, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i12/p3416_s1. See erratum [576].

Gilmore:1970:DTC

[426] Robert Gilmore. Diagrammatic technique for constructing matrix elements. Journal of Mathematical Physics, 11(12):3420–3427, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v11/i12/p3420_s1.

Robinson:1970:GRG

[427] G. de B. Robinson. Group representations and geometry. Journal of Mathematical Physics, 11(12):3428–3432, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v11/i12/p3428_s1.

Svetlichny:1970:GO

[428] George Svetlichny. Generalized operators. Journal of Mathematical Physics, 11(12):3433–3486, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v11/i12/p3433_s1.

Borie:1970:MHR

[429] Edith Borie. S matrix in the Heisenberg representation. Journal of Mathematical Physics, 11(12):3487–3496, December 1970. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i12/p3487_s1. See er-

rata [1045].

Pietenpol:1971:SPD

[430] J. L. Pietenpol and D. Speiser. Simple procedure for determining the number of components of an irreducible tensor. Journal of Mathematical Physics, 12(1):1–4, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i1/p1_s1. See addendum [636].

Kummer:1971:CPT

[431] Martin Kummer. Class of perturbation theories of ordinary differential equations. Journal of Mathematical Physics, 12(1):4–23, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i1/p4_s1.

Robinson:1971:EPE

[432] Peter D. Robinson. Extremum principles for the equation (∇^{2})ϕ = −ϕ − ϕ^{3}. Journal of Mathematical Physics, 12(1):23–28, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/
p23_s1.

Muraskin:1971:MPD

[433] M. Muraskin. A mathematical property of Dirac plane waves. Journal of Mathematical Physics, 12(1):28–31, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i1/p28_s1.

Finley:1971:TGA

[434] J. D. Finley. Tidal gravitational accelerations near an arbitrary timelike geodesic in Schwarzschild space. Journal of Mathematical Physics, 12(1): 32–36, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i1/p32_s1.

Cukierda:1971:FTD

[435] T. Cukierda. Field-theoretic description of massless particles with higher spin and definite parity. II. Half-integer spin. Journal of Mathematical Physics, 12(1):36–39, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i1/p36_s1.

Gottlieb:1971:WCG

[436] H. P. W. Gottlieb. Wigner coefficients for the group E(2). Journal of Mathematical Physics, 12(1):39–44, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i1/p39_s1.

Cleary:1971:SGT

[437] J. G. Cleary and B. G. Wybourne. Statistical group theory and the distribution of angular momentum states. Journal of Mathematical Physics, 12(1):45–52, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i1/p45_s1.

Babbitt:1971:PIC

[438] Donald G. Babbitt. Probabilistic interpretation of the classical scattering cross section. Journal of Mathematical Physics, 12(1):53–57, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i1/p53_s1.

Plybon:1971:NAN

[439] Benjamin F. Plybon. New approach to the Noether theorems. Journal of Mathematical Physics, 12(1):57–60, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i1/p57_s1.

Lee:1971:HTE

[440] M. Howard Lee. High-temperature expansion of the spin-1/2 XY model. Journal of Mathematical Physics, 12(1):61–69, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/p61_s1.

Wong:1971:ARC

[441] Cheuk-Yin Wong. Application of Regge calculus to the Schwarzschild and Reissner Nordstrom geometries at the moment of time symmetry.

Journal of Mathematical Physics, 12(1):70–78, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/p70_s1.

Shere:1971:IMA

[442] Kenneth D. Shere. Introduction to multiple asymptotic series with an application to elastic scattering. Journal of Mathematical Physics, 12(1): 78–82, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i1/p78_s1.

Ukai:1971:ADE

[443] Seiji Ukai. Asymptotic distribution of eigenvalues of the kernel in the Kirkwood–Riseman integral equation. Journal of Mathematical Physics, 12(1):83–92, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i1/p83_s1.

Claro:1971:RSP

[444] Francisco H. Claro and Gregory H. Wannier. Relaxation spectrum of phonons: A solvable model. Journal of Mathematical Physics, 12(1): 92–95, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i1/p92_s1.

Mariwalla:1971:DST

[445] K. H. Mariwalla. Dimensionality of space–time. Journal of Mathematical Physics, 12(1):96–99, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i1/p96_s1.

Hall:1971:SPT

[446] F. G. Hall and R. E. Collins. Stochastic processes and their representations in Hilbert space. Journal of Mathematical Physics, 12(1):100– 112, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i1/p100_s1.

Ilamed:1971:PMG

[447] Yehiel Ilamed. Powers of a matrix and the generalized Lucas polynomials. Journal of Mathematical Physics, 12(1):113, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/p113_s1.

Yngve:1971:NCH

[448] Staffan Yngve. Normalization of certain higher-order phase-integral approximations for wavefunctions of bound states in a potential well. Journal of Mathematical Physics, 12(1):114–117, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/p114_s1.

Hristev:1971:CCC

[449] Anatole P. Hristev. Conformal charge conjugation. Journal of Mathematical Physics, 12(1):118–124, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i1/p118_s1.

Epstein:1971:PSF

[450] Kenneth J. Epstein. Photon spin formalism. Journal of Mathematical Physics, 12(1):125–130, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i1/p125_s1.

Schwarz:1971:UIR

[451] Fritz Schwarz. Unitary irreducible representations of the groups SO_{0}(n, 1). Journal of Mathematical Physics, 12(1):131–139, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i1/p131_s1.

Simon:1971:DTH

[452] Barry Simon. Distributions and their Hermite expansions. Journal of Mathematical Physics, 12(1):140–148, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i1/p140_s1.

DellAntonio:1971:CLG

[453] G. F. Dell’Antonio. Can local gauge transformations be implemented? Journal of Mathematical Physics, 12(1):148–156, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/p148_s1.

Hajicek:1971:BST

[454] P. Hajicek. Bifurcate space–times. Journal of Mathematical Physics, 12(1):157–160, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i1/p157_s1.

Hamilton:1971:PIP

[455] John F. Hamilton and L. S. Schulman. Path integrals and product integrals. Journal of Mathematical Physics, 12(1):160–164, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i1/
p160_s1.

Girardeau:1971:COP

[456] M. D. Girardeau. Completely orthogonalized plane waves. Journal of Mathematical Physics, 12(1):165–167, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i1/p165_s1.

Hegerfeldt:1971:EBD

[457] Gerhard C. Hegerfeldt. Equivalence of basis-dependent and basisindependent approach to canonical field operators. Journal of Mathematical Physics, 12(1):167–172, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i1/p167_s1.

Louck:1971:ISR

[458] J. D. Louck and L. C. Biedenharn. Identity satisfied by the Racah coefficients of u(n). Journal of Mathematical Physics, 12(2):173–177, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i2/p173_s1.

Witriol:1971:BCP

[459] Norman M. Witriol. Boundary conditions in the pairwise point transformation method. Journal of Mathematical Physics, 12(2):177–179, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i2/p177_s1.

Bahar:1971:NGB

[460] E. Bahar. New generalized Bessel transform and its relationship to the Fourier, Watson, and Kontorowich–Lebedev transforms. Journal of Mathematical Physics, 12(2):179–185, February 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i2/p179_s1.

Bahar:1971:DEW

[461] E. Bahar. Diffraction of electromagnetic waves by cylindrical structures characterized by variable curvature and surface impedance. Journal of Mathematical Physics, 12(2):186–196, February 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i2/p186_s1.

Wolf:1971:RMC

[462] Kurt Bernardo Wolf. Recursive method for the computation of the SO_{n}, SO_{n,1}, and ISO_{n}, representation matrix elements. Journal of Mathematical Physics, 12(2):197–206, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i2/p197_s1.

Somorjai:1971:IIT

[463] R. L. Somorjai. Iterated integral-transform trial functions. Journal of Mathematical Physics, 12(2):206–210, February 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i2/p206_s1.

Padgett:1971:ESS

[464] W. J. Padgett and C. P. Tsokos. Existence of a solution of a stochastic integral equation in turbulence theory. Journal of Mathematical Physics, 12(2):210–212, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i2/p210_s1.

Halpern:1971:QCT

[465] M. Halpern and S. Malin. Quasigroups of coordinate transformations in curved space–time. Journal of Mathematical Physics, 12(2):213–225, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i2/p213_s1.

Carmeli:1971:IDR

[466] M. Carmeli and S. Malin. Infinite-dimensional representations of the Lorentz group: Complementary series of representations. Journal of Mathematical Physics, 12(2):225–229, February 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i2/p225_s1.

Katsura:1971:LGFa

[467] Shigetoshi Katsura and Tsuyoshi Horiguchi. Lattice Green’s function for the body-centered cubic lattice. Journal of Mathematical Physics, 12(2):230–231, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i2/p230_s1.

Das:1971:SRS

[468] A. Das. “short-ranged” scalar gravity in Einstein’s theory. Journal of Mathematical Physics, 12(2):232–234, February 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i2/p232_s1.

Suzuki:1971:ZPF

[469] Masuo Suzuki and Michael E. Fisher. Zeros of the partition function for the Heisenberg, ferroelectric, and general Ising models. Journal of Mathematical Physics, 12(2):235–246, February 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i2/p235_s1.

Sutherland:1971:QMBa

[470] Bill Sutherland. Quantum many-body problem in one dimension: Ground state. Journal of Mathematical Physics, 12(2):246–250, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i2/p246_s1.

Sutherland:1971:QMBb

[471] Bill Sutherland. Quantum many-body problem in one dimension: Thermodynamics. Journal of Mathematical Physics, 12(2):251–256, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i2/p251_s1.

Safko:1971:SPC

[472] John L. Safko and Louis Witten. Some properties of cylindrically symmetric Einstein–Maxwell fields. Journal of Mathematical Physics, 12(2):257–271, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i2/p257_s1.

Karlsson:1971:HFI

[473] Per W. Karlsson. Hypergeometric functions with integral parameter differences. Journal of Mathematical Physics, 12(2):270–271, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i2/p270_s1.

Dufty:1971:GME

[474] James W. Dufty. Generalized master equations for inhomogeneous systems. Journal of Mathematical Physics, 12(2):272–280, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i2/
p272_s1.

Macfadyen:1971:CSE

[475] N. W. Macfadyen and P. Winternitz. Crossing symmetric expansions of physical scattering amplitudes; the O(2, 1) group and Lamé functions. Journal of Mathematical Physics, 12(2):281–293, February 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i2/p281_
s1.

Constantinescu:1971:APN

[476] F. Constantinescu. Analytic properties of nonstrictly localizable fields. Journal of Mathematical Physics, 12(2):293–298, February 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i2/p293_
s1.

Lam:1971:MEM

[477] Lui Lam. Motion in electric and magnetic fields. I. Klein–Gordon particles. Journal of Mathematical Physics, 12(2):299–303, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i2/
p299_s1.

Schulman:1971:AT

[478] L. S. Schulman. Approximate topologies. Journal of Mathematical Physics, 12(2):304–308, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i2/p304_s1.

Williams:1971:CPS

[479] J. G. Williams. Conserved particle structures in general relativity. Journal of Mathematical Physics, 12(2):308–313, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i2/p308_s1.

Geroch:1971:ASP

[480] Robert Geroch and E. T. Newman. Application of the semidirect product of groups. Journal of Mathematical Physics, 12(2):314, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i2/p314_s1.

Yao:1971:UIR

[481] Tsu Yao. Unitary irreducible representations of SU(2, 2). III. Reduction with respect to an iso-Poincaré subgroup. Journal of Mathematical Physics, 12(3):315–342, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i3/p315_s1.

Gutzwiller:1971:POC

[482] Martin C. Gutzwiller. Periodic orbits and classical quantization conditions. Journal of Mathematical Physics, 12(3):343–358, March 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i3/p343_
s1.

Adawi:1971:SWM

[483] I. Adawi. Scattering of waves in many dimensions. Journal of Mathematical Physics, 12(3):358–359, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i3/p358_s1.

Benioff:1971:SAR

[484] Paul A. Benioff. Some aspects of the relationship between mathematical logic and physics. II. Journal of Mathematical Physics, 12(3):360– 376, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i3/p360_s1.

Santhanam:1971:GIT

[485] T. S. Santhanam, P. S. Chandrasekaran, and Nalini B. Menon. General involutional transformations and the representation of GL(n). Journal of Mathematical Physics, 12(3):377–383, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i3/p377_s1.

Patera:1971:LSC

[486] J. Patera. Labeling states and constructing matrix representations of F_{4}. Journal of Mathematical Physics, 12(3):384–389, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i3/p384_s1.

Solomon:1971:GTS

[487] A. I. Solomon. Group theory of superfluidity. Journal of Mathematical Physics, 12(3):390–394, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i3/p390_s1.

Nanda:1971:TMS

[488] S. Nanda. Topology for Minkowski space. Journal of Mathematical Physics, 12(3):394–401, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i3/p394_s1.

Burke:1971:GRD

[489] William L. Burke. Gravitational radiation damping of slowly moving systems calculated using matched asymptotic expansions. Journal of Mathematical Physics, 12(3):401–418, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i3/p401_s1.

Calogero:1971:SOD

[490] F. Calogero. Solution of the one-dimensional N-body problems with quadratic and/or inversely quadratic pair potentials. Journal of Mathematical Physics, 12(3):419–436, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i3/p419_s1. See erratum [1784].

Wu:1971:SCG

[491] Alfred C. T. Wu. Structure of the combinatorial generalization of hypergeometric functions for SU(n) states. Journal of Mathematical Physics, 12(3):437–440, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i3/p437_s1.

Mittag:1971:DTM

[492] L. Mittag and M. J. Stephen. Dual transformations in many-component Ising models. Journal of Mathematical Physics, 12(3):441–450, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i3/p441_s1.

Richardson:1971:EMC

[493] Thomas G. Richardson and Lawrence Sirovich. Effect of molecular collision frequency on solutions of the linearized Boltzmann equation. Journal of Mathematical Physics, 12(3):450–453, March 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i3/p450_s1.

delaPena-Auerbach:1971:STQ

[494] L. de la Peña-Auerbach. Stochastic theory of quantum mechanics for particles with spin. Journal of Mathematical Physics, 12(3):453–461, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i3/p453_s1.

Goldin:1971:NCA

[495] Gerald A. Goldin. Nonrelativistic current algebras as unitary representations of groups. Journal of Mathematical Physics, 12(3):462–487, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i3/p462_s1.

Guyard:1971:EHE

[496] Jacques Guyard, André Nadeau, Germain Baumann, and Marc R. Feix. Eigenvalues of the Hill equation to any order in the adiabatic limit. Journal of Mathematical Physics, 12(3):488–492, March 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i3/p488_s1.

Macfadyen:1971:R

[497] N. W. Macfadyen. The reduction O(3, 1) ⊃ O(2, 1) ⊃ O(1, 1). Journal of Mathematical Physics, 12(3):492–498, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i3/p492_s1.

Lovelock:1971:ETG

[498] David Lovelock. The Einstein tensor and its generalizations. Journal of Mathematical Physics, 12(3):498–501, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i3/p498_s1.

Hazeltine:1971:DRL

[499] R. D. Hazeltine, M. N. Rosenbluth, and A. M. Sessler. Diffraction radiation by a line charge moving past a comb: a model of radiation losses in an electron ring accelerator. Journal of Mathematical Physics, 12(3): 502–514, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i3/p502_s1.

Keam:1971:PTB

[500] R. F. Keam. Perturbation treatment of the Bethe–Salpeter equation for tightly bound fermion-antifermion systems. Journal of Mathematical Physics, 12(3):515–528, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i3/p515_s1.

Chakraborty:1971:CWN

[501] B. Chakraborty. Coupling of waves in nonuniform plasma. Journal of Mathematical Physics, 12(3):529–533, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i3/p529_s1.

Sanders:1971:SRT

[502] J. W. Sanders, Th. W. Ruijgrok, and J. J. ten Bosch. Some remarks on the theory of trapping of excitons in the photosynthetic unit. Journal of Mathematical Physics, 12(3):534–541, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i3/p534_s1.

McDonald:1971:EME

[503] J. F. McDonald. Eigenvalue moments for ensembles of random matrices. Journal of Mathematical Physics, 12(3):542–549, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i3/p542_s1.

Rafiqullah:1971:IRT

[504] A. K. Rafiqullah. Integral representation technique for expansion of arbitrary analytic functions of the distance between two points. Journal of Mathematical Physics, 12(3):549–552, March 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i3/p549_s1.

Romerio:1971:APF

[505] M. V. Romerio. Almost periodic functions and the theory of disordered systems. Journal of Mathematical Physics, 12(3):552–562, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i3/p552_s1.

Robertson:1971:EMN

[506] Baldwin Robertson and William C. Mitchell. Equations of motion in nonequilibrium statistical mechanics III. Open systems. Journal of Mathematical Physics, 12(3):563–568, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i3/p563_s1.

Leff:1971:CIC

[507] Harvey S. Leff. Correlation inequalities for coupled oscillators. Journal of Mathematical Physics, 12(4):569–578, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i4/p569_s1.

Sen:1971:STT

[508] D. K. Sen and K. A. Dunn. A scalar-tensor theory of gravitation in a modified Riemannian manifold. Journal of Mathematical Physics, 12(4): 578–586, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i4/p578_s1.

Thomas:1971:MCB

[509] R. B. Thomas, Jr. Magnetic corrections to the Boltzmann distribution function. Journal of Mathematical Physics, 12(4):586–591, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i4/p586_s1.

Chen:1971:CAQ

[510] Etang Chen. Continuity of automorphisms of quasilocal algebra. Journal of Mathematical Physics, 12(4):591–593, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i4/p591_s1.

Alisauskas:1971:WLO

[511] S. J. Alisauskas and A. P. Jucys. Weight lowering operators and the multiplicity-free isoscalar factors for the group R_{5}. Journal of Mathematical Physics, 12(4):594–605, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i4/p594_s1. See erratum [879].

Godfrey:1971:HAE

[512] Brendan B. Godfrey. Horizons and analytic extensions in static twodimensional space–times. Journal of Mathematical Physics, 12(4):606– 611, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i4/p606_s1.

Guertin:1971:TIA

[513] Ralph F. Guertin. Theorem on invariant amplitudes. Journal of Mathematical Physics, 12(4):612–630, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i4/p612_s1.

Deser:1971:CST

[514] Stanley Deser and L. Kent Morrison. Current-stress tensor commutators. Journal of Mathematical Physics, 12(4):631–633, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i4/p631_s1.

Lakatos-Lindenberg:1971:RWN

[515] Katja Lakatos-Lindenberg and Kurt E. Shuler. Random walks with nonnearest neighbor transitions. I. Analytic 1-D theory for next-nearest neighbor and exponentially distributed steps. Journal of Mathematical Physics, 12(4):633–652, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i4/p633_s1.

Harrison:1971:GAI

[516] B. Kent Harrison and Frank B. Estabrook. Geometric approach to invariance groups and solution of partial differential systems. Journal of Mathematical Physics, 12(4):653–666, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i4/p653_s1.

Oleson:1971:CTP

[517] M. Oleson. A class of type [4] perfect fluid space–times. Journal of Mathematical Physics, 12(4):666–672, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i4/p666_s1.

Rosen:1971:FTV

[518] S. P. Rosen. Finite transformations in various representations of SU(3). Journal of Mathematical Physics, 12(4):673–681, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i4/p673_s1.

Raman:1971:CFS

[519] K. Raman. Consequences of a field-source identity in a spin-2 meson theory. Journal of Mathematical Physics, 12(4):682–688, April 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i4/p682_
s1.

Dashen:1971:SFM

[520] Roger Dashen and Shang-Keng Ma. Singularities in forward multiparticle scattering amplitudes and the S-matrix interpretation of higher virial coefficients. Journal of Mathematical Physics, 12(4):689–715, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i4/p689_s1.

Uzes:1971:GFP

[521] C. A. Uzes. Gauge fields with positive-definite energy density. Journal of Mathematical Physics, 12(4):716–723, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i4/p716_s1.

Jain:1971:EDT

[522] D. L. Jain and R. P. Kanwal. Electromagnetic diffraction by a thin conducting annular disk. Journal of Mathematical Physics, 12(4):723– 736, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i4/p723_s1.

Nelson:1971:RBS

[523] C. A. Nelson, A. K. Rajagopal, and C. S. Shastry. Relation between the singularities of the S matrix and the L^{2} class of solutions in potential theory. Journal of Mathematical Physics, 12(4):737–741, April 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i4/p737_
s1.

Muraskin:1971:EME

[524] M. Muraskin. Erratum: “Maxwell’s Equations Having a Gradient as Source”. Journal of Mathematical Physics, 12(4):742, April 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i4/p742_

s2. See [266].

Rowe:1971:EPO

[525] D. J. Rowe. Erratum: “Properties of Overcomplete and Nonorthogonal Basis Vectors”. Journal of Mathematical Physics, 12(4):742, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i4/p742_s1. See [2].

Stern:1971:EDC

[526] David P. Stern. Erratum: “Direct Canonical Transformations”. Journal of Mathematical Physics, 12(4):742, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i4/p742_s3. See [342].

Chen:1971:ATW

[527] Kenneth C. Chen. Asymptotic theory of wave propagation in spatial and temporal dispersive inhomogeneous media. Journal of Mathematical Physics, 12(5):743–753, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i5/p743_s1.

Leonard:1971:TDQ

[528] A. Leonard. Two-dimensional quarter space problems in one-speed transport theory. Journal of Mathematical Physics, 12(5):754–766, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i5/
p754_s1.

Gates:1971:ESC

[529] D. J. Gates. Exactly solvable cell model with a melting transition. Journal of Mathematical Physics, 12(5):766–778, May 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i5/p766_s1.

Dobson:1971:PRV

[530] P. N. Dobson and W. A. Simmons. Properties of the residues in the Veneziano model. Journal of Mathematical Physics, 12(5):778–782, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i5/p778_s1.

Galbraith:1971:FBP

[531] Harold William Galbraith. Four-body problem: A complete system of angular functions. Journal of Mathematical Physics, 12(5):782–794, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i5/p782_s1.

Asvestas:1971:LFS

[532] J. S. Asvestas and R. E. Kleinman. Low-frequency scattering by perfectly conducting obstacles. Journal of Mathematical Physics, 12(5): 795–811, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i5/p795_s1.

Rosen:1971:FCT

[533] Gerald Rosen. Functional calculus theory for incompressible fluid turbulence. Journal of Mathematical Physics, 12(5):812–820, May 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i5/p812_
s1.

Gruber:1971:MCG

[534] B. Gruber, B. H. Han, and J. A. Saldanha. Multiplicities in the classical groups. II. Journal of Mathematical Physics, 12(5):821–827, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i5/p821_s1.

Wainwright:1971:GPN

[535] J. Wainwright. Geometric properties of neutrino fields in curved space– time. Journal of Mathematical Physics, 12(5):828–835, May 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i5/p828_
s1.

Seetharaman:1971:ASW

[536] M. Seetharaman, J. Jayaraman, and P. M. Mathews. Arbitraryspin wave equations and Lorentz invariance. Journal of Mathematical Physics, 12(5):835–840, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i5/p835_s1.

Barut:1971:FSB

[537] A. O. Barut and G. L. Bornzin. SO(4, 2)-formulation of the symmetry breaking in relativistic Kepler problems with or without magnetic charges. Journal of Mathematical Physics, 12(5):841–846, May 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i5/p841_
s1.

Szego:1971:REEa

[538] K. Szegő and K. Tóth. SL(2, C) representations in explicitly ‘energydependent’ basis. I. Journal of Mathematical Physics, 12(5):846–852, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i5/p846_s1.

Szego:1971:REEb

[539] K. Szegő and K. Tóth. SL(2, C) representations in explicitly ‘energydependent’ basis. II. Journal of Mathematical Physics, 12(5):853–860, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i5/p853_s1.

Deliyannis:1971:DS

[540] Platon C. Deliyannis. Density of states. Journal of Mathematical Physics, 12(5):860–862, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i5/p860_s1.

Anderson:1971:SF

[541] J. T. Anderson. Strong forces. Journal of Mathematical Physics, 12(5): 862–867, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i5/p862_s1.

Gomatam:1971:CRK

[542] Jagannathan Gomatam and C. W. Gardiner. Continuous representations and the kernel function. Journal of Mathematical Physics, 12(5): 868–873, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i5/p868_s1.

Essam:1971:PPI

[543] John W. Essam. Percolation processes. II. The pair connectedness. Journal of Mathematical Physics, 12(5):874–882, May 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i5/p874_s1.

Jaffe:1971:DLS

[544] Leon Jaffe. Discrete Lorentz and Singleton representations of the universal covering group of the 3 + 2 de Sitter group. Journal of Mathematical Physics, 12(5):882–891, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i5/p882_s1.

Katsura:1971:LGFb

[545] Shigetoshi Katsura, Tohru Morita, Sakari Inawashiro, Tsuyoshi Horiguchi, and Yoshihiko Abe. Lattice Green’s function. Introduction. Journal of Mathematical Physics, 12(5):892–895, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i5/p892_s1.

Katsura:1971:LGFc

[546] Shigetoshi Katsura, Sakari Inawashiro, and Yoshihiko Abe. Lattice Green’s function for the simple cubic lattice in terms of a Mellin–Barnes type integral. Journal of Mathematical Physics, 12(5):895–899, May 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i5/p895_s1.

Ruelle:1971:AGF

[547] David Ruelle. Analyticity of Green’s functions of dilute quantum gases. Journal of Mathematical Physics, 12(6):901–903, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i6/p901_s1.

Davidson:1971:RQS

[548] Russell Davidson and John J. Kozak. Relaxation to quantum-statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation field. III. The quantum-mechanical solution. Journal of Mathematical Physics, 12(6):903–917, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i6/p903_s1.

Geroch:1971:MGS

[549] Robert Geroch. A method for generating solutions of Einstein’s equations. Journal of Mathematical Physics, 12(6):918–924, June 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i6/p918_
s1.

Gersten:1971:WET

[550] A. Gersten. Wavefunction expansion in terms of spherical Bessel functions. Journal of Mathematical Physics, 12(6):924–929, June 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i6/p924_
s1.

Mahanta:1971:ASV

[551] M. N. Mahanta and D. R. K. Reddy. An approximate solution of the vacuum static case of spherical symmetry in Brans–Dicke theory. Journal of Mathematical Physics, 12(6):929–932, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i6/p929_s1.

Galgani:1971:AEE

[552] Luigi Galgani, Luciano Manzoni, and Antonio Scotti. Asymptotic equivalence of equilibrium ensembles of classical statistical mechanics. Journal of Mathematical Physics, 12(6):933–935, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i6/p933_s1.

Li:1971:SNP

[553] Ming Chiang Li. Scattering by a nonspherical potential. Journal of Mathematical Physics, 12(6):936–944, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i6/p936_s1.

Anderson:1971:STD

[554] David L. T. Anderson. Stability of time-dependent particlelike solutions in nonlinear field theories. II. Journal of Mathematical Physics, 12(6): 945–952, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i6/p945_s1.

Armstrong:1971:HOR

[555] Lloyd Armstrong, Jr. O(2, 1) and the harmonic oscillator radial function. Journal of Mathematical Physics, 12(6):953–957, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i6/p953_s1.

Barnsley:1971:SRF

[556] Michael Barnsley. Sum rule functions. Journal of Mathematical Physics, 12(6):957–964, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i6/p957_s1.

Dragt:1971:APP

[557] Alex J. Dragt. Analytic properties of the phase matrix. Journal of Mathematical Physics, 12(6):965–969, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i6/p965_s1.

Bolsterli:1971:SAM

[558] M. Bolsterli and J. L. Norton. Separable approximations to matrices and functions of two variables. Journal of Mathematical Physics, 12(6): 969–970, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i6/p969_s1.

Chang:1971:DCD

[559] B. Tsu-Shen Chang. Density and current density as coordinates for a system of interacting bosons at absolute zero. Journal of Mathematical Physics, 12(6):971–980, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i6/p971_s1.

Morita:1971:LGF

[560] Tohru Morita and Tsuyoshi Horiguchi. Lattice Green’s functions for the cubic lattices in terms of the complete elliptic integral. Journal of Mathematical Physics, 12(6):981–986, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i6/p981_s1.

Morita:1971:CLG

[561] Tohru Morita and Tsuyoshi Horiguchi. Calculation of the lattice Green’s function for the bcc, fcc, and rectangular lattices. Journal of Mathematical Physics, 12(6):986–992, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i6/p986_s1.

Schmidt:1971:ASB

[562] V. Hugo Schmidt. AC susceptibility of biased one-dimensional stochastic Ising model. Journal of Mathematical Physics, 12(6):992–995, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i6/p992_s1.

Robinson:1971:SBG

[563] David C. Robinson and Jeffrey Winicour. Scaling behavior of gravitational energy. Journal of Mathematical Physics, 12(6):995–999, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i6/p995_s1.

Millard:1971:ISL

[564] Kenneth Millard and Harvey S. Leff. Infinite-spin limit of the quantum Heisenberg model. Journal of Mathematical Physics, 12(6):1000– 1005, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i6/p1000_s1.

Cantrell:1971:GFF

[565] C. D. Cantrell. Generating function of the N-fold photoelectric counting distribution for Gaussian light. Journal of Mathematical Physics, 12(6): 1005–1008, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i6/p1005_s1.

Bose:1971:LBS

[566] S. K. Bose and R. Parker. Lorentz basis for the 3 + 2 de Sitter group. Journal of Mathematical Physics, 12(6):1009–1013, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i6/p1009_s1.

Deliyannis:1971:GHV

[567] Platon C. Deliyannis. Generalized hidden variables theorem. Journal of Mathematical Physics, 12(6):1013–1017, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i6/p1013_s1.

Oreg:1971:GRW

[568] J. Oreg. Generalized Rajnak–Wybourne identity. Journal of Mathematical Physics, 12(6):1018–1021, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i6/p1018_s1.

Ifantis:1971:AFQ

[569] Evangelos K. Ifantis. Abstract formulation of the quantum mechanical oscillator phase problem. Journal of Mathematical Physics, 12(6):1021– 1026, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i6/p1021_s1.

Scott:1971:CPB

[570] H. L. Scott. Critical-point behavior of a class of model Hamiltonians. Journal of Mathematical Physics, 12(6):1026–1029, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i6/p1026_s1.

Neubert:1971:AEL

[571] Jerome A. Neubert. Approach to the Eulerian–Lagrangian problem for sound propagation in continuous stochastic media. Journal of Mathematical Physics, 12(6):1030–1034, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i6/p1030_s1.

Chen:1971:ETQ

[572] Tien Chi Chen. Evaluation techniques for quantum mechanical integrals involving orthogonal polynomials. Journal of Mathematical Physics, 12(6):1035–1038, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i6/p1035_s1.

Novello:1971:GCS

[573] Mario Novello. Gravitation as a consequence of the self-interaction of the Gamma fields. Journal of Mathematical Physics, 12(6):1039–1041, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i6/p1039_s1.

Belinfante:1971:OPS

[574] J. G. F. Belinfante and P. Winternitz. One-parameter subgroups of unitary groups with indefinite metric and in particular of the conformal group. Journal of Mathematical Physics, 12(6):1041–1054, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i6/p1041_s1.

Chakrabarti:1971:RCR

[575] A. Chakrabarti. Remarks on a class of representations of current density algebras. Journal of Mathematical Physics, 12(6):1055–1057, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i6/p1055_s1.

Burniston:1971:EHR

[576] E. E. Burniston and C. E. Siewert. Erratum: “Half-Range Expansion Theorems in Studies of Polarized Light”. Journal of Mathematical Physics, 12(6):1057, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i6/p1057_s1. See [425].

Miller:1971:EBE

[577] Melvin N. Miller. Erratum: “Bounds for Effective Bulk Modulus of Heterogeneous Materials”. Journal of Mathematical Physics, 12(6):1057, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i6/p1057_s2. See [3].

Khan:1971:I

[578] M. N. Khan and Arif-Uz-Zaman. Invariants of πN → πB_{s}. Journal of Mathematical Physics, 12(7):1059–1066, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1059_s1.

Pechlaner:1971:SGF

[579] E. Pechlaner and A. Das. Static gravitational field of multipoles in the second approximation. Journal of Mathematical Physics, 12(7):1066– 1068, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1066_s1.

Mortell:1971:ROR

[580] Michael P. Mortell. Resonant oscillations: a regular perturbation approach. Journal of Mathematical Physics, 12(7):1069–1075, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/
p1069_s1.

Mashhoon:1971:PSG

[581] Bahram Mashhoon. Particles with spin in a gravitational field. Journal of Mathematical Physics, 12(7):1075–1077, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1075_s1.

Ozsvath:1971:SHR

[582] István Ozsváth. Spatially homogeneous rotating world models. Journal of Mathematical Physics, 12(7):1078–1082, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1078_s1.

Bergstrom:1971:ETC

[583] K. Bergström and S. Yngström. Equal-time commutation relations between components of a current and the energy-momentum tensor. Journal of Mathematical Physics, 12(7):1083–1087, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1083_s1.

Debney:1971:IAS

[584] George C. Debney, Jr. Invariant approach to a space–time symmetry. Journal of Mathematical Physics, 12(7):1088–1098, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1088_s1.

Post:1971:NEA

[585] E. J. Post and D. D. Bahulikar. Note on the electrodynamics of accelerated systems. Journal of Mathematical Physics, 12(7):1098–1099, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1098_s1.

Hintz:1971:NBV

[586] Gerald R. Hintz. A nonhomogeneous boundary-value problem for the linear transport equation for a slab geometry. Journal of Mathematical Physics, 12(7):1100–1112, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i7/p1100_s1.

Luehr:1971:SIP

[587] C. P. Luehr and M. Rosenbaum. Simplified intrinsic proof of the Rainich differential relation. Journal of Mathematical Physics, 12(7):1113–1115, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1113_s1.

Barston:1971:EPLa

[588] E. M. Barston. Eigenvalue problem for Lagrangian systems. IV. Journal of Mathematical Physics, 12(7):1116–1122, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1116_s1.

Okubo:1971:SGI

[589] Susumu Okubo. Some general inequalities in quantum statistical mechanics. Journal of Mathematical Physics, 12(7):1123–1128, July 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1123_
s1.

Lebowitz:1971:PTB

[590] J. L. Lebowitz and G. Gallavotti. Phase transitions in binary lattice gases. Journal of Mathematical Physics, 12(7):1129–1133, July 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1129_
s1.

Yakimiw:1971:NNS

[591] E. Yakimiw. Note on the new symmetry of the Racah coefficients. Journal of Mathematical Physics, 12(7):1134, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1134_s1.

Joshi:1971:RMS

[592] Vinaya Joshi. Remark on Minton’s symmetry for Racah functions. Journal of Mathematical Physics, 12(7):1134–1136, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1134_s2.

Das:1971:SGF

[593] A. Das. Static gravitational fields. I. Eight theorems. Journal of Mathematical Physics, 12(7):1136–1142, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i7/p1136_s1.

Kelemen:1971:ACR

[594] P. J. Kelemen. Analytic continuation of the regular solution of the Schrödinger equation in the complex angular momentum plane. Journal of Mathematical Physics, 12(7):1142–1155, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1142_s1.

Matcha:1971:PSE

[595] R. L. Matcha, R. H. Pritchard, and C. W. Kern. Prolate-spheroidal expansions of the spin-orbit, spin-spin, and orbit-orbit operators. Journal of Mathematical Physics, 12(7):1155–1159, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1155_s1.

Stettner:1971:EER

[596] R. Stettner. Effect of electromagnetic radiation on the secular stability of a homogeneous, charged, rotating drop. Journal of Mathematical Physics, 12(7):1159–1162, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i7/p1159_s1.

Fuda:1971:MEP

[597] Michael G. Fuda. T matrix for the exponential potential. Journal of Mathematical Physics, 12(7):1163–1166, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1163_s1.

Coudray:1971:CRP

[598] Christiane Coudray and Marcel Coz. Construction of relativistic potentials when the energy is fixed. Journal of Mathematical Physics, 12(7):

1166–1178, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1166_s1.

Steiger:1971:PITa

[599] Arno D. Steiger. Poincaré-irreducible tensor operators for positive mass one-particle states. I. Journal of Mathematical Physics, 12(7):1178–1191, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1178_s1.

Rosen:1971:ESS

[600] Gerald Rosen. Exact stationary states in |ϕ|^{2N} field theories. Journal of Mathematical Physics, 12(7):1192–1194, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1192_s1.

Wilcox:1971:TEW

[601] Ralph M. Wilcox. Transmission of electromagnetic waves through a conducting plasma slab. Journal of Mathematical Physics, 12(7):1195–1207, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1195_s1.

Dobson:1971:SAO

[602] Peter N. Dobson, Jr. Self-adjoint operators in indefinite metric spaces. Journal of Mathematical Physics, 12(7):1207–1210, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1207_s1.

Torrieri:1971:PRE

[603] Don J. Torrieri and Leonard S. Taylor. Propagation of random electromagnetic fields. Journal of Mathematical Physics, 12(7):1211–1214, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1211_s1.

Wolfe:1971:EIE

[604] Peter Wolfe. Eigenfunctions of the integral equation for the potential of the charged disk. Journal of Mathematical Physics, 12(7):1215–1218, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1215_s1.

Ludwig:1971:GMC

[605] Garry Ludwig. Geometrization of a massless complex scalar field. Journal of Mathematical Physics, 12(7):1218–1219, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1218_s1.

Ludwig:1971:GMS

[606] Garry Ludwig. Geometrization of a massive scalar field. Journal of Mathematical Physics, 12(7):1220–1221, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1220_s1.

Glass:1971:REA

[607] Edward N. Glass. Radiation and energy in an asymptotically Friedmann space–time. Journal of Mathematical Physics, 12(7):1222–1230, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1222_s1.

Alexanian:1971:MFQ

[608] Moorad Alexanian. Myriotic fields in quantum statistical mechanics. Journal of Mathematical Physics, 12(7):1230–1242, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1230_s1.

Auberson:1971:DRF

[609] G. Auberson and G. Mahoux. Dispersion relations, Froissart–Gribov transforms, and distributions. Journal of Mathematical Physics, 12(7): 1242–1257, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1242_s1.

DInverno:1971:CHM

[610] R. A. D’Inverno and R. A. Russell-Clark. Classification of the Harrison metrics. Journal of Mathematical Physics, 12(7):1258–1263, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1258_s1.

Shen:1971:RSI

[611] C. Y. Shen and D. S. Carter. Representations of states of infinite systems in statistical mechanics. Journal of Mathematical Physics, 12(7):1263– 1269, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1263_s1.

Bradbury:1971:PSN

[612] Ted Clay Bradbury and Alton Brintzenhoff. Periodic solutions of a nonlinear differential equation. Journal of Mathematical Physics, 12(7): 1269–1275, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1269_s1.

Casti:1971:IIV

[613] J. Casti, R. Kalaba, and S. Ueno. Invariant imbedding and the variational treatment of Fredholm integral equations with displacement kernels. Journal of Mathematical Physics, 12(7):1276–1278, July 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1276_
s1.

Ueno:1971:RTS

[614] S. Ueno, H. Kagiwada, and R. Kalaba. Radiative transfer in spherical shell atmospheres with radial symmetry. Journal of Mathematical Physics, 12(7):1279–1286, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i7/p1279_s1.

Arif-Uz-Zaman:1971:CSB

[615] Arif-Uz-Zaman. Contributions of spin-J baryon poles to πN → πN and πN → γN amplitudes. Journal of Mathematical Physics, 12(7):1286– 1295, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1286_s1.

Storer:1971:DFE

[616] R. G. Storer. Distribution functions for an exactly soluble quantum statistical system. Journal of Mathematical Physics, 12(7):1296–1302, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1296_s1.

Sabatier:1971:NTS

[617] Pierre C. Sabatier. New tool for scattering studies. Journal of Mathematical Physics, 12(7):1303–1326, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i7/p1303_s1.

Hammer:1971:MQR

[618] C. L. Hammer and R. H. Tucker. A method of quantization for relativistic fields. Journal of Mathematical Physics, 12(7):1327–1333, July 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1327_
s1.

Hehl:1971:NSE

[619] F. W. Hehl and B. K. Datta. Nonlinear spinor equation and asymmetric connection in general relativity. Journal of Mathematical Physics, 12(7): 1334–1339, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1334_s1.

Dresner:1971:PPA

[620] Lawrence Dresner. Phase-plane analysis of nonlinear, second-order, ordinary differential equations. Journal of Mathematical Physics, 12(7): 1339–1348, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1339_s1.

Johannsen:1971:ESM

[621] Günther Johannsen. Electromagnetic sources in a moving conducting medium. Journal of Mathematical Physics, 12(7):1349–1351, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/
p1349_s1.

Katriel:1971:EDR

[622] J. Katriel and G. Adam. Equivalence of different representations of the generalized Bose operator. Journal of Mathematical Physics, 12(7):1352– 1353, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1352_s1.

Eringen:1971:MS

[623] A. Cemal Eringen. Micromagnetism and superconductivity. Journal of Mathematical Physics, 12(7):1353–1358, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1353_s1.

Pfaffelhuber:1971:VEV

[624] E. Pfaffelhuber. Vacuum expectation values as indicatrices of maximal regularity. Journal of Mathematical Physics, 12(7):1359–1369, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1359_s1.

Pfaffelhuber:1971:PUF

[625] E. Pfaffelhuber. Propagators of unrenormalizable fields. Journal of Mathematical Physics, 12(7):1370–1378, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i7/p1370_s1.

Gould:1971:RPS

[626] H. W. Gould. Remarks on a paper of Srivastava. Journal of Mathematical Physics, 12(7):1378–1379, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i7/p1378_s1.

Zachary:1971:SMC

[627] W. W. Zachary. Off-shell T matrix corresponding to a sum of Coulomb and separable potentials by expansion in O(4) harmonics. Journal of Mathematical Physics, 12(7):1379–1386, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1379_s1. See er-

rata [1414].

Sullivan:1971:COM

[628] John J. Sullivan. On the construction of orthogonal matrix basis elements of finite group algebras symmetry adapted to a subgroup. Journal of Mathematical Physics, 12(7):1387–1389, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1387_s1.

Joyce:1971:ERB

[629] G. S. Joyce. Exact results for a body-centered cubic lattice Green’s function with applications in lattice statistics. I. Journal of Mathematical Physics, 12(7):1390–1414, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i7/p1390_s1.

Gidas:1971:QFI

[630] Basilis Gidas. Quadratic fermion interaction Hamiltonian. Journal of Mathematical Physics, 12(7):1414–1419, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i7/p1414_s1.

Mohling:1971:MTM

[631] Franz Mohling and Charles Ray Smith. Microscopic theory of a multicomponent system of charged and neutral particles. I. General quantum statistical formulation. Journal of Mathematical Physics, 12(7):1419– 1427, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1419_s1.

Bhowmik:1971:MTM

[632] Anup Bhowmik and Charles Ray Smith. Microscopic theory of a multicomponent system of charged and neutral particles. II. Investigation of properties of photons. Journal of Mathematical Physics, 12(7):1427– 1436, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i7/p1427_s1.

Macfadyen:1971:CGP

[633] N. W. Macfadyen. Conformal group in a Poincaré basis. I. Principal continuous series. Journal of Mathematical Physics, 12(7):1436–1445, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i7/p1436_s1.

Adomian:1971:EROa

[634] G. Adomian. Erratum: “Random Operator Equations in Mathematical Physics. I”. Journal of Mathematical Physics, 12(7):1446, July 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1446_

s3. See [134, 715].

Basu:1971:EHA

[635] Debabrata Basu and S. N. Biswas. Erratum: “O(5) Harmonics and Abnormal Solutions in the Bethe–Salpeter Equation”. Journal of Mathematical Physics, 12(7):1446, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i7/p1446_s2. See [152].

Pietenpol:1971:ASP

[636] J. L. Pietenpol and D. Speiser. Addendum: “Simple Procedure for Determining the Number of Components of an Irreducible Tensor”. Journal of Mathematical Physics, 12(7):1446, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i7/p1446_s1. See [430].

Dashen:1971:MLE

[637] Roger Dashen and Shang keng Ma. S matrix and low energy theorem in the theory of correlation functions. Journal of Mathematical Physics, 12(8):1449–1471, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i8/p1449_s1.

Testa:1971:QOO

[638] F. J. Testa. Quantum operator ordering and the Feynman formulation. Journal of Mathematical Physics, 12(8):1471–1474, August 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1471_
s1.

Basu:1971:FEH

[639] Debabrata Basu. Formal equivalence of the hydrogen atom and harmonic oscillator and factorization of the Bethe–Salpeter equations. Journal of Mathematical Physics, 12(8):1474–1482, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1474_s1.

Roels:1971:FWP

[640] Jacques Roels and Alan Weinstein. Functions whose Poisson brackets are constants. Journal of Mathematical Physics, 12(8):1482–1486, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1482_s1.

Niemeijer:1971:LSC

[641] Th. Niemeijer. Linear spin chain with nearest and next-nearest neighbor interactions. Journal of Mathematical Physics, 12(8):1487–1491, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1487_s1.

Falk:1971:SHE

[642] H. Falk. Specific heat and entropy bounds for Ising spin systems. Journal of Mathematical Physics, 12(8):1492–1494, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1492_s1.

Papanicolaou:1971:MPR

[643] George C. Papanicolaou. Motion of a particle in a random field. Journal of Mathematical Physics, 12(8):1494–1496, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1494_s1.

Steiger:1971:PITb

[644] Arno D. Steiger. Poincaré-irreducible tensor operators for positive mass one-particle states. II. Journal of Mathematical Physics, 12(8):1497– 1507, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1497_s1.

Schneider:1971:TCR

[645] Thomas R. Schneider. Theorem on the convergence rate of generalized Fourier series. Journal of Mathematical Physics, 12(8):1508–1509, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1508_s1.

Cooper:1971:GTD

[646] C. S. Cooper and M. R. Pennington. Group-theoretic derivation of crossing sum rules. Journal of Mathematical Physics, 12(8):1509–1515, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1509_s1.

Rajagopal:1971:RES

[647] A. K. Rajagopal and C. S. Shastry. Remarks on the existence of solutions of the two-particle Lippmann–Schwinger equation. Journal of Mathematical Physics, 12(8):1515–1520, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1515_s1.

Howgate:1971:ECS

[648] David W. Howgate. Evaluation of certain statistical averages in the Hamiltonian form. Journal of Mathematical Physics, 12(8):1521–1529, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1521_s1.

Wong:1971:UAT

[649] M. K. F. Wong. Unit adjoint tensor operators in SO(n). Journal of Mathematical Physics, 12(8):1530–1535, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1530_s1.

Kulhanek:1971:STB

[650] J. Kulhánek. Space–time of the de Broglie wave field. Journal of Mathematical Physics, 12(8):1535–1538, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i8/p1535_s1.

Lerche:1971:KDT

[651] Ian Lerche. Kinematic dynamo theory: The Dyson equation and the large-scale field; the Bethe–Salpeter equation and the fluctuation intensity. Journal of Mathematical Physics, 12(8):1538–1547, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/
p1538_s1.

Gardner:1971:KVE

[652] Clifford S. Gardner. Korteweg–de Vries equation and generalizations. IV. The Korteweg–de Vries equation as a Hamiltonian system. Journal of Mathematical Physics, 12(8):1548–1551, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1548_s1.

Newton:1971:FMT

[653] Roger G. Newton. Fredholm methods in the three-body problem. I. Journal of Mathematical Physics, 12(8):1552–1567, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1552_s1.

Thomas:1971:GAC

[654] Brian K. Thomas and E. Gerjuoy. Glauber amplitudes in charged particle hydrogen atom collisions. Journal of Mathematical Physics, 12(8):1567– 1576, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1567_s1.

Barouch:1971:TDI

[655] Eytan Barouch. On the two-dimensional Ising model with random impurities. Journal of Mathematical Physics, 12(8):1577–1578, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1577_s1.

Nakanishi:1971:VSS

[656] Noboru Nakanishi. Vector-scalar sector solutions to the spinor-spinor Bethe–Salpeter equation. Journal of Mathematical Physics, 12(8):1578– 1582, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1578_s1.

Gudder:1971:OFQ

[657] S. Gudder and C. Piron. Observables and the field in quantum mechanics. Journal of Mathematical Physics, 12(8):1583–1588, August 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1583_
s1.

King:1971:MRP

[658] R. C. King. Modification rules and products of irreducible representations of the unitary, orthogonal, and symplectic groups. Journal of Mathematical Physics, 12(8):1588–1598, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1588_s1.

Boyer:1971:MEM

[659] Charles P. Boyer. Matrix elements for the most degenerate continuous principal series of representations of SO(p, 1). Journal of Mathematical Physics, 12(8):1599–1603, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i8/p1599_s1.

DeSilva:1971:TTD

[660] C. N. DeSilva and A. B. Whitman. Thermodynamical theory of directed curves. Journal of Mathematical Physics, 12(8):1603–1609, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1603_s1.

Johnson:1971:RTH

[661] Porter W. Johnson. Regge trajectories at high energies for a class of analytic potentials. Journal of Mathematical Physics, 12(8):1610–1616, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1610_s1.

Brans:1971:CFR

[662] Carl H. Brans. Complex 2-form representation of the Einstein equations: The Petrov type III solutions. Journal of Mathematical Physics, 12(8):1616–1619, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i8/p1616_s1.

Seetharaman:1971:RWE

[663] M. Seetharaman, J. Jayaraman, and P. M. Mathews. Relativistic wave equations: proper Lorentz invariance and invariance under discrete transformations. Journal of Mathematical Physics, 12(8):1620–1622, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1620_s1.

Katsura:1971:LGFd

[664] Shigetoshi Katsura and Sakari Inawashiro. Lattice Green’s functions for the rectangular and the square lattices at arbitrary points. Journal of Mathematical Physics, 12(8):1622–1630, August 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1622_s1.

RamaRao:1971:PTN

[665] I. RamaRao and Franz Mohling. Perturbation theory of a nonideal Bose gas. I. Journal of Mathematical Physics, 12(8):1631–1651, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1631_s1.

Tait:1971:CSS

[666] W. Tait and J. F. Cornwell. Coupling of de Sitter space–time and internal symmetry. Journal of Mathematical Physics, 12(8):1651–1654, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1651_s1.

Grypeos:1971:ASR

[667] M. E. Grypeos. Analytic solutions of the reference spectrum equation. Journal of Mathematical Physics, 12(8):1655–1662, August 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1655_
s1.

Zickendraht:1971:CSP

[668] W. Zickendraht. Collective and single-particle coordinates in nuclear physics. Journal of Mathematical Physics, 12(8):1663–1674, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1663_s1.

Gaudin:1971:BGOa

[669] M. Gaudin. Bose gas in one dimension. I. The closure property of the scattering wavefunctions. Journal of Mathematical Physics, 12(8):1674– 1676, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1674_s1.

Gaudin:1971:BGOb

[670] M. Gaudin. Bose gas in one dimension. II. Orthogonality of the scattering states. Journal of Mathematical Physics, 12(8):1677–1680, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1677_s1.

Maduemezia:1971:EFG

[671] Awele Maduemezia. Euler factorization of SO(p, q) and SU(p, q) groups. Journal of Mathematical Physics, 12(8):1681–1685, August 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1681_
s1.

Rubin:1971:ALT

[672] Robert J. Rubin and William L. Greer. Abnormal lattice thermal conductivity of a one-dimensional, harmonic, isotopically disordered crystal. Journal of Mathematical Physics, 12(8):1686–1701, August 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/p1686_
s1.

Casher:1971:HFR

[673] A. Casher and J. L. Lebowitz. Heat flow in regular and disordered harmonic chains. Journal of Mathematical Physics, 12(8):1701–1711, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1701_s1.

Tutsch:1971:MMP

[674] J. H. Tutsch. Mathematics of the measurement problem in quantum mechanics. Journal of Mathematical Physics, 12(8):1711–1718, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1711_s1.

Combes:1971:PSC

[675] J. M. Combes. Properties of some connected kernels in multiparticle systems. Journal of Mathematical Physics, 12(8):1719–1731, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1719_s1.

Devi:1971:PBR

[676] V. Syamala Devi. Polynomial bases for representations of semisimple Lie algebras. Journal of Mathematical Physics, 12(8):1732–1734, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1732_s1.

Syamala:1971:PBR

[677] V. Devi Syamala. Polynomial bases for representations of semisimple Lie Algebras. Journal of Mathematical Physics, 12(8):1732–1734, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skyrme:1971:KDE

[678] T. H. R. Skyrme. Kinks and the Dirac equation. Journal of Mathematical Physics, 12(8):1735–1743, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i8/p1735_s1.

Gould:1971:CPC

[679] H. W. Gould. Comment on a paper of Chiang. Journal of Mathematical Physics, 12(8):1743, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i8/p1743_s1. See [4].

Morita:1971:UPC

[680] Tohru Morita. Useful procedure for computing the lattice Green’s function-square, tetragonal, and b.c.c. lattices. Journal of Mathematical Physics, 12(8):1744–1747, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i8/p1744_s1.

Carhart:1971:CPT

[681] R. A. Carhart. Canonical perturbation theory for nonlinear quantum oscillators and fields. Journal of Mathematical Physics, 12(8):1748– 1762, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1748_s1.

Campbell:1971:TSS

[682] William B. Campbell. Tensor and spinor spherical harmonics and the spin-s harmonics _{s}Y_{lm}(θ, ϕ). Journal of Mathematical Physics, 12 (8):1763–1770, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i8/p1763_s1.

Greenberg:1971:QPL

[683] William Greenberg. Quantum potentials on lattices. Journal of Mathematical Physics, 12(8):1770–1771, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i8/p1770_s1.

Moshinsky:1971:LCT

[684] M. Moshinsky and C. Quesne. Linear canonical transformations and their unitary representations. Journal of Mathematical Physics, 12(8):1772– 1780, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1772_s1.

Quesne:1971:CTM

[685] C. Quesne and M. Moshinsky. Canonical transformations and matrix elements. Journal of Mathematical Physics, 12(8):1780–1783, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i8/p1780_s1.

Richardson:1971:SWB

[686] T. G. Richardson and L. Sirovich. The sound wave boundary value problem in kinetic theory. I. Journal of Mathematical Physics, 12(8):1784– 1798, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1784_s1.

Girardeau:1971:FMB

[687] M. D. Girardeau. Formulation of the many-body problem for composite particles. III. The projected Hamiltonian. Journal of Mathematical Physics, 12(8):1799–1810, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i8/p1799_s1.

Chakrabarti:1971:RLC

[688] A. Chakrabarti. Remarks on lightlike continuous spin and spacelike representations of the Poincaré group. Journal of Mathematical Physics, 12 (9):1813–1822, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i9/p1813_s1.

Chakrabarti:1971:LBP

[689] A. Chakrabarti. Lorentz basis of the Poincaré group. II. Journal of Mathematical Physics, 12(9):1822–1840, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i9/p1822_s1.

Nishimura:1971:PEM

[690] Keiichi Nishimura. Polynomial expansion method for problems involving a relativistic Maxwellian distribution. Journal of Mathematical Physics, 12(9):1840–1845, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i9/p1840_s1.

Cohen:1971:PCV

[691] Jeffrey M. Cohen and Robert M. Wald. Point charge in the vicinity of a Schwarzschild black hole. Journal of Mathematical Physics, 12(9): 1845–1849, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i9/p1845_s1.

Schroeck:1971:GCF

[692] F. E. Schroeck, Jr. Generalization of the Cook formalism for Fock space. Journal of Mathematical Physics, 12(9):1849–1857, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i9/p1849_s1.

Lerche:1971:SES

[693] I. Lerche. Some exact statistical properties of turbulent kinematic dynamo equations. Journal of Mathematical Physics, 12(9):1858–1867, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i9/p1858_s1.

Barston:1971:EPLb

[694] E. M. Barston. Eigenvalue problems for Lagrangian systems. V. Journal of Mathematical Physics, 12(9):1867–1872, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i9/p1867_s1.

Corbella:1971:WIS

[695] Oscar D. Corbella. L-wave inverse scattering problem. Journal of Mathematical Physics, 12(9):1873–1882, September 1, 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i9/p1873_s1.

Lohe:1971:BCO

[696] M. A. Lohe and C. A. Hurst. The boson calculus for the orthogonal and symplectic groups. Journal of Mathematical Physics, 12(9):1882–1889, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i9/p1882_s1.

Blount:1971:JTT

[697] E. I. Blount. The Jahn–Teller theorem. Journal of Mathematical Physics, 12(9):1890–1896, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i9/p1890_s1.

Nuttall:1971:AFT

[698] J. Nuttall. Asymptotic form of the three-particle scattering wavefunction for free incident particles. Journal of Mathematical Physics, 12(9):

1896–1899, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i9/p1896_s1.

Soward:1971:NSK

[699] A. M. Soward. Nearly symmetric kinematic and hydromagnetic dynamos. Journal of Mathematical Physics, 12(9):1900–1906, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i9/p1900_s1.

Chen:1971:SNM

[700] L. C. Chen, J. Guo, and J. E. Russell. Structure of neutral mesonic atoms formed in liquid helium III. More accurate treatment of the electron wavefunction. Journal of Mathematical Physics, 12(9):1906–1913, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i9/p1906_s1.

DeSanto:1971:SPC

[701] John A. DeSanto. Scattering from a periodic corrugated structure: thin comb with soft boundaries. Journal of Mathematical Physics, 12(9): 1913–1923, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i9/p1913_s1.

Wu:1971:MDC

[702] F. Y. Wu. Multiple density correlations in a many-particle system. Journal of Mathematical Physics, 12(9):1923–1929, September 1, 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i9/p1923_
s1.

Pajerski:1971:TSD

[703] David Walter Pajerski and Ezra T. Newman. Trapped surfaces and the development of singularities. Journal of Mathematical Physics, 12(9): 1929–1937, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i9/p1929_s1.

Silver:1971:UTD

[704] David M. Silver. Unified treatment of diatomic electron interaction integrals over Slater-type atomic orbitals. Journal of Mathematical Physics, 12(9):1937–1943, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i9/p1937_s1.

Adomian:1971:LROa

[705] G. Adomian. Linear random operator equations in mathematical physics. II. Journal of Mathematical Physics, 12(9):1944–1948, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i9/p1944_s1.

Adomian:1971:LROb

[706] G. Adomian. Linear random operator equations in mathematical physics. III. Journal of Mathematical Physics, 12(9):1948–1955, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v12/i9/p1948_s1. See addendum [833].

Harper:1971:SOJ

[707] E. Y. Harper, I-Dee Chang, and G. W. Grube. A second-order JWKB approximation with one turning point and two singular points: Stability of an accelerating liquid sphere. Journal of Mathematical Physics, 12 (9):1955–1960, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i9/p1955_s1.

Ifantis:1971:SSE

[708] Evangelos K. Ifantis. Solution of the Schrödinger equation in the Hardy– Lebesgue space. Journal of Mathematical Physics, 12(9):1961–1965, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i9/p1961_s1. See erratum [880].

Gabriel:1971:RGR

[709] John R. Gabriel. Reduction of group representations. V. Journal of Mathematical Physics, 12(9):1966–1970, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i9/p1966_s1.

Case:1971:FTM

[710] K. M. Case and R. D. Hazeltine. Fourier transform methods in linear transport theory. Journal of Mathematical Physics, 12(9):1970–1980, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i9/p1970_s1.

vanVliet:1971:MADa

[711] K. M. van Vliet. Markov approach to density fluctuations due to transport and scattering. I. Mathematical formalism. Journal of Mathematical Physics, 12(9):1981–1998, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i9/p1981_s1.

vanVliet:1971:MADb

[712] K. M. van Vliet. Markov approach to density fluctuations due to transport and scattering. II. Applications. Journal of Mathematical Physics, 12(9):1998–2012, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v12/i9/p1998_s1.

Evans:1971:DLT

[713] Kenneth Evans, Jr. and E. Atlee Jackson. Double Laplace transformation in mixed boundary-initial value problems. Journal of Mathematical Physics, 12(9):2012–2023, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i9/p2012_s1.

Chatwin:1971:ASS

[714] R. A. Chatwin and J. E. Purcell. Approximate solution of a Sturm– Liouville system using nonorthogonal expansions: Application to α–α nuclear scattering. Journal of Mathematical Physics, 12(9):2024–2030, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i9/p2024_s1.

Adomian:1971:EROb

[715] G. Adomian. Erratum: “Random Operator Equations in Mathematical Physics. I”. Journal of Mathematical Physics, 12(9):2031, September 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i9/p2031_s1. See [134, 634].

Wu:1971:SGVa

[716] Alfred C. T. Wu. Structure of the generalized Veneziano amplitudes. I. Radon transforms. Journal of Mathematical Physics, 12(10):2035–2036, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2035_s1.

Wu:1971:SGVb

[717] Alfred C. T. Wu. Structure of the generalized Veneziano amplitudes. II. Possible group theoretic content. Journal of Mathematical Physics, 12 (10):2036–2039, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2036_s1.

Camiz:1971:EST

[718] P. Camiz, A. Gerardi, C. Marchioro, E. Presutti, and E. Scacciatelli. Exact solution of a time-dependent quantal harmonic oscillator with a singular perturbation. Journal of Mathematical Physics, 12(10):2040– 2043, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2040_s1.

Emch:1971:RLT

[719] Gérard G. Emch and Charles Radin. Relaxation of local thermal deviations from equilibrium. Journal of Mathematical Physics, 12(10):2043– 2046, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2043_s1.

Eftimiu:1971:MNC

[720] C. Eftimiu. On the Martin–Newton condition for finite-dimensional unitary matrices. Journal of Mathematical Physics, 12(10):2047–2050, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2047_s1.

Federbush:1971:CER

[721] Paul Federbush. A convergent expansion for the resolvent of: ϕ^{4}: i + 1. Journal of Mathematical Physics, 12(10):2050–2051, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i10/
p2050_s1.

Soward:1971:NSA

[722] A. M. Soward. Nearly symmetric advection. Journal of Mathematical Physics, 12(10):2052–2062, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i10/p2052_s1.

Morita:1971:TTS

[723] Tohru Morita. Two-time spin-pair correlation function of the Heisenberg magnet at infinite temperature. Journal of Mathematical Physics, 12 (10):2062–2069, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2062_s1.

Boyer:1971:DMD

[724] Charles P. Boyer and Farhad Ardalan. On the decomposition SO(p, 1) ⊃ SO(p−1, 1) for most degenerate representations. Journal of Mathematical Physics, 12(10):2070–2075, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i10/p2070_s1. See erratum [1183].

Bahethi:1971:MHP

[725] O. P. Bahethi and M. G. Fuda. The T matrix for the hulthén potential. Journal of Mathematical Physics, 12(10):2076–2080, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i10/
p2076_s1.

Bailey:1971:SDC

[726] Paul B. Bailey and George E. Barr. Scalar diffraction by a curved edge. Journal of Mathematical Physics, 12(10):2080–2092, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i10/
p2080_s1.

Savit:1971:ABA

[727] R. Savit, R. Blankenbecler, and M. B. Einhorn. Another bound on the absorptive part of elastic scattering amplitudes. Journal of Mathematical Physics, 12(10):2092–2098, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i10/p2092_s1.

Bracken:1971:VOP

[728] A. J. Bracken and H. S. Green. Vector operators and a polynomial identity for SO(n). Journal of Mathematical Physics, 12(10):2099–2106, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2099_s1.

Green:1971:CIG

[729] H. S. Green. Characteristic identities for generators of GL(n), O(n) and SP(n). Journal of Mathematical Physics, 12(10):2106–2113, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i10/p2106_s1.

McQuistan:1971:ENN

[730] Richmond B. McQuistan. Exact nearest-neighbor statistics for twodimensional rectangular lattices. Journal of Mathematical Physics, 12 (10):2113–2115, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2113_s1.

Bedeaux:1971:RBM

[731] Dick Bedeaux, Katja Lakatos-Lindenberg, and Kurt E. Shuler. On the relation between master equations and random walks and their solutions. Journal of Mathematical Physics, 12(10):2116–2123, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i10/p2116_s1.

Shih:1971:KVE

[732] L. Y. Shih. The Korteweg–de Vries equation: Transition period. Journal of Mathematical Physics, 12(10):2124–2126, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i10/p2124_s1.

McKenna:1971:MSC

[733] J. McKenna and J. A. Morrison. Moments of solutions of a class of stochastic differential equations. Journal of Mathematical Physics, 12 (10):2126–2136, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2126_s1.

Tripathy:1971:R

[734] Kishor C. Tripathy. Representations of O(3). Journal of Mathematical Physics, 12(10):2136–2143, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i10/p2136_s1.

Liu:1971:ARS

[735] C. J. Liu and Yutze Chow. An algebraic realization of spin-wave theory. Journal of Mathematical Physics, 12(10):2144–2150, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i10/
p2144_s1.

Bender:1971:CDI

[736] Carl M. Bender and David J. Griffiths. Covariant derivatives of infinitedimensional fields. Journal of Mathematical Physics, 12(10):2151–2156, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2151_s1.

Ma:1971:SSL

[737] Shang keng Ma. Second sound in a low-temperature weakly interacting Bose gas. Journal of Mathematical Physics, 12(10):2157–2170, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i10/p2157_s1.

Couch:1971:RSE

[738] W. E. Couch and W. H. Hallidy. Radiation scattering in Einstein– Maxwell theory. Journal of Mathematical Physics, 12(10):2170–2175, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2170_s1.

Miller:1971:SMD

[739] Bruce N. Miller. The statistical mechanical derivation of the stress tensor and heat flux for a system of spherical molecules. Journal of Mathematical Physics, 12(10):2175–2187, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i10/p2175_s1.

Kanal:1971:MST

[740] Madhoo Kanal. Multiple scattering theory of radiative transfer in the light of singular modes. Journal of Mathematical Physics, 12(10):2187– 2194, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2187_s1.

Shastry:1971:RES

[741] C. S. Shastry and A. K. Rajagopal. Remarks on the existence of solutions of the two-particle Lippmann–Schwinger equation. II. Journal of Mathematical Physics, 12(10):2194–2198, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i10/p2194_s1.

Doolen:1971:AFW

[742] G. Doolen and J. Nuttall. Asymptotic form of the wavefunction for neutral particle-hydrogen atom scattering. Journal of Mathematical Physics, 12(10):2198–2203, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2198_s1.

Castagnino:1971:RSS

[743] Mario A. Castagnino. The Riemannian structure of space–time as a consequence of a measurement method. Journal of Mathematical Physics, 12(10):2203–2211, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2203_s1.

Robinson:1971:ARC

[744] G. de B. Robinson. The algebras of representations and classes of finite groups. Journal of Mathematical Physics, 12(10):2212–2215, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i10/p2212_s1.

Mizushima:1971:PF

[745] Masataka Mizushima. Perturbation formulas. Journal of Mathematical Physics, 12(10):2216–2220, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i10/p2216_s1.

Li:1971:EDG

[746] Shou yung Li. Explicit decomposition of a general two-body Hamiltonian into irreducible symplectic tensors. Journal of Mathematical Physics, 12 (10):2221–2226, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2221_s1.

Stern:1971:CNV

[747] David P. Stern. The canonization of nice variables. Journal of Mathematical Physics, 12(10):2226–2231, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i10/p2226_s1.

Stern:1971:CAP

[748] David P. Stern. Classical adiabatic perturbation theory. Journal of Mathematical Physics, 12(10):2231–2242, October 1, 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i10/p2231_s1.

Hooshyar:1971:ISP

[749] M. A. Hooshyar. On the inverse scattering problem at fixed energy for the tensor and spin-orbit potentials. Journal of Mathematical Physics, 12(10):2243–2258, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2243_s1.

Wegner:1971:DGI

[750] Franz J. Wegner. Duality in generalized Ising models and phase transitions without local order parameters. Journal of Mathematical Physics, 12(10):2259–2272, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2259_s1.

Burt:1971:IAP

[751] P. B. Burt and F. J. Woods. Integral approximations in potential scattering. Journal of Mathematical Physics, 12(11):2275–2280, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i11/p2275_s1.

Heskia:1971:UGS

[752] Said Heskia and S. A. Sofroniou. The unitary group U(m) as the symmetry group of curved phase space. Journal of Mathematical Physics, 12(11):2280–2286, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i11/p2280_s1.

Pomeau:1971:CLN

[753] Y. Pomeau. Correlation length in a nonequilibrium gas. Journal of Mathematical Physics, 12(11):2286–2295, November 1971. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i11/p2286_s1.

Stacey:1971:VFT

[754] Weston M. Stacey, Jr. Variational field theory. Journal of Mathematical Physics, 12(11):2296–2301, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i11/p2296_s1.

Shaffer:1971:MAU

[755] Gedaliah R. Shaffer. T -matrix analyticity using Fredholm theory. Journal of Mathematical Physics, 12(11):2302–2305, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i11/p2302_s1.

Huerta:1971:EEH

[756] M. A. Huerta, H. S. Robertson, and J. C. Nearing. Exact equilibration of harmonically bound oscillator chains. Journal of Mathematical Physics, 12(11):2305–2311, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i11/p2305_s1.

Hostler:1971:ESR

[757] Levere Hostler. Excited state reduced Coulomb Green’s function and f-dimensional Coulomb Green’s function in momentum space. Journal of Mathematical Physics, 12(11):2311–2319, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i11/p2311_s1.

Newman:1971:MSS

[758] Ezra T. Newman and Roger Posadas. Motion and structure of singularities in general relativity II. Journal of Mathematical Physics, 12 (11):2319–2327, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i11/p2319_s1.

Twersky:1971:RSE

[759] Victor Twersky. Relativistic scattering of electromagnetic waves by moving obstacles. Journal of Mathematical Physics, 12(11):2328–2341, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i11/p2328_s1.

Twersky:1971:CEM

[760] Victor Twersky. Conservation of energy and momentum in relativistic electromagnetic scattering. Journal of Mathematical Physics, 12 (11):2341–2348, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i11/p2341_s1.

Frank:1971:PSB

[761] William M. Frank. The phase shift. I. As a bounded functional. Journal of Mathematical Physics, 12(11):2348–2354, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i11/p2348_s1.

Persides:1971:NAM

[762] S. Persides. A new approximation method for wave theories. Journal of Mathematical Physics, 12(11):2355–2361, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i11/p2355_s1.

Mano:1971:RSP

[763] Koichi Mano. _{3}F_{2} representations for spin projection coefficients. Journal of Mathematical Physics, 12(11):2361–2364, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i11/p2361_s1.

Chen:1971:OAA

[764] Etang Chen. Operator algebras and axioms of measurements. Journal of Mathematical Physics, 12(11):2364–2371, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i11/p2364_s1.

Debney:1971:VST

[765] George C. Debney, Jr. On vacuum space–times admitting a null Killing bivector. Journal of Mathematical Physics, 12(11):2372–2377, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i11/p2372_s1.

Mickelsson:1971:EBR

[766] Jouko Mickelsson. An explicit basis for the reduction U(n + m) ↓ U(n) × U(m). Journal of Mathematical Physics, 12(11):2378–2382, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i11/p2378_s1.

Galbraith:1971:CDB

[767] Harold W. Galbraith. Coordinates and democracy in the N-body problem. Journal of Mathematical Physics, 12(11):2382–2386, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i11/p2382_s1.

Rajagopal:1971:LGS

[768] A. K. Rajagopal and C. S. Shastry. The Lorentz group symmetry of the hydrogen atom and the Coulomb T matrix. Journal of Mathematical Physics, 12(11):2387–2394, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i11/p2387_s1.

Ernst:1971:EAD

[769] Frederick J. Ernst. Exterior-algebraic derivation of Einstein field equations employing a generalized basis. Journal of Mathematical Physics, 12(11):2395–2397, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i11/p2395_s1.

Yeh:1971:BAV

[770] Robert H. T. Yeh. On the bounds of the average value of a function. Journal of Mathematical Physics, 12(11):2397–2398, November 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i11/
p2397_s1.

Lopuszanski:1971:SPP

[771] Jan T. Lopuszanski. On some properties of physical symmetries. Journal of Mathematical Physics, 12(12):2401–2412, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i12/p2401_s1.

Servadio:1971:CCC

[772] Sergio Servadio. Computation of the cluster coefficients of the δ-function gases through the U functions. Journal of Mathematical Physics, 12 (12):2413–2415, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i12/p2413_s1.

Arens:1971:CLI

[773] Richard Arens. Classical Lorentz invariant particles. Journal of Mathematical Physics, 12(12):2415–2422, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i12/p2415_s1.

Lonke:1971:CMP

[774] A. Lonke. The classical moment problem and the calculation of thermal averages. Journal of Mathematical Physics, 12(12):2422–2438, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i12/p2422_s1.

Neville:1971:TSR

[775] Donald Neville. A technique for solving recurrence relations approximately and its application to the 3-J and 6-J symbols. Journal of Mathematical Physics, 12(12):2438–2453, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i12/p2438_s1.

Chen:1971:IVR

[776] Li-Ching Chen and Joseph L. Birman. Irreducible vector and ray representations of some cubic crystal point groups in four dimensions. Journal of Mathematical Physics, 12(12):2454–2461, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i12/p2454_s1.

Aronson:1971:NAF

[777] B. Aronson, R. Lind, J. Messmer, and E. Newman. A note on asymptotically flat spaces. Journal of Mathematical Physics, 12(12):2462–2467, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i12/p2462_s1.

Witriol:1971:MBP

[778] Norman M. Witriol. Many-body point transforms. II. An exact noncluster approach to the hard-core many-body problem. Journal of Mathematical Physics, 12(12):2467–2480, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i12/p2467_s1.

Herbst:1971:OPO

[779] Ira W. Herbst. One-particle operators and local internal symmetries. Journal of Mathematical Physics, 12(12):2480–2490, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i12/
p2480_s1.

Sigal:1971:ERC

[780] Richard Sigal. Electromagnetic radiation in curved spaces. Journal of Mathematical Physics, 12(12):2490–2494, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i12/p2490_s1.

Trim:1971:CNG

[781] D. Trim and J. Wainwright. Combined neutrino-gravitational fields in general relativity. Journal of Mathematical Physics, 12(12):2494–2498, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i12/p2494_s1.

Okubo:1971:SCE

[782] S. Okubo and A. Isihara. Some considerations of entropy change. Journal of Mathematical Physics, 12(12):2498–2503, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i12/p2498_s1.

Uhlenbrock:1971:PSS

[783] D. A. Uhlenbrock. Perturbation of statistical semigroups in quantum statistical mechanics. Journal of Mathematical Physics, 12(12):2503– 2512, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i12/p2503_s1.

Ifantis:1971:MUS

[784] Evangelos K. Ifantis. Minimal uncertainty states for bounded observables. Journal of Mathematical Physics, 12(12):2512–2516, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i12/p2512_s1.

Percus:1971:CND

[785] Jerome K. Percus and Nikitas L. Petrakopoulos. On Clifford numbers, Dirac and relativistic Hamilton–Jacobi equations. Journal of Mathematical Physics, 12(12):2516–2520, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v12/i12/p2516_s1.

McLaughlin:1971:PIC

[786] David W. McLaughlin and L. S. Schulman. Path integrals in curved spaces. Journal of Mathematical Physics, 12(12):2520–2524, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i12/p2520_s1.

Frank:1971:PSIa

[787] William M. Frank and David W. McLaughlin. The phase shift. II. As a continuous functional. Journal of Mathematical Physics, 12(12):2525– 2529, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i12/p2525_s1.

Frank:1971:PSIb

[788] William M. Frank. The phase shift. III. As a continuous functional. Journal of Mathematical Physics, 12(12):2529–2537, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i12/p2529_s1.

Horwitz:1971:IDP

[789] L. P. Horwitz, J. A. LaVita, and J.-P. Marchand. The inverse decay problem. Journal of Mathematical Physics, 12(12):2537–2543, December 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i12/p2537_s1.

Warke:1972:JFM

[790] Chindhu S. Warke. The Jost function of a momentum-dependent potential. Journal of Mathematical Physics, 13(1):1–3, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i1/p1_s1.

Chung:1972:GDE

[791] Sun Ho Chung and Yung Ming Chen. Green’s dyadics for elastic and electromagnetic waves in a random medium. Journal of Mathematical Physics, 13(1):4–11, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i1/p4_s1.

Nanda:1972:WVZ

[792] S. Nanda. Weaker versions of Zeeman’s conjectures on topologies for Minkowski space. Journal of Mathematical Physics, 13(1):12–15, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i1/p12_s1.

Horiguchi:1972:LGFa

[793] Tsuyoshi Horiguchi, Yoshitake Yamazaki, and Tohru Morita. Lattice Green’s function for the orthorhombic lattice in terms of the complete elliptic integral. Journal of Mathematical Physics, 13(1):16–22, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i1/p16_s1.

Mickelsson:1972:RCG

[794] J. Mickelsson and J. Niederle. On representations of the conformal group which when restricted to its Poincaré or Weyl subgroups remain irreducible. Journal of Mathematical Physics, 13(1):23–27, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i1/
p23_s1.

Devi:1972:ATI

[795] V. Syamala Devi. On an application of the theory of invariants. Journal of Mathematical Physics, 13(1):28–33, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i1/p28_s1.

Cunningham:1972:RME

[796] M. J. Cunningham. Radial matrix elements of the radial-angular factorized hydrogen atom. Journal of Mathematical Physics, 13(1):33–39, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i1/p33_s1.

Davies:1972:ERF

[797] E. B. Davies. Example related to the foundations of quantum theory. Journal of Mathematical Physics, 13(1):39–41, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i1/p39_s1.

Zauderer:1972:UAS

[798] Erich Zauderer. Uniform asymptotic solution of the radial Schrödinger equation for an electron bound in a central field potential. Journal of Mathematical Physics, 13(1):42–44, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i1/p42_s1.

Hegerfeldt:1972:CCR

[799] Gerhard C. Hegerfeldt. On canonical commutation relations and infinitedimensional measures. Journal of Mathematical Physics, 13(1):45–50, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i1/p45_s1.

Cheng:1972:ATO

[800] Ing-Yih S. Cheng and John J. Kozak. Application of the theory of Orlicz spaces to statistical mechanics. I. Integral equations. Journal of Mathematical Physics, 13(1):51–58, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i1/p51_s1.

Molyneux:1972:WPC

[801] John E. Molyneux. Wave propagation in certain one-dimensional random media. Journal of Mathematical Physics, 13(1):58–69, January 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i1/p58_s1.

Couch:1972:ABV

[802] W. E. Couch and R. J. Torrence. Asymptotic behavior of vacuum space– times. Journal of Mathematical Physics, 13(1):69–73, January 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i1/p69_s1.

Nelson:1972:NSE

[803] C. A. Nelson and A. K. Rajagopal. Non-L^{2} solutions of exactly soluble relativistic wave equations and the Lee model. Journal of Mathematical Physics, 13(1):73–78, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i1/p73_s1.

Johannsen:1972:ERU

[804] Günther Johannsen. Electromagnetic radiation in a uniformly moving, homogeneous medium. Journal of Mathematical Physics, 13(1): 78–83, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i1/p78_s1.

Wu:1972:SWC

[805] Alfred C. T. Wu. Structure of the Wigner 9-j coefficients in the Bargmann approach. Journal of Mathematical Physics, 13(1):84–90, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i1/p84_s1.

Dyson:1972:CME

[806] Freeman J. Dyson. A class of matrix ensembles. Journal of Mathematical Physics, 13(1):90–97, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i1/p90_s1.

Hermann:1972:CCF

[807] Robert Hermann. Currents in classical field theories. Journal of Mathematical Physics, 13(1):97–99, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i1/p97_s1.

Eardley:1972:VDS

[808] D. Eardley, E. Liang, and Rainer Sachs. Velocity-dominated singularities in irrotational dust cosmologies. Journal of Mathematical Physics, 13(1): 99–107, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i1/p99_s1.

Lerche:1972:SEK

[809] I. Lerche. Statistically exact kinematic dynamo action in a box. Journal of Mathematical Physics, 13(1):107–115, January 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i1/p107_s1.

Morita:1972:GSD

[810] Tohru Morita. General structure of the distribution functions for the Heisenberg model and the Ising model. Journal of Mathematical Physics, 13(1):115–123, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i1/p115_s1.

York:1972:MSG

[811] James W. York, Jr. Mapping onto solutions of the gravitational initial value problem. Journal of Mathematical Physics, 13(2):125–130, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i2/p125_s1.

Bars:1972:OTG

[812] Itzhak Bars and Feza Gürsey. Operator treatment of the Gel’fand– Naimark basis for SL(2, C). Journal of Mathematical Physics, 13(2):131– 143, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p131_s1.

Maugin:1972:DMSa

[813] Gérard A. Maugin and A. Cemal Eringen. Deformable magnetically saturated media. I. Field equations. Journal of Mathematical Physics, 13(2):143–155, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i2/p143_s1.

Brandas:1972:VMW

[814] Erkki J. Brändas and David A. Micha. Variational methods in the wave operator formalism: applications in variation-perturbation theory and the theory of energy bounds. Journal of Mathematical Physics, 13(2):155–160, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i2/p155_s1.

Ramaker:1972:OCT

[815] David E. Ramaker. One-center two-electron integrals arising in electronion scattering calculations. Journal of Mathematical Physics, 13(2):161– 169, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p161_s1.

Neubert:1972:ASL

[816] Jerome A. Neubert and John L. Lumley. Approach to stochastic Lagrangian integrals and their asymptotic evaluation for sound propagation in continuous stochastic media. Journal of Mathematical Physics, 13(2):169–183, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i2/p169_s1.

Sharp:1972:B

[817] R. T. Sharp. SU(n − 2) × SU(2) × U(1) bases for SU(n). Journal of Mathematical Physics, 13(2):183–186, February 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i2/p183_s1.

Prosser:1972:ABC

[818] Reese T. Prosser. On the asymptotic behavior of certain dynamical systems. Journal of Mathematical Physics, 13(2):186–196, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i2/
p186_s1.

Conatser:1972:CLD

[819] C. W. Conatser. Contractions of the low-dimensional real Lie algebras. Journal of Mathematical Physics, 13(2):196–203, February 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i2/p196_
s1.

Sheaks:1972:PTS

[820] O. J. Sheaks. Particle transport in spherical media with a central black cavity. Journal of Mathematical Physics, 13(2):203–209, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i2/
p203_s1.

Cobb:1972:DPS

[821] Donald D. Cobb and Edward R. McCliment. Decomposition of the principal series of unitary irreducible representations of SU(2, 2) restricted to the subgroup SU(1, 1) ⊗ SU(2). Journal of Mathematical Physics, 13(2):209–214, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i2/p209_s1.

Coleman:1972:NCR

[822] A. J. Coleman. Necessary conditions for N-representability of reduced density matrices. Journal of Mathematical Physics, 13(2):214– 222, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p214_s1.

Millard:1972:SMA

[823] Kenneth Millard. A statistical mechanical approach to the problem of a fluid in an external field. Journal of Mathematical Physics, 13(2):222– 226, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p222_s1.

Yeh:1972:LBE

[824] Robert H. T. Yeh. Lower bounds of the energy eigenvalues of systems containing identical particles. Journal of Mathematical Physics, 13(2): 227, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p227_s1.

Yeh:1972:ULB

[825] Robert H. T. Yeh. Upper and lower bounds of the eigenvalues of a secondorder linear self-adjoint differential equation. Journal of Mathematical Physics, 13(2):227–228, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v13/i2/p227_s2.

Kagiwada:1972:RCN

[826] Harriet Kagiwada, Robert Kalaba, and Chung-Chun Yang. Reduction of a class of nonlinear integral equations to a Cauchy system. Journal of Mathematical Physics, 13(2):228–231, February 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i2/p228_s1.

Benioff:1972:OVM

[827] Paul A. Benioff. Operator valued measures in quantum mechanics: finite and infinite processes. Journal of Mathematical Physics, 13(2):231– 242, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p231_s1.

McQuistan:1972:OSP

[828] R. B. McQuistan, S. J. Lichtman, and L. P. Levine. Occupation statistics for parallel dumbbells on a 2 × N lattice space. Journal of Mathematical Physics, 13(2):242–248, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v13/i2/p242_s1.

Liu:1972:PCA

[829] C. J. Liu. Perturbation calculation of an algebraic realization of spin waves. Journal of Mathematical Physics, 13(2):248–252, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i2/
p248_s1.

Wilhelm:1972:NIB

[830] H. E. Wilhelm. Nonlinear initial-boundary-value problem for convection, diffusion, ionization, and recombination processes. Journal of Mathematical Physics, 13(2):252–256, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i2/p252_s1.

Volkel:1972:TAS

[831] A. H. Völkel. Timelike asymptotic series of almost local fields. Journal of Mathematical Physics, 13(2):257–268, February 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i2/p257_s1.

OCarroll:1972:ECG

[832] Michael O’Carroll. The existence and completeness of generalized wave operators for spherically symmetric, asymptotically Coulomb potentials. Journal of Mathematical Physics, 13(2):268–271, February 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i2/p268_
s1.

Adomian:1972:ALR

[833] G. Adomian. Addendum: “Linear Random Operator Equations in Mathematical Physics. III”. Journal of Mathematical Physics, 13(2): 272, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p272_s1. See [706].

Liboff:1972:ERF

[834] Richard L. Liboff. Errata: “Relativistic Fields Due to a Particle in a Grounded Cylindrical Box”. Journal of Mathematical Physics, 13(2):

272, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i2/p272_s2. See [162].

Dresden:1972:CNR

[835] M. Dresden and A. Albano. A class of nonlinear realizations of the Poincaré group. Journal of Mathematical Physics, 13(3):275–282, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i3/p275_s1.

Ryan:1972:HFG

[836] C. Ryan. Hamiltonian formalism for general Lagrangian systems in an exceptional case. Journal of Mathematical Physics, 13(3):283–285, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i3/p283_s1.

Szekeres:1972:CPG

[837] P. Szekeres. Colliding plane gravitational waves. Journal of Mathematical Physics, 13(3):286–294, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i3/p286_s1.

Thorpe:1972:RWP

[838] M. F. Thorpe. Random walks in polytype structures. Journal of Mathematical Physics, 13(3):294–299, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i3/p294_s1.

Morrison:1972:MCF

[839] J. A. Morrison. Moments and correlation functions of solutions of some stochastic matrix differential equations. Journal of Mathematical Physics, 13(3):299–306, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i3/p299_s1.

Warren:1972:PRR

[840] Russell E. Warren and William H. Klink. Projective representations and the relation of internal to space–time symmetries. Journal of Mathematical Physics, 13(3):306–312, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i3/p306_s1.

Warren:1972:POI

[841] Russell E. Warren and William H. Klink. Position operators as “internal” symmetries. Journal of Mathematical Physics, 13(3):312–316, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i3/p312_s1.

Devi:1972:CRP

[842] K. R. Sandhya Devi and Nazakat Ullah. On the correlations of the resonance parameters for the overlapping resonances. Journal of Mathematical Physics, 13(3):316–321, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i3/p316_s1.

Raghavacharyulu:1972:SDE

[843] I. V. V. Raghavacharyulu and A. R. Tekumalla. Solution of the difference equations of generalized Lucas polynomials. Journal of Mathematical Physics, 13(3):321–324, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v13/i3/p321_s1.

Yngve:1972:NCH

[844] Staffan Yngve. Normalization of certain higher-order phase-integral approximations for wavefunctions of bound states in a potential well. II. Journal of Mathematical Physics, 13(3):324–331, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i3/p324_s1.

Andrie:1972:CML

[845] M. Andrié and D. J. Simms. Constants of motion and Lie group actions. Journal of Mathematical Physics, 13(3):331–336, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i3/p331_s1.

DeSanto:1972:SPC

[846] John A. DeSanto. Scattering from a periodic corrugated structure. II. Thin comb with hard boundaries. Journal of Mathematical Physics, 13(3):336–341, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i3/p336_s1.

Janssen:1972:PUA

[847] T. Janssen. On projective unitary antiunitary representations of finite groups. Journal of Mathematical Physics, 13(3):342–351, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i3/p342_s1.

Joseph:1972:CPQ

[848] A. Joseph. Commuting polynomials in quantum canonical operators and realizations of Lie algebras. Journal of Mathematical Physics, 13(3): 351–357, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i3/p351_s1.

Besieris:1972:SME

[849] Ioannis M. Besieris. Stochastic master equations: A perturbative approach. Journal of Mathematical Physics, 13(3):358–360, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i3/
p358_s1.

Case:1972:WPI

[850] K. M. Case. On wave propagation in inhomogeneous media. Journal of Mathematical Physics, 13(3):360, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i3/p360_s1.

Tiersten:1972:IEF

[851] H. F. Tiersten and C. F. Tsai. On the interaction of the electromagnetic field with heat conducting deformable insulators. Journal of Mathematical Physics, 13(3):361–378, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i3/p361_s1.

Greer:1972:QTH

[852] William L. Greer and Robert J. Rubin. Quantum theory of heat transport in an isotopically substituted, one-dimensional, harmonic crystal. Journal of Mathematical Physics, 13(3):379–385, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i3/p379_s1.

Liang:1972:VDS

[853] E. P. T. Liang. Velocity-dominated singularities in irrotational hydrodynamic cosmological models. Journal of Mathematical Physics, 13(3): 386–393, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i3/p386_s1.

Geroch:1972:MGN

[854] Robert Geroch. A method for generating new solutions of Einstein’s equation. II. Journal of Mathematical Physics, 13(3):394–404, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i3/p394_s1.

Lepore:1972:RSZ

[855] Joseph V. Lepore and Robert J. Riddell, Jr. Relativistic spin-zero wave equation. Journal of Mathematical Physics, 13(3):405–418, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i3/
p405_s1.

Andrews:1972:NPO

[856] L. C. Andrews and T. Triffet. Nonlinear perturbation operators. Journal of Mathematical Physics, 13(3):419–433, March 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i3/p419_s1.

Huang:1972:EGF

[857] Justin C. Huang and Brian DeFacio. Errata: “On Green’s Functions to the Bethe–Salpeter Equation”, [J. Math. Phys. 11, 715 (1970)]. Journal of Mathematical Physics, 13(3):434, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i3/p434_s1. See [90].

Lalor:1972:ASRa

[858] Éamon Lalor. The angular spectrum representation of electromagnetic fields in crystals. I. Uniaxial crystals. Journal of Mathematical Physics, 13(4):437–443, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i4/p437_s1.

Lalor:1972:ASRb

[859] Éamon Lalor. The angular spectrum representation of electromagnetic fields in crystals. II. Biaxial crystals. Journal of Mathematical Physics, 13(4):443–449, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i4/p443_s1.

Lalor:1972:AAT

[860] Éamon Lalor. An analytical approach to the theory of internal conical refraction. Journal of Mathematical Physics, 13(4):449–454, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i4/p449_s1.

Wybourne:1972:ILG

[861] B. G. Wybourne. Internal labeling and the group G_{2}. Journal of Mathematical Physics, 13(4):455–457, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i4/p455_s1.

Tsilimigras:1972:ERA

[862] P. Tsilimigras. On the existence of relations among the eigenvalues of difference equations. Journal of Mathematical Physics, 13(4):457–458, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i4/p457_s1.

Eftimiu:1972:PEC

[863] C. Eftimiu. Proof of an existence condition for solutions of the unitarity system of equations. Journal of Mathematical Physics, 13(4):458– 459, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i4/p458_s1.

Hermann:1972:GCM

[864] Robert Hermann. Geodesics and classical mechanics on Lie groups. Journal of Mathematical Physics, 13(4):460–464, April 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i4/p460_s1.

Delbourgo:1972:NLI

[865] R. Delbourgo. Nonpolynomial Lagrangians with isospin. Journal of Mathematical Physics, 13(4):464–468, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i4/p464_s1.

Silver:1972:CMF

[866] H. Silver, N. E. Frankel, and B. W. Ninham. A class of mean field models. Journal of Mathematical Physics, 13(4):468–474, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i4/p468_s1.

Deliyannis:1972:ESM

[867] Platon C. Deliyannis. Exact and simultaneous measurements. Journal of Mathematical Physics, 13(4):474–477, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i4/p474_s1.

Dimock:1972:SLH

[868] Jonathan Dimock. Spectrum of local Hamiltonians in the yukawa_{2} field theory. Journal of Mathematical Physics, 13(4):477–481, April 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i4/p477_
s1.

Howard:1972:MTE

[869] A. Q. Howard. On the mathematical theory of electromagnetic radiation from flanged waveguides. Journal of Mathematical Physics, 13(4):482– 490, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i4/p482_s1.

Wald:1972:UKN

[870] Robert Wald. On uniqueness of the Kerr–Newman black holes. Journal of Mathematical Physics, 13(4):490–499, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i4/p490_s1.

Ferrara:1972:LCS

[871] S. Ferrara and G. Rossi. Light-cone singularities and Lorentz poles. Journal of Mathematical Physics, 13(4):499–504, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i4/p499_s1.

Cattaneo:1972:SEIa

[872] U. Cattaneo. Sliced extensions, irreducible extensions, and associated graphs: an analysis of Lie algebra extensions. I. General theory. Journal of Mathematical Physics, 13(4):504–518, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i4/p504_s1.

Cattaneo:1972:SEIb

[873] U. Cattaneo. Sliced extensions, irreducible extensions, and associated graphs: An analysis of Lie algebra extensions. II. Application to Euclidean, Poincaré, and Galilean algebras. Journal of Mathematical Physics, 13(4):518–528, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i4/p518_s1.

Hoffman:1972:GEA

[874] David K. Hoffman, Richard C. Raffenetti, and Klaus Ruedenberg. Generalization of Euler angles to N-dimensional orthogonal matrices. Journal of Mathematical Physics, 13(4):528–533, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i4/p528_s1.

Onofri:1972:DQ

[875] E. Onofri and M. Pauri. Dynamical quantization. Journal of Mathematical Physics, 13(4):533–543, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i4/p533_s1.

Cohen:1972:NAM

[876] Jeffrey M. Cohen and Robert M. Wald. Note on the angular momentum and mass of gravitational geons. Journal of Mathematical Physics, 13(4): 543–545, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i4/p543_s1.

Fischer:1972:EEE

[877] Arthur E. Fischer and Jerrold E. Marsden. The Einstein equations of evolution — a geometric approach. Journal of Mathematical Physics, 13(4): 546–568, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i4/p546_s1.

Ifantis:1972:SMU

[878] Evangelos K. Ifantis. States, minimizing the uncertainty product of the oscillator phase operators. Journal of Mathematical Physics, 13(4):568– 575, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i4/p568_s1.

Alisauskas:1972:EWL

[879] S. J. Alisauskas and A. P. Jucys. Erratum: “Weight Lowering Operators and the Multiplicity-Free Isoscalar Factors for the Group R_{5}”. Journal of Mathematical Physics, 13(4):575, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i4/p575_s1. See [511].

Ifantis:1972:ESS

[880] E. K. Ifantis. Erratum: “Solution of the Schrödinger Equation in the Hardy–Lebesgue Space”. Journal of Mathematical Physics, 13(4):575, April 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i4/p575_s2. See [708].

Chacon:1972:EMF

[881] E. Chacón, Mikael Ciftan, and L. C. Biedenharn. On the evaluation of the multiplicity-free Wigner coefficients of U(n). Journal of Mathematical Physics, 13(5):577–590, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i5/p577_s1.

Taylor:1972:SRE

[882] Leonard S. Taylor. Scintillation of randomized electromagnetic fields. Journal of Mathematical Physics, 13(5):590–595, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p590_s1.

Rosen:1972:ESY

[883] Gerald Rosen. Exact solutions to the Yang–Mills field equations. Journal of Mathematical Physics, 13(5):595–597, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i5/p595_s1.

Chadam:1972:CPC

[884] John M. Chadam. On the Cauchy problem for the coupled Maxwell– Dirac equations. Journal of Mathematical Physics, 13(5):597–604, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p597_s1.

Penrose:1972:FET

[885] O. Penrose and J. L. Lebowitz. A functional equation in the theory of fluids. Journal of Mathematical Physics, 13(5):604–607, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p604_s1.

Falk:1972:EDT

[886] H. Falk. Entropy decomposition and transfer-matrix problems. Journal of Mathematical Physics, 13(5):608–609, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i5/p608_s1.

Zachary:1972:NTD

[887] W. W. Zachary. Nonrelativistic time-dependent scattering theory and von Neumann algebras. I. Single channel scattering. Journal of Mathematical Physics, 13(5):609–615, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i5/p609_s1.

Canosa:1972:ABS

[888] José Canosa. Asymptotic behavior of Sturm–Liouville systems. Journal of Mathematical Physics, 13(5):615–621, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i5/p615_s1.

Partensky:1972:EIO

[889] A. Partensky. On the eigenvalues of the invariant operators of the unitary unimodular group SU(n). Journal of Mathematical Physics, 13(5): 621–625, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i5/p621_s1.

Fleming:1972:SSU

[890] Gordon N. Fleming. The spin spectrum of an unstable particle. Journal of Mathematical Physics, 13(5):626–637, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i5/p626_s1.

Babbitt:1972:PIQ

[891] Donald G. Babbitt. Probabilistic interpretation of the quantum scattering cross section. Journal of Mathematical Physics, 13(5):637–640, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p637_s1.

Hajj:1972:MRS

[892] Jack P. Hajj and David L. Weaver. Multipole radiation solutions for a class of massive particle wave equations. Journal of Mathematical Physics, 13(5):640–645, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i5/p640_s1.

Lu:1972:MWA

[893] Pao Lu and Sheldon S. Wald. Modified WKB approximation applied to the solution of the repulsive singular potential. Journal of Mathematical Physics, 13(5):646–648, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i5/p646_s1.

Miller:1972:CGCa

[894] Willard Miller, Jr. Clebsch–Gordan coefficients and special function identities. I. The harmonic oscillator group. Journal of Mathematical Physics, 13(5):648–655, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i5/p648_s1.

Maekawa:1972:NEF

[895] Takayoshi Maekawa. Note on the explicit form of invariant operators for O(n). Journal of Mathematical Physics, 13(5):655–656, May 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p655_
s1.

Lee:1972:RTD

[896] S. W. Lee. Ray theory of diffraction by open-ended waveguides. II. Applications. Journal of Mathematical Physics, 13(5):656–664, May 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p656_
s1.

Lugiato:1972:SRG

[897] Luigi A. Lugiato and Vittorio Gorini. On the structure of relativity groups. Journal of Mathematical Physics, 13(5):665–671, May 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p665_
s1.

Ranada:1972:ESE

[898] Antonio F. Rañada and Mario Soler. Elementary spinorial excitations in a model universe. Journal of Mathematical Physics, 13(5):671–675, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p671_s1.

Sabatier:1972:CSI

[899] Pierre C. Sabatier. Complete solution of the inverse scattering problem at fixed energy. Journal of Mathematical Physics, 13(5):675–699, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p675_s1.

Lunsford:1972:SPO

[900] Gary H. Lunsford and Joseph Ford. On the stability of periodic orbits for nonlinear oscillator systems in regions exhibiting stochastic behavior. Journal of Mathematical Physics, 13(5):700–705, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p700_s1.

Majumdar:1972:SFE

[901] Chanchal K. Majumdar. Solution of Faddeev equations for a onedimensional system. Journal of Mathematical Physics, 13(5):705–708, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p705_s1.

Mahanta:1972:ASS

[902] M. N. Mahanta and D. R. K. Reddy. An approximate solution for the static, spherically symmetric metric due to a point charged mass in Brans–Dicke theory. Journal of Mathematical Physics, 13(5):708–709, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p708_s1.

Hautot:1972:ASD

[903] André Hautot. About the solutions of Dirac’s equations in the presence of new magnetic fields. Journal of Mathematical Physics, 13(5): 710–714, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i5/p710_s1.

Morita:1972:TTSa

[904] Tohru Morita. Two-time spin-pair correlation function of the Heisenberg magnet at infinite temperature. II. Journal of Mathematical Physics, 13(5):714–720, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i5/p714_s1.

Barston:1972:EPL

[905] E. M. Barston. Eigenvalue problem for Lagrangian systems. VI. Journal of Mathematical Physics, 13(5):720–725, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i5/p720_s1.

Chen:1972:EIM

[906] H. H. Chen and R. K. Joseph. Exchange interaction model of ferromagnetism. Journal of Mathematical Physics, 13(5):725–739, May 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p725_

s1. See erratum [1073].

Kunzle:1972:CDS

[907] H. P. Künzle. Canonical dynamics of spinning particles in gravitational and electromagnetic fields. Journal of Mathematical Physics, 13(5):739– 744, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i5/p739_s1.

Gouyet:1972:SDO

[908] J. F. Gouyet. Spin-dependent operators in the occupation-branching number representation. Journal of Mathematical Physics, 13(5):745–751, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p745_s1.

Abellanas:1972:AUM

[909] L. Abellanas and A. Galindo. Analytic unimodular matrix groups containing the image D^{j} of the rotation group: a simple derivation. Journal of Mathematical Physics, 13(5):751–753, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i5/p751_s1.

Shen:1972:FCA

[910] C. Y. Shen. A functional calculus approach to the Ursell–Mayer functions. Journal of Mathematical Physics, 13(5):754–759, May 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p754_
s1.

Whitman:1972:PPD

[911] G. M. Whitman and L. B. Felsen. Pulse propagation in a dispersive medium with moving density profile. Journal of Mathematical Physics, 13(5):760–768, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i5/p760_s1.

Kuchar:1972:BTC

[912] Karel Kuchar. A bubble-time canonical formalism for geometrodynamics. Journal of Mathematical Physics, 13(5):768–781, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p768_s1.

Lieb:1972:PZP

[913] Elliott H. Lieb and David Ruelle. A property of zeros of the partition function for Ising spin systems. Journal of Mathematical Physics, 13(5): 781–784, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i5/p781_s1.

McLaughlin:1972:PIA

[914] David W. McLaughlin. Path integrals, asymptotics, and singular perturbations. Journal of Mathematical Physics, 13(5):784–796, May 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i5/p784_
s1.

Tesser:1972:GM

[915] Herbert Tesser. Generalized mechanics. Journal of Mathematical Physics, 13(5):796–799, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i5/p796_s1.

Gudder:1972:NPN

[916] S. Gudder and J.-P. Marchand. Noncommutative probability on von Neumann algebras. Journal of Mathematical Physics, 13(5):799–806, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i5/p799_s1.

Cooper:1972:PNG

[917] W. Grant Cooper and Donald J. Kouri. N-particle noninteracting Green’s function. Journal of Mathematical Physics, 13(6):809–813, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i6/p809_s1.

Schilder:1972:ASA

[918] Michael Schilder. Another set of axioms for classical gas dynamics. Journal of Mathematical Physics, 13(6):813–821, June 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i6/p813_s1.

Hegerfeldt:1972:GDA

[919] Gerhard C. Hegerfeldt. Gårding domains and analytic vectors for quantum fields. Journal of Mathematical Physics, 13(6):821–827, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/
p821_s1.

Miller:1972:CGCb

[920] Willard Miller, Jr. Clebsch–Gordan coefficients and special function identities. II. The rotation and Lorentz groups in 3-space. Journal of Mathematical Physics, 13(6):827–833, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i6/p827_s1.

Hermann:1972:SGAa

[921] Robert Hermann. Spectrum generating algebras and symmetries in mechanics. I. Journal of Mathematical Physics, 13(6):833–838, June 1972.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/
p833_s1.

Hermann:1972:SGAb

[922] Robert Hermann. Spectrum generating algebras and symmetries in mechanics. II. Journal of Mathematical Physics, 13(6):838–842, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/
p838_s1.

vanKampen:1972:EME

[923] N. G. van Kampen and Irwin Oppenheim. Expansion of the master equation for one-dimensional random walks with boundary. Journal of Mathematical Physics, 13(6):842–849, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i6/p842_s1.

Gunzig:1972:TAS

[924] E. Gunzig. Topological analysis of self-generating interactions in Lagrangian field theory. Journal of Mathematical Physics, 13(6):850–862, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i6/p850_s1.

deVries:1972:CCT

[925] E. de Vries and A. J. van Zanten. Calculation of a certain type of 6jsymbol. Journal of Mathematical Physics, 13(6):862–863, June 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/p862_
s1.

Muraskin:1972:LES

[926] M. Muraskin. Local existence of solutions to the equations Γ_{jk;l}^{i} = 0, g_{ij;k} = 0 when R_{jkl}^{i} ≠ 0. Journal of Mathematical Physics, 13(6): 863–865, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i6/p863_s1.

Israel:1972:CSE

[927] Werner Israel and Gordon A. Wilson. A class of stationary electromagnetic vacuum fields. Journal of Mathematical Physics, 13(6):865–867, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i6/p865_s1.

Shukre:1972:CER

[928] C. S. Shukre and T. F. Jordan. Constructing examples in relativistic classical particle mechanics. Journal of Mathematical Physics, 13(6): 868–871, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i6/p868_s1.

Gelman:1972:IIC

[929] Donald Gelman. Identities involving the coefficients in the asymptotic expansion of the outgoing scattering solution of the Schrödinger equation. Journal of Mathematical Physics, 13(6):871–873, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/p871_s1.

Lovelock:1972:FDS

[930] David Lovelock. The four-dimensionality of space and the Einstein tensor. Journal of Mathematical Physics, 13(6):874–876, June 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/p874_
s1.

Seligman:1972:WBU

[931] T. H. Seligman. The Weyl basis of the unitary group U(k). Journal of Mathematical Physics, 13(6):876–879, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i6/p876_s1.

Newton:1972:RBT

[932] Roger G. Newton. Relation between the three-dimensional Fredholm determinant and the Jost functions. Journal of Mathematical Physics, 13(6):880–883, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i6/p880_s1.

Gilmore:1972:RE

[933] Robert Gilmore. Rank 1 expansions. Journal of Mathematical Physics, 13(6):883–886, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i6/p883_s1.

Glasser:1972:APF

[934] M. L. Glasser, D. B. Abraham, and E. H. Lieb. Analytic properties of the free energy for the ‘ice’ models. Journal of Mathematical Physics, 13(6): 887–900, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i6/p887_s1.

Moshinsky:1972:CTR

[935] M. Moshinsky, T. H. Seligman, and K. B. Wolf. Canonical transformations and the radial oscillator and Coulomb problems. Journal of Mathematical Physics, 13(6):901–907, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i6/p901_s1.

Benioff:1972:DPQ

[936] Paul A. Benioff. Decision procedures in quantum mechanics. Journal of Mathematical Physics, 13(6):908–915, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i6/p908_s1.

Beckers:1972:HCS

[937] J. Beckers. Helicity and canonical spin operators of the Poincaré algebra. Journal of Mathematical Physics, 13(6):915–918, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/p915_s1.

Rosen:1972:RHS

[938] Lon Rosen. Renormalization of the Hilbert space in the mass shift model. Journal of Mathematical Physics, 13(6):918–927, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/p918_s1.

Baraff:1972:EGF

[939] G. A. Baraff. Erratum: “Green’s Function for the Nonlocal Wave Equation”. Journal of Mathematical Physics, 13(6):927, June 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i6/p927_s1. See

[246].

Couch:1972:GLW

[940] W. E. Couch and E. T. Newman. Generalized Lienard–Wiechert fields. Journal of Mathematical Physics, 13(7):929–931, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i7/p929_s1.

Matzner:1972:CRC

[941] Richard A. Matzner. Closed rotating cosmologies containing matter described by the kinetic theory. Entropy production in the collision time approximation. Journal of Mathematical Physics, 13(7):931–937, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i7/p931_s1.

Seetharaman:1972:PTI

[942] M. Seetharaman and P. M. Mathews. Poincaré and TCP invariance in the determination of wave equations for particles of arbitrary spin. Journal of Mathematical Physics, 13(7):938–943, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i7/p938_s1.

Wang:1972:BSN

[943] P. K. C. Wang. Bounds for solution of nonlinear wave-wave interacting systems with well-defined phase description. Journal of Mathematical Physics, 13(7):943–947, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i7/p943_s1.

Bowen:1972:SCB

[944] Ray M. Bowen and Peter J. Chen. Some comments on the behavior of acceleration waves of arbitrary shape. Journal of Mathematical Physics, 13(7):948–950, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i7/p948_s1.

Beers:1972:HAA

[945] Brian Lee Beers. Harmonic analysis of analytic functions on hyperspheres. Journal of Mathematical Physics, 13(7):950–955, July 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i7/p950_
s1.

Geroch:1972:SGF

[946] Robert Geroch. Structure of the gravitational field at spatial infinity. Journal of Mathematical Physics, 13(7):956–968, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i7/p956_s1.

Prugovecki:1972:STF

[947] Eduard Prugovecki. Scattering theory in Fock space. Journal of Mathematical Physics, 13(7):969–976, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i7/p969_s1.

Federbush:1972:EDT

[948] Paul Federbush. On the equivalence of dressing transformations. Journal of Mathematical Physics, 13(7):977–979, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i7/p977_s1.

Foster:1972:TDM

[949] J. C. Foster, Jr. and J. R. Ray. A tachyon dust metric in general relativity. Journal of Mathematical Physics, 13(7):979–982, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i7/p979_s1.

Marchesin:1972:TDT

[950] Dan Marchesin and Michael L. O’Carroll. Time-dependent and timeindependent potential scattering for asymptotically Coulomb potentials. Journal of Mathematical Physics, 13(7):982–990, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i7/p982_s1.

Sen:1972:WLM

[951] D. K. Sen and J. R. Vanstone. On Weyl and Lyra manifolds. Journal of Mathematical Physics, 13(7):990–993, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i7/p990_s1.

Uhlenbrock:1972:PGS

[952] D. A. Uhlenbrock. Perturbations of Gibbs states. I. Grand canonical formalism for self-interacting fermion systems. Journal of Mathematical Physics, 13(7):993–1001, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i7/p993_s1.

Balanis:1972:PIP

[953] George Nick Balanis. The plasma inverse problem. Journal of Mathematical Physics, 13(7):1001–1005, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i7/p1001_s1.

Gates:1972:LWTa

[954] D. J. Gates. Lattice wind-tree models. I. Absence of diffusion. Journal of Mathematical Physics, 13(7):1005–1013, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i7/p1005_s1.

Kanal:1972:GTS

[955] Madhoo Kanal. General theory of spherically symmetric boundaryvalue problems of the linear transport theory. Journal of Mathematical Physics, 13(7):1013–1025, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v13/i7/p1013_s1.

Daniel:1972:ARE

[956] M. Daniel and P. K. Mitter. Analytic renormalization of the exponential interaction: the three-point time-ordered product with minimum light cone singularity. Journal of Mathematical Physics, 13(7):1026– 1041, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i7/p1026_s1.

Hardy:1972:THM

[957] J. Hardy and Y. Pomeau. Thermodynamics and hydrodynamics for a modeled fluid. Journal of Mathematical Physics, 13(7):1042–1051, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i7/p1042_s1.

Nelson:1972:ASK

[958] C. A. Nelson. Algebraic solution for the Källén–Pauli state vectors in the V θ sector by a congruence transformation. Journal of Mathematical Physics, 13(7):1051–1056, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v13/i7/p1051_s1.

Hemenger:1972:IEQ

[959] Richard P. Hemenger, Robert M. Pearlstein, and Katja Lakatos-Lindenberg. Incoherent exciton quenching on lattices. Journal of Mathematical Physics, 13(7):1056–1063, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i7/p1056_s1.

Abellanas:1972:MCB

[960] Lorenzo Abellanas. On the mathematical characterization of the breaking of hadronic internal symmetry. Journal of Mathematical Physics, 13(7): 1064–1069, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i7/p1064_s1.

Goovaerts:1972:ATC

[961] M. J. Goovaerts and J. T. Devreese. Analytic treatment of the Coulomb potential in the path integral formalism by exact summation of a perturbation expansion. Journal of Mathematical Physics, 13(7):1070–1082, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v13/i7/p1070_s1. See erratum [1118].

Presutti:1972:SAT

[962] E. Presutti, E. Scacciatelli, G. L. Sewell, and F. Wanderlingh. Studies in the C∗-algebraic theory of nonequilibrium statistical mechanics: dynamics of open and of mechanically driven systems. Journal of Mathematical Physics, 13(8):1085–1098, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i8/p1085_s1.

McLaughlin:1972:CTC

[963] David W. McLaughlin. Complex time, contour independent path integrals, and barrier penetration. Journal of Mathematical Physics, 13 (8):1099–1108, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i8/p1099_s1.

Cunningham:1972:SGQ

[964] M. J. Cunningham. An O(2, 1) × O(3) solution to a generalized quantum mechanical Kepler problem. Journal of Mathematical Physics, 13(8):1108–1112, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i8/p1108_s1.

Swenson:1972:SCN

[965] Robert J. Swenson. Solutions of a class of nonlinear coupled partial differential equations. Journal of Mathematical Physics, 13(8):1112– 1114, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1112_s1.

Silvennoinen:1972:MTT

[966] Pekka Silvennoinen and P. F. Zweifel. On multigroup transport theory with a degenerate transfer kernel. Journal of Mathematical Physics, 13(8):1114–1117, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i8/p1114_s1.

Satchler:1972:RMF

[967] G. R. Satchler. Radial moments of folding integrals for nonspherical distributions. Journal of Mathematical Physics, 13(8):1118–1119, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i8/p1118_s1.

Stacey:1972:VEG

[968] Weston M. Stacey, Jr. Variational estimates and generalized perturbation theory for the ratios of linear and bilinear functionals. Journal of Mathematical Physics, 13(8):1119–1125, August 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i8/p1119_s1.

Cap:1972:ADS

[969] Ferdinand F. Cap. Amplitude dispersion and stability of nonlinear weakly dissipative waves. Journal of Mathematical Physics, 13(8):1126– 1130, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1126_s1.

Chen:1972:SSE

[970] Etang Chen. Structure of simplexes of equilibrium states in quantum statistical mechanics. Journal of Mathematical Physics, 13(8):1130– 1135, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1130_s1.

Mano:1972:HSG

[971] Koichi Mano. Hypergeometric structure of the generalized Veneziano amplitudes. Journal of Mathematical Physics, 13(8):1136–1139, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i8/p1136_s1.

Avishai:1972:MFL

[972] Y. Avishai, H. Ekstein, and J. E. Moyal. Is the Maxwell field local? Journal of Mathematical Physics, 13(8):1139–1145, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i8/p1139_s1.

Glasser:1972:WSC

[973] M. L. Glasser. A Watson sum for a cubic lattice. Journal of Mathematical Physics, 13(8):1145–1146, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v13/i8/p1145_s1.

Belleni-Morante:1972:MNT

[974] A. Belleni-Morante and G. Busoni. Multigroup neutron transport. Journal of Mathematical Physics, 13(8):1146–1150, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i8/p1146_s1.

Bracci:1972:LCQ

[975] L. Bracci and F. Strocchi. Local and covariant quantization of linearized Einstein’s equations. Journal of Mathematical Physics, 13(8):1151– 1163, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1151_s1.

Ullah:1972:RGR

[976] Nazakat Ullah and K. K. Gupta. On the reduction of the generalized RPA eigenvalue problem. Journal of Mathematical Physics, 13(8):1163– 1165, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1163_s1.

Williams:1972:CKS

[977] J. G. Williams. Conservation of kinks and spin in nonlinear quantum field theories. Journal of Mathematical Physics, 13(8):1165–1168, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i8/p1165_s1.

Garrod:1972:RSM

[978] Claude Garrod and Carver Simmons. Rigorous statistical mechanics for nonuniform systems. Journal of Mathematical Physics, 13(8):1168– 1176, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1168_s1.

Manley:1972:SFS

[979] O. P. Manley and Y. M. Treve. Stability of finite-sized radiating plasmas. Journal of Mathematical Physics, 13(8):1176–1185, August 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i8/p1176_
s1.

Mattis:1972:ESM

[980] D. C. Mattis and S. B. Nam. Exactly soluble model of interacting electrons. Journal of Mathematical Physics, 13(8):1185–1189, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i8/p1185_s1.

Jackson:1972:USC

[981] E. Atlee Jackson. Unbounded solutions of coupled mode equations. Journal of Mathematical Physics, 13(8):1189–1192, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i8/p1189_s1.

Ifantis:1972:SDE

[982] Evangelos K. Ifantis and Christos B. Kouris. Study of differential equations of physics in the Hardy–Lebesgue space. Journal of Mathematical Physics, 13(8):1193–1195, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i8/p1193_s1.

Fox:1972:CTM

[983] Ronald Forrest Fox. Contributions to the theory of multiplicative stochastic processes. Journal of Mathematical Physics, 13(8):1196– 1207, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1196_s1.

Davies:1972:LWN

[984] John A. Davies. Long-wavelength normal modes of crystals with Coulomb interactions. Journal of Mathematical Physics, 13(8):1207– 1218, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1207_s1.

Akyeampong:1972:FT

[985] D. A. Akyeampong and M. A. Rashid. On the finite transformations of SU(3). Journal of Mathematical Physics, 13(8):1218–1224, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i8/p1218_s1.

Chow:1972:AFS

[986] Pao-Liu Chow. Applications of function space integrals to problems in wave propagation in random media. Journal of Mathematical Physics, 13(8):1224–1236, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i8/p1224_s1.

Emch:1972:MDB

[987] Gérard G. Emch and John C. Wolfe. A model for dissipative behavior in nonlinear quantum optics. Journal of Mathematical Physics, 13(8):1236– 1243, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1236_s1.

Morita:1972:LGF

[988] Tohru Morita and Tsuyoshi Horiguchi. Lattice Green’s function for the diatomic lattices. Journal of Mathematical Physics, 13(8):1243–1252, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i8/p1243_s1.

Devooght:1972:RKMa

[989] J. Devooght. The reproducing kernel method. I. Journal of Mathematical Physics, 13(8):1253–1259, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v13/i8/p1253_s1.

Devooght:1972:RKMb

[990] J. Devooght. The reproducing kernel method. II. Journal of Mathematical Physics, 13(8):1259–1268, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i8/p1259_s1.

Boon:1972:RIG

[991] M. H. Boon. Representations of the invariance group for a Bloch electron in a magnetic field. Journal of Mathematical Physics, 13(8):1268– 1285, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1268_s1.

Faris:1972:ICU

[992] William G. Faris. Invariant cones and uniqueness of the ground state for fermion systems. Journal of Mathematical Physics, 13(8):1285–1290, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i8/p1285_s1.

Hecht:1972:AAM

[993] W. Hecht. An analytic approximation method for the one-dimensional Schrödinger equation. I. Journal of Mathematical Physics, 13(8):1291– 1294, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i8/p1291_s1.

Love:1972:DRU

[994] J. D. Love. The dielectric ring in a uniform, axial, electrostatic field. Journal of Mathematical Physics, 13(9):1297–1304, September 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i9/p1297_
s1.

Kalnins:1972:URH

[995] E. G. Kalnins. Unitary representations of the homogeneous Lorentz group in an O(1, 1) ⊗ O(2) basis and some applications to relativistic equations. Journal of Mathematical Physics, 13(9):1304–1312, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i9/p1304_s1.

Lee:1972:DHB

[996] K. H. Lee. Dynamics of harmonically bound semi-infinite and infinite chains with friction and applied forces. Journal of Mathematical Physics, 13(9):1312–1315, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i9/p1312_s1.

Gates:1972:LWTb

[997] D. J. Gates. Lattice wind-tree models. II. Analytic property. Journal of Mathematical Physics, 13(9):1315–1317, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i9/p1315_s1.

McQuistan:1972:ENN

[998] R. B. McQuistan. Exact nearest neighbor statistics for one-dimensional lattice spaces. Journal of Mathematical Physics, 13(9):1317–1319, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i9/p1317_s1.

Banks:1972:AOP

[999] Thomas I. Banks and Carl M. Bender. Anharmonic oscillator with polynomial self-interaction. Journal of Mathematical Physics, 13(9):1320– 1324, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i9/p1320_s1.

Davies:1972:QEZ

[1000] B. Davies. Quantum electromagnetic zero-point energy of a conducting spherical shell. Journal of Mathematical Physics, 13(9):1324–1329, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i9/p1324_s1.

Alisauskas:1972:STO

[1001] S. J. Alisauskas, A.-A. A. Jucys, and A. P. Jucys. On the symmetric tensor operators of the unitary groups. Journal of Mathematical Physics, 13(9):1329–1333, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i9/p1329_s1.

Maugin:1972:DMSb

[1002] Gérard A. Maugin and A. Cemal Eringen. Deformable magnetically saturated media. II. Constitutive theory. Journal of Mathematical Physics, 13(9):1334–1347, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i9/p1334_s1.

Benioff:1972:PQM

[1003] Paul A. Benioff. Procedures in quantum mechanics without Von Neumann’s projection Axiom. Journal of Mathematical Physics, 13(9):1347– 1355, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i9/p1347_s1.

Barnebey:1972:WTR

[1004] T. A. Barnebey. Ward–Takahashi relations in massive Yang–Mills theory. Journal of Mathematical Physics, 13(9):1355–1368, September 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i9/p1355_
s1.

Riewe:1972:QDH

[1005] F. Riewe and A. E. S. Green. Quantum dynamics of higher-derivative fields. Journal of Mathematical Physics, 13(9):1368–1374, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i9/p1368_s1.

Riewe:1972:PIH

[1006] F. Riewe and A. E. S. Green. Physical interpretation of higherderivative field theories. Journal of Mathematical Physics, 13(9):1374– 1380, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i9/p1374_s1.

Li:1972:STC

[1007] Ming Chiang Li. Scattering by two charged centers. Journal of Mathematical Physics, 13(9):1381–1386, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i9/p1381_s1.

Litvin:1972:DDP

[1008] Daniel B. Litvin. On the decomposition of direct products of irreducible representations. Journal of Mathematical Physics, 13(9):1386– 1389, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i9/p1386_s1.

Lezuo:1972:SGG

[1009] K. J. Lezuo. The symmetric group and the Gel’fand basis of U(3). generalizations of the Dirac identity. Journal of Mathematical Physics, 13 (9):1389–1393, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i9/p1389_s1.

Miller:1972:LTL

[1010] Willard Miller, Jr. Lie theory and the Lauricella functions F_{D}. Journal of Mathematical Physics, 13(9):1393–1399, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i9/p1393_s1.

McDonald:1972:ACR

[1011] J. F. McDonald. Averages of the components of random unit vectors. Journal of Mathematical Physics, 13(9):1399–1404, September 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i9/p1399_
s1.

Chow:1972:WPR

[1012] Pao-Liu Chow and Joseph B. Keller. Wave propagation in a random lattice. I. Journal of Mathematical Physics, 13(9):1404–1411, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i9/p1404_s1.

Horiguchi:1972:LGFb

[1013] Tsuyoshi Horiguchi. Lattice Green’s functions for the triangular and honeycomb lattices. Journal of Mathematical Physics, 13(9):1411–1419, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i9/p1411_s1.

Liu:1972:ATS

[1014] Tsu huei Liu. An approximation technique for solving multiparticle and bound-state scattering processes. Journal of Mathematical Physics, 13 (9):1419–1428, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i9/p1419_s1.

Morita:1972:TTSb

[1015] Tohru Morita. Two-time spin-pair correlation function of the Heisenberg magnet at infinite temperature. III. BCC lattice and arbitrary spin. Journal of Mathematical Physics, 13(9):1428–1435, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i9/p1428_s1.

Robinson:1972:EGS

[1016] David C. Robinson and Jeffrey Winicour. Energy of gravitational shock waves. Journal of Mathematical Physics, 13(9):1435–1441, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i9/p1435_s1.

Weston:1972:KTS

[1017] V. H. Weston. Kinetic theory of scattering by a plasma cylinder. Journal of Mathematical Physics, 13(10):1443–1450, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1443_s1.

Ray:1972:LDP

[1018] John R. Ray. Lagrangian density for perfect fluids in general relativity. Journal of Mathematical Physics, 13(10):1451–1453, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/
p1451_s1.

Kubo:1972:KKR

[1019] Ryogo Kubo and Munetake Ichimura. Kramers–Kronig relations and sum rules. Journal of Mathematical Physics, 13(10):1454–1461, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i10/p1454_s1.

Burniston:1972:SSS

[1020] E. E. Burniston, T. W. Mullikin, and C. E. Siewert. Steady-state solutions in the two-group theory of neutron diffusion. Journal of Mathematical Physics, 13(10):1461–1465, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i10/p1461_s1.

Streifer:1972:PMT

[1021] William Streifer and K. Bradley Paxton. A perturbation method for two synchronously tuned, coupled, autonomous, nonlinear oscillators. Journal of Mathematical Physics, 13(10):1466–1469, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1466_s1.

Debney:1972:SEM

[1022] George Debney. Symmetry in Einstein–Maxwell space–time. Journal of Mathematical Physics, 13(10):1469–1477, October 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1469_s1.

Tchrakian:1972:FFS

[1023] D. H. Tchrakian. Form factors for any spin and charge coupling. Journal of Mathematical Physics, 13(10):1477–1484, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1477_s1.

OBrien:1972:NSS

[1024] T. J. O’Brien and T. T. Holloway. Neumann series solution for the atom-rigid rotor collision. Journal of Mathematical Physics, 13(10):1485– 1488, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i10/p1485_s1.

Ranada:1972:GTE

[1025] Antonio F. Rañada. A geometrical theory of the electromagnetic field and the gauge transformation. Journal of Mathematical Physics, 13 (10):1488–1490, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i10/p1488_s1.

Pearson:1972:CEG

[1026] D. B. Pearson. Conditions for the existence of the generalized wave operators. Journal of Mathematical Physics, 13(10):1490–1499, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i10/p1490_s1.

Becchi:1972:CBL

[1027] C. M. Becchi and R. Collina. On the connection between Levinson’s theorem and singular integral equations. Journal of Mathematical Physics, 13(10):1499–1502, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i10/p1499_s1.

Partensky:1972:GEO

[1028] A. Partensky. On the generalized exchange operators for SU(n). Journal of Mathematical Physics, 13(10):1503–1510, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1503_s1.

Neubert:1972:SSS

[1029] Jerome A. Neubert. Steady-state sound propagation in continuous, statistically isotropic media. Journal of Mathematical Physics, 13(10):1510– 1524, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i10/p1510_s1.

Mano:1972:RFV

[1030] Koichi Mano. Recurrence formula for the Veneziano model N-point functions. Journal of Mathematical Physics, 13(10):1525–1526, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/
p1525_s1.

McRae:1972:LID

[1031] Walter B. McRae and Ernest R. Davidson. Linear inequalities for density matrices. II. Journal of Mathematical Physics, 13(10):1527–1538, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i10/p1527_s1.

Misra:1972:SAS

[1032] R. M. Misra and Deo Bhushan Pandey. Stationary axially symmetric fields and the Kerr metric. Journal of Mathematical Physics, 13 (10):1538–1540, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i10/p1538_s1.

Bushell:1972:CBS

[1033] P. J. Bushell. On the convergence of the Born series for all energies. Journal of Mathematical Physics, 13(10):1540–1542, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1540_s1.

Pereira:1972:NSS

[1034] C. Martin Pereira. Nonmetrical specification of space–time sources. Journal of Mathematical Physics, 13(10):1542–1551, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1542_s1.

Stapp:1972:PLC

[1035] Henry P. Stapp. Properties of Lorentz covariant analytic functions. Journal of Mathematical Physics, 13(10):1551–1568, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/p1551_s1.

Glimm:1972:QFT

[1036] James Glimm and Arthur Jaffe. The λφ_{2}^{4} quantum field theory without cutoffs. IV. Perturbations of the Hamiltonian. Journal of Mathematical Physics, 13(10):1568–1584, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i10/p1568_s1.

Casati:1972:SRT

[1037] G. Casati and E. Lazzaro. Some remarks on three-waves interaction. Journal of Mathematical Physics, 13(10):1584–1585, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/
p1584_s1.

MacDowell:1972:RPG

[1038] W. W. MacDowell and Ralph Roskies. Reduction of the Poincaré group with respect to the Lorentz group. Journal of Mathematical Physics, 13(10):1585–1592, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i10/p1585_s1. See errata [1412].

Beers:1972:ASD

[1039] Brian Beers and H. H. Nickle. Algebraic solution for a Dirac electron in a plane-wave electromagnetic field. Journal of Mathematical Physics, 13(10):1592–1595, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i10/p1592_s1.

McGuire:1972:STI

[1040] J. B. McGuire and C. A. Hurst. The scattering of three impenetrable particles in one dimension. Journal of Mathematical Physics, 13(10):1595– 1607, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i10/p1595_s1.

Erdahl:1972:CSS

[1041] R. M. Erdahl. The convex structure of the set of N-representable reduced 2- matrices. Journal of Mathematical Physics, 13(10):1608–1621, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v13/i10/p1608_s1. See comment [1786].

Schwarz:1972:AEF

[1042] Fritz Schwarz. Asymptotic expansions of Fourier integrals with light cone singularities. Journal of Mathematical Physics, 13(10):1621–1634, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i10/p1621_s1.

Wolf:1972:RME

[1043] Kurt Bernardo Wolf. The U_{n,1} and IU_{n} representation matrix elements. Journal of Mathematical Physics, 13(10):1634–1638, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i10/p1634_s1.

Lorente:1972:CSS

[1044] M. Lorente and B. Gruber. Classification of semisimple subalgebras of simple Lie algebras. Journal of Mathematical Physics, 13(10):1639– 1663, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i10/p1639_s1.

Borie:1972:EMH

[1045] Edith Borie. Errata: “S Matrix in the Heisenberg Representation”. Journal of Mathematical Physics, 13(10):1664, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i10/

p1664_s1. See [429].

Foulis:1972:OSB

[1046] D. J. Foulis and C. H. Randall. Operational statistics. I. Basic concepts. Journal of Mathematical Physics, 13(11):1667–1675, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i11/p1667_s1.

Jenc:1972:STA

[1047] F. Jenc. Some theorems on atomicity in axiomatic quantum mechanics. Journal of Mathematical Physics, 13(11):1675–1681, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i11/
p1675_s1.

Siegel:1972:AFI

[1048] Armand Siegel and Terence Burke. Approximate functional integral methods in statistical mechanics. I. Moment expansions. Journal of Mathematical Physics, 13(11):1681–1694, November 1972. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i11/

p1681_s1. See erratum [1413].

Kota:1972:ASV

[1049] J. Kóta and Z. Perjés. All stationary vacuum metrics with shearing geodesic eigenrays. Journal of Mathematical Physics, 13(11):1695–1699, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1695_s1. See erratum [1785].

Halford:1972:STT

[1050] W. D. Halford. Scalar-tensor theory of gravitation in a Lyra manifold. Journal of Mathematical Physics, 13(11):1699–1703, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i11/p1699_s1.

Treat:1972:GID

[1051] Richard P. Treat. Gauge invariant decomposition of Yang–Mills potentials. Journal of Mathematical Physics, 13(11):1704–1714, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v13/i11/p1704_s1. See erratum [1310].

Brandt:1972:NPR

[1052] Richard A. Brandt and Paul Otterson. Null plane restrictions of current commutators. Journal of Mathematical Physics, 13(11):1714–1722, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1714_s1.

Cheng:1972:QGD

[1053] K. S. Cheng. Quantization of a general dynamical system by Feynman’s path integration formulation. Journal of Mathematical Physics, 13(11):1723–1726, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i11/p1723_s1.

Fox:1972:TMS

[1054] Ronald Forrest Fox. An “H-theorem” for multiplicative stochastic processes. Journal of Mathematical Physics, 13(11):1726–1729, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i11/p1726_s1.

Kyriakopoulos:1972:AEB

[1055] E. Kyriakopoulos. Algebraic equations for Bethe–Salpeter and Coulomb Green’s functions. Journal of Mathematical Physics, 13(11):1729–1735, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1729_s1.

Garrison:1972:ALR

[1056] John C. Garrison, Jack Wong, and Harry L. Morrison. Absence of longrange order in thin films. Journal of Mathematical Physics, 13(11):1735– 1742, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1735_s1. See erratum [1226].

Lee:1972:EFT

[1057] K. H. Lee. Exponential Fourier transforms for coupled harmonic oscillator chains. Journal of Mathematical Physics, 13(11):1742–1745, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i11/p1742_s1.

desCloizeaux:1972:SAE

[1058] J. des Cloizeaux and M. L. Mehta. Some asymptotic expressions for prolate spheroidal functions and for the eigenvalues of differential and integral equations of which they are solutions. Journal of Mathematical Physics, 13(11):1745–1754, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i11/p1745_s1.

Sijacki:1972:GSP

[1059] Dj. Sijacki, M. Vujicić, and F. Herbut. Generalized semidirect product in group extensions. Journal of Mathematical Physics, 13(11):1755– 1759, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1755_s1.

Agarwal:1972:RBS

[1060] G. S. Agarwal and E. Wolf. Relation between the statistical representations of real and associated complex fields in optical coherence theory. Journal of Mathematical Physics, 13(11):1759–1764, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i11/
p1759_s1.

Tortorella:1972:EUQ

[1061] Michael Tortorella. Existence and uniqueness questions for the unitarity equation. Journal of Mathematical Physics, 13(11):1764–1767, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i11/p1764_s1.

Raghavacharyulu:1972:NNR

[1062] I. V. V. Raghavacharyulu. Note on nonlinear representations. Journal of Mathematical Physics, 13(11):1768, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i11/p1768_s1.

Hsieh:1972:SRW

[1063] D. Y. Hsieh. A study of relaxing waves. Journal of Mathematical Physics, 13(11):1769–1776, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i11/p1769_s1.

Maugin:1972:PEM

[1064] Gérard A. Maugin and A. Cemal Eringen. Polarized elastic materials with electronic spin — a relativistic approach. Journal of Mathematical Physics, 13(11):1777–1788, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i11/p1777_s1.

Maugin:1972:RCD

[1065] Gérard A. Maugin and A. Cemal Eringen. Relativistic continua with directors. Journal of Mathematical Physics, 13(11):1788–1797, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i11/p1788_s1.

Gerjuoy:1972:IRV

[1066] Edward Gerjuoy, A. R. P. Rau, and Larry Spruch. Identities related to variational principles. Journal of Mathematical Physics, 13(11):1797– 1804, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1797_s1.

McCoy:1972:CFQ

[1067] John J. McCoy. On the correlation of field quantities in a random linearly elastic solid. Journal of Mathematical Physics, 13(11):1804–1814, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1804_s1.

Merlini:1972:SLS

[1068] D. Merlini and C. Gruber. Spin-1/2 lattice system: group structure and duality relation. Journal of Mathematical Physics, 13(11):1814–1823, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1814_s1.

Santhanam:1972:RCB

[1069] T. S. Santhanam. Remarkable connection between the multi-Veneziano integrand and the character of special unitary groups. Journal of Mathematical Physics, 13(11):1823–1825, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i11/p1823_s1.

Piepenbrink:1972:NYH

[1070] John Piepenbrink. Nonnegativity of the Yukawa Hamiltonian. Journal of Mathematical Physics, 13(11):1825–1827, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i11/p1825_s1.

Liboff:1972:REC

[1071] Richard L. Liboff. Relativistic excitation of cavity modes. Journal of Mathematical Physics, 13(11):1828–1837, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i11/p1828_s1.

McCarthy:1972:SBM

[1072] Patrick J. McCarthy. Structure of the Bondi–Metzner–Sachs group. Journal of Mathematical Physics, 13(11):1837–1842, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i11/
p1837_s1.

Chen:1972:EEI

[1073] H. H. Chen and R. I. Joseph. Erratum: “Exchange Interaction Model of Ferromagnetism”. Journal of Mathematical Physics, 13(11):1842, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i11/p1842_s1. See [906].

Aronson:1972:CSA

[1074] B. Aronson and E. T. Newman. Coordinate systems associated with asymptotically shear-free null congruences. Journal of Mathematical Physics, 13(12):1847–1851, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i12/p1847_s1.

Roman:1972:RDS

[1075] P. Roman, J. J. Aghassi, and P. L. Huddleston. Relativistic dynamical symmetries and dilatations. Journal of Mathematical Physics, 13 (12):1852–1861, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i12/p1852_s1.

Baker:1972:CBF

[1076] George A. Baker, Jr. Converging bounds for the free energy in certain statistical mechanical problems. Journal of Mathematical Physics, 13 (12):1862–1864, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i12/p1862_s1.

Townsend:1972:SRC

[1077] J. Townsend, G. Feldman, and T. Fulton. A spectral representation for Coulomb wavefunctions. Journal of Mathematical Physics, 13(12):1865– 1868, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i12/p1865_s1.

Trubatch:1972:EBS

[1078] S. L. Trubatch. Erroneous bound state conditions from an algebraic misrepresentation of spin wave theory. Journal of Mathematical Physics, 13(12):1869, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i12/p1869_s1.

Kelemen:1972:NMT

[1079] Peter J. Kelemen and Abraham Robinson. The nonstandard λ: ϕ_{2}^{4}(x): model. I. The technique of nonstandard analysis in theoretical physics. Journal of Mathematical Physics, 13(12):1870–1874, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i12/
p1870_s1.

Kelemen:1972:NMI

[1080] Peter J. Kelemen and Abraham Robinson. The nonstandard λ: ϕ_{2}^{4}(x): model. II. The standard model from a nonstandard point of view. Journal of Mathematical Physics, 13(12):1875–1878, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i12/p1875_s1.

Lind:1972:AIO

[1081] Robert W. Lind, James Messmer, and Ezra T. Newman. Applications of the intertwining operators for representations of the restricted Lorentz group. Journal of Mathematical Physics, 13(12):1879–1883, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i12/p1879_s1.

Lind:1972:EMS

[1082] Robert W. Lind, James Messmer, and Ezra T. Newman. Equations of motion for the sources of asymptotically flat spaces. Journal of Mathematical Physics, 13(12):1884–1891, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i12/p1884_s1.

Buch:1972:ITO

[1083] T. Buch. Irreducible tensor operators for finite groups. Journal of Mathematical Physics, 13(12):1892–1895, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i12/p1892_s1.

Bart:1972:SNI

[1084] G. R. Bart and R. L. Warnock. Solutions of a nonlinear integral equation for high energy scattering. III. Analyticity of solutions in a parameter, explored numerically. Journal of Mathematical Physics, 13(12):1896– 1902, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i12/p1896_s1.

Zipfel:1972:SSW

[1085] G. G. Zipfel, Jr. and John A. DeSanto. Scattering of a scalar wave from a random rough surface: A diagrammatic approach. Journal of Mathematical Physics, 13(12):1903–1911, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v13/i12/p1903_s1.

Papanicolaou:1972:KTP

[1086] George C. Papanicolaou. A kinetic theory for power transfer in stochastic systems. Journal of Mathematical Physics, 13(12):1912–1918, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i12/p1912_s1.

Rocher:1972:NEE

[1087] E. Y. Rocher. Noumenon: Elementary entity of a new mechanics. Journal of Mathematical Physics, 13(12):1919–1925, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i12/p1919_s1.

Kent:1972:DIA

[1088] W. H. Kent and S. W. Lee. Diffraction by an infinite array of parallel strips. Journal of Mathematical Physics, 13(12):1926–1930, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i12/p1926_s1.

Hooshyar:1972:ISP

[1089] M. A. Hooshyar. The inverse scattering problem at fixed energy for L^{2}-dependent potentials. Journal of Mathematical Physics, 13(12):1931– 1933, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i12/p1931_s1.

Glass:1972:EGR

[1090] Edward N. Glass and Jeffrey Winicour. Elastic general relativistic systems. Journal of Mathematical Physics, 13(12):1934–1940, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v13/i12/p1934_s1. See errata [1535].

Dwivedi:1972:POF

[1091] I. Dwivedi and R. Kantowski. On the possibility of observing firstorder corrections to geometrical optics in a curved space–time. Journal of Mathematical Physics, 13(12):1941–1943, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v13/i12/p1941_s1.

Stephen:1972:NRS

[1092] M. J. Stephen and L. Mittag. A new representation of the solution of the Ising model. Journal of Mathematical Physics, 13(12):1944–1951, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i12/p1944_s1.

Weston:1972:IPH

[1093] V. H. Weston. On the inverse problem for a hyperbolic dispersive partial differential equation. Journal of Mathematical Physics, 13(12):1952– 1956, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i12/p1952_s1. See erratum [1311].

Biedenharn:1972:SCTa

[1094] L. C. Biedenharn, J. D. Louck, E. Chacón, and M. Ciftan. On the structure of the canonical tensor operators in the unitary groups. I. An extension of the pattern calculus rules and the canonical splitting in U(3). Journal of Mathematical Physics, 13(12):1957–1984, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v13/i12/p1957_s1.

Biedenharn:1972:SCTb

[1095] L. C. Biedenharn and J. D. Louck. On the structure of the canonical tensor operators in the unitary groups. II. The tensor operators in U(3) characterized by maximal null space. Journal of Mathematical Physics, 13(12):1985–2001, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i12/p1985_s1.

Gille:1972:GTS

[1096] J. F. Gille and J. Manuceau. Gauge transformations of second type and their implementation. I. Fermions. Journal of Mathematical Physics, 13 (12):2002–2007, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i12/p2002_s1. See errata [1281].

Graffi:1972:DPE

[1097] S. Graffi and V. Grecchi. On the divergent perturbation expansion for the vacuum polarization by an external field. Journal of Mathematical Physics, 13(12):2008–2012, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i12/p2008_s1.

Yener:1972:RTA

[1098] Y. Yener and M. N. Özisik. Radiative transfer in adjacent half-spaces with specular reflection. Journal of Mathematical Physics, 13(12):2013– 2019, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i12/p2013_s1.

Campbell:1973:MPG

[1099] Gerald A. Campbell and Richard A. Matzner. A model for peaking of galactic gravitational radiation. Journal of Mathematical Physics, 14(1): 1–6, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i1/p1_s1.

Bardeen:1973:RFS

[1100] James M. Bardeen and William H. Press. Radiation fields in the Schwarzschild background. Journal of Mathematical Physics, 14(1):7– 19, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i1/p7_s1.

Fox:1973:PAM

[1101] Ronald Forrest Fox. Physical applications of multiplicative stochastic processes. Journal of Mathematical Physics, 14(1):20–25, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i1/p20_s1.

Andersson:1973:LCF

[1102] Bo Andersson. Locality conditions on form factors. Journal of Mathematical Physics, 14(1):26–41, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i1/p26_s1.

Love:1973:DSB

[1103] J. D. Love. Differential solutions of the biharmonic Poisson and first order Stokes equations. Journal of Mathematical Physics, 14(1):42– 44, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i1/p42_s1.

Mallett:1973:POS

[1104] R. L. Mallett and G. N. Fleming. Position operators in a (3+1) de Sitter space. Journal of Mathematical Physics, 14(1):45–51, January 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i1/p45_s1.

Schiffer:1973:KGC

[1105] Menahem M. Schiffer, Ronald J. Adler, James Mark, and Charles Sheffield. Kerr geometry as complexified Schwarzschild geometry. Journal of Mathematical Physics, 14(1):52–56, January 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i1/p52_s1.

Macfadyen:1973:CGPa

[1106] N. W. Macfadyen. Conformal group in a Poincaré basis. II. Principal discrete series. Journal of Mathematical Physics, 14(1):57–64, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i1/p57_s1.

Lindell:1973:CID

[1107] Ismo V. Lindell. Coordinate independent dyadic formulation of wave normal and ray surfaces of general anisotropic media. Journal of Mathematical Physics, 14(1):65–67, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i1/p65_s1.

Taylor:1973:PTN

[1108] J. G. Taylor. Perturbation theory for nonderivative nonpolynomial Lagrangians. Journal of Mathematical Physics, 14(1):68–78, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i1/
p68_s1.

Bogolubov:1973:MCQ

[1109] N. N. Bogolubov, Jr. Method of calculating quasiaverages. Journal of Mathematical Physics, 14(1):79–83, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i1/p79_s1.

Papanicolaou:1973:SGB

[1110] G. C. Papanicolaou, D. McLaughlin, and R. Burridge. A stochastic Gaussian beam. Journal of Mathematical Physics, 14(1):84–89, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v14/i1/p84_s1. See erratum [1282].

Choudhury:1973:SBL

[1111] A. L. Choudhury. U and V sectors in the Bronzan–Lee model. Journal of Mathematical Physics, 14(1):90–96, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i1/p90_s1.

Tannenwald:1973:NTE

[1112] L. M. Tannenwald. Neutron transport equations with spin-orbit coupling. Journal of Mathematical Physics, 14(1):97–101, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i1/p97_s1. See

erratum [1283].

Newman:1973:MEC

[1113] Ezra T. Newman. Maxwell’s equations and complex Minkowski space. Journal of Mathematical Physics, 14(1):102–103, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i1/p102_s1.

Ellis:1973:EFT

[1114] Homer G. Ellis. Ether flow through a drainhole: a particle model in general relativity. Journal of Mathematical Physics, 14(1):104–118, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v14/i1/p104_s1. See errata [1507].

Merchant:1973:SPL

[1115] Allan H. Merchant and Richard L. Liboff. Spectral properties of the linearized Balescu–Lenard operator. Journal of Mathematical Physics, 14(1):119–129, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i1/p119_s1.

Schroeck:1973:MFT

[1116] Franklin E. Schroeck, Jr. A model of field theory treated in the Fock– Cook formalism. Journal of Mathematical Physics, 14(1):130–138, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i1/p130_s1.

Gutzwiller:1973:AKP

[1117] Martin C. Gutzwiller. The anisotropic Kepler problem in two dimensions. Journal of Mathematical Physics, 14(1):139–152, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i1/p139_s1.

Goovaerts:1973:EAT

[1118] M. J. Goovaerts and J. T. Devreese. Erratum: “Analytic treatment of the Coulomb potential in the path integral formalism by exact summation of a perturbation expansion”. Journal of Mathematical Physics, 14(1): 153, January 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i1/p153_s1. See [961].

Benard:1973:DEP

[1119] Christine Benard and Odile Macchi. Detection and “emission” processes of quantum particles in a “chaotic state”. Journal of Mathematical Physics, 14(2):155–167, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i2/p155_s1.

Zund:1973:SAG

[1120] Joseph Zund and William Maher. A spinor approach to the generalized singular electromagnetic field theory of Bel, Lapiedra, and Montserrat. Journal of Mathematical Physics, 14(2):168–172, February 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i2/p168_
s1.

deAzcarraga:1973:NVM

[1121] J. A. de Azcárraga. A note on vector-meson dominance and the compactness of current algebra. Journal of Mathematical Physics, 14(2):173– 175, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i2/p173_s1.

Ashmore:1973:MS

[1122] J. Ashmore and R. Delbourgo. On matrix superpropagators. I. Journal of Mathematical Physics, 14(2):176–181, February 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i2/p176_s1.

Calogero:1973:EBS

[1123] F. Calogero and C. Marchioro. Exact bound states of some N-body systems with two- and three-body forces. Journal of Mathematical Physics, 14(2):182–184, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i2/p182_s1.

Killingbeck:1973:CSO

[1124] J. Killingbeck. The class sum operator approach for the point group O and D_{4}. Journal of Mathematical Physics, 14(2):185–187, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i2/p185_s1.

Chakraborty:1973:MPR

[1125] B. Chakraborty. The mathematical problem of reflection solved by an extension of the WKB method. Journal of Mathematical Physics, 14(2):188–190, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i2/p188_s1.

Holzwarth:1973:MFS

[1126] N. A. W. Holzwarth. Mathieu function solutions to the radial Schrödinger equation for the −f^{2}/r^{4} interaction. Journal of Mathematical Physics, 14(2):191–204, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i2/p191_s1.

Zohni:1973:ESM

[1127] O. Zohni. Off-energy-shell t matrix for local potentials with nonlocal square well core interaction. Journal of Mathematical Physics, 14(2):205– 208, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i2/p205_s1. See erratum [1341].

Eardley:1973:STF

[1128] D. Eardley and R. K. Sachs. Space–times with a future projective infinity. Journal of Mathematical Physics, 14(2):209–212, February 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i2/p209_
s1.

Thomas:1973:CFG

[1129] Brian K. Thomas and E. Gerjuoy. Closed-form Glauber cross sections in p–H and e^{−}–H collisions. I. Journal of Mathematical Physics, 14(2):213– 218, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i2/p213_s1.

Halpern:1973:QMA

[1130] Francis R. Halpern. Quantum mechanics of the anharmonic oscillator. Journal of Mathematical Physics, 14(2):219–227, February 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i2/p219_
s1.

Marathe:1973:SAG

[1131] Kishore B. Marathe. Spaces admitting gravitational fields. Journal of Mathematical Physics, 14(2):228–233, February 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i2/p228_s1.

Gudder:1973:MTC

[1132] Stanley P. Gudder and Harry C. Mullikin. Measure theoretic convergences of observables and operators. Journal of Mathematical Physics, 14(2):234–242, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i2/p234_s1.

Parisi:1973:ADO

[1133] G. Parisi and F. Zirilli. Anomalous dimensions in one-dimensional quantum field theory. Journal of Mathematical Physics, 14(2):243–245, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i2/p243_s1.

Fleischer:1973:NPA

[1134] J. Fleischer. Nonlinear Padé approximants for Legendre series. Journal of Mathematical Physics, 14(2):246–248, February 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i2/p246_s1.

Deliyannis:1973:VSM

[1135] Platon C. Deliyannis. Vector space models of abstract quantum logics. Journal of Mathematical Physics, 14(2):249–253, February 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i2/p249_
s1.

Mukunda:1973:NFR

[1136] N. Mukunda and B. Radhakrishnan. New forms for the representations of the three-dimensional Lorentz group. Journal of Mathematical Physics, 14(2):254–258, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i2/p254_s1.

Weger:1973:DSI

[1137] M. Weger, S. Alexander, and G. Della Riccia. Description of a system of interacting linear chains by means of the electron network model. Journal of Mathematical Physics, 14(2):259–262, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i2/p259_s1.

Huang:1973:SCG

[1138] Cheng-Sheng Huang and Alfred C. T. Wu. Structure of the combinatorial generalization of hypergeometric functions for SU(n) states. II. Journal of Mathematical Physics, 14(2):263–266, February 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i2/p263_s1.

Callegari:1973:NSP

[1139] Andrew J. Callegari and Edward L. Reiss. Nonlinear stability problems for the sine-Gordon equation. Journal of Mathematical Physics, 14(2):267–276, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i2/p267_s1.

Gilmore:1973:CSO

[1140] Robert Gilmore. Coset symmetrization operators. Journal of Mathematical Physics, 14(2):277–278, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i2/p277_s1.

Patera:1973:NMO

[1141] J. Patera. The Nagel–Moshinsky operators for U(p, 1) ⊃ U(p). Journal of Mathematical Physics, 14(2):279–284, February 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i2/p279_s1.

Couch:1973:ASP

[1142] W. E. Couch and E. T. Newman. Algebraically special perturbations of the Schwarzschild metric. Journal of Mathematical Physics, 14(2):285– 286, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i2/p285_s1.

Chen:1973:AEF

[1143] T. W. Chen. Asymptotic estimation of Fourier transforms and light cone dominance. Journal of Mathematical Physics, 14(2):287–291, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i2/p287_s1.

Voros:1973:INA

[1144] A. Voros. Introduction to nonstandard analysis. Journal of Mathematical Physics, 14(2):292–296, February 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v14/i2/p292_s1.

Barnsley:1973:BPM

[1145] Michael Barnsley. The bounding properties of the multipoint Padé approximant to a series of Stieltjes on the real line. Journal of Mathematical Physics, 14(3):299–313, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i3/p299_s1.

Epstein:1973:VAT

[1146] Saul T. Epstein and Michael F. Barnsley. A variational approach to the theory of multipoint Padé approximants. Journal of Mathematical Physics, 14(3):314–325, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i3/p314_s1.

Denman:1973:SHJ

[1147] Harry H. Denman and Lawrence H. Buch. Solution of the Hamilton– Jacobi equation for certain dissipative classical mechanical systems. Journal of Mathematical Physics, 14(3):326–329, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/p326_s1.

Holman:1973:ARS

[1148] Wayne J. Holman III. The application of Regge symmetry to the 9-j symbol. Journal of Mathematical Physics, 14(3):330–336, March 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/p330_
s1.

Miller:1973:SME

[1149] Marvin Miller and Stanly Steinberg. The solution of the moment equations associated with a partial differential equation with polynomial coefficients. Journal of Mathematical Physics, 14(3):337–339, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/
p337_s1.

LoDato:1973:RPO

[1150] Vincent A. LoDato. The renormalized projection operator technique for linear stochastic differential equations. Journal of Mathematical Physics, 14(3):340–345, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v14/i3/p340_s1.

Boffi:1973:CMS

[1151] V. C. Boffi, F. Premuda, and G. Spiga. Convergence in the mean of solutions to the neutron integral Boltzmann equation in three-dimensional systems. Journal of Mathematical Physics, 14(3):346–352, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/
p346_s1.

Girard:1973:CAF

[1152] André Girard. Current algebras of free systems at finite temperature. Journal of Mathematical Physics, 14(3):353–365, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/p353_s1.

Quesne:1973:BRU

[1153] C. Quesne. Boson realization of the unitary symplectic group. Journal of Mathematical Physics, 14(3):366–372, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i3/p366_s1.

Osborn:1973:CSE

[1154] T. A. Osborn. On the convergence of separable expansions for the t matrix. Journal of Mathematical Physics, 14(3):373–375, March 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/p373_

s1. See erratum [1340].

Lavine:1973:CWO

[1155] Richard Lavine. Completeness of the wave operators in the repulsive Nbody problem. Journal of Mathematical Physics, 14(3):376–379, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i3/p376_s1.

Cawley:1973:MCL

[1156] Robert G. Cawley. Motion of a charged lightlike particle in an external field. Journal of Mathematical Physics, 14(3):380–386, March 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/p380_
s1.

Sullivan:1973:TRT

[1157] John J. Sullivan. Nth rank tensor representations of U(n) symmetry adapted to subgroups of the symmetric group S_{N}. Journal of Mathematical Physics, 14(3):387–395, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i3/p387_s1.

Barucchi:1973:DEO

[1158] G. Barucchi and G. Ponzano. Differential equations for one-loop generalized Feynman integrals. Journal of Mathematical Physics, 14(3):396– 401, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i3/p396_s1.

Bassanini:1973:SS

[1159] Piero Bassanini. On stochastic scattering. Journal of Mathematical Physics, 14(3):402–405, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v14/i3/p402_s1.

Weston:1973:IPH

[1160] V. H. Weston and R. J. Krueger. On the inverse problem for a hyperbolic dispersive partial differential equation. II. Journal of Mathematical Physics, 14(3):406–409, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i3/p406_s1.

Glasser:1973:ELSa

[1161] M. L. Glasser. The evaluation of lattice sums. I. Analytic procedures. Journal of Mathematical Physics, 14(3):409–413, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/p409_s1. See

comments [1466].

Davidson:1973:RQSa

[1162] Russell Davidson and John J. Kozak. On the relaxation to quantumstatistical equilibrium of the Wigner–Weisskopf atom in a onedimensional radiation field. IV. Exact solution for finite systems. Journal of Mathematical Physics, 14(3):414–421, March 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i3/p414_s1.

Davidson:1973:RQSb

[1163] Russell Davidson and John J. Kozak. On the relaxation to quantumstatistical equilibrium of the Wigner–Weisskopf atom in a onedimensional radiation-field. V. Exact solution for infinite systems. Journal of Mathematical Physics, 14(4):423–431, April 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i4/p423_s1.

Eaves:1973:PVG

[1164] R. E. Eaves. On the phase velocity and group velocity of guided waves. Journal of Mathematical Physics, 14(4):432–433, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i4/p432_s1.

Mayes:1973:HOF

[1165] I. W. Mayes and J. S. Dowker. Hamiltonian orderings and functional integrals. Journal of Mathematical Physics, 14(4):434–439, April 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i4/p434_
s1.

Holman:1973:WC

[1166] Wayne J. Holman III. On the Wigner coefficients of Sp(4) and SO(5). Journal of Mathematical Physics, 14(4):440–447, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i4/p440_s1.

Holman:1973:RT

[1167] Wayne J. Holman III. The representation theory of SU(4) and Sp(4). Journal of Mathematical Physics, 14(4):448–455, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i4/p448_s1.

York:1973:CIO

[1168] James W. York, Jr. Conformally invariant orthogonal decomposition of symmetric tensors on Riemannian manifolds and the initial-value problem of general relativity. Journal of Mathematical Physics, 14(4):456– 464, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i4/p456_s1.

Adler:1973:CST

[1169] Ronald J. Adler and Charles Sheffield. Classification of space–times in general relativity. Journal of Mathematical Physics, 14(4):465–469, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i4/p465_s1.

Chern:1973:ENS

[1170] De-Ching Chern and Peter Havas. Exact nonplanar solutions of the classical relativistic three-body problem. Journal of Mathematical Physics, 14(4):470–478, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i4/p470_s1.

Mo:1973:GSM

[1171] Tse Chin Mo, C. H. Papas, and Carl E. Baum. General scaling method for electromagnetic fields with application to a matching problem. Journal of Mathematical Physics, 14(4):479–483, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i4/p479_s1.

Miller:1973:ECL

[1172] J. G. Miller and M. D. Kruskal. Extension of a compact Lorentz manifold. Journal of Mathematical Physics, 14(4):484–485, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i4/p484_s1.

Miller:1973:GAK

[1173] J. G. Miller. Global analysis of the Kerr–Taub NUT metric. Journal of Mathematical Physics, 14(4):486–494, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i4/p486_s1.

Woodhouse:1973:DCS

[1174] N. M. J. Woodhouse. The differentiable and causal structures of space– time. Journal of Mathematical Physics, 14(4):495–501, April 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i4/p495_
s1.

Newman:1973:GEG

[1175] C. E. Newman. A generalization of the equations governing the evolution of a particle distribution in a random force field. Journal of Mathematical Physics, 14(4):502–508, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i4/p502_s1.

Korringa:1973:TEC

[1176] J. Korringa. Theory of elastic constants of heterogeneous media. Journal of Mathematical Physics, 14(4):509–513, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i4/p509_s1.

Hori:1973:STEa

[1177] Motoo Hori. Statistical theory of effective electrical, thermal, and magnetic properties of random heterogeneous materials. I. Perturbation expansions for the effective permittivity of cell materials. Journal of Mathematical Physics, 14(4):514–523, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i4/p514_s1.

Galas:1973:CEM

[1178] D. J. Galas, C. A. Jensen, and H. L. Sahlin. A computable expansion for multiparticle propagators. Journal of Mathematical Physics, 14(4): 524–530, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i4/p524_s1.

Thomson:1973:GFM

[1179] J. J. Thomson and Gregory Benford. Green’s function for Markovian systems (application to plasma turbulence). Journal of Mathematical Physics, 14(4):531–536, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i4/p531_s1.

Nauenberg:1973:HDE

[1180] Michael Nauenberg. High density expansions in the Thomas–Fermi approximation. Journal of Mathematical Physics, 14(4):537–539, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i4/p537_s1.

Butler:1973:LL

[1181] P. H. Butler. On Litvin’s lemma. Journal of Mathematical Physics, 14 (4):540, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i4/p540_s1.

Dobson:1973:SZC

[1182] Peter N. Dobson, Jr. Spin zero couplings in the Veneziano model. Journal of Mathematical Physics, 14(4):541–547, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i4/p541_s1.

Boyer:1973:EDM

[1183] Charles P. Boyer and Farhad Ardalan. Erratum: “On the decomposition SO(p, 1) ⊃ SO(p − 1, 1) for most degenerate representations”. Journal of Mathematical Physics, 14(4):548, April 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i4/p548_s1. See [724].

Band:1973:RAE

[1184] William Band and James L. Park. Relations among elements of the density matrix. I. Definiteness inequalities. Journal of Mathematical Physics, 14(5):551–553, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i5/p551_s1.

Goovaerts:1973:NEM

[1185] M. J. Goovaerts, A. Babcenco, and J. T. Devreese. A new expansion method in the Feynman path integral formalism: application to a one-dimensional delta-function potential. Journal of Mathematical Physics, 14(5):554–559, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i5/p554_s1.

Inawashiro:1973:LGF

[1186] Sakari Inawashiro, Shigetoshi Katsura, and Yoshihiko Abe. Lattice Green’s function for the simple cubic lattice in terms of a Mellin–Barnes type integral. II. Journal of Mathematical Physics, 14(5):560–562, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p560_s1.

Hormozdiari:1973:GFB

[1187] Iraj Hormozdiari, Justin C. Huang, and Brian DeFacio. On the Green’s functions for the Bethe–Salpeter equation. II. Journal of Mathematical Physics, 14(5):563–568, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i5/p563_s1.

Ashmore:1973:MSI

[1188] J. Ashmore and R. Delbourgo. On matrix superpropagators. II. Journal of Mathematical Physics, 14(5):569–571, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i5/p569_s1.

Enosh:1973:GFT

[1189] M. Enosh and A. Kovetz. Generalized Fermi transport in generalized gravitation theories. Journal of Mathematical Physics, 14(5):572–575, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p572_s1.

Malkin:1973:LAI

[1190] I. A. Malkin, V. I. Man’ko, and D. A. Trifonov. Linear adiabatic invariants and coherent states. Journal of Mathematical Physics, 14(5): 576–582, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i5/p576_s1.

Kac:1973:FPS

[1191] M. Kac and J. M. Luttinger. A formula for the pressure in statistical mechanics. Journal of Mathematical Physics, 14(5):583–585, May 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i5/p583_
s1.

Luttinger:1973:GIIa

[1192] J. M. Luttinger. Generalized isoperimetric inequalities. Journal of Mathematical Physics, 14(5):586–593, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i5/p586_s1.

Case:1973:DVI

[1193] K. M. Case and M. Kac. A discrete version of the inverse scattering problem. Journal of Mathematical Physics, 14(5):594–603, May 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i5/p594_
s1.

Popp:1973:WQ

[1194] Fritz A. Popp. Weak quantization. Journal of Mathematical Physics, 14(5):604–608, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i5/p604_s1.

Boyer:1973:SS

[1195] Charles P. Boyer. On the supplementary series of SO_{0}(p, 1). Journal of Mathematical Physics, 14(5):609–617, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i5/p609_s1.

Tanabe:1973:EDL

[1196] Kosai Tanabe. The explicit determination of the linear boson transformation coefficients. Journal of Mathematical Physics, 14(5):618–622, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p618_s1.

Shtokhamer:1973:GFC

[1197] R. Shtokhamer. Generating functional for covariant time-ordered products of currents. Journal of Mathematical Physics, 14(5):623–629, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p623_s1.

Rashid:1973:IST

[1198] M. A. Rashid and Saifuddin. Identity satisfied by the d-type coefficients of SU(n). Journal of Mathematical Physics, 14(5):630–631, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p630_s1.

Cheng:1973:RBB

[1199] Ing-Yih S. Cheng and John J. Kozak. On the relationship between bifurcation points of the Kirkwood–Salsburg equation and phase transitions. Journal of Mathematical Physics, 14(5):632–637, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i5/p632_s1.

Macfadyen:1973:CGPb

[1200] N. W. Macfadyen. Conformal group in a Poincaré basis. III. Degenerate series. Journal of Mathematical Physics, 14(5):638–646, May 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i5/p638_
s1.

Casey:1973:HFO

[1201] K. F. Casey and L. Weaver. Hill’s functions and the one-dimensional Schrödinger equation. Journal of Mathematical Physics, 14(5):647–650, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p647_s1.

Kinnersley:1973:GSE

[1202] William Kinnersley. Generation of stationary Einstein–Maxwell fields. Journal of Mathematical Physics, 14(5):651–653, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i5/p651_s1.

Kalnins:1973:MBM

[1203] E. G. Kalnins. Mixed basis matrix elements for the subgroup reductions of SO(2, 1). Journal of Mathematical Physics, 14(5):654–657, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p654_s1.

Horwitz:1973:LCE

[1204] Gerald Horwitz. A linked cluster evaluation of contour integrals in statistical mechanics. Journal of Mathematical Physics, 14(5):658–665, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p658_s1.

Liu:1973:UPS

[1205] Tsu huei Liu and Robert L. Zimmerman. Unstable particle scattering in the LSZ formalism. Journal of Mathematical Physics, 14(5):666–675, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i5/p666_s1.

Shanmugadhasan:1973:CFD

[1206] S. Shanmugadhasan. Canonical formalism for degenerate Lagrangians. Journal of Mathematical Physics, 14(6):677–687, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i6/p677_s1.

Park:1973:RAE

[1207] James L. Park and William Band. Relations among elements of the density matrix. II. Exotic conservation laws. Journal of Mathematical Physics, 14(6):688–691, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i6/p688_s1.

Louck:1973:CTAa

[1208] J. D. Louck, M. Moshinsky, and K. B. Wolf. Canonical transformations and accidental degeneracy. I. The anisotropic oscillator. Journal of Mathematical Physics, 14(6):692–695, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i6/p692_s1.

Louck:1973:CTAb

[1209] J. D. Louck, M. Moshinsky, and K. B. Wolf. Canonical transformations and accidental degeneracy. II. The isotropic oscillator in a sector. Journal of Mathematical Physics, 14(6):696–700, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i6/p696_s1.

Glasser:1973:ELSb

[1210] M. L. Glasser. The evaluation of lattice sums. II. Number-theoretic approach. Journal of Mathematical Physics, 14(6):701–703, June 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i6/p701_

s1. See erratum [1508].

Besieris:1973:PFM

[1211] Ioannis M. Besieris and Frederick D. Tappert. Propagation of frequencymodulated pulses in a randomly stratified plasma. Journal of Mathematical Physics, 14(6):704–707, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i6/p704_s1.

Dollard:1973:SCI

[1212] John D. Dollard. Scattering into cones. II. n-body problems. Journal of Mathematical Physics, 14(6):708–718, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i6/p708_s1.

Reinhart:1973:MFE

[1213] Bruce L. Reinhart. Maximal foliations of extended Schwarzschild space. Journal of Mathematical Physics, 14(6):719, June 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i6/p719_s1.

Case:1973:ODM

[1214] K. M. Case and C. W. Lau. The one-dimensional X–Y model in inhomogeneous magnetic fields. Journal of Mathematical Physics, 14(6): 720–732, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i6/p720_s1.

Eringen:1973:TNE

[1215] A. Cemal Eringen. Theory of nonlocal electromagnetic elastic solids. Journal of Mathematical Physics, 14(6):733–740, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i6/p733_s1.

Butler:1973:BRE

[1216] P. H. Butler and R. C. King. Branching rules for U(N) ⊃ U(M) and the evaluation of outer plethysms. Journal of Mathematical Physics, 14(6): 741–745, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i6/p741_s1.

Aharonov:1973:OPS

[1217] Y. Aharonov, E. C. Lerner, H. W. Huang, and J. M. Knight. Oscillator phase states, thermal equilibrium and group representations. Journal of Mathematical Physics, 14(6):746–756, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i6/p746_s1.

Tsokos:1973:SPT

[1218] C. P. Tsokos and D. P. Telionis. Stochastic particle trajectories in turbulent flow. Journal of Mathematical Physics, 14(6):757–758, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i6/p757_s1.

Bars:1973:PRO

[1219] Itzhak Bars. Projective representations of SL(3, C) in the Z-operator formalism. Journal of Mathematical Physics, 14(6):759–769, June 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i6/p759_
s1.

Celik:1973:FCD

[1220] Tarik Celik and David R. Palmer. Finite-contour dispersive inequalities. Journal of Mathematical Physics, 14(6):770–773, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i6/p770_s1.

Newman:1973:CCT

[1221] E. T. Newman. Complex coordinate transformations and the Schwarzschild–Kerr metrics. Journal of Mathematical Physics, 14(6): 774–776, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i6/p774_s1.

Loupias:1973:CDI

[1222] G. Loupias and M. Mebkhout. Central decomposition of invariant states: applications to the Galilei group. Journal of Mathematical Physics, 14(6): 777–786, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i6/p777_s1.

Sommers:1973:KTC

[1223] Paul Sommers. On Killing tensors and constants of motion. Journal of Mathematical Physics, 14(6):787–790, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i6/p787_s1.

Wilde:1973:HF

[1224] Ivan F. Wilde. Hypercontractivity for fermions. Journal of Mathematical Physics, 14(6):791–792, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i6/p791_s1.

Demmie:1973:CDT

[1225] Paul N. Demmie and Allen I. Janis. The characteristic development of trapped surfaces. Journal of Mathematical Physics, 14(6):793–802, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i6/p793_s1.

Garrison:1973:EAL

[1226] John C. Garrison. Erratum: “Absence of long-range order in thin films”. Journal of Mathematical Physics, 14(6):803, June 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i6/p803_s1. See

[1056].

Hirota:1973:EES

[1227] Ryogo Hirota. Exact envelope-soliton solutions of a nonlinear wave equation. Journal of Mathematical Physics, 14(7):805–809, July 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i7/p805_
s1.

Hirota:1973:ESS

[1228] Ryogo Hirota. Exact N-soliton solutions of the wave equation of long waves in shallow-water and in nonlinear lattices. Journal of Mathematical Physics, 14(7):810–814, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i7/p810_s1.

Cordero:1973:DCR

[1229] P. Cordero and G.-C. Ghirardi. Dynamics for classical relativistic particles: circular orbit solutions and the nonrelativistic limit. Journal of Mathematical Physics, 14(7):815–822, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p815_s1.

Spencer:1973:PQF

[1230] Thomas Spencer. Perturbation of the P (φ)_{2} quantum field Hamiltonian. Journal of Mathematical Physics, 14(7):823–828, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i7/p823_s1.

Schroeck:1973:NKP

[1231] F. E. Schroeck, Jr. A note on Kato’s perturbation theory. Journal of Mathematical Physics, 14(7):829–830, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p829_s1.

White:1973:PCM

[1232] Paul C. White. C^{∞}-perturbations of a cosmological model. Journal of Mathematical Physics, 14(7):831–836, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p831_s1.

Suzuki:1973:CIP

[1233] Masuo Suzuki. Correlation inequalities and phase transition in the generalized X − Y model. Journal of Mathematical Physics, 14(7):837–838, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i7/p837_s1.

Caron:1973:THS

[1234] Laurent G. Caron. The transformation from a harmonic single-particle basis to the self-consistent harmonic approximation. Journal of Mathematical Physics, 14(7):839–848, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i7/p839_s1.

Buoncristiani:1973:AYM

[1235] A. M. Buoncristiani. An algebra of the Yang–Mills field. Journal of Mathematical Physics, 14(7):849–854, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p849_s1.

Wang:1973:LOT

[1236] Alan P. Wang. Linear operators and transfer of radiation with spherical symmetry. Journal of Mathematical Physics, 14(7):855–862, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i7/p855_s1.

vanTrigt:1973:ASI

[1237] C. van Trigt. Asymptotic solutions of integral equations with a convolution kernel. I. Journal of Mathematical Physics, 14(7):863–873, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i7/p863_s1.

Geroch:1973:STC

[1238] R. Geroch, A. Held, and R. Penrose. A space–time calculus based on pairs of null directions. Journal of Mathematical Physics, 14(7):874–881, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i7/p874_s1.

Maia:1973:I

[1239] M. D. Maia. Isospinors. Journal of Mathematical Physics, 14(7):882–887, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i7/p882_s1.

Kanal:1973:MST

[1240] Madhoo Kanal. Multiple scattering theory of radiative transfer in inhomogeneous atmospheres. Journal of Mathematical Physics, 14(7):888– 892, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i7/p888_s1.

Hestenes:1973:LOD

[1241] David Hestenes. Local observables in the Dirac theory. Journal of Mathematical Physics, 14(7):893–905, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i7/p893_s1.

Agarwal:1973:SPD

[1242] G. S. Agarwal, A. J. Devaney, and D. N. Pattanayak. On the solution of the partial differential equation ∏_{i=1}^{N}(∇^{2} + k_{i}^{2})Ψ = 0. Journal of Mathematical Physics, 14(7):906–908, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p906_s1.

Fabrey:1973:WSM

[1243] James D. Fabrey. Weyl systems for the (φ^{4})_{3} model. Journal of Mathematical Physics, 14(7):909–910, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i7/p909_s1.

Wang:1973:UET

[1244] P. K. C. Wang. Unidirectional energy transfer in nonlinear wave-wave interactions. Journal of Mathematical Physics, 14(7):911–915, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i7/

p911_s1. See erratum and addendum [1537].

Case:1973:DIS

[1245] K. M. Case. On discrete inverse scattering problems. II. Journal of Mathematical Physics, 14(7):916–920, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p916_s1.

Durand:1973:POR

[1246] Bernice Durand. Position operators in relativistic quantum theory: Analysis of algebraic operator relationships. Journal of Mathematical Physics, 14(7):921–933, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i7/p921_s1.

Anderson:1973:ELG

[1247] D. Anderson. Equivalent Lagrangians in generalized mechanics. Journal of Mathematical Physics, 14(7):934–936, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p934_s1.

Au-Yang:1973:TAB

[1248] Helen Au-Yang. Thermodynamics of an anisotropic boundary of a twodimensional Ising model. Journal of Mathematical Physics, 14(7):937– 946, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i7/p937_s1.

Cheng:1973:SFT

[1249] Chiang S. Cheng. Scattering function of a two-component fully ionized gas. Journal of Mathematical Physics, 14(7):947–956, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i7/p947_s1.

Prugovecki:1973:MSS

[1250] E. Prugovecki. Multichannel stationary scattering theory in two-Hilbert space formulation. Journal of Mathematical Physics, 14(7):957–962, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v14/i7/p957_s1. See erratum [1776].

Banos:1973:CRT

[1251] Alfredo Baños, Jr. Calculation of reflection and transmission coefficients for a class of one-dimensional wave propagation problems in inhomogeneous media. Journal of Mathematical Physics, 14(7):963–972, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i7/p963_s1.

Boillat:1973:CDE

[1252] Guy Boillat. Covariant disturbances and exceptional waves. Journal of Mathematical Physics, 14(7):973–976, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i7/p973_s1.

Chaffin:1973:GSR

[1253] Eugene F. Chaffin. Group structure of the relativistic equivalent oscillator. Journal of Mathematical Physics, 14(7):977–979, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i7/p977_s1.

Cheng:1973:QFP

[1254] Kuo-Shung Cheng. Quantization by Feynman’s path integration method. Journal of Mathematical Physics, 14(7):980–981, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i7/p980_s1.

Screaton:1973:ASL

[1255] G. R. Screaton and A. Truman. Analytic structure of the Laplace transforms of distributions with restricted support. Journal of Mathematical Physics, 14(7):982–985, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i7/p982_s1.

Goorjian:1973:UTC

[1256] Peter Goorjian. Uniqueness of the time coordinate for certain singular space–times in general relativity. Journal of Mathematical Physics, 14(7): 986–987, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i7/p986_s1.

Leaf:1973:PVB

[1257] Boris Leaf. Product-vector basis and occupation-number basis in Fock space for bosons and fermions. Journal of Mathematical Physics, 14(7): 988–994, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i7/p988_s1.

Martin:1973:TCO

[1258] Ph. Martin and B. Misra. On trace-class operators of scattering theory, and the asymptotic behavior of scattering cross section at high energy. Journal of Mathematical Physics, 14(8):997–1005, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i8/p997_s1.

McDonald:1973:PMR

[1259] A. McDonald and S. P. Rosen. Properties of matrix representations of SU(3). Journal of Mathematical Physics, 14(8):1006–1016, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i8/p1006_s1.

Persides:1973:RWE

[1260] S. Persides. On the radial wave equation in Schwarzschild’s space–time. Journal of Mathematical Physics, 14(8):1017–1021, August 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i8/p1017_
s1.

Montet:1973:NGF

[1261] G. L. Montet. Note on Green’s functions for open lattices. Journal of Mathematical Physics, 14(8):1022–1023, August 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i8/p1022_s1.

Bahar:1973:EWPa

[1262] E. Bahar. Electromagnetic wave propagation in inhomogeneous multilayered structures of arbitrarily varying thickness- Generalized field transforms. Journal of Mathematical Physics, 14(8):1024–1029, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i8/p1024_s1.

Bahar:1973:EWPb

[1263] E. Bahar. Electromagnetic wave propagation in inhomogeneous multilayered structures of arbitrary thickness — full wave solutions. Journal of Mathematical Physics, 14(8):1030–1036, August 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i8/p1030_s1.

Miller:1973:NNB

[1264] K. Miller and G. A. Viano. On the necessity of nearly-best-possible methods for analytic continuation of scattering data. Journal of Mathematical Physics, 14(8):1037–1048, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i8/p1037_s1.

Klein:1973:PSK

[1265] William Klein. Perturbation solution of the Kirkwood–Salsburg equation. Journal of Mathematical Physics, 14(8):1049–1059, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i8/
p1049_s1.

Sykes:1973:DLTa

[1266] M. F. Sykes, D. S. Gaunt, J. W. Essam, and D. L. Hunter. Derivation of low-temperature expansions for Ising model. II. General theory. Journal of Mathematical Physics, 14(8):1060–1065, August 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i8/p1060_s1.

Sykes:1973:DLTb

[1267] M. F. Sykes, D. S. Gaunt, S. R. Mattingly, J. W. Essam, and C. J. Elliott. Derivation of low-temperature expansions for Ising model. III. Twodimensional lattices-field grouping. Journal of Mathematical Physics, 14(8):1066–1070, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i8/p1066_s1.

Sykes:1973:DLTc

[1268] M. F. Sykes, D. S. Gaunt, J. L. Martin, S. R. Mattingly, and J. W. Essam. Derivation of low-temperature expansions for Ising model. IV. Two-dimensional lattices-temperature grouping. Journal of Mathematical Physics, 14(8):1071–1074, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i8/p1071_s1.

Simmons:1973:DNS

[1269] Carver S. Simmons and Claude Garrod. The density of a nonuniform system in the thermodynamic limit. Journal of Mathematical Physics, 14(8):1075–1087, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i8/p1075_s1.

Suzuki:1973:RRI

[1270] Masuo Suzuki. Rigorous results for Ising ferromagnets of general spin with degeneracy or symmetric potentials. Journal of Mathematical Physics, 14(8):1088–1093, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v14/i8/p1088_s1.

Robinson:1973:EVF

[1271] Peter D. Robinson. Equivalent variational functionals in potential scattering. Journal of Mathematical Physics, 14(8):1094–1096, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i8/
p1094_s1.

Testa:1973:FIC

[1272] F. J. Testa. Finite and infinitesimal canonical transformations. II. Journal of Mathematical Physics, 14(8):1097–1098, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i8/p1097_s1.

Das:1973:SGF

[1273] A. Das. Static gravitational fields. II. Ricci rotation coefficients. Journal of Mathematical Physics, 14(8):1099–1105, August 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i8/p1099_s1.

Onofri:1973:CMD

[1274] E. Onofri and M. Pauri. Constants of motion and degeneration in Hamiltonian systems. Journal of Mathematical Physics, 14(8):1106–1115, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i8/p1106_s1.

Kouri:1973:NSI

[1275] Donald J. Kouri. On the noniterative solution of integral equations for scattering of electromagnetic waves. Journal of Mathematical Physics, 14(8):1116–1120, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i8/p1116_s1.

Herbut:1973:MEG

[1276] F. Herbut, M. Vujicić, and Dj. Sijacki. Minimal extensions as generalized semidirect products. Journal of Mathematical Physics, 14(8):1121– 1124, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i8/p1121_s1.

Rogers:1973:STC

[1277] Harold H. Rogers. Symmetry transformations of the classical Kepler problem. Journal of Mathematical Physics, 14(8):1125–1129, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i8/p1125_s1.

Patera:1973:NBR

[1278] J. Patera and P. Winternitz. A new basis for the representations of the rotation group Lame and Heun polynomials. Journal of Mathematical Physics, 14(8):1130–1139, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v14/i8/p1130_s1.

Hemmer:1973:NYH

[1279] P. C. Hemmer. Nonnegativity of the Yukawa Hamiltonian. Journal of Mathematical Physics, 14(8):1140, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i8/p1140_s1.

Yao:1973:ANP

[1280] Te Hai Yao. On analytic nonlocal potentials. I. A forward dispersion relation. Journal of Mathematical Physics, 14(8):1141–1151, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v14/i8/p1141_s1. See erratum [1775].

Gille:1973:EGT

[1281] J. F. Gille and J. Manuceau. Errata: “Gauge transformations of second type and their implementations. I. Fermions”. Journal of Mathematical Physics, 14(8):1152, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i8/p1152_s3. See [1096].

Papanicolaou:1973:ESG

[1282] G. C. Papanicolaou, D. McLaughlin, and R. Burridge. Erratum: “A stochastic Gaussian beam”. Journal of Mathematical Physics, 14(8): 1152, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i8/p1152_s2. See [1110].

Tannenwald:1973:ENT

[1283] L. M. Tannenwald. Erratum: “Neutron transport equations with spinorbit coupling”. Journal of Mathematical Physics, 14(8):1152, August 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i8/p1152_s1. See [1112].

Krejs:1973:TGS

[1284] Franz Krejs and Abraham Klein. Theory of ground state correlations of closed shell nuclei: a density-matrix formulation. Journal of Mathematical Physics, 14(9):1155–1163, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i9/p1155_s1.

Bayen:1973:CCC

[1285] F. Bayen and P. Hillion. Covariance of the conserved current equations and transformations of Noether’s current. Journal of Mathematical Physics, 14(9):1164–1169, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i9/p1164_s1.

Richard:1973:SMP

[1286] J. L. Richard and E. C. G. Sudarshan. Statistical mechanics of parafermi systems. Journal of Mathematical Physics, 14(9):1170–1175, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i9/p1170_s1.

Butler:1973:SGC

[1287] P. H. Butler and R. C. King. The symmetric group: characters, products and plethysms. Journal of Mathematical Physics, 14(9):1176–1183, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i9/p1176_s1.

Graffi:1973:SSC

[1288] Sandro Graffi. Stieltjes summability and convergence of the Padé approximants for the vacuum polarization by an external field. Journal of Mathematical Physics, 14(9):1184–1186, September 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1184_s1.

Fox:1973:SSB

[1289] Ronald Forrest Fox. Stochastic symmetry breaking of time reversal invariance. Journal of Mathematical Physics, 14(9):1187–1189, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i9/p1187_s1.

Biswas:1973:EAO

[1290] S. N. Biswas, K. Datta, R. P. Saxena, P. K. Srivastava, and V. S. Varma. Eigenvalues of λx^{2m} anharmonic oscillators. Journal of Mathematical Physics, 14(9):1190–1195, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i9/p1190_s1.

Murtaza:1973:DYD

[1291] G. Murtaza and M. A. Rashid. Duality of a Young diagram describing a representation and dimensionality formulas. Journal of Mathematical Physics, 14(9):1196–1198, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i9/p1196_s1.

Ruskai:1973:LPO

[1292] Mary Beth Ruskai and John R. Klauder. Local products as operators. Journal of Mathematical Physics, 14(9):1199–1201, September 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1199_
s1.

Shen:1973:CCT

[1293] C. Y. Shen. On a certain class of transformations in statistical mechanics. Journal of Mathematical Physics, 14(9):1202–1204, September 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1202_
s1.

Klein:1973:NFT

[1294] James R. Klein and Irving I. Zinnes. On nonrelativistic field theory: interpolating fields and long-range forces. Journal of Mathematical Physics, 14(9):1205–1212, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i9/p1205_s1.

Katzin:1973:RIT

[1295] Gerald H. Katzin. Related integral theorem. II. A method for obtaining quadratic constants of the motion for conservative dynamical systems admitting symmetries. Journal of Mathematical Physics, 14(9):1213– 1217, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i9/p1213_s1.

Snellman:1973:CGU

[1296] H. Snellman. Charges as generators of unitary symmetry groups. Journal of Mathematical Physics, 14(9):1218–1221, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1218_s1.

Wu:1973:SAJ

[1297] Alfred C. T. Wu. Structure of Alisauskas–Jucys form of the 9-j coefficients. Journal of Mathematical Physics, 14(9):1222–1223, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i9/p1222_s1.

Boon:1973:CTA

[1298] M. H. Boon and T. H. Seligman. Canonical transformations applied to the free Landau electron. Journal of Mathematical Physics, 14(9):1224– 1227, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i9/p1224_s1.

Goldstein:1973:EPS

[1299] Sheldon Goldstein, Oscar E. Lanford III, and Joel L. Lebowitz. Ergodic properties of simple model system with collisions. Journal of Mathematical Physics, 14(9):1228–1230, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i9/p1228_s1.

Cutkosky:1973:CRK

[1300] R. E. Cutkosky. Construction of reproducing kernels for analytic Hilbert spaces. Journal of Mathematical Physics, 14(9):1231–1238, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i9/p1231_s1.

Nackoney:1973:GIB

[1301] Raymond W. Nackoney. The gravitational influence of a beam of light. I. Journal of Mathematical Physics, 14(9):1239–1247, September 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1239_
s1.

Majumdar:1973:CGC

[1302] S. Datta Majumdar and Biplab Kumar Basu. The Clebsch–Gordan coefficients of SU(3) and the orthogonalization problem. Journal of Mathematical Physics, 14(9):1248–1253, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i9/p1248_s1.

Berezdivin:1973:MSG

[1303] R. Berezdivin and R. K. Sachs. Matter symmetries in general relativistic kinetic theory. Journal of Mathematical Physics, 14(9):1254–1257, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i9/p1254_s1.

McQuistan:1973:EPO

[1304] R. B. McQuistan. Exact parastatistics for one-dimensional lattice space. Journal of Mathematical Physics, 14(9):1258–1261, September 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1258_
s1.

Shapiro:1973:POT

[1305] Joseph Y. Shapiro. Projection operator techniques for compact groups. Journal of Mathematical Physics, 14(9):1262–1270, September 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1262_
s1.

Wyss:1973:QFT

[1306] Walter Wyss. Quantum field theory in terms of sesquilinear forms. Journal of Mathematical Physics, 14(9):1271–1276, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1271_s1.

Ablowitz:1973:DCS

[1307] Mark J. Ablowitz and Alan C. Newell. The decay of the continuous spectrum for solutions of the Korteweg–de Vries equation. Journal of Mathematical Physics, 14(9):1277–1284, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i9/p1277_s1.

Glass:1973:GGH

[1308] Edward N. Glass and Jeffrey Winicour. A geometric generalization of Hooke’s law. Journal of Mathematical Physics, 14(9):1285–1290, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v14/i9/p1285_s1. See errata [1536].

Muller-Kirsten:1973:LCE

[1309] H. J. W. Müller-Kirsten and N. Vahedi-Faridi. Large coupling expansions for eigenenergies and Regge trajectories of the Yukawa potential. Journal of Mathematical Physics, 14(9):1291–1295, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1291_s1.

Treat:1973:EGI

[1310] Richard P. Treat. Erratum: “Gauge Invariant Decomposition of Yang– Mills Potentials”. Journal of Mathematical Physics, 14(9):1296, September 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i9/p1296_s2. See [1051].

Weston:1973:EAI

[1311] V. H. Weston and R. J. Krueger. Erratum and addendum: “On the inverse problem for a hyperbolic dispersive partial differential equation. II”. Journal of Mathematical Physics, 14(9):1296, September 1973. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i9/p1296_

s1. See [1093].

Schafer:1973:DMR

[1312] Lothar Schäfer. Diagrammatic methods and renormalization in the Iwamoto–Yamada cluster expansion. Journal of Mathematical Physics, 14(10):1299–1307, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i10/p1299_s1.

Eggarter:1973:SER

[1313] Thomas P. Eggarter. Schrödinger equation with a random potential: A functional approach. Journal of Mathematical Physics, 14(10):1308– 1313, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i10/p1308_s1.

Levinger:1973:USE

[1314] J. S. Levinger. On the utility of separable expansions for the t matrix. Journal of Mathematical Physics, 14(10):1314–1315, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i10/
p1314_s1.

Kalnins:1973:RIE

[1315] E. G. Kalnins. The relativistically invariant expansion of a scalar function on imaginary Lobachevski space. Journal of Mathematical Physics, 14(10):1316–1319, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i10/p1316_s1.

Hautot:1973:EMN

[1316] André Hautot. Exact motion in noncentral electric fields. Journal of Mathematical Physics, 14(10):1320–1327, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1320_s1.

Chandler:1973:TTD

[1317] Colston Chandler and A. G. Gibson. Transition from time-dependent to time-independent multichannel quantum scattering theory. Journal of Mathematical Physics, 14(10):1328–1335, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1328_s1.

Louck:1973:SCT

[1318] J. D. Louck and L. C. Biedenharn. On the structure of the canonical tensor operators in the unitary groups. III. Further developments of the boson polynomials and their implications. Journal of Mathematical Physics, 14(10):1336–1357, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i10/p1336_s1.

Li:1973:SSD

[1319] Ming Chiang Li. Scattering by a spin-dependent spheroidal potential. Journal of Mathematical Physics, 14(10):1358–1363, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i10/
p1358_s1.

Chow:1973:WPR

[1320] Pao-Liu L. Chow. Wave propagation in a random lattice. II. Journal of Mathematical Physics, 14(10):1364–1373, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1364_s1.

Hosokawa:1973:EPF

[1321] I. Hosokawa. Entropy production in the functional random-walk model. Journal of Mathematical Physics, 14(10):1374–1377, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i10/
p1374_s1.

Friedman:1973:ROE

[1322] Charles N. Friedman. Renormalized oscillator equations. Journal of Mathematical Physics, 14(10):1378–1380, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1378_s1.

Lerche:1973:AKDa

[1323] I. Lerche. Application of Kraichnan’s direct interaction approximation to kinematic dynamo theory. I. Incompressible isotropic turbulence and a singular integral equation. Journal of Mathematical Physics, 14 (10):1381–1388, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i10/p1381_s1.

Taniuti:1973:PMN

[1324] Tosiya Taniuti and Nobuo Yajima. Perturbation method for a nonlinear wave modulation. III. Journal of Mathematical Physics, 14(10):1389– 1397, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i10/p1389_s1.

Prugovecki:1973:MLS

[1325] E. Prugovecki and J. Zorbas. Modified Lippmann–Schwinger equations for two-body scattering theory with long-range interactions. Journal of Mathematical Physics, 14(10):1398–1409, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1398_s1. See er-

ratum [1467].

Prugovecki:1973:BKF

[1326] Eduard Prugovecki. The bra and ket formalism in extended Hilbert space. Journal of Mathematical Physics, 14(10):1410–1422, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i10/p1410_s1.

vanZanten:1973:TOF

[1327] A. J. van Zanten and E. de Vries. Tensor operators of finite groups. Journal of Mathematical Physics, 14(10):1423–1429, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1423_s1.

Wilhelm:1973:STC

[1328] H. E. Wilhelm. Similarity transformation for the compressible Jeffery– Hamel flow of dissipative plasmas. Journal of Mathematical Physics, 14(10):1430–1433, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i10/p1430_s1.

Sigal:1973:SPT

[1329] R. Sigal. A symmetry of Petrov-type N spaces. Journal of Mathematical Physics, 14(10):1434–1436, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i10/p1434_s1.

Anstis:1973:KTO

[1330] G. R. Anstis, H. S. Green, and D. K. Hoffman. Kinetic theory of a one-dimensional model. Journal of Mathematical Physics, 14(10):1437– 1443, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i10/p1437_s1.

Luttinger:1973:GIIb

[1331] J. M. Luttinger. Generalized isoperimetric inequalities. II. Journal of Mathematical Physics, 14(10):1444–1447, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1444_s1.

Luttinger:1973:GIIc

[1332] J. M. Luttinger. Generalized isoperimetric inequalities. III. Journal of Mathematical Physics, 14(10):1448–1450, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1448_s1.

Beck:1973:CAB

[1333] Robert E. Beck. A condition for the applicability of Biedenharn’s lemma. Journal of Mathematical Physics, 14(10):1451–1452, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i10/p1451_s1.

Wald:1973:PKB

[1334] Robert M. Wald. On perturbations of a Kerr black hole. Journal of Mathematical Physics, 14(10):1453–1461, October 1973. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i10/p1453_s1.

Chen:1973:FAS

[1335] Etang Chen. Facial aspect of superposition principle in algebraic quantum theory. Journal of Mathematical Physics, 14(10):1462–1465, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i10/p1462_s1.

Chakraborty:1973:MRP

[1336] B. Chakraborty. The mathematical reflection problem. Journal of Mathematical Physics, 14(10):1466–1468, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i10/p1466_s1.

Maczynski:1973:FRN

[1337] M. J. Maczyński. The field of real numbers in axiomatic quantum mechanics. Journal of Mathematical Physics, 14(10):1469–1471, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i10/p1469_s1.

Randall:1973:OSI

[1338] C. H. Randall and D. J. Foulis. Operational statistics. II. Manuals of operations and their logics. Journal of Mathematical Physics, 14(10):1472– 1480, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i10/p1472_s1.

Berson:1973:EFT

[1339] I. Berson and J. Valdmanis. Electron in the field of two monochromatic electromagnetic waves. Journal of Mathematical Physics, 14(10):1481– 1484, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i10/p1481_s1.

Osborn:1973:ECS

[1340] T. A. Osborn. Erratum: “On the convergence of separable expansions for the t matrix”. Journal of Mathematical Physics, 14(10):1485, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i10/p1485_s1. See [1154].

Zohni:1973:EES

[1341] O. Zohni. Erratum: “Off-energy-shell t matrix for local potentials with nonlocal square well core interaction”. Journal of Mathematical Physics, 14(10):1486, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i10/p1486_s1. See [1127].

Bellman:1973:IIR

[1342] Richard Bellman and Sueo Ueno. Invariant imbedding and the resolvent of Fredholm integral equation with composite symmetric kernel. Journal of Mathematical Physics, 14(11):1489–1490, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i11/p1489_s1.

Aviran:1973:CGC

[1343] A. Aviran and D. B. Litvin. Clebsch–Gordan coefficients for nonunitary groups. Journal of Mathematical Physics, 14(11):1491–1494, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1491_s1.

Lorente:1973:SPM

[1344] M. Lorente, P. L. Huddleston, and P. Roman. Some properties of matrices occurring in SL(2, C) invariant wave equations. Journal of Mathematical Physics, 14(11):1495–1497, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i11/p1495_s1.

Pereira:1973:TSC

[1345] C. Martin Pereira. Thin sandwich conjecture and a fifth initial value equation for the gravitational field. Journal of Mathematical Physics, 14 (11):1498–1501, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i11/p1498_s1.

Bahar:1973:DEWa

[1346] E. Bahar. Depolarization of electromagnetic waves excited by distributions of electric and magnetic sources in inhomogeneous multilayered structures of arbitrarily varying thickness. Generalized field transforms. Journal of Mathematical Physics, 14(11):1502–1509, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1502_s1.

Bahar:1973:DEWb

[1347] E. Bahar. Depolarization of electromagnetic waves excited by distributions of electric and magnetic sources in inhomogeneous multilayered structures of arbitrarily varying thickness. Full wave solutions. Journal of Mathematical Physics, 14(11):1510–1515, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i11/p1510_s1.

Vujicic:1973:GSP

[1348] M. Vujicić, F. Herbut, and Dj. Sijacki. Generalized semidirect product in group unifications. Journal of Mathematical Physics, 14(11):1516– 1518, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1516_s1.

Hecht:1973:AAM

[1349] W. Hecht. An analytic approximation method for the one-dimensional Schrödinger equation. II. Journal of Mathematical Physics, 14(11):1519– 1521, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1519_s1.

Nesbet:1973:RJF

[1350] R. K. Nesbet. Radial Jost functions in scattering theory. Journal of Mathematical Physics, 14(11):1522–1526, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v14/i11/p1522_s1.

Anderson:1973:IEW

[1351] Robert L. Anderson, Sukeyuki Kumei, and Carl E. Wulfman. Invariants of the equations of wave mechanics: Rigid rotator and symmetric top. Journal of Mathematical Physics, 14(11):1527–1531, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1527_s1.

Federbush:1973:PSG

[1352] Paul Federbush. Positivity for some generalized Yukawa models in one space dimension. Journal of Mathematical Physics, 14(11):1532–1542, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1532_s1.

Weissman:1973:MEW

[1353] Y. Weissman and J. A. Sussmann. Motion and expansion of wave packets in periodic potentials. Journal of Mathematical Physics, 14(11):1543– 1545, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1543_s1.

Blasi:1973:LAC

[1354] A. Blasi and R. Collina. Lie algebras connected with infinite momentum kinematics. Journal of Mathematical Physics, 14(11):1546–1550, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1546_s1.

OMurchadha:1973:EUS

[1355] Niall O’Murchadha and James W. York, Jr. Existence and uniqueness of solutions of the Hamiltonian constraint of general relativity on compact manifolds. Journal of Mathematical Physics, 14(11):1551–1557, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1551_s1.

Bart:1973:CAC

[1356] G. R. Bart, P. W. Johnson, and R. L. Warnock. Continuum ambiguity in the construction of unitary analytic amplitudes from fixed-energy scattering data. Journal of Mathematical Physics, 14(11):1558–1565, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1558_s1.

DeSanto:1973:SRR

[1357] John A. DeSanto. Scattering from a random rough surface: Diagram methods for elastic media. Journal of Mathematical Physics, 14 (11):1566–1573, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i11/p1566_s1.

Coz:1973:EGT

[1358] Marcel Coz and Christiane Coudray. Existence of generalized translation operators from the Agranovitch–Marchenko transformation (Jost solutions). Journal of Mathematical Physics, 14(11):1574–1578, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1574_s1.

Lerche:1973:AKDb

[1359] I. Lerche. Application of Kraichnan’s direct interaction approximation to kinematic dynamo theory. II. incompressible, helical velocity turbulence and a pair of coupled, singular, nonlinear integral equations. Journal of Mathematical Physics, 14(11):1579–1587, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i11/p1579_s1.

Lerche:1973:AKDc

[1360] I. Lerche. Application of Kraichnan’s direct interaction approximation to kinematic dynamo theory. III. Solution of the Kraichnan equations under Parker’s ‘short-sudden’ conditions. Journal of Mathematical Physics, 14 (11):1588–1591, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i11/p1588_s1.

Hanson:1973:UVA

[1361] F. B. Hanson, A. Klimas, G. V. Ramanathan, and G. Sandri. Uniformly valid asymptotic solution to a Volterra equation on an infinite interval. Journal of Mathematical Physics, 14(11):1592–1600, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1592_s1.

Mason:1973:DIS

[1362] A. L. Mason. Deep inelastic scattering in a renormalized perturbative model. Journal of Mathematical Physics, 14(11):1601–1619, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1601_s1.

Huszar:1973:URG

[1363] M. Huszár. Unitary representations of the SL(2, C) group in horospheric basis. Journal of Mathematical Physics, 14(11):1620–1625, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i11/p1620_s1.

Kac:1973:BEC

[1364] M. Kac and J. M. Luttinger. Bose–Einstein condensation in the presence of impurities. Journal of Mathematical Physics, 14(11):1626–1628, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1626_s1.

Tjotta:1973:KTW

[1365] Jacqueline Naze Tjøtta and Alf H. Øien. Kinetic theory of a weakly coupled and weakly inhomogeneous plasma in a magnetic field. Journal of Mathematical Physics, 14(11):1629–1636, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i11/p1629_s1.

Baer:1973:CCO

[1366] Michael Baer and Donald J. Kouri. Coupled channel operator approach to e–H scattering. Journal of Mathematical Physics, 14(11):1637–1642, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1637_s1.

Case:1973:DVM

[1367] K. M. Case and S. C. Chiu. The discrete version of the Marchenko equations in the inverse scattering problem. Journal of Mathematical Physics, 14(11):1643–1647, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i11/p1643_s1.

desCloizeaux:1973:ABS

[1368] J. des Cloizeaux and M. L. Mehta. Asymptotic behavior of spacing distributions for the eigenvalues of random matrices. Journal of Mathematical Physics, 14(11):1648–1650, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i11/p1648_s1.

Gunaydin:1973:QSO

[1369] Murat Günaydin and Feza Gürsey. Quark structure and octonions. Journal of Mathematical Physics, 14(11):1651–1667, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i11/
p1651_s1.

Evans:1973:PMP

[1370] G. T. Evans and D. K. Sen. Projective manifolds and projective theory of relativity. Journal of Mathematical Physics, 14(11):1668–1674, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1668_s1.

Ciulli:1973:OAE

[1371] S. Ciulli and G. Nenciu. Optimal analytic extrapolation for the scattering amplitude from cuts to interior points. Journal of Mathematical Physics, 14(11):1675–1691, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i11/p1675_s1.

Calogero:1973:NLM

[1372] F. Calogero. Nonlinear Lee model. II. Journal of Mathematical Physics, 14(11):1692–1705, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i11/p1692_s1.

Rae:1973:CTT

[1373] J. Rae and R. Davidson. A comparison of two transformation theories of classical mechanics. Journal of Mathematical Physics, 14(11):1706– 1714, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1706_s1.

Wood:1973:STS

[1374] D. W. Wood and H. P. Griffiths. Simplex transformations of the spin-1/2 Ising model. I. Journal of Mathematical Physics, 14(11):1715–1722, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i11/p1715_s1.

Torrence:1973:GIC

[1375] R. J. Torrence and W. M. Tulczyjew. Gauge invariant canonical mechanics for charged particles. Journal of Mathematical Physics, 14(12):1725– 1732, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1725_s1.

Kohler:1973:PSW

[1376] W. Kohler and G. C. Papanicolaou. Power statistics for wave propagation in one dimension and comparison with radiative transport theory. Journal of Mathematical Physics, 14(12):1733–1745, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1733_s1.

Hardy:1973:TET

[1377] J. Hardy, Y. Pomeau, and O. de Pazzis. Time evolution of a twodimensional model system. I. Invariant states and time correlation functions. Journal of Mathematical Physics, 14(12):1746–1759, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1746_s1.

Fischbach:1973:DKP

[1378] Ephraim Fischbach, Michael Martin Nieto, and C. K. Scott. Diffin– Kemmer–Petiau subalgebras: representations and applications. Journal of Mathematical Physics, 14(12):1760–1774, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1760_s1.

Emch:1973:DEF

[1379] Gérard G. Emch. Diffusion, Einstein formula and mechanics. Journal of Mathematical Physics, 14(12):1775–1783, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1775_s1.

Bracken:1973:PQM

[1380] A. J. Bracken and H. S. Green. Parastatistics and the quark model. Journal of Mathematical Physics, 14(12):1784–1793, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1784_s1.

Hudgin:1973:GRG

[1381] Richard H. Hudgin. Generalizing Riemannian geometry. Journal of Mathematical Physics, 14(12):1794–1799, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1794_s1.

Albeverio:1973:STM

[1382] Sergio Albeverio. Scattering theory in a model of quantum fields. I. Journal of Mathematical Physics, 14(12):1800–1816, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1800_s1.

McLaughlin:1973:RBT

[1383] David W. McLaughlin and Alwyn C. Scott. A restricted Bäcklund transformation. Journal of Mathematical Physics, 14(12):1817–1828, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1817_s1.

Besieris:1973:KEQ

[1384] Ioannis M. Besieris and Frederick D. Tappert. Kinetic equations for the quantized motion of a particle in a randomly perturbed potential field. Journal of Mathematical Physics, 14(12):1829–1836, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1829_s1.

Saenz:1973:TMB

[1385] A. W. Sáenz and W. W. Zachary. Two-magnon bound states in Heisenberg ferromagnets. Journal of Mathematical Physics, 14(12):1837–1852, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1837_s1. See erratum [1711].

Boyer:1973:DIC

[1386] Charles P. Boyer and Kurt Bernardo Wolf. Deformations of inhomogeneous classical Lie algebras to the algebras of the linear groups. Journal of Mathematical Physics, 14(12):1853–1859, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1853_s1.

Ahlrichs:1973:ABA

[1387] Reinhart Ahlrichs. Asymptotic behavior of atomic bound state wave functions. Journal of Mathematical Physics, 14(12):1860–1863, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1860_s1.

Vawter:1973:EEB

[1388] Richard Vawter. Energy eigenvalues of a bounded centrally located harmonic oscillator. Journal of Mathematical Physics, 14(12):1864–1870, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1864_s1.

Lee:1973:GIA

[1389] Ching Tsung Lee. Griffiths’ inequalities for Ashkin–Teller model. Journal of Mathematical Physics, 14(12):1871–1875, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1871_s1.

Fernandez:1973:IBT

[1390] Julio F. Fernández. Irreversible behavior of a thermally insulated system. Journal of Mathematical Physics, 14(12):1876–1882, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1876_s1.

Freeman:1973:IES

[1391] D. F. Freeman and J. Nuttall. Integral equations for scattering in potentials of infinite range. Journal of Mathematical Physics, 14(12):1883– 1885, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1883_s1.

Levine:1973:SMC

[1392] Jack Levine and Gerald H. Katzin. Symmetry mappings of constrained dynamical systems and an associated related integral theorem. Journal of Mathematical Physics, 14(12):1886–1891, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1886_s1.

Ohanian:1973:GNI

[1393] Hans C. Ohanian. Gravitation and the new improved energy-momentum tensor. Journal of Mathematical Physics, 14(12):1892–1897, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1892_s1.

Taylor:1973:SRE

[1394] Leonard S. Taylor and Danai Lekhyananda. Scintillations of randomized electromagnetic fields. II. Journal of Mathematical Physics, 14(12):1898– 1903, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1898_s1.

Draayer:1973:WRC

[1395] J. P. Draayer and Yoshimi Akiyama. Wigner and Racah coefficients for su_{3}. Journal of Mathematical Physics, 14(12):1904–1912, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1904_s1.

Boffi:1973:LIT

[1396] V. C. Boffi and G. Spiga. Linear integral transformations generated by the three-dimensional neutron transport kernel. Journal of Mathematical Physics, 14(12):1913–1917, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i12/p1913_s1.

Knops:1973:ISD

[1397] Hubert J. F. Knops. Infinite spin dimensionality limit for nontranslationally invariant interactions. Journal of Mathematical Physics, 14(12):1918–1920, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i12/p1918_s1.

Lang:1973:PDF

[1398] R. H. Lang. Probability density function and moments of the field in a slab of one-dimensional random medium. Journal of Mathematical Physics, 14(12):1921–1926, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i12/p1921_s1.

Brascamp:1973:SRR

[1399] H. J. Brascamp, H. Kunz, and F. Y. Wu. Some rigorous results for the vertex model in statistical mechanics. Journal of Mathematical Physics, 14(12):1927–1932, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i12/p1927_s1.

Ihrig:1973:ESS

[1400] E. Ihrig and D. K. Sen. An exact solution of a scalar-tensor theory of gravitation. Journal of Mathematical Physics, 14(12):1933–1934, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1933_s1.

Ben-Abraham:1973:QGD

[1401] S. I. Ben-Abraham and A. Lonke. Quantization of a general dynamical system. Journal of Mathematical Physics, 14(12):1935–1937, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1935_s1.

Lieb:1973:PSS

[1402] Elliott H. Lieb and Mary Beth Ruskai. Proof of the strong subadditivity of quantum-mechanical entropy. Journal of Mathematical Physics, 14 (12):1938–1941, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i12/p1938_s1.

Hori:1973:STEb

[1403] Motoo Hori. Statistical theory of effective electrical, thermal and magnetic properties of random heterogeneous materials. II. Bounds for the effective permittivity of statistically anisotropic materials. Journal of Mathematical Physics, 14(12):1942–1948, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1942_s1.

Lerche:1973:INC

[1404] I. Lerche and E. N. Parker. On the improper neglect of certain terms in random function theory. Journal of Mathematical Physics, 14(12):1949– 1954, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1949_s1.

Lohe:1973:RBB

[1405] M. A. Lohe and C. A. Hurst. Relation between the boson calculus and Zhelobenko’s method. Journal of Mathematical Physics, 14(12):1955– 1958, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1955_s1.

Lohe:1973:SRO

[1406] M. A. Lohe. Spinor representations of the orthogonal groups. Journal of Mathematical Physics, 14(12):1959–1964, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/p1959_s1.

Volkin:1973:POP

[1407] Howard C. Volkin. Phase operators and phase relations for photon states. Journal of Mathematical Physics, 14(12):1965–1976, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v14/i12/p1965_s1.

Patera:1973:ITC

[1408] J. Patera and P. Winternitz. Invariant trilinear couplings involving both SU(2) and SU(1, 1) states. Journal of Mathematical Physics, 14 (12):1977–1983, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i12/p1977_s1.

Piasecki:1973:TPV

[1409] J. Piasecki and P. Résibois. On the transport properties of a van der Waals fluid. I. Formal theory. Journal of Mathematical Physics, 14 (12):1984–2004, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v14/i12/p1984_s1.

Mukunda:1973:MFL

[1410] N. Mukunda. Matrices of finite Lorentz transformations in a noncompact basis. III. Completeness relation for O(2, 1). Journal of Mathematical Physics, 14(12):2005–2010, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i12/p2005_s1.

Hautot:1973:MRM

[1411] André Hautot. Motion in radial magnetic fields. Journal of Mathematical Physics, 14(12):2011–2017, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i12/p2011_s1.

MacDowell:1973:ERP

[1412] S. W. MacDowell and Ralph Roskies. Errata: “Reduction of the Poincaré group with respect to the Lorentz group”. Journal of Mathematical Physics, 14(12):2018, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v14/i12/p2018_s2. See [1038].

Siegel:1973:EAF

[1413] Armand Siegel and Terence Burke. Erratum: “Approximate functional integral methods in statistical mechanics. I. Moment expansions”. Journal of Mathematical Physics, 14(12):2018, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/

p2018_s3. See [1048].

Zachary:1973:ESM

[1414] W. W. Zachary. Errata: “Off-shell T matrix corresponding to a sum of Coulomb and separable potentials by expansion in O(4) harmonics”. Journal of Mathematical Physics, 14(12):2018, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i12/

p2018_s1. See [627].

Speer:1974:RWI

[1415] Eugene R. Speer. Renormalization and Ward identities using complex space–time dimension. Journal of Mathematical Physics, 15(1):1–6, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i1/p1_s1.

Eggarter:1974:CRL

[1416] T. P. Eggarter. On a complex representation of Lorentzian random variables. Journal of Mathematical Physics, 15(1):7–8, January 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p7_s1.

Kiehn:1974:EHP

[1417] R. M. Kiehn. An extension of Hamilton’s principle to include dissipative systems. Journal of Mathematical Physics, 15(1):9–13, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i1/p9_s1.

Stephani:1974:DPR

[1418] H. Stephani. Debye potentials in Riemannian spaces. Journal of Mathematical Physics, 15(1):14–16, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i1/p14_s1.

Bellman:1974:IIR

[1419] R. Bellman and S. Ueno. Invariant imbedding and the resolvent of Fredholm integral equations with variable parameters. Journal of Mathematical Physics, 15(1):17–19, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i1/p17_s1.

Weder:1974:LMD

[1420] R. A. Weder. On the Lee model with dilatation analytic cutoff function. Journal of Mathematical Physics, 15(1):20–24, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p20_s1.

Wong:1974:UR

[1421] M. K. F. Wong. Unitary representations of SO(n, 1). Journal of Mathematical Physics, 15(1):25–30, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i1/p25_s1.

Reents:1974:SPE

[1422] G. Reents. Scattering of photons by an external current. Journal of Mathematical Physics, 15(1):31–34, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i1/p31_s1.

Grmela:1974:AEK

[1423] Miroslav Grmela. On the approach to equilibrium in kinetic theory. Journal of Mathematical Physics, 15(1):35–40, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p35_s1.

Dodd:1974:CCO

[1424] L. R. Dodd and A. M. Gibbs. Cluster coefficients for the one-dimensional Bose gas with point interactions. Journal of Mathematical Physics, 15(1): 41–43, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i1/p41_s1.

Newman:1974:LWF

[1425] E. T. Newman. Lienard–Wiechert fields and general relativity. Journal of Mathematical Physics, 15(1):44–45, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i1/p44_s1.

Hansen:1974:MMS

[1426] R. O. Hansen. Multipole moments of stationary space–times. Journal of Mathematical Physics, 15(1):46–52, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i1/p46_s1.

Wu:1974:CGF

[1427] Alfred C. T. Wu and Tai Tsun Wu. Classical SU(3) gauge field equations. Journal of Mathematical Physics, 15(1):53–54, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p53_s1.

Cunningham:1974:ECE

[1428] Gordon W. Cunningham and Paul H. E. Meijer. Exact calculation of the energy and heat capacity for the triangular lattice with three different coupling constants. Journal of Mathematical Physics, 15(1):55– 59, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i1/p55_s1.

Fischbach:1974:LAD

[1429] Ephraim Fischbach, James D. Louck, Michael Martin Nieto, and C. K. Scott. The Lie algebra SO(N) and the Duffin–Kemmer–Petiau ring. Journal of Mathematical Physics, 15(1):60–64, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p60_s1.

Brascamp:1974:ZPF

[1430] H. J. Brascamp and H. Kunz. Zeroes of the partition function for the Ising model in the complex temperature plane. Journal of Mathematical Physics, 15(1):65–66, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i1/p65_s1.

Srinivas:1974:SMV

[1431] M. D. Srinivas and Y. S. Prahalad. Statistical mechanics for velocity dependent interactions. Journal of Mathematical Physics, 15(1):67– 69, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i1/p67_s1.

Lorente:1974:GEP

[1432] M. Lorente and P. Roman. General expressions for the position and spin operators of relativistic systems. Journal of Mathematical Physics, 15(1): 70–74, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i1/p70_s1.

Larsen:1974:ASN

[1433] Edward W. Larsen and Joseph B. Keller. Asymptotic solution of neutron transport problems for small mean free paths. Journal of Mathematical Physics, 15(1):75–81, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i1/p75_s1.

Capper:1974:DRZ

[1434] D. M. Capper and G. Leibbrandt. Dimensional regularization for zero-mass particles in quantum field theory. Journal of Mathematical Physics, 15(1):82–85, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i1/p82_s1.

Capper:1974:CHV

[1435] D. M. Capper and G. Leibbrandt. On a conjecture by ’t Hooft and Veltman. Journal of Mathematical Physics, 15(1):86–87, January 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p86_s1.

Goldin:1974:NCA

[1436] G. A. Goldin, J. Grodnik, R. T. Powers, and D. H. Sharp. Nonrelativistic current algebra in the N/V limit. Journal of Mathematical Physics, 15(1):88–100, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i1/p88_s1.

Zipfel:1974:SSW

[1437] G. G. Zipfel, Jr. Scattering of scalar waves from a random irregular interface. Journal of Mathematical Physics, 15(1):101–113, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i1/p101_s1.

Matcha:1974:NBE

[1438] Robert L. Matcha and Kenneth C. Daiker. Natural basis expansion of functions of the distance between two points: Generalization of Laplace’s and Neumann’s expansions. Journal of Mathematical Physics, 15(1):114– 118, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i1/p114_s1.

Backhouse:1974:ESR

[1439] N. B. Backhouse. Extra selection rules (in magnetic systems). Journal of Mathematical Physics, 15(1):119–124, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i1/p119_s1.

Bollini:1974:ES

[1440] C. G. Bollini and J. J. Giambiagi. On the exponential superpropagator. Journal of Mathematical Physics, 15(1):125–128, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p125_s1.

Urbantke:1974:KRC

[1441] Helmuth K. Urbantke. KS-related f–g couples as exact vacuum solutions of Salam’s two-tensor theory. Journal of Mathematical Physics, 15(1):129–132, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i1/p129_s1. See erratum [1773].

Lipszyc:1974:ODM

[1442] K. Lipszyc. One-dimensional model of the rearrangement process and the Faddeev equations. Journal of Mathematical Physics, 15(1):133– 138, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i1/p133_s1.

Calogero:1974:CPL

[1443] F. Calogero and C. Marchioro. A comment on the paper “Lower Bounds on the Energy Eigenvalues of Systems Containing Identical Particles”. Journal of Mathematical Physics, 15(1):139–140, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i1/p139_s1. See

[1].

Case:1974:DIS

[1444] K. M. Case. The discrete inverse scattering problem in one dimension. Journal of Mathematical Physics, 15(2):143–146, February 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i2/p143_
s1.

StJohn:1974:LOG

[1445] Richard H. St. John and J. D. Finley III. Local observations of geodesics in the extended Kerr manifold. Journal of Mathematical Physics, 15 (2):147–148, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i2/p147_s1.

Varga:1974:MTS

[1446] B. B. Varga and S. Ø. Aks. Multiple time scales and the ϕ^{4} model of quantum field theory. Journal of Mathematical Physics, 15(2):149– 154, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i2/p149_s1.

Angelopoulos:1974:RLS

[1447] E. Angelopoulos. Reduction on the Lorentz subgroup of U.I.R.’s of the Poincaré group induced by a semisimple little group. Journal of Mathematical Physics, 15(2):155–165, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i2/p155_s1.

Ebner:1974:GES

[1448] Dieter W. Ebner. The general exact solution of the equation of geodesics of the general exact homogeneous plane gravitational wave. Journal of Mathematical Physics, 15(2):166–171, February 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i2/p166_s1.

Kyriakopoulos:1974:LAH

[1449] E. Kyriakopoulos. Lie algebra of the hypergeometric functions. Journal of Mathematical Physics, 15(2):172–175, February 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i2/p172_s1.

Ali:1974:FOQ

[1450] S. Twareque Ali and Gérard G. Emch. Fuzzy observables in quantum mechanics. Journal of Mathematical Physics, 15(2):176–182, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i2/p176_s1.

Kac:1974:BEC

[1451] M. Kac and J. M. Luttinger. Bose–Einstein condensation in the presence of impurities. II. Journal of Mathematical Physics, 15(2):183–186, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i2/p183_s1.

Zucker:1974:NLS

[1452] I. J. Zucker. A note on lattice sums in two dimensions. Journal of Mathematical Physics, 15(2):187, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i2/p187_s1.

Glasser:1974:ELS

[1453] M. L. Glasser. The evaluation of lattice sums. III. Phase modulated sums. Journal of Mathematical Physics, 15(2):188–189, February 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i2/p188_
s1.

Sloan:1974:PCT

[1454] Alan D. Sloan. The polaron without cutoffs in two space dimensions. Journal of Mathematical Physics, 15(2):190–201, February 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i2/p190_
s1.

Bahar:1974:DNM

[1455] E. Bahar. Depolarization in nonuniform multilayered structures-full wave solutions. Journal of Mathematical Physics, 15(2):202–208, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i2/p202_s1.

Weston:1974:IS

[1456] V. H. Weston. On inverse scattering. Journal of Mathematical Physics, 15(2):209–213, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i2/p209_s1.

Grimson:1974:EFA

[1457] R. C. Grimson. Exact formulas for 2 × n arrays of dumbbells. Journal of Mathematical Physics, 15(2):214–216, February 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i2/p214_s1.

Fox:1974:PAM

[1458] Ronald Forrest Fox. Physical applications of multiplicative stochastic processes. II. Derivation of the Bloch equations for magnetic relaxation. Journal of Mathematical Physics, 15(2):217–219, February 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i2/p217_
s1.

Um:1974:NPP

[1459] Gregory S. Um. On normalization problems of the path integral method. Journal of Mathematical Physics, 15(2):220–224, February 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i2/p220_
s1.

Bosco:1974:CVF

[1460] B. Bosco and M. T. Sacchi. On the conditions that a vector field vanish outside a given radius. Journal of Mathematical Physics, 15(2):225– 229, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i2/p225_s1.

Graves-Morris:1974:ACS

[1461] P. R. Graves-Morris and J. F. Rennison. Analyticity in the coupling strength. Journal of Mathematical Physics, 15(2):230–233, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i2/p230_s1.

Devaney:1974:MEP

[1462] A. J. Devaney and E. Wolf. Multipole expansions and plane wave representations of the electromagnetic field. Journal of Mathematical Physics, 15(2):234–244, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i2/p234_s1.

Nagel:1974:AVP

[1463] Bengt Nagel and Håkan Snellman. Analytic vectors for the Poincaré group in quantum field theory. Journal of Mathematical Physics, 15 (2):245–246, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i2/p245_s1.

Balakrishnan:1974:RTS

[1464] V. Balakrishnan. Regge trajectory structure of the Amati–Bertocchi– Fubini–Stranghellini–Tonin multiperipheral model. Journal of Mathematical Physics, 15(2):247–257, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i2/p247_s1.

King:1974:BRE

[1465] R. C. King. Branching rules for GL(N) ⊃ Σ_{m} and the evaluation of inner plethysms. Journal of Mathematical Physics, 15(2):258–267, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i2/p258_s1.

Hautot:1974:SCR

[1466] André Hautot. Some comments on a recent paper of Glasser (J. Math. Phys. 14, 409 (1973)). Journal of Mathematical Physics, 15(2):268, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i2/p268_s1. See [1161].

Prugovecki:1974:EML

[1467] E. Prugovecki and J. Zorbas. Erratum: “Modified Lippmann–Schwinger equations for two-body scattering theory with long-range interactions”. Journal of Mathematical Physics, 15(2):268, February 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i2/p268_s2. See

[1325].

Pereira:1974:NCI

[1468] C. Martin Pereira. New commutator identities on the Riemann tensor. Journal of Mathematical Physics, 15(3):269–272, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i3/p269_s1.

Nieto:1974:LNS

[1469] Michael Martin Nieto. Lowest nonnegative solution of the Yukawa Hamiltonian. Journal of Mathematical Physics, 15(3):273–274, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i3/
p273_s1.

Prugovecki:1974:ELN

[1470] E. Prugovecki and A. Tip. On ergodic limits of normal states in quantum statistical mechanics. Journal of Mathematical Physics, 15(3):275– 282, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p275_s1.

DeSanto:1974:GFE

[1471] John A. DeSanto. Green’s function for electromagnetic scattering from a random rough surface. Journal of Mathematical Physics, 15(3):283– 288, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p283_s1.

Kalaba:1974:IVM

[1472] R. Kalaba and E. Zagustin. An initial value method for an inverse problem in wave propagation. Journal of Mathematical Physics, 15(3):289– 290, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p289_s1.

Chandler:1974:TDM

[1473] Colston Chandler and A. G. Gibson. Time-dependent multichannel Coulomb scattering theory. Journal of Mathematical Physics, 15(3):291– 294, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p291_s1.

Schulman:1974:SDD

[1474] L. S. Schulman. Some differential-difference equations containing both advance and retardation. Journal of Mathematical Physics, 15(3):295– 298, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p295_s1.

Larsen:1974:SSO

[1475] Edward W. Larsen. Solutions of the steady, one-speed neutron transport equation for small mean free paths. Journal of Mathematical Physics, 15(3):299–305, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i3/p299_s1.

Leibowitz:1974:SCF

[1476] Elhanan Leibowitz. Spinor calculus in five-dimensional relativity. Journal of Mathematical Physics, 15(3):306–312, March 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i3/p306_s1.

Leibowitz:1974:FDA

[1477] Elhanan Leibowitz. Five-dimensional approach to the neutrino and electron theories. Journal of Mathematical Physics, 15(3):313–316, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i3/p313_s1.

Wenzel:1974:SBC

[1478] Alan R. Wenzel. Smoothed boundary conditions for randomly rough surfaces. Journal of Mathematical Physics, 15(3):317–323, March 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i3/p317_
s1.

Moscheo:1974:ODF

[1479] R. Moscheo and R. P. Moya. A one-dimensional fermion model exhibiting an anomalous type of phase transition. Journal of Mathematical Physics, 15(3):324–328, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i3/p324_s1.

duPlessis:1974:CKV

[1480] J. C. du Plessis. Conformal Killing vector fields on timelike two-surfaces. Journal of Mathematical Physics, 15(3):329–333, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i3/p329_s1.

Fabrey:1974:RNO

[1481] James D. Fabrey. Renormalized number operators. Journal of Mathematical Physics, 15(3):334–337, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i3/p334_s1.

Newton:1974:TPM

[1482] Roger G. Newton. The three-particle S-matrix. Journal of Mathematical Physics, 15(3):338–343, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i3/p338_s1.

Lukierski:1974:GFF

[1483] J. Lukierski and K. Sienkiewicz. Generalized free fields and the representations of Weyl group. Journal of Mathematical Physics, 15(3): 344–349, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p344_s1.

Arnal:1974:AIR

[1484] D. Arnal and G. Pinczon. On algebraically irreducible representations of the Lie algebra sl(2). Journal of Mathematical Physics, 15(3):350– 359, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p350_s1.

Tilgner:1974:NTV

[1485] Hans Tilgner. On nonlinear transformations in vector spaces. Colineation and conformal groups. Journal of Mathematical Physics, 15(3):

360–366, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p360_s1.

Gregorio:1974:FLE

[1486] M. A. Gregorio and D. R. Avalos. Faddeev-like equations with multibody forces. Journal of Mathematical Physics, 15(3):367–373, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i3/p367_s1.

Mirman:1974:MUF

[1487] R. Mirman. A model unified field theory. Journal of Mathematical Physics, 15(3):374–379, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i3/p374_s1.

Montgomery:1974:NLA

[1488] W. Montgomery and L. O’Raifeartaigh. Noncompact Lie-algebraic approach to the unitary representations of SŨ(1, 1): Role of the confluent hypergeometric equation. Journal of Mathematical Physics, 15(3): 380–382, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p380_s1.

Stuck:1974:SMA

[1489] B. W. Stuck. On some mathematical aspects of deterministic classical electrodynamics. Journal of Mathematical Physics, 15(3):383– 393, March 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i3/p383_s1.

Greenberger:1974:SUP

[1490] Daniel M. Greenberger. Some useful properties of a theory of variable mass particles. Journal of Mathematical Physics, 15(4):395–405, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i4/p395_s1.

Greenberger:1974:WPI

[1491] Daniel M. Greenberger. Wavepackets for particles of indefinite mass. Journal of Mathematical Physics, 15(4):406–412, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i4/p406_s1.

Hallidy:1974:PAF

[1492] William Hallidy. A property of asymptotically flat, static, vacuum space– times. Journal of Mathematical Physics, 15(4):413–414, April 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i4/p413_
s1.

Liebman:1974:MTA

[1493] Alan J. Liebman and Jean-Yves Parlange. Mayer’s theory analysis of repulsive and weak, long range potentials. Journal of Mathematical Physics, 15(4):415–419, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i4/p415_s1.

Maia:1974:CSF

[1494] M. D. Maia. Conformal spinor fields in general relativity. Journal of Mathematical Physics, 15(4):420–425, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i4/p420_s1.

Newman:1974:KC

[1495] Ezra T. Newman and Jeffrey Winicour. The Kerr congruence. Journal of Mathematical Physics, 15(4):426–428, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i4/p426_s1.

Barcilon:1974:ISI

[1496] Victor Barcilon. Iterative solution of the inverse Sturm–Liouville problem. Journal of Mathematical Physics, 15(4):429–436, April 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i4/p429_
s1.

Bailey:1974:PD

[1497] Paul B. Bailey. On the problem of diffraction. Journal of Mathematical Physics, 15(4):437–444, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i4/p437_s1.

Hioe:1974:ASQ

[1498] F. T. Hioe. An approach to the study of quantum systems. Journal of Mathematical Physics, 15(4):445–452, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i4/p445_s1.

Watanabe:1974:NGI

[1499] Yoiti Watanabe. Note on gauge invariance and conservation laws for a class of nonlinear partial differential equations. Journal of Mathematical Physics, 15(4):453–457, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i4/p453_s1.

Johnson:1974:TPA

[1500] R. E. Johnson, N. K. Pope, and A. J. Barrett. Two-particle approximation for the quantum third virial coefficient: Comparison of two approaches. Journal of Mathematical Physics, 15(4):458–467, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i4/
p458_s1.

Mariwalla:1974:ACS

[1501] K. H. Mariwalla. Applications of the concept of strength of a system of partial differential equations. Journal of Mathematical Physics, 15(4): 468–473, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i4/p468_s1.

Rashid:1974:OSE

[1502] M. A. Rashid. Occupation statistics from exact recursion relations for occupation by dumbbells of a 2 × N array. Journal of Mathematical Physics, 15(4):474–476, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i4/p474_s1.

Radhakrishnan:1974:SRI

[1503] B. Radhakrishnan and N. Mukunda. Spacelike representations of the inhomogeneous Lorentz group in a Lorentz basis. Journal of Mathematical Physics, 15(4):477–490, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i4/p477_s1.

Yang:1974:RQS

[1504] Janilla Jyy wen Yang, Russell Davidson, and John J. Kozak. On the relaxation to quantum-statistical equilibrium of the Wigner–Weisskopf atom in a one-dimensional radiation field. VI. Influence of the coupling function on the dynamics. Journal of Mathematical Physics, 15(4):491– 501, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i4/p491_s1.

Greenman:1974:ESM

[1505] J. V. Greenman. On the existence of solutions to S-matrix models. Journal of Mathematical Physics, 15(4):502–507, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i4/p502_s1.

Kutz:1974:ETR

[1506] Henry Douglas Kutz and John J. Kozak. Energy transfer in radiation chemistry. I. Dynamics of the electron-oscillators system. Journal of Mathematical Physics, 15(4):508–519, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i4/p508_s1.

Ellis:1974:EEF

[1507] Homer G. Ellis. Errata: “Ether flow through a drainhole: A particle model in general relativity”. Journal of Mathematical Physics, 15(4): 520, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i4/p520_s1. See [1114].

Glasser:1974:EEL

[1508] M. Lawrence Glasser. Erratum: “The evaluation of lattice sums. II. Number theoretic approach”. Journal of Mathematical Physics, 15(4): 520, April 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i4/p520_s2. See [1210].

Graffi:1974:RPB

[1509] S. Graffi. Regular perturbations, Brillouin–Wigner expansion, and continued fractions. Journal of Mathematical Physics, 15(5):521–523, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i5/p521_s1.

Brissaud:1974:SLS

[1510] A. Brissaud and U. Frisch. Solving linear stochastic differential equations. Journal of Mathematical Physics, 15(5):524–534, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p524_s1.

Trim:1974:NAS

[1511] D. W. Trim and J. Wainwright. Nonradiative algebraically special space– times. Journal of Mathematical Physics, 15(5):535–546, May 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p535_
s1.

Chela-Flores:1974:APS

[1512] J. Chela-Flores. An action principle for superconductivity. Journal of Mathematical Physics, 15(5):547–551, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i5/p547_s1.

Benioff:1974:SCS

[1513] Paul Benioff. Some consequences of the strengthened interpretative rules of quantum mechanics. Journal of Mathematical Physics, 15(5):552–559, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i5/p552_s1.

Boyer:1974:GCC

[1514] Charles P. Boyer and Kurt Bernardo Wolf. The group chain SO_{n,1} ⊃ SO_{1,1} ⊗ SO_{n−1}, a complete solution to the “missing label” problem. Journal of Mathematical Physics, 15(5):560–564, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p560_s1.

Klink:1974:EBW

[1515] William H. Klink. Expansions in Breit–Wigner amplitudes and biorthogonal functions. Journal of Mathematical Physics, 15(5):565–569, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i5/p565_s1.

Li:1974:SAC

[1516] Ming Chiang Li. Spheroidal analysis of Coulomb scattering. Journal of Mathematical Physics, 15(5):570–575, May 1974. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i5/p570_s1.

Dombrowski:1974:SPP

[1517] J. Dombrowski. Spectral properties of phase operators. Journal of Mathematical Physics, 15(5):576–577, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i5/p576_s1.

Kim:1974:BMA

[1518] Sungwoon Kim. Brownian motion in assemblies of coupled harmonic oscillators. Journal of Mathematical Physics, 15(5):578–582, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/
p578_s1.

Rajagopal:1974:NAE

[1519] A. K. Rajagopal and G. S. Grest. A novel approach to the exact calculation of correlation functions of a one-dimensional random Ising chain. Journal of Mathematical Physics, 15(5):583–588, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p583_s1.

Grest:1974:PST

[1520] G. S. Grest and A. K. Rajagopal. Perpendicular susceptibility of two one-dimensional Ising chains. Journal of Mathematical Physics, 15(5): 589–593, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i5/p589_s1.

Letelier:1974:GSE

[1521] P. S. Letelier and R. R. Tabensky. The general solution to Einstein– Maxwell equations with plane symmetry. Journal of Mathematical Physics, 15(5):594–595, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i5/p594_s1.

McDonald:1974:MDW

[1522] J. F. McDonald. On the moments of the distribution of widths corresponding to an ensemble of random matrices. Journal of Mathematical Physics, 15(5):596–601, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i5/p596_s1.

Doebner:1974:DGC

[1523] H. D. Doebner and J. Hennig. On dynamical groups: Classification of Lie algebras with Galilei subalgebras. Journal of Mathematical Physics, 15(5):602–608, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i5/p602_s1.

Mattis:1974:NWO

[1524] Daniel C. Mattis. New wave-operator identity applied to the study of persistent currents in 1D. Journal of Mathematical Physics, 15(5): 609–612, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i5/p609_s1.

Thiebaux:1974:AQL

[1525] M. L. Thiebaux. On the application of quasi-Lagrangian coordinates to random shear flows. Journal of Mathematical Physics, 15(5):613–618, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i5/p613_s1.

Sniatycki:1974:PC

[1526] Jedrzej Śniatycki. Prequantization of charge. Journal of Mathematical Physics, 15(5):619–620, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i5/p619_s1.

Hegerfeldt:1974:SOF

[1527] Gerhard C. Hegerfeldt. Smoothing operators for field domains. Journal of Mathematical Physics, 15(5):621–624, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i5/p621_s1.

Ardalan:1974:LMD

[1528] F. Ardalan. Limit of the most degenerate representations of SO(p, 1). Journal of Mathematical Physics, 15(5):625–630, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p625_s1.

Uslenghi:1974:HFB

[1529] P. L. E. Uslenghi and S. W. Lee. High-frequency backscattering from an elliptic metal plate. Journal of Mathematical Physics, 15(5):631–639, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i5/p631_s1.

Petru:1974:QCL

[1530] Zygmunt K. Petru. Quantum conservation laws for charged particle systems in magnetic fields. Journal of Mathematical Physics, 15(5):640–642, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i5/p640_s1.

Aizenman:1974:MDT

[1531] M. Aizenman and H. Gutfreund. Momentum distribution in the Tomonaga model at finite temperatures. Journal of Mathematical Physics, 15(5):643–647, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i5/p643_s1.

Semwal:1974:TCI

[1532] B. S. Semwal and P. K. Sharma. Thermal conductivity of an isotopically disordered harmonic crystal. Journal of Mathematical Physics, 15(5): 648–653, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i5/p648_s1.

Huszar:1974:DSL

[1533] M. Huszár. Deformation of the SO(2, C) subgroup of the Lorentz group. Journal of Mathematical Physics, 15(5):654–658, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p654_s1.

Horiguchi:1974:LGF

[1534] T. Horiguchi and C. C. Chen. Lattice Green’s function for the diced lattice. Journal of Mathematical Physics, 15(5):659–660, May 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p659_
s1.

Glass:1974:EEG

[1535] E. N. Glass and J. Winicour. Errata: “Elastic general relativistic systems”. Journal of Mathematical Physics, 15(5):661, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/p661_s1. See

[1090].

Glass:1974:EGG

[1536] E. N. Glass and J. Winicour. Errata: “A geometric generalization of Hooke’s law”. Journal of Mathematical Physics, 15(5):661, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i5/

p661_s2. See [1308].

Wang:1974:EAU

[1537] P. K. C. Wang. Erratum and addendum: “Unidirectional energy transfer in nonlinear wave-wave interactions”. Journal of Mathematical Physics, 15(5):662, May 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i5/p662_s1. See [1244].

Lubkin:1974:TMT

[1538] Elihu Lubkin. Theory of multibin tests: Definition and existence of extraneous tests. Journal of Mathematical Physics, 15(6):663–672, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i6/p663_s1.

Lubkin:1974:PSW

[1539] Elihu Lubkin. A physical system which can be forced to execute an arbitrary unitary transformation, and its use to perform arbitrary tests. Journal of Mathematical Physics, 15(6):673–674, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p673_s1.

Barston:1974:EPL

[1540] E. M. Barston. Eigenvalue problem for Lagrangian systems. VII. Journal of Mathematical Physics, 15(6):675–682, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i6/p675_s1.

Plybon:1974:CCL

[1541] Benjamin F. Plybon. Concerning conservation laws resulting from geometric invariance groups for field theories. Journal of Mathematical Physics, 15(6):683–686, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i6/p683_s1.

Wu:1974:EVM

[1542] F. Y. Wu. Eight-vertex model on the honeycomb lattice. Journal of Mathematical Physics, 15(6):687–691, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i6/p687_s1.

OConnor:1974:HCS

[1543] A. J. O’Connor and J. L. Lebowitz. Heat conduction and sound transmission in isotopically disordered harmonic crystals. Journal of Mathematical Physics, 15(6):692–703, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i6/p692_s1.

Inoue:1974:LGF

[1544] Michiko Inoue. Lattice Green’s function for the face centered cubic lattice. Journal of Mathematical Physics, 15(6):704–707, June 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p704_
s1.

Kuchar:1974:GRL

[1545] Karel Kuchar. Geometrodynamics regained: A Lagrangian approach. Journal of Mathematical Physics, 15(6):708–715, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p708_s1.

Hadinger:1974:CFE

[1546] G. Hadinger, N. Bessis, and G. Bessis. Closed-form expressions of matrix elements and eigenfunctions from ladder-operator considerations. Journal of Mathematical Physics, 15(6):716–726, June 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i6/p716_s1.

Adler:1974:FSG

[1547] Ronald J. Adler. A fluid sphere in general relativity. Journal of Mathematical Physics, 15(6):727–729, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i6/p727_s1. See erratum [1781].

Brydges:1974:SEA

[1548] David Brydges and Paul Federbush. A semi-Euclidean approach to boson-fermion model theories. Journal of Mathematical Physics, 15(6):

730–732, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i6/p730_s1.

Halpern:1974:PCQ

[1549] Francis R. Halpern. A pair of coupled quantum anharmonic oscillators. Journal of Mathematical Physics, 15(6):733–737, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p733_s1.

Ladde:1974:DIS

[1550] G. S. Ladde. Differential inequalities and stochastic functional differential equations. Journal of Mathematical Physics, 15(6):738–743, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i6/p738_s1.

Louat:1974:EKO

[1551] Norman Louat. The evaluation of ‘Kondo’ and other integrals of arbitrary range. Journal of Mathematical Physics, 15(6):744, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p744_s1.

Tortorella:1974:UES

[1552] Michael Tortorella. The unitarity equation for scattering in the absence of spherical symmetry. Journal of Mathematical Physics, 15(6): 745–747, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i6/p745_s1. See addendum [1614].

Moses:1974:PSM

[1553] H. E. Moses and A. F. Quesada. The power spectrum of the Mellin transformation with applications to scaling of physical quantities. Journal of Mathematical Physics, 15(6):748–752, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i6/p748_s1.

Kyriakopoulos:1974:GFH

[1554] E. Kyriakopoulos. Generating functions of the hypergeometric functions. Journal of Mathematical Physics, 15(6):753–759, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p753_s1.

Klauder:1974:OET

[1555] J. R. Klauder and L. Streit. Optical equivalence theorem for unbounded observables. Journal of Mathematical Physics, 15(6):760–763, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/
p760_s1.

Grgin:1974:DOG

[1556] Emile Grgin and Aage Petersen. Duality of observables and generators in classical and quantum mechanics. Journal of Mathematical Physics, 15(6):764–769, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i6/p764_s1.

Krause:1974:CGF

[1557] Thomas O. Krause and Bernard Mulligan. Cancellation of the Green’s function in the generation of continuum bound states by nonlocal potentials. Journal of Mathematical Physics, 15(6):770–773, June 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p770_
s1.

Honig:1974:MCP

[1558] Eli Honig, Engelbert L. Schucking, and C. V. Vishveshwara. Motion of charged particles in homogeneous electromagnetic fields. Journal of Mathematical Physics, 15(6):774–781, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i6/p774_s1.

Hongoh:1974:EWC

[1559] M. Hongoh, R. T. Sharp, and D. E. Tilley. Explicit O(5) Wigner coefficients. Journal of Mathematical Physics, 15(6):782–788, June 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p782_
s1.

Goedecke:1974:GE

[1560] G. H. Goedecke. On global embedding. Journal of Mathematical Physics, 15(6):789–791, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i6/p789_s1.

Goedecke:1974:SET

[1561] G. H. Goedecke. On stress-energy tensors. Journal of Mathematical Physics, 15(6):792–794, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i6/p792_s1.

Capper:1974:HOD

[1562] D. M. Capper and G. Leibbrandt. Higher order diagrams in quantum gravity and the continuous dimension method. Journal of Mathematical Physics, 15(6):795–797, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i6/p795_s1.

Kershaw:1974:NMS

[1563] David S. Kershaw. A new method for the study of Feynman perturbation series. Journal of Mathematical Physics, 15(6):798–807, June 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p798_
s1.

Cheng:1974:ELP

[1564] Kuo-Shung Cheng. Equivalent Lagrangians and path integration for generalized mechanics. Journal of Mathematical Physics, 15(6):808–811, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i6/p808_s1.

Ryan:1974:HCD

[1565] Michael P. Ryan, Jr. Hamiltonian cosmology: Death and transfiguration. Journal of Mathematical Physics, 15(6):812–815, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p812_s1.

Hauser:1974:SSS

[1566] I. Hauser and R. J. Malhiot. Spherically symmetric static space–times which admit stationary Killing tensors of rank two. Journal of Mathematical Physics, 15(6):816–823, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i6/p816_s1.

Constantinescu:1974:CNF

[1567] F. Constantinescu and J. G. Taylor. Causality and nonlocalizable fields. Journal of Mathematical Physics, 15(6):824–830, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p824_s1.

Fan:1974:SPLa

[1568] W. Fan and Y. M. Chen. Scattering of a plane longitudinal elastic wave by a large convex rigid object with a statistically corrugated surface. I. Perturbation solutions. Journal of Mathematical Physics, 15(6): 831–841, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i6/p831_s1.

Cornette:1974:MQS

[1569] William M. Cornette and Stanley P. Gudder. The mixture of quantum states. Journal of Mathematical Physics, 15(6):842–850, June 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p842_
s1.

Altschul:1974:RCT

[1570] Martin S. Altschul. Representations of the coordinate transformation group. Journal of Mathematical Physics, 15(6):851–855, June 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/p851_
s1.

Raychaudhuri:1974:TG

[1571] A. K. Raychaudhuri. Tachyons and gravitation. Journal of Mathematical Physics, 15(6):856, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i6/p856_s1.

Flato:1974:RLL

[1572] Moshé Flato and Håkan Snellman. Regge lattice and Lie-algebra representations. Journal of Mathematical Physics, 15(6):857–860, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/
p857_s1.

Gidas:1974:PIH

[1573] Basilis Gidas. Properties of the (: ϕ^{4}: )_{1+1} interaction Hamiltonian. Journal of Mathematical Physics, 15(6):861–866, June 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i6/p861_s1.

Gidas:1974:SAL

[1574] Basilis Gidas. On the self-adjointness of the Lorentz generator for (: ϕ^{4}: )_{1+1}. Journal of Mathematical Physics, 15(6):867–869, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i6/
p867_s1.

Marcuvitz:1974:TPT

[1575] N. Marcuvitz. On the theory of plasma turbulence. Journal of Mathematical Physics, 15(6):870–879, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i6/p870_s1.

Ullah:1974:DLE

[1576] Nazakat Ullah. Distribution of the lowest eigenvalue for a class of random matrices. Journal of Mathematical Physics, 15(6):880–883, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i6/p880_s1.

Persides:1974:SWE

[1577] S. Persides. Scalar waves in the exterior of a Schwarzschild black hole. Journal of Mathematical Physics, 15(7):885–891, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p885_s1.

Mika:1974:SPM

[1578] Janusz Mika. Singular perturbation method in neutron transport theory. Journal of Mathematical Physics, 15(7):892–898, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p892_s1.

Mathews:1974:MZI

[1579] P. M. Mathews, M. Seetharaman, and M. T. Simon. On mass zero indecomposable representations of the Poincaré group. Journal of Mathematical Physics, 15(7):899–901, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i7/p899_s1.

Melsheimer:1974:RHSa

[1580] Olaf Melsheimer. Rigged Hilbert space formalism as an extended mathematical formalism for quantum systems. I. General theory. Journal of Mathematical Physics, 15(7):902–916, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p902_s1.

Melsheimer:1974:RHSb

[1581] O. Melsheimer. Rigged Hilbert space formalism as an extended mathematical formalism for quantum systems. II. Transformation theory in non-relativistic quantum mechanics. Journal of Mathematical Physics, 15(7):917–925, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i7/p917_s1.

McCormick:1974:IPR

[1582] N. J. McCormick and I. Kuscer. On the inverse problem in radiative transfer. Journal of Mathematical Physics, 15(7):926–927, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/
p926_s1.

Ludwig:1974:GNFa

[1583] Garry Ludwig. On the geometrization of neutrino fields. I. Journal of Mathematical Physics, 15(7):928–932, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p928_s1.

Ludwig:1974:GNFb

[1584] Garry Ludwig. On the geometrization of neutrino fields. II. Journal of Mathematical Physics, 15(7):933–937, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p933_s1.

Wu:1974:AIOa

[1585] S. Y. Wu, C. C. Tung, and M. Schwartz. Applications of infinite order perturbation theory in linear systems. I. Journal of Mathematical Physics, 15(7):938–946, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i7/p938_s1.

Wu:1974:AIOb

[1586] Shi yu Wu. Application of infinite order perturbation theory in linear systems. II. The frequency spectrum of disordered chains. Journal of Mathematical Physics, 15(7):947–949, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i7/p947_s1. See erratum [1774].

Fan:1974:SPLb

[1587] W. Fan and Y. M. Chen. Scattering of plane longitudinal elastic wave by a large convex rigid object with a statistically corrugated surface. II. Far field solution. Journal of Mathematical Physics, 15(7):950–953, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i7/p950_s1.

Lai:1974:TMF

[1588] C. K. Lai. Thermodynamics of a mixture of fermions and bosons in one dimension with a repulsive delta-function potential. Journal of Mathematical Physics, 15(7):954–956, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i7/p954_s1.

Bose:1974:ESK

[1589] S. K. Bose and M. Y. Wang. Elastic scattering in the Kerr metric. Journal of Mathematical Physics, 15(7):957–962, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p957_s1.

Okubo:1974:TSR

[1590] Susumu Okubo. Theory of self-reproducing kernel and dispersion inequalities. Journal of Mathematical Physics, 15(7):963–973, July 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p963_
s1.

Case:1974:STO

[1591] K. M. Case. Scattering theory, orthogonal polynomials, and the transport equation. Journal of Mathematical Physics, 15(7):974–983, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i7/p974_s1.

Runnels:1974:IMD

[1592] L. K. Runnels. Ising models derived from binary lattice gases. Journal of Mathematical Physics, 15(7):984–987, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p984_s1.

Ranger:1974:SPS

[1593] K. B. Ranger. A scattering problem for a straight line segment. Journal of Mathematical Physics, 15(7):988–991, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p988_s1.

Debney:1974:EFE

[1594] George Debney. Expansion-free electromagnetic solutions of the Kerr– Schild class. Journal of Mathematical Physics, 15(7):992–997, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/
p992_s1.

Monroe:1974:SPC

[1595] James L. Monroe. A simple proof of certain FKG inequalities. Journal of Mathematical Physics, 15(7):998–999, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p998_s1.

Barut:1974:UEC

[1596] A. O. Barut and G. L. Bornzin. Unification of the external conformal symmetry group and the internal conformal dynamical group. Journal of Mathematical Physics, 15(7):1000–1006, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p1000_s1.

Boyer:1974:QFT

[1597] C. P. Boyer and G. N. Fleming. Quantum field theory on a sevendimensional homogeneous space of the Poincaré group. Journal of Mathematical Physics, 15(7):1007–1024, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i7/p1007_s1.

Kalnins:1974:LTSa

[1598] E. G. Kalnins and W. Miller, Jr. Lie theory and separation of variables. 3. The equation f_{tt} −f_{ss} = γ^{2}f. Journal of Mathematical Physics, 15(7): 1025–1032, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i7/p1025_s1. See erratum [1780].

Kunzle:1974:HDR

[1599] H. P. Künzle. Hamiltonian description of relativistically interacting twoparticle systems. Journal of Mathematical Physics, 15(7):1033–1044, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i7/p1033_s1.

Lev:1974:ESS

[1600] A. Paliov (Lev) and S. Rosendorff. Exponentially small scattering amplitude in high energy potential scattering. II. Journal of Mathematical Physics, 15(7):1045–1053, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i7/p1045_s1.

Paliov:1974:ESS

[1601] A. Paliov and S. Rosendorff. Exponentially small scattering amplitude in high energy potential scattering. II. Journal of Mathematical Physics, 15 (7):1045–1053, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Snellman:1974:PAV

[1602] Håkan Snellman. Poincaré analytic vectors in axiomatic quantum field theory. Journal of Mathematical Physics, 15(7):1054–1059, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/
p1054_s1.

Falkowski:1974:FID

[1603] B. J. Falkowski. Factorizable and infinitely divisible PUA representations of locally compact groups. Journal of Mathematical Physics, 15(7):1060– 1066, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i7/p1060_s1.

Hudson:1974:NSC

[1604] R. L. Hudson. A new system, of Casimir operators for U(n). Journal of Mathematical Physics, 15(7):1067–1070, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i7/p1067_s1.

DeFacio:1974:RKP

[1605] B. DeFacio and C. L. Hammer. Remarks on the Klauder phenomenon. Journal of Mathematical Physics, 15(7):1071–1077, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p1071_s1.

Ridener:1974:PES

[1606] F. L. Ridener, Jr. and R. H. Good, Jr. Polarization of electrons in screened Coulomb scattering. Journal of Mathematical Physics, 15(7): 1078–1084, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i7/p1078_s1.

Yeh:1974:RDG

[1607] Hsin-Yang Yeh. Representation and differential geometry of the semi simple Lie groups. Journal of Mathematical Physics, 15(7):1085–1095, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i7/p1085_s1.

Kloster:1974:SGF

[1608] S. Kloster, M. M. Som, and A. Das. On the stationary gravitational fields. Journal of Mathematical Physics, 15(7):1096–1102, July 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p1096_

s1. See erratum [1782].

Lind:1974:CAS

[1609] Robert W. Lind and Ezra T. Newman. Complexification of the algebraically special gravitational fields. Journal of Mathematical Physics, 15(7):1103–1112, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i7/p1103_s1.

Newman:1974:CCA

[1610] Ezra T. Newman and Jeffrey Winicour. A curiosity concerning angular momentum. Journal of Mathematical Physics, 15(7):1113–1115, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i7/p1113_s1.

Anderson:1974:RGP

[1611] James L. Anderson. Relativistic grad polynomials. Journal of Mathematical Physics, 15(7):1116–1119, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i7/p1116_s1.

Luehr:1974:SCG

[1612] C. P. Luehr and M. Rosenbaum. Spinor connections in general relativity. Journal of Mathematical Physics, 15(7):1120–1137, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i7/p1120_s1.

Menikoff:1974:HGF

[1613] Ralph Menikoff. The Hamiltonian and generating functional for a nonrelativistic local current algebra. Journal of Mathematical Physics, 15(7): 1138–1152, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i7/p1138_s1.

Tortorella:1974:AUE

[1614] Michael Tortorella. Addendum: “The unitarity equation for scattering in the absence of spherical symmetry”. Journal of Mathematical Physics, 15(7):1153, July 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i7/p1153_s1. See [1552].

Cattaneo:1974:MMC

[1615] U. Cattaneo and A. Janner. Mackey–Moore cohomology and topological extensions of Polish groups. Journal of Mathematical Physics, 15(8):1155–1165, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i8/p1155_s1.

Cattaneo:1974:CST

[1616] U. Cattaneo and A. Janner. Coupling of space–time and electromagnetic gauge transformations. Journal of Mathematical Physics, 15(8):1166– 1173, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i8/p1166_s1.

Hioe:1974:CSL

[1617] F. T. Hioe. Coherent states and Lie algebras. Journal of Mathematical Physics, 15(8):1174–1177, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i8/p1174_s1.

Shamaly:1974:IBR

[1618] A. Shamaly. Invariants of Born reciprocity theory. Journal of Mathematical Physics, 15(8):1178–1180, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i8/p1178_s1.

Klein:1974:BDF

[1619] William Klein. Behavior of distribution functions in the thermodynamic limit. Journal of Mathematical Physics, 15(8):1181–1185, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/
p1181_s1.

Wulf:1974:PZR

[1620] A. Wulf. Proof of Zwanzig’s rule of “planar” graphs. Journal of Mathematical Physics, 15(8):1186–1189, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i8/p1186_s1.

Eardley:1974:NST

[1621] Douglas M. Eardley. Note on space–times that admit constant electromagnetic fields. Journal of Mathematical Physics, 15(8):1190–1191, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i8/p1190_s1.

McQuistan:1974:ENNa

[1622] Richmond B. McQuistan. Exact nearest neighbor degeneracy for dumbbells on a one-dimensional lattice space. Journal of Mathematical Physics, 15(8):1192–1196, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i8/p1192_s1.

Strand:1974:GPT

[1623] Alf E. Strand. Gel’fand-pattern technique applied to the physically important SU(3) ⊗ SU(2) ⊂ SU(6) decomposition of SU(6). Journal of Mathematical Physics, 15(8):1197–1201, August 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1197_s1.

Zirilli:1974:SOO

[1624] F. Zirilli. Some observations on the operator H = −(1/2)d^{2}/dx^{2} + m^{2}x^{2}/2 + g/x^{2}. Journal of Mathematical Physics, 15(8):1202, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i8/p1202_s1.

Kasperkovitz:1974:ITS

[1625] P. Kasperkovitz and R. Dirl. Irreducible tensorial sets within the group algebra of a compact group. Journal of Mathematical Physics, 15(8):1203–1210, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i8/p1203_s1.

Yao:1974:ANPa

[1626] Te Hai Yao. On analytic nonlocal potentials. II. Analyticity of the S matrix, for fixed l, its representations, and a dispersion relation for fixed t. Journal of Mathematical Physics, 15(8):1211–1224, August 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1211_
s1.

Yao:1974:ANPb

[1627] Te Hai Yao. On analytic nonlocal potentials. III. Local correlations. Journal of Mathematical Physics, 15(8):1225–1226, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1225_s1.

Yao:1974:OIP

[1628] Te Hai Yao. The T operator and an inverse problem for nonlocal potentials. Journal of Mathematical Physics, 15(8):1227–1231, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v15/i8/p1227_s1. See erratum [1778].

Yao:1974:RJS

[1629] Te Hai Yao. Representations of the Jost solutions and the S matrix for a class of analytic nonlocal potentials. Journal of Mathematical Physics, 15(8):1232–1234, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i8/p1232_s1.

Mahanta:1974:AIS

[1630] M. N. Mahanta and D. R. K. Reddy. An approximate interior solution in Brans–Dicke theory. Journal of Mathematical Physics, 15(8):1235– 1237, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i8/p1235_s1.

Resibois:1974:TPV

[1631] P. Résibois, Y. Pomeau, and J. Piasecki. On the transport properties of van der Waals fluids. II. Explicit calculations. Journal of Mathematical Physics, 15(8):1238–1254, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i8/p1238_s1.

Raghavacharyulu:1974:PAI

[1632] I. V. V. Raghavacharyulu. Polynomial algebras. II. Commutation relations. Journal of Mathematical Physics, 15(8):1255–1262, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/
p1255_s1.

Kalnins:1974:LTSb

[1633] E. G. Kalnins and Willard Miller, Jr. Lie theory and separation of variables. 4. The groups SO(2, 1) and SO(3). Journal of Mathematical Physics, 15(8):1263–1274, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i8/p1263_s1.

Bellissard:1974:CCS

[1634] J. Bellissard and R. Holtz. Composition of coherent spin states. Journal of Mathematical Physics, 15(8):1275–1276, August 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1275_s1.

Raychaudhuri:1974:NEF

[1635] A. K. Raychaudhuri and A. K. Dutta. Null electromagnetic fields in spaces of high symmetry. Journal of Mathematical Physics, 15(8):1277– 1280, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i8/p1277_s1.

Lazarides:1974:SBV

[1636] G. Lazarides and A. A. Patani. Superpropagators as boundary values of analytic distributions. Journal of Mathematical Physics, 15(8):1281– 1284, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i8/p1281_s1.

Huss:1974:BVP

[1637] R. Huss, R. Kalaba, and R. Vasudevan. On a boundary value problem for integro-differential equations. Journal of Mathematical Physics, 15(8):1285–1287, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i8/p1285_s1.

Flato:1974:GCP

[1638] M. Flato and A. Piard. Generalized covariance principles and neutrino physics. Journal of Mathematical Physics, 15(8):1288–1294, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i8/p1288_s1.

Wolf:1974:CTC

[1639] Kurt Bernardo Wolf. Canonical transforms. I. Complex linear transforms. Journal of Mathematical Physics, 15(8):1295–1301, August 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1295_
s1.

Budic:1974:CBG

[1640] R. Budic and R. K. Sachs. Causal boundaries for general relativistic space–times. Journal of Mathematical Physics, 15(8):1302–1309, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i8/p1302_s1.

Hakim:1974:ESM

[1641] Rémi Hakim. Equilibrium statistical mechanics of relativistic particles with variable masses. I. Nonquantal theory. Journal of Mathematical Physics, 15(8):1310–1319, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i8/p1310_s1.

Mukunda:1974:CGPa

[1642] N. Mukunda and B. Radhakrishnan. Clebsch–Gordan problem and coefficients for the three-dimensional Lorentz group in a continuous basis. I. Journal of Mathematical Physics, 15(8):1320–1331, August 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1320_
s1.

Mukunda:1974:CGPb

[1643] N. Mukunda and B. Radhakrishnan. Clebsch–Gordan problem and coefficients for the three-dimensional Lorentz group in a continuous basis. II. Journal of Mathematical Physics, 15(8):1332–1342, August 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1332_
s1.

Wolfe:1974:AFCa

[1644] John C. Wolfe and Gérard G. Emch. C∗-algebraic formalism for coarse graining. I. General theory. Journal of Mathematical Physics, 15(8):1343– 1347, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i8/p1343_s1.

Wolfe:1974:AFCb

[1645] John C. Wolfe and Gérard G. Emch. C∗-algebraic formalism for coarse graining. II. Momentum coarse graining for Fermi systems in finite volume. Journal of Mathematical Physics, 15(8):1348–1350, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i8/
p1348_s1.

Wolfe:1974:AFCc

[1646] John C. Wolfe and Gérard G. Emch. C∗-algebraic formalism for coarse graining. III. Momentum coarse graining for Fermi systems in infinite free space. Journal of Mathematical Physics, 15(8):1351–1365, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i8/p1351_s1.

Gibson:1974:TIM

[1647] A. G. Gibson and Colston Chandler. Time independent multichannel scattering theory for charged particles. Journal of Mathematical Physics, 15(8):1366–1377, August 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i8/p1366_s1.

Patera:1974:MSSa

[1648] J. Patera, P. Winternitz, and H. Zassenhaus. The maximal solvable subgroups of the SU(p, q) groups and all subgroups of SU(2, 1). Journal of Mathematical Physics, 15(8):1378–1393, August 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1378_s1.

Menikoff:1974:GFD

[1649] Ralph Menikoff. Generating functionals determining representations of a non relativistic local current algebra in the N/V limit. Journal of Mathematical Physics, 15(8):1394–1408, August 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i8/p1394_s1.

Ernst:1974:CPF

[1650] Frederick J. Ernst. Complex potential formulation of the axially symmetric gravitational field problem. Journal of Mathematical Physics, 15 (9):1409–1412, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i9/p1409_s1. See erratum [1777].

Lerner:1974:WGF

[1651] D. Lerner and J. R. Porter. Weak gravitational fields. Journal of Mathematical Physics, 15(9):1413–1415, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i9/p1413_s1.

Lerner:1974:ASS

[1652] D. Lerner and J. R. Porter. Asymptotically simple space–time manifolds. Journal of Mathematical Physics, 15(9):1416–1419, September 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1416_
s1.

Wolfes:1974:TBL

[1653] J. Wolfes. On the three-body linear problem with three-body interaction. Journal of Mathematical Physics, 15(9):1420–1424, September 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1420_
s1.

Calogero:1974:ESO

[1654] F. Calogero and C. Marchioro. Exact solution of a one-dimensional three-body scattering problem with two-body and/or three-body inverse square potentials. Journal of Mathematical Physics, 15(9):1425–1430, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i9/p1425_s1.

Lee:1974:WPR

[1655] L. C. Lee. Wave propagation in a random medium: A complete set of the moment equations with different wavenumbers. Journal of Mathematical Physics, 15(9):1431–1435, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i9/p1431_s1.

Burdet:1974:ARL

[1656] G. Burdet, M. Perrin, and P. Sorba. On the automorphisms of real Lie algebras. Journal of Mathematical Physics, 15(9):1436–1442, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i9/p1436_s1.

Frederiksen:1974:SRP

[1657] J. S. Frederiksen. Spectral representation of the pentagon diagram amplitude. Journal of Mathematical Physics, 15(9):1443–1450, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i9/p1443_s1.

Buhring:1974:SEI

[1658] Wolfgang Bühring. Schrödinger equation with inverse fourth-power potential, a differential equation with two irregular singular points. Journal of Mathematical Physics, 15(9):1451–1459, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1451_s1.

Katzin:1974:DSC

[1659] Gerald H. Katzin and Jack Levine. Dynamical symmetries and constants of the motion for classical particle systems. Journal of Mathematical Physics, 15(9):1460–1470, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i9/p1460_s1.

Hirondel:1974:SRG

[1660] D. Hirondel. Special relativity in general relativity? Journal of Mathematical Physics, 15(9):1471–1473, September 1974. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i9/p1471_s1.

Chang:1974:SSF

[1661] M. H. Chang, G. Ladde, and P. T. Liu. Stability of stochastic functional differential equations. Journal of Mathematical Physics, 15(9):1474– 1478, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i9/p1474_s1.

Fox:1974:ACT

[1662] Ronald Forrest Fox. Application of cumulant techniques to multiplicative stochastic processes. Journal of Mathematical Physics, 15(9):1479–1483, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i9/p1479_s1.

Boyer:1974:CSO

[1663] Charles P. Boyer and Willard Miller, Jr. A classification of second-order raising operators for Hamiltonians in two variables. Journal of Mathematical Physics, 15(9):1484–1489, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i9/p1484_s1.

Huang:1974:SCB

[1664] Cheng-Sheng Huang and Alfred C. T. Wu. Structure of the 12-j and 15-j coefficients in the Bargmann approach. Journal of Mathematical Physics, 15(9):1490–1493, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i9/p1490_s1.

Maugin:1974:VFR

[1665] Gérard A. Maugin and A. Cemal Eringen. Variational formulation of the relativistic theory of microelectromagnetism. Journal of Mathematical Physics, 15(9):1494–1499, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i9/p1494_s1.

Katsura:1974:LGF

[1666] Shigetoshi Katsura and Hiroshi Watanabe. Lattice Green’s function of the body-centered cubic lattice at arbitrary points. Journal of Mathematical Physics, 15(9):1500–1504, September 1974. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i9/p1500_s1.

Zelazny:1974:MRP

[1667] R. Zelazny and T. Duracz. Multigroup replication property for external, spherically symmetric problems of transport theory. Journal of Mathematical Physics, 15(9):1505–1507, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i9/p1505_s1.

Nedelkov:1974:EUS

[1668] I. P. Nedelkov. Existence and uniqueness of solutions to Low’s problem. Journal of Mathematical Physics, 15(9):1508–1515, September 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1508_
s1.

Brown:1974:PTP

[1669] William Fuller Brown, Jr. Properties of two-phase “cell materials”. Journal of Mathematical Physics, 15(9):1516–1524, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1516_s1.

Mennessier:1974:SUB

[1670] G. Mennessier and J. Nuyts. Some unitarity bounds for finite matrices. Journal of Mathematical Physics, 15(9):1525–1537, September 1974. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1525_
s1.

Cicogna:1974:SBC

[1671] G. Cicogna and P. de Mottoni. On the spontaneous breakdown of compound symmetries. Journal of Mathematical Physics, 15(9):1538–1542, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i9/p1538_s1.

DAdda:1974:SEC

[1672] A. D’Adda, R. D’Auria, and G. Ponzano. Symmetries of extended 3-j coefficients. Journal of Mathematical Physics, 15(9):1543–1553, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i9/p1543_s1.

Hajicek:1974:COF

[1673] Petr Hájícek. Can outside fields destroy black holes? Journal of Mathematical Physics, 15(9):1554–1558, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i9/p1554_s1.

Brans:1974:CSR

[1674] Carl H. Brans. Complex structures and representations of the Einstein equations. Journal of Mathematical Physics, 15(9):1559–1566, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i9/p1559_s1.

Fulling:1974:ATS

[1675] S. A. Fulling. Absence of trivial subrepresentations from tensor products of unitary representations of pseudo-orthogonal groups. Journal of Mathematical Physics, 15(9):1567–1570, September 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1567_s1.

Lee:1974:SMA

[1676] Jon Lee. Statistical mechanical approaches to fluid turbulence. Journal of Mathematical Physics, 15(9):1571–1586, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1571_s1.

Okuniewicz:1974:QPS

[1677] Joyce M. Okuniewicz. Quasiperiodic pointwise solutions of the periodic, time-dependent Schrödinger equation. Journal of Mathematical Physics, 15(9):1587–1595, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i9/p1587_s1.

Hill:1974:ECS

[1678] Robert Nyden Hill. Elimination of a coordinate singularity in the three-body problem. Journal of Mathematical Physics, 15(9):1596–1598, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i9/p1596_s1.

Brink:1974:SHS

[1679] L. Brink, P. H. Frampton, and H. B. Nielsen. A study of the Hilbert space properties of the Veneziano model operator formalism. Journal of Mathematical Physics, 15(9):1599–1605, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i9/p1599_s1.

Marin:1974:EST

[1680] Lennart Marin. Electromagnetic scattering from two circular, coaxial disks. Journal of Mathematical Physics, 15(9):1606–1613, September 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i9/p1606_s1.

Hioe:1974:RPM

[1681] F. T. Hioe. The recursive perturbation method and its application to the study of nuclear shell models. Journal of Mathematical Physics, 15(10):1615–1619, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1615_s1.

Presutti:1974:LBP

[1682] E. Presutti and E. Scacciatelli. Local bounded perturbations of KMS states. Journal of Mathematical Physics, 15(10):1620–1625, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i10/p1620_s1.

Narcowich:1974:MTMa

[1683] F. J. Narcowich. Mathematical theory of the R matrix. I. The eigenvalue problem. Journal of Mathematical Physics, 15(10):1626–1634, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i10/p1626_s1.

Narcowich:1974:MTMb

[1684] Francis J. Narcowich. The mathematical theory of the R-matrix. II. The R-matrix and its properties. Journal of Mathematical Physics, 15 (10):1635–1642, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1635_s1.

Mukunda:1974:CGPc

[1685] N. Mukunda and B. Radhakrishnan. The Clebsch–Gordan problem and coefficients for the three-dimensional Lorentz group in a continuous basis. III. Journal of Mathematical Physics, 15(10):1643–1655, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i10/p1643_s1.

Mukunda:1974:CGPd

[1686] N. Mukunda and B. Radhakrishnan. The Clebsch–Gordan problem and coefficients for the three-dimensional Lorentz group in a continuous basis. IV. Journal of Mathematical Physics, 15(10):1656–1668, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i10/p1656_s1.

Sheaks:1974:MCG

[1687] O. J. Sheaks and T. W. Schnatz. Multiregion criticality in general geometries. Journal of Mathematical Physics, 15(10):1669–1674, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i10/p1669_s1.

Lai:1974:LGN

[1688] C. K. Lai. Lattice gas with nearest neighbor interaction in one dimension with arbitrary statistics. Journal of Mathematical Physics, 15(10):1675– 1676, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i10/p1675_s1.

Bernstein:1974:MQT

[1689] Herbert J. Bernstein. Must quantum theory assume unrestricted superposition? Journal of Mathematical Physics, 15(10):1677–1679, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i10/p1677_s1.

Truman:1974:SCD

[1690] A. Truman. Spectrality, cluster decomposition, and small distance properties in Wightman field theory. Journal of Mathematical Physics, 15 (10):1680–1685, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1680_s1.

Bacry:1974:ORG

[1691] H. Bacry. Orbits of the rotation group on spin states. Journal of Mathematical Physics, 15(10):1686–1688, October 1974. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i10/p1686_s1.

Hirondel:1974:FPL

[1692] D. Hirondel. Free-particle-like formulation of Newtonian instantaneous action-at-a-distance electrodynamics. Journal of Mathematical Physics, 15(10):1689–1697, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1689_s1.

Finley:1974:RCP

[1693] J. D. Finley III. Radial charged particle trajectories in the extended Reissner–Nordstrom manifold. Journal of Mathematical Physics, 15 (10):1698–1701, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1698_s1.

Sakata:1974:GMOa

[1694] Isao Sakata. A general method for obtaining Clebsch–Gordan coefficients of finite groups. I. Its application to point and space groups. Journal of Mathematical Physics, 15(10):1702–1709, October 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i10/p1702_s1.

Sakata:1974:GMOb

[1695] Isao Sakata. A general method for obtaining Clebsch–Gordan coefficients of finite groups. II. Extension to antiunitary groups. Journal of Mathematical Physics, 15(10):1710–1711, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i10/p1710_s1.

Runnels:1974:PTM

[1696] L. K. Runnels and J. L. Lebowitz. Phase transitions of a multicomponent Widom–Rowlinson model. Journal of Mathematical Physics, 15 (10):1712–1717, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1712_s1.

Halpern:1974:MMA

[1697] Francis R. Halpern and Thomas W. Yonkman. Matrix mechanics approach to a nonlinear oscillator. Journal of Mathematical Physics, 15 (10):1718–1721, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1718_s1.

Hautot:1974:NME

[1698] André Hautot. A new method for the evaluation of slowly convergent series. Journal of Mathematical Physics, 15(10):1722–1727, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i10/
p1722_s1.

Kalnins:1974:LTSc

[1699] E. G. Kalnins and W. Miller, Jr. Lie theory and separation of variables. 5. The equations iU_{t} + U_{xx} = 0 and iU_{t} + U_{xx} − c/x^{2}U = 0. Journal of Mathematical Physics, 15(10):1728–1737, October 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i10/p1728_s1.

Lavine:1974:DZJ

[1700] J. P. Lavine. The distribution of the zeroes of the Jost function: The swave attractive exponential potential. Journal of Mathematical Physics, 15(10):1738–1739, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1738_s1.

LoDato:1974:RPO

[1701] Vincent A. LoDato. The renormalized projection operator technique for quadratic stochastic differential equations. II. Journal of Mathematical Physics, 15(10):1740–1744, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i10/p1740_s1.

Albeverio:1974:RCB

[1702] Sergio Albeverio and Raphael Høegh-Krohn. A remark on the connection between stochastic mechanics and the heat equation. Journal of Mathematical Physics, 15(10):1745–1747, October 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i10/p1745_s1.

Hu:1974:STE

[1703] B. L. Hu. Separation of tensor equations in a homogeneous space by group theoretical methods. Journal of Mathematical Physics, 15 (10):1748–1755, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1748_s1.

Teixeira:1974:GSE

[1704] A. F. da-F. da F. Teixeira, Idel Wolk, and M. M. Som. Generalized static electromagnetic fields in relativity. Journal of Mathematical Physics, 15(10):1756–1759, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1756_s1.

Roman:1974:GTG

[1705] P. Roman and J. P. Leveille. Gauge theories and Galilean symmetry. Journal of Mathematical Physics, 15(10):1760–1767, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i10/
p1760_s1.

Hestenes:1974:PPM

[1706] David Hestenes. Proper particle mechanics. Journal of Mathematical Physics, 15(10):1768–1777, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i10/p1768_s1.

Hestenes:1974:PDR

[1707] David Hestenes. Proper dynamics of a rigid point particle. Journal of Mathematical Physics, 15(10):1778–1786, October 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i10/p1778_s1.

Judd:1974:CSC

[1708] B. R. Judd, W. Miller, Jr., J. Patera, and P. Winternitz. Complete sets of commuting operators and O(3) scalars in the enveloping algebra of SU(3). Journal of Mathematical Physics, 15(10):1787–1799, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i10/p1787_s1.

Shafee:1974:STG

[1709] A. Shafee. Some topological and graphical aspects of phase contours. Journal of Mathematical Physics, 15(10):1800–1807, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i10/
p1800_s1.

Truex:1974:LSE

[1710] William L. Truex and Joseph L. Strecker. Limits in systems exhibiting a one-dimensional phase transition. Journal of Mathematical Physics, 15(10):1808–1818, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i10/p1808_s1.

Saenz:1974:ETM

[1711] A. W. Saénz and W. W. Zachary. Erratum: “Two-magnon bound states in Heisenberg ferromagnets”. Journal of Mathematical Physics, 15(10): 1819, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i10/p1819_s1. See [1385].

Millard:1974:ZPF

[1712] K. Y. Millard and K. S. Viswanathan. Zeros of the partition function for higher-order spin systems. Journal of Mathematical Physics, 15(11):1821–1825, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i11/p1821_s1.

Frederiksen:1974:DSR

[1713] J. S. Frederiksen. Double spectral representations of single loop amplitudes with k vertices: k ≥ 4. Journal of Mathematical Physics, 15 (11):1826–1834, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i11/p1826_s1.

Kobe:1974:PCD

[1714] Donald H. Kobe and Gerald W. Goble. Principle of compensation of dangerous diagrams for boson systems. III. Finite temperature. Journal of Mathematical Physics, 15(11):1835–1844, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/p1835_s1.

McQuistan:1974:ENNb

[1715] R. B. McQuistan. Exact next nearest neighbor degeneracy. Journal of Mathematical Physics, 15(11):1845–1848, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/p1845_s1.

King:1974:SEQ

[1716] Jean I. F. King. Schrödinger equation and quantum state codons in discrete transform space. Journal of Mathematical Physics, 15(11):1849– 1851, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i11/p1849_s1.

Ablowitz:1974:CPP

[1717] Mark J. Ablowitz, David J. Kaup, and Alan C. Newell. Coherent pulse propagation, a dispersive irreversible phenomenon. Journal of Mathematical Physics, 15(11):1852–1858, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i11/p1852_s1.

Smith:1974:FRP

[1718] J. Denmead Smith. Factorizability of resonance poles in multiparticle amplitudes. Journal of Mathematical Physics, 15(11):1859–1865, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i11/p1859_s1.

Ma:1974:RGL

[1719] Shang keng Ma. The renormalization group and the large n limit. Journal of Mathematical Physics, 15(11):1866–1891, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/p1866_s1.

diSessa:1974:QHF

[1720] A. di Sessa. Quantization on hyperboloids and full space–time field expansion. Journal of Mathematical Physics, 15(11):1892–1900, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i11/p1892_s1.

Beran:1974:PTA

[1721] M. J. Beran and J. J. McCoy. Propagation through an anisotropic random medium. Journal of Mathematical Physics, 15(11):1901–1912, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i11/p1901_s1.

Griffin:1974:PWE

[1722] P. Griffin, P. Nagel, and R. D. Koshel. The plane-wave expansion method. Journal of Mathematical Physics, 15(11):1913–1917, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i11/p1913_s1.

Fox:1974:MSP

[1723] Ronald Forrest Fox. Multiplicative stochastic processes, Fokker–Planck equations, and a possible dynamical mechanism for critical behavior. Journal of Mathematical Physics, 15(11):1918–1929, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/
p1918_s1.

Glass:1974:WTD

[1724] E. N. Glass. Weyl tensor decomposition in stationary vacuum space– times. Journal of Mathematical Physics, 15(11):1930–1931, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i11/p1930_s1.

Patera:1974:MSSb

[1725] J. Patera, P. Winternitz, and H. Zassenhaus. The maximal solvable subgroups of SO(p, q) groups. Journal of Mathematical Physics, 15 (11):1932–1938, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i11/p1932_s1.

Ueno:1974:IIF

[1726] S. Ueno. Invariant imbedding and Fredholm integral equations with degenerate kernels. Journal of Mathematical Physics, 15(11):1939–1941, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i11/p1939_s1.

Weyland:1974:KTL

[1727] A. Weyland. Kinetic theory and the Lorentz gas. Journal of Mathematical Physics, 15(11):1942–1943, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i11/p1942_s1.

Kashyap:1974:DCS

[1728] S. C. Kashyap. Diffraction characteristics of a slit formed by two staggered parallel planes. Journal of Mathematical Physics, 15(11):1944– 1949, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i11/p1944_s1.

Matsuno:1974:ABM

[1729] Koichiro Matsuno. Asymptotic behavior of Markoffian kinetics. Journal of Mathematical Physics, 15(11):1950–1953, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/p1950_s1.

Santos:1974:QFS

[1730] Emilio Santos. Quantumlike formulation of stochastic problems. Journal of Mathematical Physics, 15(11):1954–1962, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/p1954_s1.

Berezdivin:1974:ANC

[1731] R. Berezdivin. The analytic noncharacteristic Cauchy problem for nonlightlike isometries in vacuum space–times. Journal of Mathematical Physics, 15(11):1963–1966, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i11/p1963_s1.

Lerche:1974:CCV

[1732] I. Lerche. Concerning a criterion for the validity of the first order smoothing approximation. Journal of Mathematical Physics, 15(11):1967–1971, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i11/p1967_s1.

Tokuoka:1974:ITE

[1733] Tatsuo Tokuoka. Interaction and transformation of electromagnetic and gravitational waves in Einstein–Maxwell field. Journal of Mathematical Physics, 15(11):1972–1976, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i11/p1972_s1.

Bahar:1974:RWPa

[1734] E. Bahar. Radio wave propagation in nonuniform multilayered cylindrical structures-generalized field transforms. Journal of Mathematical Physics, 15(11):1977–1981, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i11/p1977_s1.

Bahar:1974:RWPb

[1735] E. Bahar. Radio wave propagation in nonuniform multilayered cylindrical structures-full wave solutions. Journal of Mathematical Physics, 15 (11):1982–1986, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i11/p1982_s1.

Larsen:1974:SLT

[1736] Edward W. Larsen and Paul F. Zweifel. On the spectrum of the linear transport operator. Journal of Mathematical Physics, 15(11):1987– 1997, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i11/p1987_s1.

DeVega:1974:DR

[1737] H. J. De Vega and F. A. Schaposnik. Dimensional renormalization. Journal of Mathematical Physics, 15(11):1998–2000, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/
p1998_s1.

Elsayed:1974:BET

[1738] Mostafa A. Elsayed. Bounds for effective thermal, electrical, and magnetic properties of heterogeneous materials using high order statistical information. Journal of Mathematical Physics, 15(11):2001–2015, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i11/p2001_s1.

Harter:1974:RMS

[1739] W. G. Harter. Representations of multidimensional symmetries in networks. Journal of Mathematical Physics, 15(11):2016–2021, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i11/p2016_s1.

Chen:1974:CSS

[1740] Etang Chen. Choquet simplexes of states in physical systems. Journal of Mathematical Physics, 15(11):2022–2025, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i11/p2022_s1.

Ernst:1974:WCT

[1741] Frederick J. Ernst. Weyl conform tensor for stationary gravitational fields. Journal of Mathematical Physics, 15(12):2027–2030, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2027_s1.

Rudra:1974:ICF

[1742] P. Rudra. On irreducible corepresentations of finite magnetic groups. Journal of Mathematical Physics, 15(12):2031–2035, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i12/
p2031_s1.

Davies:1974:DMW

[1743] E. B. Davies. Dynamics of a multilevel Wigner–Weisskopf atom. Journal of Mathematical Physics, 15(12):2036–2041, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i12/p2036_s1.

Rno:1974:CGC

[1744] Jung Sik Rno. Clebsch–Gordan coefficients and special functions related to the Euclidean group in three space. Journal of Mathematical Physics, 15(12):2042–2047, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i12/p2042_s1.

Troianiello:1974:STS

[1745] Giovanni M. Troianiello. Scattering theory for Schrödinger operators with L^{∞} potentials and distorted Bloch waves. Journal of Mathematical Physics, 15(12):2048–2052, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i12/p2048_s1.

Roman:1974:RQM

[1746] P. Roman and J. P. Leveille. Relativistic quantum mechanics and local gauge symmetry. Journal of Mathematical Physics, 15(12):2053–2062, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2053_s1.

Nayfeh:1974:ASS

[1747] Ali H. Nayfeh. Asymptotic solutions of second-order linear equations with three transition points. Journal of Mathematical Physics, 15 (12):2063–2066, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i12/p2063_s1.

Petti:1974:GIC

[1748] Richard Petti. A geometric interpretation of classical relativistic electrodynamics. Journal of Mathematical Physics, 15(12):2067–2070, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2067_s1.

Bessis:1974:SRP

[1749] D. Bessis, L. Epele, and M. Villani. Summation of regularized perturbative expansions for singular interactions. Journal of Mathematical Physics, 15(12):2071–2078, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i12/p2071_s1.

Provost:1974:PPS

[1750] J. P. Provost, F. Rocca, G. Vallee, and M. Sirugue. Phase properties of some photon states with nonzero energy density. Journal of Mathematical Physics, 15(12):2079–2085, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i12/p2079_s1.

Balazs:1974:SFS

[1751] N. L. Balazs and G. G. Zipfel, Jr. The semiclassical fermion μ-space density in three dimensions. Journal of Mathematical Physics, 15(12):2086– 2089, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2086_s1.

Gilmore:1974:BCH

[1752] R. Gilmore. Baker–Campbell–Hausdorff formulas. Journal of Mathematical Physics, 15(12):2090–2092, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i12/p2090_s1.

Capri:1974:EWR

[1753] Anton Z. Capri. On the existence of weakly retarded and advanced Green’s functions. Journal of Mathematical Physics, 15(12):2093–2095, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2093_s1. See erratum [1779].

Wolf:1974:AGD

[1754] Kurt Bernardo Wolf and Charles P. Boyer. The algebra and group deformations I^{m}[SO(n) ⊗ SO(m)] ⇒ SO(n, m), I^{m}[U(n) ⊗ U(m)] ⇒ U(n, m), and I^{m}[S_{p}(n) ⊗ S_{p}(m)] ⇒ S_{p}(n, m) for 1 ≤ m ≤ n. Journal of Mathematical Physics, 15(12):2096–2101, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i12/p2096_s1.

Wolf:1974:CTI

[1755] Kurt Bernardo Wolf. Canonical transforms. II. Complex radial transforms. Journal of Mathematical Physics, 15(12):2102–2111, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2102_s1.

Peskoff:1974:GFL

[1756] A. Peskoff. Green’s function for Laplace’s equation in an infinite cylindrical cell. Journal of Mathematical Physics, 15(12):2112–2120, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2112_s1.

Kinnersley:1974:LTS

[1757] William Kinnersley and Edward F. Kelley. Limits of the Tomimatsu– Sato gravitational field. Journal of Mathematical Physics, 15(12):2121– 2126, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2121_s1.

Phillipson:1974:ELR

[1758] Paul E. Phillipson. Effects of long range interactions in harmonically coupled systems. I. Equilibrium fluctuations and diffusion. Journal of Mathematical Physics, 15(12):2127–2147, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i12/p2127_s1.

Hecht:1974:SRR

[1759] K. T. Hecht. Summation relation for U(N) Racah coefficients. Journal of Mathematical Physics, 15(12):2148–2156, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v15/i12/p2148_s1.

Lamb:1974:BTC

[1760] G. L. Lamb, Jr. Bäcklund transformations for certain nonlinear evolution equations. Journal of Mathematical Physics, 15(12):2157–2165, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2157_s1.

Case:1974:OPV

[1761] K. M. Case. Orthogonal polynomials from the viewpoint of scattering theory. Journal of Mathematical Physics, 15(12):2166–2174, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2166_s1.

Mano:1974:RGF

[1762] Koichi Mano. A remark on the Green’s function for the face-centered cubic lattices. Journal of Mathematical Physics, 15(12):2175–2176, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2175_s1.

Hori:1974:STE

[1763] Motoo Hori and Fumiko Yonezawa. Statistical theory of effective electrical, thermal, and magnetic properties of random heterogeneous materials. III. Perturbation treatment of the effective permittivity in completely random heterogeneous materials. Journal of Mathematical Physics, 15 (12):2177–2185, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i12/p2177_s1.

Kohler:1974:PSW

[1764] W. Kohler and G. C. Papanicolaou. Power statistics for wave propagation in one-dimension and comparison with radiative transport theory. II. Journal of Mathematical Physics, 15(12):2186–2197, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i12/
p2186_s1.

Strocchi:1974:PCS

[1765] F. Strocchi and A. S. Wightman. Proof of the charge superselection rule in local relativistic quantum field theory. Journal of Mathematical Physics, 15(12):2198–2224, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v15/i12/p2198_s1. See erratum [1783].

RamaRao:1974:PTN

[1766] I. RamaRao and Franz Mohling. Perturbation theory of a nonideal Bose gas. II. Journal of Mathematical Physics, 15(12):2225–2228, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2225_s1.

Dunn:1974:STT

[1767] K. A. Dunn. A scalar-tensor theory of gravitation. Journal of Mathematical Physics, 15(12):2229–2231, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i12/p2229_s1.

Tariq:1974:UBR

[1768] N. Tariq and B. O. J. Tupper. The uniqueness of the Bertotti–Robinson electromagnetic universe. Journal of Mathematical Physics, 15(12):2232– 2235, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v15/i12/p2232_s1.

Marlow:1974:PA

[1769] A. R. Marlow. Physical axiomatics. Journal of Mathematical Physics, 15(12):2236–2241, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v15/i12/p2236_s1.

Klotz:1974:TCG

[1770] Frederick S. Klotz. Twistors and the conformal group. Journal of Mathematical Physics, 15(12):2242–2247, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i12/p2242_s1.

Dokmeci:1974:HOT

[1771] M. Cengiz Dökmeci. On the higher order theories of piezoelectric crystal surfaces. Journal of Mathematical Physics, 15(12):2248–2252, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v15/i12/p2248_s1.

Burdet:1974:ETD

[1772] G. Burdet, M. Perrin, and P. Sorba. On the explicit-time-dependent invariance properties of quantum mechanical systems. Journal of Mathematical Physics, 15(12):2253–2256, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v15/i12/p2253_s1.

Urbantke:1974:EKR

[1773] Helmuth K. Urbantke. Erratum: “KS-related f–g couples as exact vacuum solutions of Salam’s two-tensor theory”. Journal of Mathematical Physics, 15(12):2257, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v15/i12/p2257_s1. See [1441].

Wu:1975:EAI

[1774] Shi yu Wu. Erratum: “Application of infinite order perturbation theory in linear systems. II”. Journal of Mathematical Physics, 16(1):171, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p171_s1. See [1586].

Yao:1975:EAN

[1775] Te Hai Yao. Erratum: “On analytic nonlocal potentials. I. A forward dispersion relation”. Journal of Mathematical Physics, 16(1):171, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i1/p171_s2. See [1280].

Prugovecki:1975:EMS

[1776] E. Prugovečki. Erratum: “Multichannel stationary scattering theory in two-Hilbert space formulation”. Journal of Mathematical Physics, 16(2): 442, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i2/p442_s1. See [1250].

Ernst:1975:ECP

[1777] Frederick J. Ernst. Erratum: “Complex potential formulation of the axially symmetric gravitational field problem”. Journal of Mathematical Physics, 16(4):1011, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i4/p1011_s1. See [1650].

Yao:1975:EOI

[1778] Te Hai Yao. Erratum: “The T operator and an inverse problem for nonlocal potentials”. Journal of Mathematical Physics, 16(4):1011, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v16/i4/p1011_s2. See [1628].

Capri:1975:EEW

[1779] A. Z. Capri. Erratum: “On the existence of weakly retarded and advanced Green’s functions”. Journal of Mathematical Physics, 16(5): 1201, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i5/p1201_s1. See [1753].

Kalnins:1975:ELT

[1780] E. G. Kalnins and W. Miller. Erratum: “Lie theory and separation of variables. 3. The equation f_{tt} − f_{ss} = γ^{2}f”. Journal of Mathematical Physics, 16(7):1531, July 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i7/p1531_s1. See [1598].

Adler:1976:EFS

[1781] Ronald J. Adler. Erratum: “A fluid sphere in general relativity”. Journal of Mathematical Physics, 17(1):158, January 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i1/p158_s1. See [1547].

Kloster:1976:ESG

[1782] S. Kloster, M. M. Som, and A. Das. Erratum: “On the stationary gravitational fields”. Journal of Mathematical Physics, 17(3):441, March 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v17/i3/p441_s2. See [1608].

Strocchi:1976:EPC

[1783] F. Strocchi and A. S. Wightman. Erratum: “Proof of the charge superselection rule in local relativistic quantum field theory”. Journal of Mathematical Physics, 17(10):1930–1931, October 1976. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v17/i10/p1930_s1. See [1765].

Calogero:1996:ESO

[1784] F. Calogero. Erratum: “Solution of the one-dimensional N-body problems with quadratic and/or inversely quadratic pair potentials”, [J. Math. Phys. 12 (1971), 419–436; MR 43 #5824]. Journal of Mathematical Physics, 37(7):3646, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [490].

Kota:1998:EAS

[1785] J. Kóta and Z. Perjés. Erratum: “All stationary vacuum metrics with shearing geodesic eigenrays”, [J. Math. Phys. 13, 1695 (1972)]. Journal of Mathematical Physics, 39(11):6241, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1049].

Chen:2012:CSR

[1786] Jianxin Chen, Zhengfeng Ji, Mary Beth Ruskai, Bei Zeng, and Duan-Lu Zhou. Comment on some results of Erdahl and the convex structure of reduced density matrices. Journal of Mathematical Physics, 53(7): 072203, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v53/i7/p072203_s1. See [1041].